A Bibliography of Publications in Centaurus: An International Journal of the History of

Science and its Cultural Aspects

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.24

Title word cross-reference

22^{∘} [588]. 2: n [566]. $49.95 [1574]. $85.00 [1577]. ^{c} [850, 1027, 558]. ^{th} [1087]. ^{v} [876]. _{2} [239, 1322]. _{4} [239]. ΔIAIT A [164]. ϵκϕανης [1031]. ΓEP ONT ΩN [164]. ≫ [517, 86, 112, 113, 741, 281, 411, 335]. ≪ [517, 86, 112, 113, 741, 281, 411, 335]. Φ [734, 801, 735]. π [703]. ′ [1215]. S [673]. \sqrt{2} [1025].

-derived [1558]. -Matrix [673].

0 [1542, 1645, 1666, 1685, 1640, 1576, 1603, 1578, 1648]. 0-8229-4547-9 [1578]. 0-8229-4556-8 [1603].

1 [1545, 1262, 1173, 1643, 1636, 1638, 1579, 1671, 1605]. 1/2 [580]. 10 [1408].

11 [335]. 12th [107]. 139780226352480 [1651]. 13th [52]. 14th [1002, 1006]. 14th-Century [1002]. 1550 [157]. 1599 [1008]. 15th [395, 32, 167]. 1650 [1039]. 1669 [636]. 1670 [1335]. 1670-1760 [1335]. 1685 [1013]. 16th [1478, 1477, 1011, 32, 421, 563, 148, 84, 826, 1054, 339, 200, 555]. 16th-century [1477]. 1700 [1279]. 1715 [1581]. 1750 [1639, 1748]. 1760 [1335]. 1764 [1277]. 1777 [55]. 1786 [182]. 17th [1478, 1570, 1571, 1638]. 17th-century [1570, 1571]. 18 [916]. 1800 [1585]. 1810 [1337]. 1820s [1755, 1055]. 1822 [438]. 1822-1972 [438]. 1830s [1362]. 1834 [1646]. 1840 [1600]. 1840s [1362, 1755]. 1850 [1021, 1713, 1338]. 1851-1916 [1654]. 1860s [1484]. 1870s [1363]. 1880’s [556, 559, 1774]. 1890 [1164]. 1890s [1210]. 18th [421, 1482, 1480, 433, 1351, 1223, 146, 194, 1224, 172, 108]. 18th-Century [433, 1351]. 19 [1747, 1559, 1547, 1549, 1561, 1556, 1553, 1546]. 1900 [442, 1154]. 1901 [1251]. 1905 [515]. 1910s [1774]. 1915 [503]. 1918 [1329]. 1920 [1583]. 1920s [1462]. 1927 [630]. 1935 [1648]. 1940 [1507, 1441, 1217, 1443]. 1942 [1401]. 1944 [799]. 1945 [1073]. 1949 [1132]. 1950 [52, 1750]. 1950s [1201]. 1953 [1098, 1099, 1095, 1097, 1100]. 1960 [1422, 1429]. 1960s [1737]. 1970 [1095, 1097, 1291]. 1970s [1434, 1433, 1432, 1435]. 1972 [1371, 438]. 1973 [1098, 1099, 1100]. 1980s [1206]. 19th [1532, 1482, 1534, 1121, 1390, 1396, 1077, 1555, 1735, 1590, 1011, 1604, 1223, 871, 1307, 1365, 41, 592, 212, 1090, 1245]. 19th-Century [1396, 1390, 1555, 1590, 1307, 1245]. 1st [488].

2 [1584, 1524, 1275, 1683, 1602, 1187, 1647, 1539, 1686, 1555]. 2. [511]. 2000 [1188, 1137]. 2003 [1115]. 2006 [1070]. 2007 [1071]. 2009 [1700]. 2010 [1220]. 2012 [1303]. 2014 [1330, 1403, 1373]. 2016 [1430]. 2020 [1739, 1620]. 2022 [1740]. 20th [1451, 1534, 1480, 1308, 1182, 1365, 1734, 1311, 1140]. 20th-century [1451]. 27 [124]. 28/26 [994]. 2nd [1642].

3 [1641, 1586, 1541, 1517, 1543, 1642, 1587]. 31 [1008]. 322 [572]. 33552 [754].

4 [1600, 1523, 1669, 1582, 1668, 1687, 1522, 1665]. 41004 [916]. 4th [86, 1171, 1185, 1196, 1220, 754].

5 [1683, 1540, 1644, 1604, 1544].

6 [1574, 1667, 1585, 1601, 1518, 1583, 1684]. 621 [343]. 6th [1330].

7 [1664, 1580, 1639, 1688, 1581].

8 [1607, 1650, 1577, 1603, 1637, 1606, 1521, 1689, 1646, 1670, 1649]. 86 [815].

9 [1519, 1520, 1578, 1573, 1575]. 921/2 [1684]. 96 [74]. 978 [1607, 1650, 1577, 1542, 1545, 1584, 1664, 1645, 1524, 1580, 1683, 1602, 1666, 1685, 1641, 1640, 1586, 1519, 1576, 1540, 1541, 1520, 1600, 1574, 1667, 1643, 1647, 1517, 1585, 1636, 1543, 1639, 1601, 1637, 1638, 1606, 1523, 1539, 1669, 1579, 1688, 1521, 1689, 1582, 1573, 1575, 1668, 1642, 1671, 1518, 1587, 1646, 1605, 1670, 1583, 1644, 1604, 1649, 1648, 1581, 1687, 1544, 1686, 1522, 1684, 1665]. 978-0-19-091436-3 [1642]. 978-0-19-093378-4 [1668]. 978-0-19-879725-8 [1577]. 978-0-19-882352-0 [1542]. 978-0-19-883795-4 [1687]. 978-0-19-884907-0 [1648]. 978-0-226-50929-7 [1581]. 978-0-226-55337-5 [1544]. 978-0-226-60599-9 [1573]. 978-0-226-70078-6 [1667]. 978-0-262-03726-6 [1518]. 978-0-262-53905-0 [1666]. 978-0-520-29698-5 [1586]. 978-0-520-30549-6 [1601]. 978-0-520-96951-3 [1586]. 978-0-520-97357-2 [1601]. 978-0-674-97201-8 [1670]. 978-0-691-17198-2 [1647]. 978-0-8153-7985-0 [1580]. 978-0-8229-4541-3 [1541]. 978-0-8229-4547-5 [1578]. 978-0-8229-4551-2 [1602]. 978-0-8229-4553-6 [1585]. 978-0-8229-4556-7 [1603]. 978-0-8229-4559-8 [1649]. 978-0-8229-4560-4 [1582]. 978-0-8229-4562-8 [1607]. 978-1-108-40600-0 [1685]. 978-1-108-41830-0 [1576]. 978-1-108-48243-1 [1643]. 978-1-351-21482-7 [1580]. 978-1-4214-2177-3 [1517]. 978-1-4214-3846-7 [1639]. 978-1-4744-4578-8 [1646]. 978-1-55481-387-2 [1584]. 978-1-55753-837-6 [1574]. 978-1-78920-113-0 [1645]. 978-1-899209-24-8 [1650]. 978-2-13-080065-1 [1579]. 978-2-271-11500-3 [1543]. 978-2-271-12300-8 [1606]. 978-2-503-57445-5 [1540]. 978-2-503-58191-0 [1640]. 978-2-503-58603-8 [1689]. 978-2-7132-2726-4 [1523]. 978-3-030-01637-1 [1638]. 978-3-030-04175-5 [1683]. 978-3-030-20598-0 [1664]. 978-3-030-20599-7 [1664]. 978-3-319-54654-4 [1522]. 978-3-319-93652-9 [1519]. 978-3-319-95251-2 [1524]. 978-3-8047-3797-6 [1583]. 978-606-543-951-1 [1545]. 978-84-17945-01-5 [1644]. 978-84-9134-439-1 [1605]. 978-84-9168-164-9 [1575]. 978-87-7304-417-9 [1520]. 978-90-04-24336-1 [1636]. 978-90-04-28481-4 [1665]. 978-90-04-32549-4 [1600]. 978-90-04-33629-2 [1686]. 978-90-04-35263-6 [1587]. 978-90-04-35264-3 [1587]. 978-90-04-36064-8 [1521]. 978-90-04-36574-2 [1539]. 978-90-04-38143-8 [1637]. 978-90-04-38866-6 [1684]. 978-90-04-40056-6 [1636]. 978-90-6299-468-7 [1688]. 978-963-693-300-5 [1604]. 978-981-12-0628-3 [1641]. 978-981-12-1848-4 [1669]. 978-989-8363-41-1 [1671]. 9th [1002]. 9th-Century [1002].

ādh [1027].

începutul [1545].

a-t-il [718]. A. [552]. A.D. [287]. Aaboe [1071]. Aage [648]. Abū [349, 558]. Abacus [449]. Abandoned [1091]. abbess [188]. Abel [1331, 1169]. Aberration [591]. Abhandlung [711]. Abingdon [1507]. Above [1366]. Abraham [1011, 527]. Absolute [633, 1169]. Absorption [831]. Abstract [218, 62]. Äbtissin [188]. Abu [523]. Abwegen [830]. Academe [1141]. Academia [1419]. Academic [1438, 1362, 1484]. Académie [1667]. Academy [1532, 488, 1455, 1484, 1305]. Access [1691]. accident [1630]. Accommodating [12]. According [224, 274, 964, 267, 717, 375]. Accoucheur [1527]. Account [1160, 330, 1134]. Accountability [1754]. Accuracies [640]. accuracy [1535]. Achaemenid [317]. Achievement [1294]. Achievements [372]. Achromatic [756]. Achtzehnten [27]. Ackermann [1637, 1398, 1726]. Acolytes [1338]. Acquaintance [479, 38]. Acronycal [1014]. Acronychal [310]. Act [1368, 1820]. Acta [1583]. Action [1403, 1429, 806]. Activity [554]. Actor [859]. Aczel [1218]. AD [1074, 1166, 74, 442]. Adaptation [630, 1336]. Addenda [360]. Adding [1050]. Address [1403, 52]. Adelene [1355]. Adelheid [1340]. Adjustable [464]. Adler [1670]. Adoption [1336]. Adrian [1073, 1518]. Adrien [1337]. Adrienne [1399]. Advantage [1153]. adventure [1671]. Advertising [1269]. Advisor [1155]. Aegyptische [354]. Aesthetics [1712, 1481]. Aetiologies [1754]. Aetna [554]. Affair [1299, 1823]. Affluence [1065]. Aflah [428, 465]. Africa [1675, 1239, 499]. African [202, 499, 1756, 1612]. After [1152, 1715, 1314, 534, 1297, 1821, 1512]. Aftermath [1132, 1245]. against [1559, 623, 1107, 1387, 568]. Age [1422, 1746, 1392, 1269, 1278, 173, 1291, 1604, 1194, 463, 1213, 1469, 1786, 12, 1830, 1457, 1808, 554, 1383, 1725, 1216, 1218]. Age-old [173]. Agenda [1201]. Agendas [1197, 1447]. Agents [461]. Ages [1253, 81, 64, 150, 159, 158, 291, 24, 427]. Aggregating [1264]. Agricultural [1365]. Agustí [1643, 1081]. Agustín [1507]. ägyptischen [638]. Ahargana [45]. Ahmad [1215]. AIDS [1393, 1701, 1700, 1801, 449]. Air [105, 610, 165]. Aire [589]. Akademien [1279]. Al [459, 816, 1215, 1027, 632, 384, 768, 380, 811, 816, 996, 1537, 1027, 237, 325, 728, 154, 155, 374, 73, 644, 896, 1215, 776, 224, 821, 533, 558, 409, 649, 384, 736, 768, 493, 868, 558]. Al-Andalus [1027]. al-Bīrūnī [644]. al-Bag̀dâdî [768]. Al-Battānī [374]. al-Battani [409]. Al-Bitruji [154, 155]. al-Dīn [632]. al-Dīn-Tūsī [816]. al-Din [384]. al-Hāsīb [649]. Al-Haitham [73]. Al-Hajjaj [1537]. al-Hassan [1215]. al-Haytham [736, 493]. al-jabr [380]. al-Jayyān [1027]. al-Kāmil [868]. al-Kāshī [776]. al-Kūh [996]. Al-Kūhī [816]. al-Kammâd [821]. al-Khāzin [558]. Al-Khwārizmī [237]. Al-Khwarizmi [325]. al-Kimya [1215]. al-Latîf [768]. al-Misāha [811]. al-Munajjūn [533]. al-Nayrızı [896]. al-Safāa’ih [558]. Al-Samaw’al [996]. al-Tadhkirah [384]. al-Tusı [384]. al-Zarqāllu [224]. Al-Zarqālluh [728]. Alan [1636]. Albert [1241, 74, 1050, 608]. Alberti [1126, 74]. Alberto [1227]. Albrecht [55, 92, 55]. alchemical [618]. alchemistische [618]. Alchemy [1215, 439, 1640]. Alex [1747, 1544]. Alexander [1749, 1249, 1541, 1175, 1060, 1064, 165]. Alexandria [1211]. Alfonsine [998, 1678, 1493, 527]. Alfred [1382, 1790, 508]. Algebra [1168, 855, 595, 815, 1038, 876, 514]. algebraic [355]. algebraischer [355]. algorismi [876]. Algorithm [1050]. Alhacen [995]. Alhaseni [311]. Alike [805]. Allan [1381]. Alleged [37, 6]. Allegory [1389]. Allies [983, 1284]. Almagest [457, 992, 880, 635, 576]. Almanac [262, 1685]. Almanach [1011, 527]. almost [1791, 1370]. along [805]. Alper [1396]. Altar [966]. Altchinesischer [137]. Altdeutsche [291]. alten [113]. Altenstein [1065]. alter [254]. Alternative [1144]. Alternatives [817, 1747]. Altertum [33]. Altes [78]. ältesten [618]. Altindischen [139, 220]. altindischer [54, 181]. Altmedizin [12]. Álvarez [1586]. am [893, 555]. Amakin [459]. Amanuenses [1179]. Amateur [172]. Amateurs [1311]. Amazonia [1420]. Ameghino [1414]. Amelia [1602]. America [1429, 1603, 1164, 1371, 1197, 1560, 1307, 1593, 1815]. American [1746, 1751, 1423, 1418, 1747, 1033, 1774, 3, 1390, 1470, 1628, 1309, 1321, 1681, 1707, 1286, 246, 1202]. Amir [1218]. Among [62, 1141, 1638, 148]. Amontons [350]. Amputation [181]. Amputations [181]. Amputations-Lehren [181]. Amra [850]. Amsterdam [1021, 1688, 16, 102, 1024]. Ana [1671]. Analemma [1035]. analemmatic [506]. analemmatique [506]. Analogy [1173, 1187]. Analyse [419]. Analysis [795, 1033, 1013, 1121, 105, 563, 630, 1169, 915, 505, 873, 523, 68, 880, 419, 1464]. Anamnesis [1759]. Anaphoric [72]. Anaritius [896]. Anatole [1543]. Anatomen [148]. anatomical [47]. anatomically [53]. Anatomie [137, 24]. Anatomies [1773, 1519]. anatomisch [53]. anatomische [47]. anatomischen [46]. anatomists [148]. Anatomy [567, 1278, 46, 41, 24]. anciens [25]. Ancient [1769, 833, 525, 1683, 320, 60, 603, 966, 94, 85, 120, 145, 135, 640, 575, 1724, 1384, 784, 106, 782, 181, 1818, 448, 1173, 1187, 579, 1111, 113, 25, 139, 54, 220, 164, 461]. and/or [1222]. Andalus [1027]. Andersen [1126, 1167]. Anderson [1645]. Andes [1775]. Andrault [1521]. André [1238]. Andrea [1714]. Andreas [1651, 1686]. Andrei [1545]. Andrew [1302, 1296, 1278, 1601]. Androids [1340]. Anfängen [119]. angehängt [400]. Angel [1575]. Angela [1150]. Angeles [1021]. Angelsächsischen [47]. Angle [1168]. Anglo [47]. Anglo-Saxon [47]. Aníbal [175]. Animal [1734, 1735, 425]. Animals [1731, 210]. animate [1563]. Ann [1639]. Anne [1584, 1685]. Annette [1641, 1369]. Annie [1261]. Anniversary [1283]. anno [1069]. annotations [1686]. Announcement [1171, 1185, 1196, 1220, 1229]. annual [52]. Anomaly [990, 1005]. Anonymous [343]. Anpassung [1133]. Ansichten [540]. Anteaters [1730]. Antecedents [350]. Anthony [1632]. Anthropocene [1647]. Anthropological [1364]. Anthropology [1364, 1604, 1483]. anti [1492]. anti-nuclear [1492]. Antike [164]. Antikythera [483, 664]. Antiquarianism [1194]. Antique [500]. Antiquities [1775]. Antiquity [1325, 330, 461, 484, 1640, 1688, 33]. Anton [1313]. Antonio [555]. Antony [1234, 1670]. Anwendungen [560]. Anxious [1602]. Any [1775]. Apatite [452]. Apatits [452]. apertures [1570]. Aphrodisias [330]. Apogee [267]. apogees [1678]. Apollonius [751, 603, 479]. Apostate [661]. Apostle [661]. Appearance [968, 1389]. Appendix [583, 527]. Appetite [1750]. Appleby [1346]. Application [1227, 1039, 646, 652, 324]. applications [560]. Applied [1027]. Appraisal [515, 1743]. Approach [1403, 676, 349, 1703, 1679, 55]. Approaches [697]. Appropriation [1332, 1342, 1459]. approximation [1025]. Approximations [308, 4]. April [1071]. Aquarama [1737]. Aquinas [1519]. Arab [717]. Arabic [1215, 896, 776, 217, 343, 83, 1011, 114, 107]. arabischen [717]. Arabischer [107]. Arabus [694]. Arbeit [1279, 335]. Arbeiten [102]. Arbeitsweise [55]. Arboretum [1486]. arc [355, 1017]. Archaeological [320, 1768, 618]. archaeology [1570]. Archimedean [417, 577]. Archimedes [820, 816, 168, 127, 261, 2, 537, 726, 1016]. Architectural [1803, 1441]. archival [1453]. Archives [1104, 1114, 16, 1107, 1111, 1094, 1101, 1286, 1289, 1096, 1287, 1093]. Archytas [391, 391]. Arclength [336]. Area [577, 1757]. Argentina [1414]. Argon [601]. Argument [401]. Argyre [532, 532]. aria [1755]. Arianrhod [1352]. Aristotele [86]. Aristotelian [1238, 494]. Aristotelischen [494]. Aristotle [1685, 1519, 1074, 1250, 341, 238, 1122, 964, 86]. Aristyllus [640]. Arithmetic [514, 493, 1464, 1006, 629, 455]. arithmétique [493, 629]. arithmétiques [1006]. Armillary [1409]. Army [1317]. Arne [1137]. Arons [867, 867]. Arribas [1726]. arrows [1664]. Arsenal [1198]. Art [1126, 1128, 1170, 1194, 1339, 1822, 8, 1586, 53, 1687, 14, 1440, 1333]. art-historical [8]. Arthur [1401]. Article [889, 890, 1265, 881, 1263, 1266, 1264]. articles [1458]. Artifacts [1769, 1625]. Artificial [1570, 1125]. Artisanal [1472, 1477, 1476, 1505]. Artisans [1340]. Artistic [1170]. Arts [1540, 1481, 1813, 1295]. Aryan [1439]. Arzneibuch [376]. Ärzte [220, 14, 1059, 54, 180]. Arztes [84]. Ärzteschulen [12]. Ärztin [188]. Ascending [744]. Ascensions [198, 1005]. Ascham [1632]. ascribed [896]. Asger [1071]. Ash [1329]. Ashley [1194]. Asia [1336, 1197]. Asian [921, 1556]. Asif [1296]. Aspects [369, 1027]. Assamawal [455]. Assembly [1220]. Assessment [491, 552, 293]. assessments [193]. assimilation [107]. Assimilations [107]. Assimilations-Zentren [107]. assistance [1610]. assistants [1179]. Association [1078]. Assyrian [326]. Astatic [852]. Astral [1405]. Astrolabe [764, 72, 551, 394]. astrolabes [429, 1726]. Astrolabium [394, 1411]. Astrological [1718, 1123, 583, 722, 1027, 582, 868, 1637]. astrologie [999]. Astrology [1376, 1394, 1076, 584, 1347, 1339, 1782, 1799, 999, 1148]. astronomica [1575]. Astronomical [318, 1153, 801, 887, 583, 1013, 605, 1497, 582, 632, 1002, 384, 465, 736, 1410, 739, 1035, 178, 197, 868, 720, 124, 693, 289, 1529, 167]. astronomie [1689, 231, 335]. astronomische [289]. astronomischen [167]. Astronomy [223, 573, 525, 1161, 1032, 716, 851, 1284, 1406, 1003, 579, 727, 1111, 1014, 349, 738, 825, 834, 846, 879, 152, 1407, 428, 1005, 575, 281, 100, 1617, 1001, 543, 1259, 101, 1678, 1636, 1498, 1529, 231, 335, 498]. Athanasius [1043]. Atlantic [1607, 1772, 1472]. Atlas [432]. atmosphere [1500]. atmospheres [1595]. Atmospheric [478, 282, 995]. Atom [965, 1489, 1492]. Atomic [1392, 1609, 1611, 1786]. atomism [699]. atomistic [875]. Atomistik [699]. atoms [1627]. Attack [1439]. Attempt [338]. Attempting [1290, 1613]. Attempts [861]. Attraction [443]. attribuée [5]. Attributed [401, 896, 5]. Aubin [1689]. audience [1663]. Aufgabe [119, 357]. aufgefundenes [92]. Aufnahme [12]. Aufrecht [1131]. August [9, 110, 594]. Augustin [24, 711, 711]. Augustine [1519, 24]. ausgehenden [291, 108]. ausseren [46]. Australian [1728]. Austria [1574]. Auswirkungen [847]. Auszüge [167]. Author [1062, 1463]. Authority [1692, 1552]. Authors [998]. Authorship [1544]. Autobiography [1213]. Automaton [1271]. Autonomie [1133]. autonomy [1552]. autorité [1523]. autre [1606]. Aux [244]. Auxiliaries [94]. avant [1240]. Avempace [238]. Avian [227]. Awakening [511]. Axiom [2]. axiomatic [552]. Axioms [417, 467]. Axolotl [1731]. az-Zarqal [333]. Azarquiel [262]. Azimuth [764].

B [511, 1214, 1081, 281, 484, 498, 224, 318, 794]. B.C [578]. B.M. [754]. Bīrūnī [644]. Bīrünī [118]. Baby [1310]. Babylon [317]. Babylonian [1683, 335, 93, 126, 223, 573, 574, 734, 733, 801, 872, 1032, 716, 754, 990, 207, 887, 1014, 125, 888, 738, 886, 708, 281, 721, 989, 331, 572, 1038, 231]. Babylonischen [125, 231, 335]. Bacciagaluppi [1234]. Back [1373, 1736, 1504]. Background [515]. Bacon [23, 1565, 1562, 1563]. Baconian [1566]. bacteriology [48]. Badash [1206]. Bädekers [545]. Badge [1222]. Baedeker [545]. Baer [1438, 1483]. Baert [375]. Bag̀dâdî [768]. Baghdad [1002]. Bahnen [134]. Bahr [1140]. Bakker [1539]. Bakteriologie [48]. Ball [1402]. Ballester [487]. balloons [1500]. Balsamital [26, 26]. Balthazar [1481]. Balthazar-Georges [1481]. Baltic [843]. Baltimore [1517, 1639]. banalization [1488]. Bang [1705, 822]. Banishing [1489]. Banks [1762, 1527]. Bantu [202]. Baptiste [1439, 1068]. Barany [1716]. barber [157]. Barbers [1719]. Barbiers [157]. Barcelona [1171, 1185, 1196, 1220, 1575, 1770, 1737, 1144, 1415]. Barnabas [1147]. Barnacle [158]. Baron [423, 166]. Baroque [1796, 1343]. barrels [219]. Barrow [233, 827]. Bart [1333]. Bartholin [538, 156]. Bartholinerzeit [156]. Bartolomeo [6, 6]. Based [879, 1310, 1632]. Basel [1638]. Basis [576, 256]. Bateson [347]. Battānī [374]. Battani [409]. Battery [881]. Battista [1564, 1565, 826, 1562]. battle [1559]. Battles [1666]. BC [1074, 1039, 442, 483]. Be [1043, 1558, 1537, 220]. Beati [1152]. Became [1230]. Become [1773]. Becquerel [723]. Bedeutung [146, 562]. Been [1157]. Beer [485, 485]. Before [895, 1240, 967, 1646, 1581]. Beginn [212]. beginnenden [41]. Beginning [690, 1193, 212]. Beginnings [1083, 1455, 119, 1462]. Behandlung [194]. Being [1439, 439, 1533]. beings [1563]. Beintker [88, 88]. Beintker-Kahlenberg [88, 88]. Beitrag [41, 112, 545, 108, 159]. Beiträge [291]. bekannten [38]. Belege [254]. Belgian [1722, 1486]. Belgique [35]. Belgium [1640, 1540, 1664, 1722, 1245, 35]. Bell [1375]. Belt [32]. Bemerkungen [86, 113, 383, 182]. ben [363, 410, 680, 1005]. Benedetti [826]. Bengt [1066].

Bennett [1193, 1581]. Benoît [1495]. Bensaude [1125]. Bensaude-Vincent [1125]. Beobachtungen [704]. Beobachtungsfehler [128]. Berg [1583]. Berghahn [1645]. Berghe [1149]. Bergmans [518, 518]. Bericht [717]. berichteten [704]. Berichtigung [335]. Berlin [1356, 1279, 1800]. Bern [74]. Bernadette [1125]. Bernal [1105, 1106]. Bernard [1237, 1177, 1425, 987, 999]. Bernardino [1587]. Bernat [851]. Bernhard [1213]. Bernoulli [1128, 517, 671]. Bernoullis [671]. Berr [99]. Berthelot [1516]. Bertoloni [1578, 1085]. Bertomeu [1081]. Berzelius [226]. Beta [251]. Betim [1227]. Betrachtung [338]. Better [1369, 1829]. Between [1459, 1365, 1121, 841, 1319, 1717, 1232, 1657, 871, 896, 845, 818, 650, 1173, 1187, 1258, 1633, 1419, 264, 1676, 1556, 301, 587, 1024, 1389, 1336, 1035, 44, 1503, 293]. Beurteilungen [193]. Bewahrung [254]. Bewegungen [231, 335]. Bewegungsempfindungen [419]. Bewertung [247]. Bewertungen [234]. Beyond [1210, 1357, 1725, 1566, 1305, 1819]. Beziehungen [1279]. Bibliographie [171]. Bibliography [487, 648, 549, 171, 513]. Bibliothek [594, 157]. Bibliothèque [215]. Bid [1317]. Bidstrup [790]. Big [1722, 1317, 822]. bigamist [1483]. Bildgedicht [618]. Bill [1646]. Bills [1552]. Bingen [188, 188]. Bioarchaeological [1703]. Biocca [1420]. Biografi [1190]. Biographer [1655]. biographical [1567, 39]. Biographisches [39]. Biological [1207, 1178, 1282]. Biology [1746, 1421, 1420, 1798, 1518, 1354, 1235, 1249]. Biosocial [1422, 1703]. Bird [1762]. birefringent [524]. Birkhäuser [1638]. Birth [1714, 1732, 433]. Birunis [459]. bis [1075, 1583, 44, 518]. Bisectable [733]. bisher [618]. bison [1485]. Bitruji [154, 155]. Bittel [1649]. Black [1747, 1586, 1746]. Blackness [1278]. Blackwell [1425]. Blair [1639]. Blame [1754]. Blattes [78]. Bleichmar [1326]. Bleisalben [194]. Blomstrand [519]. blood [234]. blossoms [1480]. Blue [1314]. Blut [234]. Blut-Bewertungen [234]. BM [916, 916]. BnF [1537]. Boantza [1353]. Bock [1664]. Bodies [791, 649, 387, 1751]. Bodleian [343]. Bodoing [1135]. Body [1386, 46, 109, 53, 1665]. Boelmans [586]. Boerhaave [1295, 1813]. Boethius [598]. bogengleicher [355]. bogy [1489]. Bohm [1794, 1184, 1794]. Bohr [600, 967]. Bois [1155, 1372, 1792, 1158]. Bois-Reymond [1155, 1372, 1792, 1158]. Bokkus [282]. Bologna [1620]. Boltzmann [250, 250]. Bomb [1314]. bone [1518, 1483]. Bonea [1602]. Boner [1347]. Bonfils [1006, 334]. Book [1607, 1749, 1650, 1442, 141, 696, 750, 923, 949, 950, 951, 952, 953, 954, 955, 956, 957, 958, 959, 960, 961, 962, 963, 924, 925, 926, 927, 928, 929, 930, 931, 932, 933, 934, 935, 936, 937, 938, 939, 940, 941, 942, 943, 944, 945, 946, 947, 948, 900, 922, 1711, 1577, 1542, 1545, 1728, 1584, 1664, 1645, 1524, 1580, 1683, 1602, 1666, 1685, 1444, 1507, 1641, 1712, 1640, 1586, 1519, 1576, 1540, 1541, 1520, 342, 1600, 1574, 1748, 1751, 1667, 1643, 1647, 1517, 1603, 1585, 1636]. Book [1543, 1639, 1497, 1601, 1637, 1638, 1606, 603, 1523, 1539, 649, 1495, 1578, 1496, 1669, 1579, 1688, 1521, 1440, 1724, 1567, 1713, 1689, 1725, 458, 1582, 1573, 1714, 1506, 1575, 1555, 1715, 1668, 1642, 1441, 1671, 1518, 1587, 1646, 1605, 1726, 1670, 1583, 1644, 1604, 1649, 1648, 1651, 1750, 1581, 505, 1687, 1544, 1686, 1522, 1684, 1665, 1443, 1716, 1727, 86, 248, 816]. Books [11, 22, 30, 43, 58, 187, 191, 206, 214, 230, 243, 260, 296, 306, 316, 362, 477, 492, 509, 522, 542, 557, 613, 627, 643, 654, 670, 682, 696, 715, 725, 750, 797, 810, 824, 839, 866, 870, 891, 878, 898, 923, 949, 950, 951, 952, 953, 954, 955, 956, 957, 958, 959, 960, 961, 962, 963, 924, 925, 926, 927, 928, 929, 930, 931, 932, 933, 934, 935, 936, 937, 938, 939, 940, 941, 942, 943, 944, 945, 946, 947, 948, 980, 985]. Books [1015, 1052, 1776, 32, 651, 46]. Boost [1314]. borders [1819, 1357]. Borel [1246]. born [684, 1481, 920]. Borrowed [683]. Boscovich [591]. Bosom [1554]. Boston [1600, 1684]. Botanical [79, 1326, 27]. Botanik [421]. Botanische [27]. botany [421, 1482]. Boubée [1463]. boundaries [1585, 871]. Bowdich [1350]. Bowen [1636, 1115]. Bowler [1178, 1205]. Box [1746]. Boyle [1519, 472, 1807]. Bradwardine [31]. Braekman [447]. Brahe [360, 982, 729, 1008, 275, 358]. Brain [1302, 41, 1812, 33]. brandy [150]. Branntweintraktat [150]. Bras [1543]. Bravai [544]. Bravais [544]. Brazil [1611, 1699]. Brazilian [1463, 1771]. Breeding [546, 1736]. Brentjes [1540]. Brepols [1640, 1540, 1689]. Breslau [1567]. Breuer [419, 419]. Bricks [966]. Bridge [803]. Bridge-Builder [803]. Bridges [1445]. Bridging [1045]. Brief [257, 271, 294, 304, 396, 413, 366, 619, 706, 620, 731, 635, 636, 782, 1652]. Briefe [658, 419, 156, 44, 250]. Briefs [1522]. Briefwechsel [699]. Bright [1031, 480]. Brigitte [1641, 1134, 1369]. Brill [1600, 1636, 1637, 1539, 1521, 1587, 1686, 1684, 1665]. Brilliant [1155]. Britain [1251, 1429, 1489, 1602, 983, 1388, 1205, 1282]. britanniques [1237]. British [1237, 1576, 426, 1233, 663, 1633, 370, 388, 114]. broad [41]. Broadview [1584]. Brock [1442]. Broecke [1148]. Bromides [1747]. Brood [227]. Brooks [427]. Brotons [1605]. Brouillon [751]. Brownian [780]. Bruce [1241]. Brummelen [1174]. Bruun [1313]. Büchern [46]. Buches [86]. Buckland [1355]. Budapest [1604]. Buffon [741]. Builder [803]. Building [1198, 1445, 1522]. Bulgaria [286]. Bulletin [1681]. Buoys [1274]. Bureaucracies [1089]. Bureaucracy [1362, 1090]. Bureaucrats [1739]. Burgerbibliothek [74]. Burgermeister [102]. Buried [1262, 1275]. Burnett [1726]. Burnt [366]. Bursting [1226]. Burzenlande [182, 182]. business [1533]. Butchers [1719]. Butlerian [1053]. Büttner [1744]. Bylica [1076]. Byzantine [1743, 1743].

C [1295, 1636, 472, 515, 608, 426, 1335, 488, 427, 423, 436, 499, 847, 1073, 44, 1232, 1775, 1753, 1754, 1011]. C. [614, 44, 279]. Côme [5]. CA [1586, 1601, 1589, 483, 1039, 1595, 1567, 1245, 1735]. Cadran [364, 450, 551, 506]. cadrans [675, 302]. Caesarean [56]. Cairo [605]. Calculate [444]. Calculation [1720, 685, 1674]. Calculations [921, 1464]. Calculus [836, 745, 106, 856]. Calendar [483, 727, 1684]. Calendarium [1567]. Calendars [1123]. Calendric [1407]. calendrical [1536]. California [1021, 1586, 1601]. Callatay [432]. Called [1439, 1002, 385]. Callippic [727].

Callippopedes [854]. Callippus [854]. Caloric [211]. Calvin [1519]. Cambridge [1666, 1685, 1576, 1643, 1495, 1178, 1518, 1670]. Campanus [346]. Camprubi [1747, 1377]. Can [1695, 220]. Canada [1596]. Canales [1204]. Cancer [1746, 84, 194, 212, 277, 1681]. Canham [1669]. Capillary [443, 659]. Capital [321, 782]. Carathéodory [799, 799]. Carathéodory-Nachfolge [799]. Carathéodory-Succession [799]. Care [1763, 1734]. Career [790, 1370, 1321, 1313, 1791]. Carl [41, 41, 1012]. Carla [1649]. Carne [1671]. Carnot [444]. Carola [1180]. Carolus [117]. Carpel [174]. Carpentry [1765]. carrée [532]. Carson [1392]. Carsten [1080]. Cartesian [78]. Cartesianism [1714]. Cartography [1618, 1475]. Carus [41, 41]. Carving [1769]. Case [889, 831, 1406, 1159, 211, 1141, 567, 1151, 174, 865, 1186, 1092, 1478, 1612, 1500, 1615, 1494, 1501, 1616]. Case-History [174]. Case-Study [865]. cases [1150]. Cassini [1689, 1257]. Cassiodorus [604]. cassowaries [1733]. Casting [722]. Catalana [1171]. Catalans [1220]. Catalog [1720]. Catalogue [480]. Catarata [1506]. Catching [1820, 1368]. Cátedra [1496]. Categories [1003]. Category [384]. Catherine [1603]. Catholic [1741]. Cathryn [1392]. Catoptrics [683]. Cauchy [423]. Causality [630]. causation [1633]. Cause [1746, 65, 1094, 1261, 128]. Cavendish [1584, 1807]. CE [1410, 1684]. Cecil [19]. Cedillo [1575]. Ceglia [1665]. Celebrates [1283]. Celebrating [1460]. Celestial [1045, 988, 1300, 326, 1826]. celestiales [1575]. cellulaire [741]. cellular [741]. CEN [1722]. census [1090, 1686]. Centaurus [1063, 1093, 335, 1498, 1690, 1538, 1672, 1691]. centenary [1516]. Center [1722, 107]. Central [689, 447, 1229, 1660]. Centralized [859]. Centre [417]. Centripetal [769, 805]. Centuries [736, 1223, 1478, 1534, 421, 84, 108, 1604]. Century [1153, 754, 1191, 217, 1121, 1027, 1205, 1308, 1239, 1396, 1282, 1077, 1355, 172, 433, 1414, 1351, 1372, 1327, 1087, 1140, 1532, 1451, 1602, 1784, 1576, 1477, 1520, 1482, 1570, 1768, 1643, 1390, 1571, 1638, 1480, 1182, 1555, 1735, 1590, 1828, 1769, 842, 79, 634, 146, 871, 70, 1104, 1307, 395, 1776, 787, 144, 1119, 1365, 41, 1792, 563, 1002, 371, 592, 148, 194, 212, 1090, 1006, 1224, 827, 1734, 1245, 1311, 1139, 826, 894, 1696, 1091, 32, 1820, 1054, 1544, 27, 388, 339, 200, 5, 555, 107, 167, 1577, 1651, 1237, 1325, 1084, 1085]. Century [1368]. CERN [1614]. Certainty [1203, 1192]. César [1689]. Ch [1190]. Châtelet [1352, 1779]. Chadarevian [1746]. Chalduns [240]. Challenges [1118, 1744]. Challenging [1744]. Cham [1683, 1519, 1522]. chamber [1483]. Chang [1322]. Change [1160, 1161, 1163, 1159, 1162, 1663, 501]. Changed [1009, 1136]. Changes [201, 729, 1604]. Changing [1676, 1742]. Chapman [1381]. chapter [86, 876]. character [139]. Charakter [139]. Charlene [1586]. Charles [1229, 1726, 1510, 812]. Charts [1475]. Chemical [842, 1295, 681, 135, 1077, 660, 420, 398, 1353, 1513, 1212, 1509, 871, 1781, 1571, 1511, 1813, 1668, 1474, 165]. Chemicals [135]. Chemie [871, 544]. chemische [871, 165]. Chemist [1419, 840, 1216, 668]. Chemistries [1723]. Chemistry [1442, 1294, 1783, 844, 845, 1295, 917, 730, 841, 843, 1378, 286, 871, 1793, 1643, 129, 895, 1813, 1503, 544, 1215, 1080, 1081]. Chemists [840, 445, 1668]. Chen [1343]. Chen-Morris [1343]. chez [523]. Ch’i [665]. Ch’i-yao [665]. Chicago [1667, 1497, 1573, 1651, 1581, 1544]. Chicherin [1515]. Chikungunya [1699]. Children [1491]. Chili [1671]. Chilton [1580]. China [446, 833, 1697, 1762, 1568, 1410, 63, 1610, 1464, 470]. Chinese [900, 1556, 1550, 1407]. Chios [1019]. Choice [1311, 1235]. Cholera [256, 1555]. Chord [392]. Chr [699]. Chr. [312]. Christensen [1190]. Christiaan [52]. Christian [1184, 1190, 27, 339, 519, 599, 699]. Christianity [1261]. Christine [1649, 1683]. Christoph [1280, 1334, 1539]. Christopher [1359]. Chromosomes [347, 1746]. Chrysanthemum [26]. Chrysokokkes [334]. chüeh [665]. church [209, 24, 14]. Churches [584]. chymistry [1667]. Ciència [1171, 1644]. científicas [1506]. Circle [988, 740, 1542, 671, 1638, 1044]. Circles [764]. Circonsion [25]. Circular [1229, 802]. Circulation [1280, 1351, 1171, 1185, 1196, 1680]. circulature [1025]. Circumcision [25]. Circumference [740]. Cities [1035]. city [1507]. Civil [446]. Civilisations [60]. Civilization [525, 1039, 470, 446, 1648]. Ckep [1021]. claims [1609]. Clarifying [893]. Clark [1301]. Classic [424, 1468]. Classical [1203, 1636, 685, 554]. Classification [1382, 1790]. Classifications [1439]. Claudia [1539]. Claus [1539]. Clausen [1017, 1017]. claw [1016]. Clay [145, 462]. Clear [1314]. Cleomedes [991, 1114]. cleric [1483]. Clifford [19]. Climate [1444, 1809, 1594, 427]. clivaxo [209, 208]. Clock [72, 1023]. Clocks [886, 486]. Closed [829]. Closure [1709]. Club [1421]. Cluj [1545]. Cluj-Napoca [1545]. Clüver [745]. CNRS [1543, 1606]. Coal [1164]. Coca [1209]. Coca-Cola [1209]. cochineal [1473]. Codifications [1673]. Coen [1345]. cognition [1539]. Cognitive [1163]. Cogwheels [1255]. col [1640]. Cola [1209]. Cold [747, 1444, 1370, 1299, 1317, 1316, 1315, 1078, 1522, 1202, 1823, 1420, 1791, 1794, 1500, 1312, 1615, 1809, 1313, 1613]. Colding [170]. Cole [169]. Collaboration [780, 1556]. Collaborative [1369, 1829]. Collected [52]. Collecting [1480, 1306, 1775]. Collection [132, 1768, 395, 1668]. Collections [1770, 1763, 1773, 1776, 1482, 1758, 1762, 1771, 1772, 1759, 12, 1481]. College [840, 490]. Colleoni [6]. Collider [1814, 1399]. collyre [5]. Collyria [461, 461]. Colonial [1319, 1631, 1473]. Colonialism [1239, 1675]. Color [1293, 1348, 1807]. Colour [564]. colours [1514]. Column [318, 734, 801, 735]. com [1671]. combination [667]. combinatorial [1687]. Combustion [104, 252]. Comet [563, 1176]. Cometary [1047]. Commemorating [31]. commemoration [1516]. Commemorative [608, 1209]. Comment [1021, 342]. Commentaries [456]. Commentariolus [527]. Commentary [1098, 1113, 1099, 1105, 1109, 1108, 1112, 1106, 1102, 1110, 1103, 1115, 1095, 1097, 1100, 1687, 459]. comments [896]. Commercial [1530, 1808, 1657, 1383]. Commitment [1144]. Committee [235, 1274, 1598]. committees [1534]. Commoditization [1269]. Common [1464]. Communicating [1330]. Communication [1403, 706, 731]. Communications [257, 271, 294, 304, 396, 413, 366, 619, 620, 635, 636].

Communities [1715, 1323]. Community [1362, 1077, 1073, 1652, 1522]. Como [5]. Companion [1425, 1743, 1423]. Companions [1636, 1425]. Company [1533]. Comparative [1428]. Comparing [1197]. Comparison [336, 231, 335]. Compendium [651]. Competing [983]. Competition [406, 1556, 1202]. completion [52]. Complex [1408]. Complexity [1721]. Components [461]. Compound [1724, 1600]. Computation [1027, 45]. Computational [1408, 1407]. Computations [605]. Computer [483, 269]. Computus [604]. Concept [1118, 1122, 199, 723, 1101, 1169, 130, 116, 540, 1177]. Conception [759, 130, 746]. Concepts [1163, 1162, 430, 129]. Conceptual [1159, 1578]. Conceptualising [1755]. Concerning [710, 738, 1038]. Conchoid [324]. conclusion [1220]. conditions [8]. Conduct [425]. Conductors [592]. Conference [1070, 1171, 1185, 1196, 1220, 1330, 1304, 1620, 1234]. Confidence [71]. Confirmation [691]. Conflict [567]. Confusion [1410]. congress [20, 1375]. Congresses [1320]. Conic [479, 4]. Conical [365]. Conics [751, 603]. Conies [609]. Conjecturing [1128]. Conjunction [578]. Conjunctions [1258]. connaissance [1240]. Connection [752]. Conquest [1775]. consciousness [1589]. consequence [263, 550, 263]. consequences [550]. Conservation [451, 170, 402, 301, 1734]. Conservative [347, 1068]. Consideration [389]. Considerations [1277]. considered [220]. Conspicuous [1031]. Constitution [139]. Constructing [1221, 859]. Construction [318, 734, 1281, 1412, 603, 726, 1509, 1564, 1630, 98, 1811, 302, 355]. Constructions [689, 537, 966]. Consultants [1722]. Consulting [1164]. Consumption [881]. Contagions [1747]. Contagious [1746]. contemporaries [182]. Contemporary [1360, 1186, 1608, 1491, 293, 1696]. Content [752, 1066]. Contestations [1716]. Contested [1291, 1830]. Context [611, 919, 1754, 1547, 1617]. contexts [1631, 1636, 1637]. Conti [1493]. Continued [1246, 744]. Continuing [1303, 1691, 544]. Continuum [1036]. Contradictions [1771]. Contribution [3, 1230, 625, 41, 1458, 112, 545, 108, 159]. Contributions [196, 593, 1289, 1054, 1641, 291]. Contributors [1116]. Control [1444, 1757, 1681, 1809, 1087]. Controversial [1244, 1316]. Controversy [1418, 723, 143, 862, 920, 1613, 293, 1684]. Convergence [1169]. Conversion [402]. Cooke [490]. Cooperation [1303, 1612]. Coordinates [1720]. Copenhagen [1520, 841]. Copernic [1579]. Copernican [1300, 666, 298, 179, 268, 1617, 1826]. Copernicu [1238]. Copernicus [1074, 460, 1376, 474, 154, 155, 467, 1074, 1238]. Coptic [717]. Corn [1473]. cornered [1017]. coronavirus [1558]. corps [109]. Corpses [1665]. Corpuscles [1036]. Corpuscular [406]. correct [53]. Correction [740, 335]. Corrections [990]. Corrects [341]. Correlation [451]. Correspondence [1033, 519, 1391, 1456, 1454, 276, 699, 471]. Corrigendum [792, 1243, 1622, 1623]. cortina [1496]. Corvinus [1076]. cosmic [1593]. Cosmo [1296]. cosmographical [1478]. Cosmography [1189, 1632]. cosmología [1605]. Cosmological [1347, 1782]. cosmologique [1579]. Cosmology [525, 1045, 657, 1152, 246, 1252]. Cosmopolitanism [1795, 1443]. Cosmos [1718, 1394, 1799, 1255]. Costs [408]. Coulter [440, 489]. Council [1479, 1621]. counterfactuality [964]. Countries [1531, 1280]. Counts [499]. Couples [1369, 1829]. Court [1128, 1076]. CoV [1555]. cover [8]. Coverage [1417]. COVID [1747, 1559, 1547, 1549, 1561, 1556, 1553, 1546]. COVID-19 [1747, 1559, 1547, 1549, 1561, 1556, 1553, 1546]. Cracow [1070]. Cradle [1721, 1310]. Craft [1395, 1824]. Crafting [1614, 1761]. Craig [1194]. Crawford [1607]. Creager [1150]. Created [1218]. Creative [1351, 1562]. Creativity [892]. Crelle [1055]. Cremona [400, 400]. Crescent [237, 794]. Crespo [1575]. Cressey [1610]. crest [6]. Crete [1631]. Crime [1081]. Crippa [1638]. Crisis [1148, 1434]. Cristina [1348]. Critical [1215, 515, 1134, 614, 491, 1709, 1169, 338, 86, 383]. Critical-Assessment [491]. Criticism [630]. Critics [1158]. Croatia [1660, 1661]. Crossroads [1229, 1234]. Crown [1394, 1799]. Crucial [616]. Cryogenic [1078]. crystal [1654, 518, 530, 540]. crystal-forming [540]. crystal-structure [530]. Crystalli [524]. Crystallographers [144]. Crystallography [144]. crystals [524]. Csiszar [1544]. Cube [1168, 708]. cuerpos [1575]. Cultivating [1320]. Cultivation [1127]. Cultural [1341, 1247, 1327, 1195, 240, 1610, 1519]. Culture [1326, 1357, 1383, 630, 1388, 1282, 1415, 1772, 1209, 1542, 1477, 1676, 1808, 1625, 1472, 1648, 1819, 1296, 1233, 499]. Cultures [1769, 1428, 1340, 1726]. Cumulative [513]. Cuneiform [318]. Cunoastere [1545]. Cupola [850]. cure [270]. Curiosi [651]. Curran [1278]. Current [1743, 1452, 193]. Currie [1518]. Curt [915]. Curtain [1502]. Curvature [685]. Curved [655]. Curves [698, 324, 854, 163, 355]. Cybernetic [1302, 1812]. Cycle [763]. Cycles [735]. Cyclical [746]. Cycloid [312, 263]. Cycloïde [375, 375, 263]. Czech [1229]. Czechoslovakia [845].

D [483, 1444, 1272, 1370, 1725, 426, 432, 1218]. D. [1390, 615]. Dīn [816, 632]. d’Abū [811]. Daily [1145, 1414]. Daipha [1397]. d’al [550]. d’al-Qūī [550]. d’al-Qūhī [550]. D’Alembert [1459, 646, 652, 1458, 1452, 1456, 1453, 1454, 1455]. Dalton [491]. Damas [364]. Damascus [774, 364]. Damien [5]. Dan [1190]. d’analyse [523]. Daniel [1587, 1383, 1686]. Daniela [1326]. dänischen [645]. Danish [1146, 235, 865, 1077, 1143, 645]. dans [5]. Danske [1520]. Darin [1394]. Darkness [1367, 1169]. Darrigol [1325]. Darstellung [53]. Dartres [76]. Darwin [1646, 1417, 1074, 1590, 1261, 1268]. Daryn [1298]. Dastūr [533]. Daston [1124, 1370]. Data [1121, 1228, 989, 310, 1453, 39, 815]. Date [1211, 650]. Dates [640]. Dating [885]. Dauder [1506]. Daughter [1261]. David [1794, 1184, 1154, 1689, 1216, 1794, 1725]. Davide [1638]. Dawkins [1249]. Day [543, 1424, 437, 484]. daylight [118]. Days [286, 1023, 92]. de’Abd [768]. Deans [638]. Death [1747, 1261, 136]. Debate [1172, 818, 1741, 1678, 1638, 1503]. debates [1633]. Debating [1396, 1828].

Debeaune [538]. Deborah [1345]. Decades [842]. Decanal [767]. December [90]. decimal [1006]. décimales [1006]. Deciphering [1674]. Decipherment [525]. Decline [211]. decolonization [1612]. découvert [718]. dedications [1659]. deduction [746]. Deduktion [746]. Deep [1596]. deepest [1592]. Defeater [279]. Defection [1299, 1823]. Defending [389]. Defense [1202]. Defining [1338]. Définition [814]. Deflection [251]. Deformity [1233]. Degree [580, 1259]. Degrees [1341]. Dekane [638]. Del [1747, 1575, 1605]. Delafosse [544, 544]. Delian [391]. delischen [391]. Delisle [518]. Delisles [518]. delivered [1021, 52]. Della [1564, 1565, 1562, 1563]. Deltqvivm [359]. Demanded [339]. Demarcation [1077]. demons [1559]. demonstrations [263, 263]. Demotic [663, 327]. Démotique [663]. Denis [1249]. Denise [1338]. Denmark [1520, 590, 278, 863, 861, 862, 859]. Denny [1136]. Denounce [748]. Density [780]. Depression [996, 1805, 1301]. depuis [1237]. Derivation [527]. Derived [586, 1558]. derzeitiger [193]. Desargues [752, 751]. Descartes [1249, 1127, 301, 1168, 1579]. Descartesschen [78]. Description [1161, 1409, 606]. descriptive [1686]. Design [1053, 887]. Destined [1603]. d’Estudis [1220]. Detail [307]. Details [624]. d’Étaples [1542]. Détente [1433]. Determination [319, 892, 301, 267, 679, 1035, 374]. Determining [324, 219]. Detleff [745]. Detrás [1496]. d’Euclide [814]. deutsch [645]. deutsch-dänischen [645]. Deutsche [1133, 150, 88]. deutscher [1059]. Deutung [128, 112, 74]. Deux [302, 5, 429]. Developed [349]. Development [1365, 534, 681, 1224, 881, 672, 562, 1179, 89, 1483, 1675, 488]. Devices [1067]. Devreese [1149]. Dezso [1604]. d’Halloy [1439, 1068]. d’histoire [20]. d’Història [1171]. Dhruv [1336]. Diagnosis [1759]. Diagram [196]. Diagrams [833, 1717, 1535]. Dialog [112, 514]. Dialogue [893, 112]. Dialogues [1689]. diary [92]. Díaz [1575]. d’Ibn [493]. Dickson [1602]. Dictum [173]. Did [1166, 856, 444, 1298, 40, 1261, 748, 1537, 1806]. Didactical [820]. Dienste [394]. Dieter [1133, 1356]. Dietrich [1354]. Difference [1740, 1576]. Differences [784]. Different [860]. Differential [745]. Diffusion [502]. Digital [269]. dimension [1594]. Dimensional [676]. dimensions [1596]. Din [384]. Dinosaur [1572]. Diocles [655]. Diocletian [136]. Diokletians [136]. Dioptra [307, 1035]. diplomacy [1608, 1627, 1612, 1629, 1614, 1625, 1616]. diplomacy-knowledge [1608]. Diplomat [1175]. Diplomatic [1742, 1630, 1628]. diplomats [1516]. Dipole [1120]. Directing [1146]. direction [746]. Directions [1287]. Dirk [1664]. d’Isaac [532]. Disagreements [918]. disappearance [712]. Disaster [1345, 1788]. discharge [1176]. Disciplinary [1613]. Disciplines [1434]. Discontents [1750, 1645, 1601]. Discourse [1245]. discover [836]. Discovered [121, 718]. Discoveries [1346, 1778]. Discovery [556, 1413, 623, 1107, 1060, 967, 624, 729, 69, 453, 452, 275, 691, 1594, 1512, 544]. Discussion [1070, 614]. discussions [264]. Disdiaclastici [524]. Disease [1711, 1747, 1704, 1576, 48, 1554]. Diseases [201, 1703, 1548]. Diskussion [264]. Disorientations [1740]. dispersion [405]. Displacement [1060, 534].

Display [1734, 1777, 1488, 1490, 1398]. Dispute [389]. dissemination [182, 1680]. Dissertation [1687]. Dissolving [915]. Distance [772, 1151, 1263, 1035, 332, 374, 779, 1476]. Distances [343, 680, 649, 1266, 728]. Distillation [85]. Distorted [1500]. Distributing [1534]. Divided [1303, 440, 489]. Divination [326]. divine [1633]. Division [218, 650]. Do [1773, 1693, 1772, 1599, 1405, 1465]. Dobeil [19]. Dobrzycki [460]. Doctoral [1341]. doctors [180, 14]. Doctrine [283, 277]. Documenting [1289]. Does [1200, 882]. dogmas [254]. Dogs [1209]. Doing [1546]. Dokumente [1583]. Dollars [841]. Dollfus [432]. Dolphins [1737]. d’Omalius [1439, 1068]. Domenico [1085, 1578]. Domin [38, 39]. Dominguez [38]. dominici [209]. Dominicus [208]. Dominique [1688]. Dominis [555, 555]. Donald [1369]. Donation [320]. Done [1430]. Dongen [1270]. Donna [1382]. Doppler [778]. Dopplerschen [778]. Dordrecht [1524]. Dosimetry [1626]. Double [829, 406, 401]. Double-Argument [401]. Doubt [1751]. down [1593, 1588]. Dr. [169]. Drachmann [648]. dragon [1556]. Drake [1265]. Dravya [253]. Drawing [764]. Dream [1388, 54]. Dreams [1271]. drei [494, 1028, 255]. Dritten [1133]. Drive [1308]. Driven [1001, 1191]. drops [5]. drtava [247]. Drucker [1382]. Druckfragmente [241]. Drugs [426, 1607]. drunkenness [270]. drunkenness-cure [270]. Dry [881]. Dubna [1614]. Duc [449]. Dudley [1128]. Duhem [1262, 1275, 1003]. d’un [109, 684]. Duncan [1282]. d’une [663, 5]. Dunn [1637]. Duplication [1168]. Dupré [1280]. Durbach [1233]. durch [671, 544, 555]. During [461, 1055, 1069, 1756, 1420, 1549, 1317, 1059, 1753, 1522]. Dust [1764]. Dutch [1383, 447, 889, 1314, 1808, 1714, 445, 668, 1020]. Dwelling [1752]. Dyadik [456]. Dyeing [94]. Dymock [426]. Dynamic [920]. Dynamics [710, 1119, 238, 301, 981]. Dynamik [699]. Dynamique [646, 652, 1455]. dynamism [699].

E. [583, 1598]. Earlier [388]. Earliest [135, 1019, 1632, 618]. earliest-known [1632]. Early [1749, 1042, 1376, 1123, 1210, 716, 1358, 851, 1541, 1030, 1365, 61, 727, 617, 1086, 1280, 1405, 507, 756, 1253, 730, 303, 1205, 308, 609, 779, 1151, 1244, 1075, 1539, 1306, 1407, 1362, 1084, 745, 981, 1348, 855, 1008, 286, 412, 392, 479, 1741, 1333, 1232, 1665, 1140, 1174, 1437, 1607, 1577, 746, 1451, 1640, 1482, 41, 1551, 1182, 1807, 1635, 1734, 1498, 1472, 1475, 1505, 1825, 1474, 144, 264, 1120, 1317, 1753, 1316, 1315, 1719, 518]. Early-Modern [1120]. Earth [1178, 1174, 1637, 991, 1114, 1221, 127, 1570, 1121, 1228, 1122, 1057, 1224, 282, 1679, 585, 408, 1058]. Earthquake [1345, 1788]. East [921, 1197, 221, 1675, 486]. Eastern [1660, 14]. Eat [1750]. Eating [1335]. eccentricity [1679]. Échanges [1237]. Eclipse [1032, 716, 1030, 921, 323, 1408, 880, 1008, 916, 1671]. Ecliptic [319, 125]. Ecology [1201]. Economic [1197, 1219, 1202, 1674]. economy [1674]. Écorché [7]. ed [1743, 1642]. Eddington [1671]. Eddystone [1658]. edge [1598]. Edgerton [1154]. Edinburgh [1277, 1646, 1530]. Edited [1249, 1072, 1166, 1584, 1251, 1133, 1137, 1357, 1081, 1125, 1082, 1354, 1084, 1291, 1073, 1150, 1021, 1188, 1237, 1179, 1240, 1280, 1334, 1296, 1178, 1329, 1267, 1336, 1217, 1180, 1333, 1232, 1369, 1337, 1645, 1507, 1641, 1640, 1540, 1539, 1521, 1649]. Edith [1128]. editing [1526]. edition [1633]. Editions [205, 229, 242, 259, 266, 273, 315, 345, 353, 415, 422, 476, 510, 536, 547, 570, 1543, 1523, 186, 1686, 1606]. editor [1533, 808]. Editorial [1056, 1063, 1093, 858, 986, 1138, 1506, 1087, 1487, 1546, 1738, 1534]. Editors [1532, 1530, 1525, 986, 1138, 1534, 1538]. Editorship [1526, 1533, 1528, 1531]. Editura [1545]. Édouard [1579]. eds [1749, 1745, 1715, 1726, 1716]. Education [1070, 1238]. Educator [803]. educators [1668]. Edvard [183]. Edward [1074, 1350]. Effect [556, 559, 615, 1094, 65]. Effects [387]. efforts [405]. Efthymios [1640]. Eggs [773]. Egypt [767]. Egyptian [638, 448, 320, 130, 298, 885, 702]. Egyptians [25]. Égyptienne [130]. Égyptiens [25]. EHESS [1523]. Ehrenberg [599]. Ehrström [279]. Eichentraktate [291]. Eighteenth [1239, 1368, 1576, 79, 1696, 1820, 27, 388]. Eighteenth-Century [1239, 1576, 1696]. eighth [1678]. Eightieth [1283]. Eighty [635]. Eighty-Thirds [635]. Einem [38, 717]. Einfluss [421, 699, 309]. Einführung [847]. einige [215, 400, 128, 254]. einiger [193]. Einstein [1671, 1800, 1050, 608, 677, 1241, 1356, 1270, 1181]. Ekliptik [125]. Elaine [1649]. Elastica [759]. Elder [164]. Electric [556, 837, 863, 534, 865, 718]. Electrical [760, 1388, 712]. Electricity [1387, 1390, 1389, 1674, 1386, 1385, 859, 1176, 1675]. Electrified [1386]. électriques [718]. electroacoustics [1469]. Electrodynamics [556, 559, 791, 623, 1107]. Electromagnetic [1210]. Electron [723, 830]. Electrophorus [755]. electrotherapeutic [38]. electrotherapy [38]. elektrischen [712]. Elektrizität [1176]. Elektronentheorie [830]. Elektrotherapeutischen [38]. Elektrotherapie [38]. Element [567, 624, 453, 1641, 1511, 1668]. Elementen [400]. Elements [657, 442, 873, 1509, 1511, 896, 400, 1537, 398, 505, 1536, 814]. Elena [1668]. Elevation [1591]. Eleven [635]. Eleventh [1027]. Elimination [1698]. Elitism [1144]. Elizabeth [1150, 1751, 1750]. Ellipse [685]. Elmer [32]. Elster [251]. Elusive [1002, 1308]. Embeddig [1747]. Embodiment [1716]. Embryologie [75]. Embryology [1741, 75]. Emendation [418]. Emerald [87]. Emerge [489]. emergence [1535]. Emerging [1548]. Emil [1155, 807, 1792, 1372]. Émilie [1352, 1779]. Emily [1580]. Emma [1335, 1232]. emotion [1591]. Emperor [136]. Empire [1326, 1362, 1329, 326, 1475, 1505, 1598, 1756, 1576, 1482, 1828, 1396]. Empirical [1118, 1119, 1120, 1117, 1162, 576]. Empiricism [1437, 1194]. Enacting [1565]. Encouragements [1277]. encyclopaedias [1531]. encyclopedia [1462]. Encyclopédie [1458]. End [1427, 871, 1705, 1702, 1706, 371, 1693, 1696, 1694, 1619, 555]. Ende [555]. Endemic [1554]. Ending [1701, 1697, 1709, 871]. Endings [1700, 1707, 1710]. Endless [1699]. Ends [1708, 1692]. enemy [1551].

Energetic [1029]. Energetics [681]. Energie [1029]. Energy [1210, 451, 673, 170, 1029, 402, 1611, 1673, 1677]. enfant [1543]. Engineer [1248]. Engineering [446, 1151, 1809, 1287, 1657, 1444]. Engineers [1785, 1288, 1377]. England [1580, 144, 1571, 1002, 172, 1696]. English [503, 1194, 1632, 1306, 1044]. English-language [1632]. Enhance [1288, 1285, 1286, 1289, 1287]. enlightened [1453]. Enlightenment [1335, 1604, 463, 1340, 1457, 1793, 1046, 1745, 1058, 1326, 1193, 1278, 1378]. Entangled [1761]. Entdeckung [544]. Entdeckungsgeschichte [452]. Enter [1389, 1772]. Enterprise [1199, 1477]. Entladungsröhre [1176]. Entre [1747, 1237, 650, 1523]. entrepreneurship [1569]. Entries [379]. Entstehungsgeschichte [778]. Entwicklung [562, 89]. Entwicklungsarbeit [1179]. Environment [1444, 1434, 630, 1809, 826, 1754, 1470]. Environmental [1723, 1615]. Environments [1752, 1755]. Ephemerides [574, 794, 665]. Ephemeris [993]. Epicycle [274, 1044]. Epicycles [736]. Epicyclic [177]. Epidemic [1709, 256, 1699, 1704, 1710, 1559, 1550, 1554]. Epidemics [1700, 1747, 1692, 1705, 1695, 1693, 256, 1619, 1747]. Epidemiology [1697, 1707, 1704]. epidemy [1549]. Epigenetics [1421]. Epilogue [1766]. Épiméthée [1579]. epistemic [1577, 1566]. Epistemological [1332]. epistemology [1557]. Epistolary [1323]. epistulis [1391]. Epitome [1617]. época [1605]. Epoch [846]. equal [355]. Equation [697, 401, 616, 798]. equations [1515]. Equator [478]. Equatorie [1406]. Equatorium [1001]. Equatorium-Driven [1001]. Equiangular [805]. Equilibrium [417]. Equinoxes [478]. Equipment [1222]. Equivalent [444]. Era [1201, 1021, 1432, 1020, 1612]. Eradicating [1702]. Eratosthenes [635]. Erforschung [165]. Ergänzende [113]. Eric [1130, 1642, 1333, 1668]. Erica [1712]. Erich [895]. Erickson [1370]. Erik [1137, 1421]. Erika [1235]. Erklärung [555]. Ernährung [188]. Ernst [1155, 1079, 1483, 250, 293, 419, 292, 1033, 250, 293, 419, 255, 171]. Erratum [1156, 1624]. Error [1010, 563, 310]. Errors [333, 128]. erste [56]. ersten [46, 53, 610]. Erwin [616]. Escorial [1408]. ESHS [1330, 1430, 1303, 1620, 1070, 1304, 1479, 1621, 1403, 1373, 1652, 1466]. España [1496]. española [1644]. especially [290]. Essai [663]. Essay [1747, 700, 1746, 837, 1290, 1745, 1743, 568, 398, 848, 397, 377, 560, 625, 1744, 520, 748, 883, 900, 663]. essays [1668, 1086]. Essential [283]. Establishment [1065, 1284, 1057]. estate [215]. etc [1179]. Eternity [1079]. Ether [351, 1577]. Ethno [1420]. Ethno-biology [1420]. Études [302, 244, 263, 375]. Euclid [896, 400, 1537, 562, 964, 758, 817, 505, 815, 873, 1536]. Euclidean [400, 628]. Euclidis [651]. Eudemus [1019]. Eudoxus [341, 2, 854]. Eugen [1176]. Euklid [400]. Euklid-Scholien [400]. Euklidinterpretation [562]. Eulalia [1506]. Euler [759, 1456, 441, 625]. Europa [27]. Europäische [1583]. Europe [1229, 1639, 1137, 1336, 1232, 1197, 1070, 1445, 1660, 1541, 1303, 1500, 1745, 1614, 1501, 1351, 1504, 1690, 1694, 1483, 1466, 27, 1749, 1711, 1253, 1745]. European [1070, 1171, 1185, 1196, 1220, 1188, 1290, 1555, 1271, 1665, 1063, 1171, 1185, 1196, 1220, 1229, 1330, 1318, 1374, 1404, 1431, 1485, 1290, 1312, 1614, 1319, 1242, 1254, 1273, 1292, 1208, 1260, 1283, 1460, 486]. Eva [1140]. Eve [1401, 1802]. Events [1153, 583, 582]. Ever [1430]. Everest [1470]. everyday [1501]. Everything [965]. Evidence [836, 1322, 1537, 887, 1567, 254, 1665]. Evidences [85]. Evolution [67, 546, 1073, 1647, 1646, 1328]. Evolutionary [812, 1235]. Evolventenbildung [671]. evolving [1125]. Ex [1391, 1567]. Exact [269]. Examination [677, 338]. Example [330]. Examples [1153]. Exceeding [1050]. exceptionalism [1645]. excerpts [96, 167]. excess [1023]. exchange [1532, 1628]. exchanged [44]. Exchanging [1772]. Excluded [1436]. Excursions [1350]. exemplary [1150]. exercise [357]. Exercitatio [192, 192]. Exhaustion [835]. exhibit [1485]. Exhibiting [1489]. Exhibition [1773, 1364]. Exhibitions [1728]. Existence [964, 475]. Exit [1772]. Expanding [690]. Expansion [105, 1753]. Expected [127]. Expedition [1420, 1591, 1470, 1069]. Expeditions [1326, 1810, 1359]. Experience [1323, 1386, 1437]. Experiences [1376, 1289]. Experiment [831, 755, 847, 1626, 1626]. Experimenta [524]. Experimental [1363, 919, 1140, 1437, 887, 1182, 1566, 411]. Experimentation [892, 1562]. Experimente [847]. Experimenting [1585]. Experiments [1225, 1712, 785, 105, 1226, 1717, 285, 252, 1565, 524, 1735, 1563, 847]. Expertise [1692, 1433, 1232, 1609]. experts [1609]. Explanation [371, 588, 555, 1250]. Explanations [1072, 481, 1696]. Explode [1490]. exploitation [1616]. Exploration [1296]. Explorer [1064]. Exploring [1608, 1408]. export [1610]. Expressions [1341]. extended [400]. External [1153]. Extinction [1736]. Extraction [85, 708]. Extremal [539]. extreme [650, 1470]. Exzerpte [96]. Eye [1688].

F [502, 44, 512]. Faber [1007]. Fabian [1747]. Fabrica [8, 1575, 8, 1686]. fabrique [1579]. Fachschrifttums [89]. Facilitating [1078]. factors [68]. facts [1500]. Faculty [1540]. Fages [1523]. Faidra [1188]. Failed [623, 1107, 1252]. Failure [251, 1615]. Fair [166]. Fairs [1401, 1802]. Faith [1603, 1079]. Falkenburg [1134]. Fall [1711, 1263, 1399, 1664, 1486, 1814]. Fallacy [777]. False [829]. Fame [166]. Family [1306]. Fantasmas [1644]. Fantasy [1670]. far [558, 1725]. Fara [1236]. Faraday [791]. Fascination [917]. Fascist [1756]. Fashioning [1077]. Fashions [1414]. Fässern [219]. Fate [1739, 1101, 116]. Fates [1183]. Father [389, 244]. Fatherland [1214]. Fats [120]. Faults [1060]. Fe [1721]. fear [1488, 1670]. feather [1733]. Feature [1431]. Fechner [250, 250]. Feeling [1712]. Félix [1575]. Felski [1727]. Female [1308, 1235, 247]. Feminism [1308]. feminist [1471]. Femmes [1689]. Ferdinand [830, 830]. Fermat [357, 324, 586]. Fermats [357]. Fernando [1358]. Fernel [1230]. Feuerwerkstext [200]. Fever [279]. Fevers [1754]. Few [1235]. Feynman [1794, 1184]. Feza [1336]. Fibonacci [650, 1147]. Fiction [1355, 1784]. Field [556, 1284, 1419, 1120, 1435, 1223, 1359, 1630, 1810]. Fieldwork [1359, 1810]. Fienus [1741]. Fifteen [1466]. Fifth [1304]. fifty [1468]. Fight [863]. Fighting [1551]. Figure [7, 1689]. figurierte [46]. Filling [816, 1762]. fin [1144]. final [1603]. Finale [1706]. Find [1286]. Finding [1007, 433, 728, 1305]. Finest [1775]. Finkelstein [1372]. Finnish [525, 279]. fireworks [200]. First [842, 218, 1229, 1418, 998, 761, 765, 1768, 601, 727, 1005, 1410, 794, 40, 1627, 46, 1489, 1023, 663, 650, 1633, 1549, 1552, 53, 1456, 56, 1529, 610, 263, 1514, 1450]. First-Crescent [794]. First-millennium [1410]. Fiss [1401]. fission [1611]. Five [1375, 1136]. fixed [1678, 1493]. Fizeau [892]. fl [1001]. Flamsteed [589, 1257]. Fleeing [1183]. flesh [1733]. Fleßenkämper [1277]. Flies [1763, 606]. flipping [1560]. Florence [320]. Florentia [876]. Florentino [1414]. Flu [1700]. Fluctuations [780]. Fluid [1119, 1001]. Foerster [1176]. fog [1502]. Folkerts [1086]. followed [5]. following [1450]. fols [876]. FOM [1314]. fontium [1570]. Food [1419, 135, 1335, 188]. football [1630]. footnote [576]. Footsteps [1356, 1800]. Forbes [433]. Force [761, 769, 805, 199, 633, 109, 226]. Forced [425]. Forces [451, 402, 405]. Forecasters [1397, 1789]. Forecasting [1685]. Forensic [1272]. Forestry [1486, 1363]. Foreword [1360]. Forgan [1247]. Forging [1371, 1815]. Forgotten [1723, 713, 302]. Form [855, 1514, 1019]. Formal [1120, 1117]. Formal-Empirical [1120]. Forman [630]. Formation [1118, 1047, 671]. forming [540]. Forms [83, 1492]. Formula [385, 577]. Formulation [646, 652, 385]. Fors [1378]. Forschungen [1012]. Forschungs [1179]. Forschungs- [1179]. Forsten [1598]. Forstner [1184]. Fortplanzungsstoffes [247]. Fortunes [584]. forward [1504]. Fossil [132, 1222]. Foundation [919, 575, 1614, 1201, 1666]. Foundations [801]. Founder [599]. Four [801, 598, 608, 1236]. Fourier [370]. Fox [1237, 1327, 1443]. Fractions [1246, 744, 1006]. Frage [53]. Fragment [366]. Fragments [168, 241]. Framework [775]. frameworks [1627]. France [1251, 1543, 1606, 1523, 1579, 1754, 1327]. Francesca [1636]. Francesco [1054, 1665]. Francis [23, 1129, 1565, 1562]. Francisco [1768]. Franciscus [38, 38]. Franck [1131]. franco [1237]. franco-britanniques [1237]. Francois [1689, 56]. Francoist [1377, 1785, 1494]. Franconia [248]. Frank [1747]. Franken [248]. Frankreich [1175]. Franz [1065, 39]. Fraser [475]. Frauenblattes [26]. Freak [1233]. Frederika [1142]. Free [1021, 1199, 1501, 1500]. Freiburg [395]. Freiburger [395]. Freiherr [1065]. French [455, 1240, 1753, 1193, 423, 1581, 500, 532, 1413, 663, 109, 650, 718, 1463, 999, 364, 450, 25, 130, 741, 1006, 429, 506, 675, 99, 1455, 493, 523, 550, 629, 684, 66, 994, 35, 208, 5, 34, 302, 244, 263, 375]. Frequent [461]. Fresco [850]. Fresnel [711]. Fretwell [1712]. Freudenthal [1219]. Freundschaft [1279]. Friend [1128]. Friends [1611]. Fries [241, 241]. Frontier [246, 1603]. Frontiers [1795, 1443]. frühe [264]. Frühen [1075, 746, 518]. Frumer [1497]. Fuchs [339]. Full [575]. Fun [1488, 1492]. Function [716, 794]. functions [386]. Fundación [1644]. Fundamental [1720]. Funding [1284, 1314, 1666]. Funktionen [386]. Furada [1418]. füreinander [1804, 1400]. Further [197]. Fusion [747]. Future [1766, 1172, 1288, 1285, 1286, 1289, 1672, 1287, 1580, 1603, 1812, 1208, 1669, 1302]. Futures [1435, 1674]. futurities [1586].

G [818, 456, 601, 1323, 502, 1074, 1329, 433, 503, 1217, 514]. Gabriel [1607, 366, 1372, 366]. Gabriele [1152]. Gaia [1344, 1816]. Gal [1343]. Galactic [412]. Galán [1507, 1643, 1081]. Galen [88, 17, 88]. Galenic [1724]. Galenos [88]. Galenübersetzungen [88]. Galilean [1744]. Galilée [1240]. Galileo [1517, 1240, 1744, 1231, 785, 238, 1010, 1152, 568, 1263, 1266, 40, 893, 1827, 748, 1380, 1248]. Galison [1109, 1108, 1124, 1110]. Game [822, 1490]. games [1482]. Gamow [822]. Gaps [1762]. García [487, 1506]. Garden [1320, 421]. Gardeners [1719]. Garrod [1541]. Gartenliteratur [421]. Gärtnerlehranstalt [1065]. Gas [862]. gases [807]. Gaspard [803]. Gastheorie [807]. Gaston [302, 244, 263, 375]. Gathering [1768]. Gaukroger [1648]. Gauß‘ [633]. Gears [483]. Gebrannten [366]. Gebri [297]. Gecekondu [1757]. Gedenken [31]. gedruckten [46]. Geert [1664]. Gegenübung [651]. Gegenwart [1075, 277]. geist [86]. Geister [1583]. Geitel [251]. gelten [220]. Gender [1306, 1142, 1139, 1305, 1716]. Gendered [1649, 1386, 1182]. Gene [1393, 1801, 1467]. Genealogy [1393, 1801, 1461]. General [1220, 1089, 1522]. Generality [1403]. Generation [1281, 1181, 1811]. Generic [1797, 1379]. genesis [778]. Genetical [430]. Genetics [1367, 1467]. genius [1800, 1356]. Geniza [605]. Genome [1746]. Genotype [546]. genre [1566]. Geo [1058]. Geo-Physical [1058]. Geodesy [685, 1618]. Geoff [1669]. Geographic [1486]. geographical [1568, 1529]. geographical-astronomical [1529]. Geography [657, 1438, 1305]. Geological [1350, 1463]. Geologist [1222]. Geologists [1223]. Geology [1225, 1221, 1059, 1784, 1590, 1355]. Geomagnetic [601, 633]. Geomagnetism [1069, 1228, 1120]. Geometric [1168, 113]. geometrica [192, 192]. Geometrical [795, 1411]. Geometrie [1147]. geometrischen [113]. Geometry [752, 1126, 598, 1039, 655, 685, 1038, 552, 517]. Geopolitics [1357, 1819]. Georg [651, 1029, 1391, 1493, 27]. George [1268, 110, 1545, 1610]. Georges [690, 1226, 1481]. Georgina [1423]. Gerard [400, 400]. Gerhard [648]. geriatrics [164]. Geriatrie [164]. German [1132, 437, 1029, 456, 1179, 357, 1176, 1075, 441, 498, 1131, 438, 436, 514, 512, 463, 8, 6, 46, 146, 871, 746, 119, 252, 638, 75, 694, 270, 395, 289, 215, 346, 400, 1459, 128, 134, 182, 830, 96, 210, 313, 366, 421, 47, 1033, 41, 528, 562, 594, 1794, 658, 618, 102, 48, 86, 112, 713, 264, 53, 74, 92, 4, 78, 113, 98, 192, 671, 494, 125, 383, 545, 150, 1059, 386, 338, 148, 84, 136, 194, 212, 277, 355, 240, 17, 88, 139, 799, 232, 806, 138, 12, 54, 181, 180, 179, 193, 220]. German [254, 234, 247, 253, 391, 56, 394, 89, 15, 645, 87, 156, 276, 80, 1012, 38, 1017, 712, 159, 452, 55, 778, 717, 847, 157, 241, 44, 188, 291, 376, 250, 293, 419, 560, 108, 610, 667, 95, 309, 636, 27, 39, 24, 1028, 31, 916, 26, 33, 711, 14, 200, 219, 248, 255, 518, 530, 540, 544, 699, 393, 807, 867, 555, 231, 335, 226, 164, 659, 107, 704, 167, 225, 171, 165]. Germany [1183, 1585, 1338, 1583, 1589, 1792, 1680, 1677, 863, 1308, 1471, 1158, 1207, 1372]. Gerontologists [1422]. Gerson [363, 410, 680, 1005]. Gersonides [1003, 1004]. Gervase [657]. Geschichte [252, 41, 108, 95, 26, 1075, 560]. Gesellschaft [1133, 498]. Gesetze [494]. Gestalt [518]. Gesundheit [165]. get [1732]. Gewechselt [44]. Ghent [1209]. Ghibaudi [1668]. Gianfrancesco [1380, 1827]. Gideon [1219]. Giora [1177]. Giorgio [1637]. Giovan [1564, 1565, 1562]. Giovanni [826]. Gittertypen [544]. Give [1261]. glaciology [1594]. Gleanings [143]. Glen [1174]. Global [1608, 1183, 1365, 1154, 1057, 1319, 1611, 1671, 1475]. Globalized [1435]. gloss [395]. glossary [270]. Glossensammlung [395]. Glow [1066]. God [1253]. Goddu [1238]. Godin [1495]. gods [1559]. Goethe [1213]. Goitre [278, 1634]. Goladipika [1409]. Golden [342, 1383, 1808, 397]. Goldstein [1176, 987, 1113, 1177, 1112]. Gómez [1575]. Gondolat [1604]. Gone [830]. Good [1127, 1235]. Goose [158]. Göppingen [270]. Gordin [1370]. Gordon [999]. Gotfredsen [183]. Gottfried [599, 248, 248]. Goussier [1058]. governance [1600, 1550]. Government [847, 847]. Grading [773]. Granada [1575]. Grand [1252]. Grandinl [1082]. Granollachs [851]. Grant [1074, 847, 458, 847]. Graphical [728]. Gravesande [1332, 1342, 1714]. Gravitation [882, 1522]. Gravitational [831]. Gravity [417, 1633]. Great [988, 805, 1715, 256, 12, 1245]. grecque [994]. Greece [1142, 164]. Greek [1325, 498, 717, 335, 177, 795, 46, 1640, 641, 683, 738, 1145, 1019, 28, 709, 1143, 993, 855, 994, 392, 33, 231, 702, 164]. Greek- [717]. Greek/German [164]. Greeks [483, 1173, 1187, 62, 113]. Green [1723]. Greene [1379]. Greenland [1317, 1316, 1315]. Gregor [1203]. Gregorio [1079]. Gregory [1638, 4]. Gregorys [4]. Grenzgänge [871]. Greyerz [1639]. grid [1675]. Griechen [113]. griechisch [717]. griechisch- [717]. Griechischen [33, 46, 231, 335, 164]. Grigorian [484]. grossen [12]. Ground [1397, 1789]. Grounds [1584]. Group [859]. Grove [451]. Growing [1072]. Growth [1172]. Grundzuege [41]. Grundzüge [240]. Guereña [1496]. Guicciardini [1192]. Guidance [1464]. guide [1518]. Guidelines [1568]. Guido [1234, 1149, 1001]. Guillaume [350]. Guldberg [660]. Günergun [1336]. Gurka [1604]. Gustav [41, 1739, 1740, 250, 250].

H [1442, 1021, 1728, 456, 461, 1390, 1401, 1084, 440, 489, 1322, 1247, 424, 1291, 1267, 438, 1217, 668, 1073, 1150, 699, 1359]. H. [818, 677, 699]. H5N1 [1558]. Hākimitic [580]. Hāsīb [649]. Haarlem [52]. Habash [649]. habitat [1596]. Habsburg [1329]. Haeckel [1155, 1079]. Haitham [73]. Hajjaj [1537]. Halbleiterforschung [545]. Half [784]. Hall [471, 1490, 556, 559]. Haller [92, 55]. Halleri [74, 74]. Hallers [92, 55]. Halleux [500]. Halo [588]. Hamal [1720]. Hamilton [981]. Hammer [1222]. Han [1687]. Hand [1718, 1016]. Hände [1179]. Handle [1316]. Hands [1259, 1179]. Handschriften [594, 291, 167, 96]. Hannah [1582].

Hans [1166, 270, 1190, 270]. Hanson [1194]. Happy [871]. Happy-End [871]. Harbsmeier [1359]. Hardcover [1577]. Harman [1354]. harmonic [244, 263]. harmonique [263]. Harmoniques [244]. Harold [503]. Harre [434]. Harriot [232, 232, 225]. Harriott [142, 160, 166]. Harrison [1267, 1426]. Hartner [342]. Hartvig [1071]. Harvard [1670, 1289]. Harvey [1519]. Hasok [1322]. Hassan [1215]. Hauptsatzes [746]. Haut [338]. Haüy [518, 530, 540]. Haüys [530, 540, 518, 544]. Haüyschen [544]. Hayek [1721]. Haytham [736, 493]. Hayton [1394]. Hazard [1757]. Headlines [1416]. Healing [1803, 1441, 1607]. Health [1711, 1756, 1757, 1390, 1803, 1441, 1555, 165]. Healthy [1087]. heart [1519]. Heat [444, 211, 614, 875]. Heaven [1001, 1637]. Heavens [1174]. Heaviside [661]. Heavy [742]. héb [1537]. Hebammenwesens [108]. Hebrew [656, 158, 1536]. Hedenus [1176]. Hedin [1591]. Heiberg [817]. Heidelberg [248]. Heidelberger [248]. Heike [1078]. Heilbron [503]. heiligen [14]. Heilmittelglossar [270]. Heim [1180]. Heinrich [146, 313, 837]. Heisenberg [1786, 673, 1392]. held [1220]. Helge [1252, 1520]. Heliocentric [460, 1618, 498]. Heliocentrism [1238]. Helium [1066]. Hell [1069]. Hellenistic [1161, 727, 1636]. Helm [1029]. Helmholtz [1155, 1203, 1688, 806, 667, 1158]. Helmut [1745]. Hendrick [639]. Henle [1155]. Henri [99]. Henrik [645, 645]. Henry [131, 1276, 245]. Hentschel [1132, 1428, 1524, 1179]. Heptagon [726, 98]. Her [1310]. Herb [1320]. Herbaria [1774]. Herbert [1129]. Here [1079, 1733]. Heredity [1053, 1746]. Herman [1295, 1813]. Hermann [1155, 806, 806, 1203]. Hero [848]. Heroic [1414]. Heron [13, 168, 307, 464, 1035, 1211, 577]. Herring [1580]. Herschel [704, 704]. Hertz [718, 787, 837, 791]. Hertzsprung [196]. Herzog [594, 594]. Hessler [1247]. Heuraet [639]. heuristic [1564]. Heymans [1209]. Heyrovskýa [845]. Hierarchy [983]. High [673, 785, 376]. High-Energy [673]. Higher [1252]. Hildegard [188, 188]. Hipparchian [93, 879]. Hipparchus [1030, 1031, 640, 267, 332, 274, 392, 576]. Hippocrates [489, 1019]. Hippopedes [1005, 854]. Hirnanatomie [41]. Hirnforschung [33]. Hirschberg [1688]. Hispanic [3]. Hispanie [684]. Histoire [35, 1606, 99]. Historia [1644]. Historia/Fundación [1644]. Historiam [74]. Historian [292, 918]. Historians [1430, 1668]. Historic [323]. Historica [1583]. Historical [1072, 747, 700, 239, 818, 461, 1010, 442, 480, 504, 1311, 1336, 1704, 8, 1557, 355, 1550, 1518, 256, 68, 576, 1446, 487, 473]. Historical-Mathematical [504]. Historicize [1290]. Historicizing [1695]. Histories [889, 1747, 1572, 1305, 1608, 1507, 1541, 1600, 1548, 1597, 1566]. Historiographic [1049]. Historiographical [1690, 1452]. Historiography [1212, 1570]. historischer [355]. History [1063, 1442, 235, 1070, 1138, 1171, 1185, 1196, 1220, 1229, 1330, 1318, 1360, 1374, 1404, 1431, 1445, 1711, 1660, 1361, 1783, 1221, 1540, 829, 1325, 1163, 1746, 1118, 1173, 1187, 812, 1126, 201, 1288, 1751, 1301, 1003, 227, 1739, 278, 3, 600, 1432, 1425, 730, 269, 1803, 1154, 609, 1082, 1244, 1276, 1334, 1723, 1075, 1321, 852, 1094, 174, 1178, 1688, 1170, 1521, 1423, 440, 489, 1713, 981, 1725, 196, 1186, 1319, 426, 1707, 1242, 1254, 1273, 1292, 1498, 1441, 1373, 1740, 256, 1328, 1212, 1368, 1208, 1260, 1283, 333, 392, 1274]. History [1460, 26, 1522, 430, 1174, 1202, 1359, 1716, 1287, 1663, 1598, 1682, 1560, 1524, 252, 1580, 1547, 20, 1661, 1574, 41, 74, 1571, 1561, 1805, 1495, 65, 782, 253, 1810, 1573, 18, 99, 1649, 1656, 1662, 1820, 108, 95, 35, 1553, 1466, 1546, 1588, 452, 560, 1650, 1578, 1428, 1204, 1272, 1425, 1236, 1195, 437, 688, 490, 512]. Hit [397]. Hitler [1780, 1402]. HIV [1393, 1700, 1801]. HIV/AIDS [1393, 1700, 1801]. Hjalmar [1378]. Hoatzin [1734]. Hoax [1416]. Hochadel [1507]. Hoddeson [1399]. Hoff [668]. Hoffenberg [1728]. Hoffmann [1133, 1356]. Holland [52]. Holy [14]. Homberg [1667]. Home [1310]. Home-Based [1310]. Homocentric [224, 736, 1005]. Homosexuality [1390]. Hon [1177]. Honorary [1341]. Honour [1167]. Hooijmaijers [1166]. Hooke [769, 149, 104]. Hooper [426]. Hope [1424]. Hopkins [1517, 1639]. Horizon [996]. horizons [1596]. Horizontal [734]. horizontale [551]. hormone [17]. Hormonforschung [17]. Horns [342]. Horoscopes [989]. Horrebow [385, 404]. Horse [1736]. Hospital [1803, 1441, 1092]. Hosted [1171]. Hot [1316, 822]. Hotspot [1350]. Hour [688]. Hours [721]. Household [1306]. Houses [767, 1027]. Howard [608]. Howson [515]. hTe [712]. Hückel [895]. Hudde [102, 102]. Hughes [615, 1147]. Huigen [1239]. Human [1746, 1253, 1438, 1321, 46, 53]. humanismo [1605]. Humanities [1727]. Humboldt [1175, 165, 1064, 165]. humility [1561]. Humor [1492]. Hund [512]. Hundred [1747]. Hungarian [636]. Hungary [1604, 1660, 1076, 256, 1663]. Hunt [1746, 1316, 1241]. Hutchinsonianism [1046]. Huygen [311]. Huygens [1638, 52, 1258, 802, 588, 312]. Hydrodynamics [1118, 807]. Hydrodynamik [807]. Hydrogen [1244]. Hygiene [1753]. Hyperbola [603]. Hyperbola-Construction [603]. Hyperbolas [815]. Hypotenuse [833]. Hypothesis [1344, 507, 337, 1816, 264, 659, 1261]. Hypothetical [233, 411]. Hyung [1422].

IAEA [1626]. Iavorenciuc [1545]. Iberian [1473, 1472, 1505, 1492]. Ibn [401, 240, 439, 428, 465, 868, 821, 580, 1027, 240, 736, 493, 995, 73]. Icelandic [290, 524]. icon [1510, 1508]. Iconography [1043]. Ida [1021, 1217]. Idea [174, 67, 1511, 544, 1495]. Ideas [752, 518]. Idee [544]. Ideen [518]. identification [638]. Identifizierung [638]. Ideology [1249, 1144]. Idiosyncrasy [756]. Idol [1380, 1827]. If [1490]. Ignace [244, 302, 244, 263, 375]. Ignaz [279]. ihre [562, 540]. ihrer [128, 560]. II [1166, 1408, 1075, 816, 559, 1342, 1314, 834, 603, 263, 693]. IIET [1361]. III [846, 339, 302]. ill [1640]. illness [136]. Illustrated [32, 46]. Illustrating [437]. Ilya [608]. Im [24, 33, 107, 871, 182, 1794, 1133, 1184, 1176, 148, 84, 194, 1131, 394, 1583, 291, 108, 636]. Image [1145, 618, 1604, 1539]. Images [1144, 1037, 1395, 1824]. imaginaries [1616]. Imagination [1387, 1388, 1386, 1385, 1778, 1539, 1590, 1346, 1271]. Imagining [1491].

Immanuel [1006, 334]. Immaterial [1067]. impact [847, 1183]. imperatorium [1411]. Imperfecta [95, 95]. Imperial [1574, 1740]. Implications [224, 1082]. Importance [673, 146, 128, 562]. Important [456]. Impossibility [1168, 1638]. Impression [590]. Improvements [1277]. in/about [1662]. Inaccessible [1228]. Incarnations [1716]. Including [105, 330]. incompatibility [494]. Incubation [81]. Indebted [634]. Independent [1311, 1449]. indeterminate [523]. indeterminate-analysis [523]. indéterminée [523]. Index [1041, 1062]. India [426]. Indian [335, 525, 1025, 703, 966, 139, 138, 12, 54, 181, 180, 193, 220, 254, 234, 247, 253, 231, 702]. Indica [426]. Indican [498]. Indies [1598]. Indifference [967, 1535]. Indigenous [1706]. Indisch [253]. indische [247, 234]. indischen [138, 12, 231, 335]. indischer [180, 193, 254]. Individual [771, 1183, 129]. Individualism [1024]. individually [540]. individually-integrated [540]. individuellen [540]. Indivisibles [827]. Induction [592, 615]. Indus [525]. Industrial [502, 1247]. Industry [70, 1104, 1419, 1082, 840]. infection [48]. Infectious [201, 1703, 48]. Infektion [48]. Infektionskrankheit [48]. Infinitesimals [820]. Influence [1148, 1540, 370, 660, 1202, 421, 309, 544, 699]. Influences [1339]. influenza [1558]. Information [203, 235, 1692, 1053, 1013, 1500]. Informations [184, 189]. Infra [1764]. Infra-thin [1764]. infrastructural [1595]. Infrastructures [1764, 1137, 1673]. ingenious [1658]. Ingenium [1136]. ingenuity [1541, 1749]. Ingjald [9]. Initiative [1290, 1202]. Initiatives [1365, 1615]. Inner [381]. Innovate [1191]. Innovation [1354, 1247, 1798, 1495]. Innovative [1309]. Inquiry [1049, 1298, 1806]. Insect [1362]. Insights [1002, 1693, 1457]. Institut [1220]. Institute [1681, 841, 1283, 1209, 1484, 1721]. institutional [1663]. Institutionalised [1024]. Institutionalization [1364]. Institutionalizing [1090, 1087]. Institutions [1715]. Instrument [1002, 1222, 1528]. Instrumentalism [1003]. Instrumente [1028]. Instruments [1221, 790, 1637, 1476, 465, 581, 1224, 1410, 1760, 1398, 1223, 1777, 1628, 1563, 1028, 302]. integrated [1580, 540]. Integration [1059]. integrierenden [540]. Intellectual [630, 1521, 1414, 1650]. Intelligence [1200, 1313]. Intelligent [1053]. Inter [1201, 1612, 1773]. Inter-African [1612]. Inter-war [1201]. Interactions [1237, 1477, 1258, 1631]. intercalation [1023]. interdisciplinary [1503]. Interest [1146, 1246]. Interests [1718, 1795, 1443]. interface [895]. Internal [1153, 68]. International [1171, 1185, 1196, 1220, 1330, 1304, 1524, 1417, 502, 1742, 1073, 1202, 20, 1615, 1555, 1609, 1078]. internationale [20]. Internationalisation [1078]. Internationalism [1313, 1158, 1626]. interplay [68]. Interpolation [198, 533]. Interpretation [342, 284, 74, 1212, 405, 663, 562, 112, 1016, 663]. Interpreting [865]. interreligious [1631]. intersecting [815]. Intersections [1539]. Intersex [1751]. Interstellar [1066]. Introduce [861]. introduced [1464]. Introduction [986, 1783, 1221, 1526, 369, 473, 1312, 1412, 1729, 1405, 860, 1167, 633, 1767, 1088, 1498, 1139, 59, 1285, 862, 514, 1117, 1255, 1386, 1385, 1588, 1653, 1128, 847, 1687, 1305, 1442]. Introductory [1]. Inv [320]. Inventing [1413, 1813, 1295]. Invention [731, 633, 1355, 872, 1784, 1129, 302, 1586]. inventory [1458]. Inverse [765, 805]. Inverse-square [805]. inversely [405]. Investigation [759, 744]. Investigations [574, 1651, 338]. Invisible [1179, 1551]. involute [671]. Involvement [1246]. Ioan [1191]. iodine [1634]. Ionic [782]. Iris [1277]. iron [1016, 1502]. Irrational [818]. Irrationalitätsbeweis [113]. irrationality [113]. Irregular [379]. Isaac [1745, 532, 1173, 1187, 1192]. ISBN [1607, 1650, 1577, 1542, 1545, 1584, 1664, 1645, 1524, 1580, 1683, 1602, 1666, 1685, 1641, 1640, 1586, 1519, 1576, 1540, 1541, 1520, 1600, 1574, 1667, 1643, 1647, 1517, 1603, 1585, 1636, 1543, 1639, 1601, 1637, 1638, 1606, 1523, 1539, 1578, 1669, 1579, 1688, 1521, 1689, 1582, 1573, 1575, 1668, 1642, 1671, 1518, 1587, 1646, 1605, 1670, 1583, 1644, 1604, 1649, 1648, 1651, 1581, 1687, 1544, 1686, 1522, 1684, 1665]. Isis [513]. Islamic [435, 764, 722, 152, 1293, 365, 384]. Islamicate [1540]. Islandici [524]. Isles [114]. isolation [15]. Isolierung [15]. Israel [1445]. Issues [1215, 1332, 1342, 1518]. Istanbul [1757]. istorie [1545]. Italian [472, 1339, 1756, 622, 1223]. Italy [1682, 1551, 1755, 1719, 1754]. Itinerarium [1567]. IV [879, 375]. ix [1607, 1639, 1581]. IXe [5].

J [1749, 487, 473, 1253, 1081, 1205, 1539, 1178, 1036, 460, 1241, 426, 503, 488, 1441, 1382, 425, 1347, 1217, 436, 668, 44, 1424, 463, 475, 1359, 1716]. J. [1542, 1541, 1210, 845, 1036, 668, 44]. Jābir [428, 465]. Ja [558]. jabr [380]. Jackson [1393]. Jacob [1128]. Jacobi [981]. Jacobus [876]. Jacques [1217, 1033, 675]. Jade [1769]. Jahre [182, 157, 636]. Jahren [44, 89]. Jahrhundert [871, 395, 148, 84, 194, 108, 200, 107]. Jahrhunderts [146, 421, 41, 212, 27, 555, 167]. Jaina [543]. Jakob [1155]. Jakub [1574]. James [1607, 1606, 1296, 1241, 1131, 1191, 1216]. Jan [1329, 1574, 16]. Janelle [1349]. Janet [1395]. jang [665]. jang-tsai-chüeh [665]. January [1071, 1008]. Japan [1629, 1497]. Jaume [1577]. Jayyān [1027]. Jean [1496, 711, 1439, 1068, 646, 652, 1452, 1230, 723, 1453]. Jean-Baptiste [1439, 1068]. Jeffrey [1269]. Jenkins [1646]. Jennifer [1602, 1517]. Jens [1583]. Jeremy [1379]. Jeroen [1270]. Jerónimo [1605]. Jesper [790]. Jesuit [1152, 1427]. Jesuits [748]. Jewish [1191, 1684, 450]. Jimena [1204]. Jo [1281]. Joan [1585]. Jochen [1744]. Joffe [1686]. Johan [1711]. Johann [658, 102, 146, 658, 102, 517, 671]. Johannes [376, 1411]. Johannsen [546, 430]. John [1747, 1683, 1295, 1276, 1178, 1193, 1297, 163, 1212, 1581, 252, 1387, 1408, 1658, 1573, 591, 1707, 376, 491]. Johns [1517, 1639]. Johnson [1324]. Joint [1138, 1658]. Joly [1237]. Jonas [610, 610]. Jones [1580, 1127]. Jordan [1324]. Jordanus [346, 346]. Jorge [1644]. Jørgensen [519]. Jorink [1333]. Jos [38]. Jos. [38]. José [1749, 1541, 1081, 1011]. Josef [86, 419, 419]. Josefina [1726]. Joseph [1607, 470, 446, 1391, 591, 39, 86, 1527, 39]. Jost [1131]. Joule [847, 105, 303, 847]. Journal [1063, 1691, 26, 80, 1529, 1544, 1055]. Journals [1138, 1534, 1526, 1530, 1531]. journeys [1593]. Joyce [1346]. Jozef [1149]. Jr [608]. Juan [1575, 1644, 1747]. Jubilees [1341]. Judging [1221]. Judy [1370]. Jugoslavien [38]. juif [450]. jumatatea [1545]. June [9]. Jupiter [1256]. Jürgen [1334, 1647]. Just [530, 540].

Kāmil [811, 868]. Kāshānı [1293]. Kāshī [776]. Kūh [996]. Kūhī [816]. Kahlenberg [88, 88]. Kaijser [1137]. Kaiser [136, 1180]. Kaiserschnitt [56]. Kalten [1794, 1184]. Kamel [1391]. Kamerlingh [1078]. Kammâd [821]. Kang [1271]. kangaroo [1732]. Kangro [424]. Kant [1037]. Kapillarerscheinungen [659]. Kapitel [86]. Karen [1401, 1073]. Kargon [1401, 1247]. Kärin [1725]. Karl [1082, 1324, 1065, 545, 1438, 1483]. Kasmil [523]. Kaspar [1639]. Kater [245]. Katherine [1272]. Katzir [1334, 1083]. Keeping [1731]. Keeps [1122]. kein [871]. Kelvin [351]. Kennedy [583, 459]. Kenya [1549]. Kepler [1045, 1123, 1782, 773, 772, 775, 774, 771, 1010, 494, 404, 1679, 249, 1617, 589, 310, 485, 1347]. Kepler-problem [404]. Keplerian [381]. Keplerschen [494]. Kevin [1580]. Key [246]. Khāzin [558]. Khaldun [240]. Khalid [439]. Khwārizmī [237]. Khwarizmi [325]. Killing [1485]. Kimya [1215]. kinetic [807]. kinetics [871]. Kinetik [871]. kinetischen [807]. King [15]. Kingdom [1542]. Kinsey [1382, 1790]. Kiprijanov [1580]. Kirchenlehrere [24]. Kircher [1043]. Kirchhoff [705]. Kirsti [1126, 1167, 1170, 1716]. Kisacky [1441]. Kitāb [811, 868]. Kitab [459]. Kitab-Tahdid-Al-Amakin [459]. Kjennerud [9]. Klaus [1132, 1428, 1524, 1179]. Klein [1370, 1713, 1232]. Kleinliteratur [291]. kleinsten [806]. kleinster [518]. Klep [1217]. Klingenstierna [685]. Know [1307, 1298, 1806, 893]. Knowing [1731, 1719]. Knowledge [79, 1307, 1346, 1737, 818, 1365, 227, 1433, 1758, 1412, 1280, 630, 1198, 1306, 1362, 1329, 1395, 1351, 1731, 1373, 1020, 1117, 475, 1305, 1607, 1608, 1478, 1778, 1507, 1790, 1634, 1680, 1647, 1240, 1635, 1573, 1505, 1544, 1824, 1827, 1649, 1269, 1382, 1380, 1336, 1239]. Known [1055, 1632]. Koenig [667]. Kolb [1399]. Kombinationstöne [667]. Komet [1176]. kongelige [1520]. König [15]. Können [220]. Konstantin [1580]. Konstitution [139]. Konstruktion [98, 355]. Kontroverse [293]. Konzeption [540]. Kopelevic [488]. kopernikanische [179]. koptische [717]. Koptisches [717]. Körpers [53]. Körperteile [46]. Kräfte [540]. Kragh [1520, 1084, 1252]. Kramers [600]. Krankheit [136]. Krebsbehandlung [84, 212]. Krebskrankheit [194]. Krebslehre [277]. Kreisbogenzweiecke [1017]. Kreisrektifikation [671]. Kremer [1398]. Krieg [1794, 1184, 1583]. Krige [1573]. Krischel [1400]. kristallbildende [540]. Kristalle [530]. Kristallstrukturtheorie [518]. Kristian [1359]. Kristin [1324]. Kritische [86, 383]. kritischen [338]. Kronland [468, 529]. Krøyer [645, 645]. Ktesibios [37, 464]. Kuhn [1213, 1160, 1468, 1279]. Kulturmorphologie [240]. kunstgeschichtlichen [8]. Kurven [355]. Kuwait [1676]. Kyrz [486].

L [511, 1600, 487, 1370, 502, 447, 1421, 440, 489, 441, 503, 1398, 498, 1267, 439, 1369]. L. [1603, 372, 593]. Laar [668]. Lab [1155, 1760, 1179]. Labels [1776]. Laborassistenten [1179]. Laboratories [842, 1571]. Laboratory [1209, 1735, 1670, 1078, 1232]. labours [1658]. Lacuna [816]. Lærke [1521]. Lafayette [1574]. Laird [1266]. Lalli [1522]. Lambert [146]. Lamberts [146]. Landowners [1774]. Landscape [1226]. Language [1464, 1632, 1599, 629]. languages [594]. langue [629]. Lapidaries [1293]. Laplace [406]. Largent [1423]. Larisa [782]. Larmor [1210]. l’asile [1543]. Last [1695, 1698, 326, 505, 89]. Lat [876]. Latîf [768]. Late [1683, 1202, 1640, 1534, 291, 108, 96, 1768, 1388, 1414, 159, 1058]. Later [329, 405]. Lateral [1060]. Latin [1215, 1086, 209, 1560, 896, 74, 524, 736, 1593, 479]. Latin/German [74]. latinus [1631]. Latitude [325]. Latitudes [729]. Latour [1727]. Lattice [920, 544]. launch [1529]. Laura [1155, 1580]. Laurence [1251]. Laurent [1715]. Laurentianus [994]. Lavoisier [1294]. Law [746, 1747, 772, 539, 1263, 586, 405]. Lawlor [1301]. Lawrence [1206, 1685, 1667, 1390]. Lawrence-Mathers [1685]. Laws [1353, 595, 1781, 494, 1150]. Layers [1549]. Lazaris [1743]. lead [194]. Learned [1323, 1828, 1396]. Learning [1406, 1795, 1443, 1540]. Least [806]. Lebenskraft [226]. Lebenstein [366]. Lebensteins [366]. Lecture [1430, 1303, 1739, 1373, 1740]. Lectures [608]. Lefèvre [1542, 1748]. Legacy [982, 440, 489, 1654]. Legal [1141]. Legendre [1037, 1159]. legends [138]. Legitimation [698]. Lehner [1334]. Lehoux [1298]. Lehre [1029]. Lehren [139, 54, 181]. Lehrern [148]. Lehrsätze [254]. Leibniz [1127, 517, 1638, 386, 514, 1687, 456]. Leibnizens [386]. Leibnizian [745]. Leiden [1600, 1636, 1637, 1539, 1521, 1587, 1686, 1684, 1665, 1078]. Lemaître [690]. Leme [1227]. Lemma [2, 815]. Lemmerich [1131]. Lemov [1370]. Length [1032, 580, 543, 527]. lengths [118]. lens [1659]. Lenses [756]. Leo [867]. Leonardo [75]. Leonardos [75]. Leong [1649]. Leonhart [339]. Leopoldina [1583]. Leprosy [151]. Lepton [691]. Letter [1128, 1208, 387]. Letters [169, 321, 808, 503, 658, 15, 44, 156, 1569, 250, 419, 441]. Lettre [1689]. letzten [89]. Leunissen [1250]. leur [550]. level [1090]. Levi [363, 410, 680, 1005]. Levin [1247]. Levine [1268]. lexical [1578]. librarians [1532]. Library [215, 32, 1463, 157, 594, 74]. lies [1506]. Life [803, 1127, 1421, 1282, 1725, 503, 1165, 1324, 491, 1066, 1518, 1328, 639, 1585, 1135, 1501, 5, 1129, 423, 366]. Lifting [435]. Light [525, 676, 1173, 1187, 892, 1050, 558, 1256, 409, 1241, 1384, 1169, 1257, 1524, 1461, 1818]. Lighthouse [1658]. Lightman [1425]. Lightning [1387]. Like [1395, 1824]. Lilley [1098, 1099, 1095, 1097, 1100]. Lillian [1399]. Limits [1172, 1148, 1378, 1558, 1793]. Lindemann [830, 830]. Linear [758, 1536]. Lines [537]. Link [301]. Linnaeus [117]. Lino [1747, 1377]. Lintsen [1291]. Lisa [1165]. Lisbon [1330, 1671, 1345, 1788]. Lissa [1600]. List [821, 594]. Listening [1502]. Lists [656]. Literature [1748, 1390, 384, 158, 1451, 421, 1527, 89, 159, 291]. little [532]. Lively [1763]. Living [1657, 1045, 1763, 1765, 1271]. Llano [1747]. Local [1365, 1362, 1319, 1090]. Localities [1351]. localization [610]. Locating [1757]. Locke [169]. Lockyer [1533]. locomotion [1651]. Loeb [1033, 425]. Logic [1352, 1779]. Logodaedalus [1541]. London [1650, 1580, 1279, 1581, 1730, 1552, 1735, 840, 490]. Long [201, 779, 1151, 1364, 1321, 1353, 585, 1532, 1781, 1476, 1651]. Long-Distance [1151, 779]. Long-period [585]. Longitude [834, 433, 267]. Look [805, 1760, 1286]. Look-Alike [805]. Looking [1093, 1504, 1235]. Loop [829]. López [487]. Lord [351]. lore [618, 391, 248]. Lorentz [677]. Lorenz [1520, 241, 779, 372, 593, 241]. Lorraine [1124, 1370, 1235]. Losee [473]. Loss [1203]. Lost [1370, 1723, 1791, 711]. Lösung [391]. Louis [1496, 1055]. Louvain [1148]. Low [1280, 1401, 1247, 1531]. Loyd [608]. Lucas [389]. Lucidarius [167]. Lucier [1164]. Lucretius [763, 875]. Ludvig [170, 1520, 779]. Ludwig [250, 250]. Luft [610, 165]. Lumières [1689]. lunaire [994]. Lunar [317, 734, 801, 990, 767, 363, 410, 888, 825, 401, 442, 328, 728, 794, 989, 1008, 358, 274, 576, 916, 994]. Lunari [851]. Lunbeck [1150]. Luso [1645]. Luso-tropicalism [1645]. Lüthy [1539, 1280]. Lvnæ [359]. Lykknes [1641, 1369]. Lyman [967]. Lynn [1207].

M [1607, 1021, 1072, 437, 1584, 1747, 1667, 487, 192, 1539, 1423, 513, 1189, 1217, 490]. M. [192]. Mādhava [740]. Mānava [703]. Mūsā [868]. MA [1666, 1600, 1495, 1518, 1670]. Maarseveen [1217]. Maas [1166]. Mach [171, 292, 1033, 250, 293, 419, 171]. Machine [1297, 1821]. Machines [1225, 1271, 1136]. Mackenbach [1711]. Macmillan [1519, 1533]. Made [1065, 1230, 1740, 1626]. Madeira [1350]. Madrid [1496, 1506, 1644, 1730]. Mads [519]. Magic [1440, 202, 1822, 1149, 1149]. magistri [209]. Magnet [852]. Magnetic [1069, 251, 1120, 633]. Magnetism [1121, 1717]. Magnitudes [716]. Maimonides [1003]. main [560]. Maintenance [1764]. major [1457]. Majorana [831]. Make [40]. Making [1749, 1747, 1507, 1177, 1673, 1377, 1595, 1186, 1143, 1291, 1414, 840, 1492, 1761, 1589, 1645, 1785, 1593, 1830, 1542, 1497]. Malaria [1698]. Malcolmson [1348]. Malerei [53]. Males [1235]. Malleability [1438]. Malus [406]. Mammoth [1415]. Man [1438, 1604, 1413]. Management [1722]. Mandelbrote [1745]. Manifesto [1029, 1449]. manifold [1654]. Mansūr [349]. Manuscript [583, 582, 1054, 1631, 74, 5]. Manuscripta [92, 92]. Manuscriptes [74]. manuscripts [46, 1535, 96, 594, 291, 167]. Manuscriptum [74]. manuscrit [5]. Manuskripte [46]. Many [1700, 1692, 1416, 1654]. many-sided [1654]. Mao [1697]. Map [689, 885, 1529]. Mapping [1652]. Mara [1398]. Maragha [679]. Marc [555]. Marcaida [1749, 1541]. Marcantonio [555]. March [110, 90]. Marchia [1001]. Marci [468, 529]. Marcial [1644]. Marcianus [1631]. Marcus [468, 529]. Margaret [1584, 1253, 1371]. marginalité [1523]. marginalities [1662]. Margócsy [1383, 1686]. Mari [1586]. Mari-Tere [1586]. Maria [1182, 1189]. Marian [780]. marine [1596]. Mario [1079].

Mariska [1250]. maritime [1477]. Marivetz [1058]. Marjorie [1519]. Mark [1427, 1666, 1133, 1136, 1129, 1423, 1384, 1180, 1686]. Market [1232]. Marr [1749, 1541]. Mars [738, 968, 1679, 1824, 310, 1395]. Marsh [343]. Martin [1687, 601, 1076]. Martina [1247]. Mary [1281, 1352, 1368, 1339, 1779, 1787, 1381]. Masculus [23]. Mass [881, 405]. Massimo [1687]. Master [163]. Masters [1789, 1397]. Matemáticas [1605]. Material [1280, 1388, 1772, 1629, 1625]. Materialisten [220]. materialists [220]. Materialized [1626]. Materials [1600, 1182, 18, 1232]. Materieteilchen [518]. Mateu [1081]. Mathematical [752, 1403, 663, 1126, 761, 759, 1121, 102, 722, 1405, 835, 1693, 1192, 1019, 327, 504, 593, 1073, 505, 1054, 589, 639, 1542, 1535, 594, 1715]. Mathematician [528, 232, 1055, 1567]. Mathematicians [1191, 622, 1320, 1183]. Mathematics [590, 733, 448, 776, 321, 1331, 370, 1027, 1037, 1086, 1240, 587, 708, 894, 572, 1320, 1174, 900, 1664, 562, 964, 1462, 830, 1230, 1073]. Mathematik [830, 562]. Mathematiker [528, 232]. Mathématique [663]. Mathématiques [1240]. mathematische [594]. Mathematischen [102]. Mathers [1685]. Matilde [1671]. Matrix [673, 617]. Matteo [1248]. Matter [1749, 1764, 1781, 1123, 1210, 1683, 1200, 1036, 1378, 1633, 1793, 518, 1353]. Matters [1038]. Matthew [1580, 1127, 1607]. Matthias [1076, 1583]. Matthis [1400]. Maurolico [1054]. Mausoleum [1765]. Max [293, 920, 293, 90]. Maximilian [1394, 1799, 1069, 15]. Maximilians [15]. Maxwell [661, 676, 718, 534]. Maxwellian [556, 559, 661]. May [52]. Mayer [507, 385, 1651]. Mayor [102]. Mayow [252, 252]. Mazliak [1715]. McEvoy [1212]. McGrath [1339]. McKay [1060]. McLaughlin [1219]. MD [1517, 1639]. Me [1201, 1483]. Mean [1034, 331, 576, 650, 231, 335]. Meaning [1747, 918, 380, 1480, 882]. Means [1708]. Measles [1701]. Measure [633]. Measurement [991, 1114, 1497]. Measurements [1228, 888, 610]. Measurer [408]. Measures [1740]. Measuring [1221, 776, 1228, 633, 1675]. Mécanique [550]. Mechanical [1210, 444, 371, 163, 1259, 560]. Mechanics [168, 600, 1085, 625, 1340, 1794, 1179, 494, 550, 1581, 484]. Mechanik [494]. Mechanikern [1179]. mechanischen [560]. Mechanism [483, 664, 1203, 1578, 1440, 1347, 1822, 1782]. mechanistic [405]. médecine [20]. Media [1413, 1417, 1676]. Mediaeval [449]. Medical [437, 131, 440, 489, 1101, 1092, 1289, 1569, 1759, 1531, 395, 148, 84, 240, 12, 253, 116, 188, 291, 5, 5, 447]. Medicin [181]. Medicinae [74]. Medicine [1138, 1330, 1602, 1576, 1272, 1323, 1081, 202, 1092, 1289, 1194, 1750, 1741, 1547, 41, 1797, 74, 139, 12, 181, 193, 254, 1582, 18, 68, 1665, 1746, 1379]. Medieval [1406, 1086, 1539, 458, 1726, 1822, 1535, 1537, 1474, 1536, 764, 1685, 421, 594, 322, 579, 1111, 198, 697, 1293, 1000, 629, 684, 720, 1440]. médiévale [629, 684]. Mediterranean [1724, 1737]. Medium [1059]. Medizin [41, 1804, 1400, 139, 193, 254]. medizingeschichtlich [253]. medizinische [395]. medizinischen [148, 240, 291]. meeting [52]. Meetings [49, 161, 20, 1059]. Meets [1748, 1319]. Mega [1545]. megawatt [1675]. Mehl [1579]. meiner [335]. Melancholia [1301, 1805]. Melanchthon [179, 179, 1044]. Melanins [1367]. Meldercreutz [610]. Meli [1578, 1085]. Melissa [1602]. Mémoire [493]. memories [1501]. Memory [1262, 1275]. men [1658]. Ménagerie [1735]. Mendel [481, 611, 285]. Mendeleev [1511, 1515]. Mendelsohn [501]. Menschen [46]. menschlichen [53]. Menso [1086]. Mental [1132, 1524]. Mentality [1132]. mentiras [1506]. Menton [1413]. Mercury [134, 322]. Mere [1388, 5]. Meridian [580, 1740]. Merkspruch [47]. Merkur [134]. mesmerism [1680]. Meson [596]. Mesopotamia [135, 94, 85, 120, 145, 684]. Mésopotamie [684]. Mesotron [742]. Messengers [1280]. Messungen [610]. Metabolic [1765]. Metallurgy [64]. Metals [61, 500]. Metaphysical [411]. Metaphysics [1134]. Meteorology [1595, 1685]. Method [752, 657, 324, 308, 835, 677, 633, 300, 728, 1353, 915, 720, 1781, 182, 610, 644, 1192, 515]. Méthodes [523]. Methodological [1342]. Methodology [1123, 411, 496]. Methods [752, 764, 1035, 1080, 523]. Metre [1742]. Metric [1742]. metrology [782]. Mexican [1731]. Mexico [1769, 1768]. Meyer [807]. Miami [1737]. Michael [1344, 1214, 1370, 1176, 1354, 1399, 1267, 1424, 1359, 1716, 1811, 334, 1281]. Michel [1746]. microbes [1557]. Micropaleontology [599]. Microscope [1746]. Mid [388]. Mid-Eighteenth [388]. Middle [1253, 1436, 1226, 159, 27, 64, 150, 158, 291, 376, 24]. midpoint [4]. midwifery [108]. Might [1043]. Miguel [1575]. Mihira [393]. Milam [1235]. mile [1698]. Miles [580]. Milicent [1310]. Military [1315]. Millennium [886, 1410]. Miller [1216]. Mills [1515]. Mind [1021, 1009, 1370, 1217, 71, 1020, 1791, 1518]. Mineralogy [420, 1481]. Minerals [1227]. Minerva [1748]. Mines [1589]. Miniati [1398]. Mining [1378, 1793]. Minsoo [1271]. Minutes [349]. Miriam [1247]. Mirror [842, 1037]. Misāha [811]. misery [1462]. Misinterpretation [418]. Misinterpretations [606]. Misinterpreted [831]. Missile [1214]. missiles [1135, 1214]. Missing [1331]. mistake [882]. mistranslation [1031]. Misunderstood [683]. Mitchell [1329]. Mithraic [739]. Mitte [27]. Mittelalter [291, 24]. Mittelalterliche [594]. mittelalterlichen [421]. Mittelalters [150]. mittelhochdeutschen [376]. Mittelpunktkegelschnittes [4]. Mittelstrass [463]. Mittler [1175]. mittleren [231, 335]. mnemonic [47]. mobility [1590]. Model [177, 1225, 754, 761, 777, 363, 410, 879, 882, 1704, 405, 1150]. Modelling [1172, 1693]. Models [1210, 579, 1111, 1721, 1005, 328, 1524, 1495]. Modern [1749, 1042, 1358, 1203, 1541, 1702, 323, 515, 1280, 1253, 1803, 1075, 1539, 1306, 1407, 1753, 1178, 1233, 1441, 1217, 1719, 1384, 1080, 1379, 463, 1333, 1232, 1665, 1437, 1607, 1664, 1797, 1551, 1129, 1635, 254, 1472, 1475, 1505, 68, 1818, 1825, 1474, 1120, 408, 1207, 463]. Modernity [1401, 1247, 1577, 1602, 1802, 1648, 1640]. Modernization [1224, 1291, 1830].

Mogens [1521]. Mohr [651]. Molecular [1746, 659]. Molecule [1244, 741]. molecules [711, 540, 544, 741, 711]. Molekül [540]. Molekularhypothese [659]. Moleküle [544]. Molella [1401]. monde [1240, 1523, 1579]. Mondeklipsen [916]. Monge [1126]. monograph [56]. Monographie [56]. Monographs [1688]. Montgomery [1423]. Montucla [1168]. monument [1630]. Moon [988, 1725, 574, 575, 100, 249, 520]. Morange [1746]. Moreno [175]. Morgagni [121]. Morgan [595]. Morienus [439]. Moritz [713, 713]. Moroccan [1627]. morphology [240]. Morris [1401, 1247, 1343]. Mortality [1552]. Moschee [394]. Moscow [1361, 1283]. Moseley [503, 284]. mosque [364, 394]. Mosquée [364]. Most [461, 1752]. mother [5]. Motion [698, 775, 771, 769, 802, 805, 517, 301, 28, 100, 780, 520, 466, 702, 1493]. Motions [1034, 1264, 585, 576, 231, 335]. motus [468, 529]. Mount [1470]. mountain [1591]. Mountains [1590, 1589]. mouvement [109]. Move [127, 1771]. movement [1650, 109, 1678, 383, 419, 472]. Movements [425]. moves [1573]. movimiento [1575]. Moving [791, 592, 1351]. moyenne [650]. Mrs [1350]. Ms [343, 1408, 1537]. Mss [114]. Mu [1027]. Mu’adh [995]. Muecke [1727]. Mugnai [1687]. mujeres [1506]. Mukhula [365]. MUL.APIN [1023]. Müller [1155]. Multidisciplinarity [1123]. multilingual [1574]. multilinguistic [1631]. Multiple [1710, 1508]. multiplication [717]. Multiplikationstafeln [717]. Munajjūn [533]. München [799]. Mundane [1083]. mundo [1575]. Mundus [1043]. Munich [799]. Municipal [74]. Munns [1725, 1444]. Muñoz [1605]. Muon [623, 1107]. muqābalah [380]. Muqarnas [776]. Murphy [1582]. Murs [1408]. Musée [35]. Museum [1764, 320, 1486, 1775, 1765, 1760, 1484, 1761, 1485, 35, 132, 663]. Museums [1773, 1762, 1309, 1771, 1485]. Music [1054]. Musical [882]. my [335]. Myanmar [1698]. Myles [1393]. mystery [1016]. Myth [1439, 772, 323, 158]. Mythological [1707].

N [1150, 239]. N. [1686]. Nabseth [502]. Nach [139, 138, 717]. Nachfolge [799]. Nachlaß [215]. Nackten [618]. Nadja [1233]. Näherungen [4]. Nahrung [188]. Naked [143, 618]. name [638]. named [290]. Namen [638]. Names [1776]. Nancy [1323]. Nanohistorical [1403]. Naples [1708]. Napoca [1545]. Napoleon [1821, 1297]. narrative [1560]. narrative-flipping [1560]. narratives [1150]. Nasīr [816, 632]. Nasir [384]. Nasr [349]. Nathalie [1337]. Nation [1768, 1089, 1491]. National [235, 889, 1666, 1033, 1795, 1214, 1592, 1180, 1443, 1773, 1613, 847, 215]. Nationale [215]. nationales [847]. nationalism [1681]. Nationality [1090]. Nationalization [1329]. Nationalsozialismus [1804, 1400]. NATO [1274]. Natur [667]. Natural [201, 1332, 1342, 1059, 488, 862, 1368, 1213, 1584, 1633, 1571, 1125, 1338, 1566, 1587, 188, 1820, 1598, 688]. Naturalist [1324]. NATURE [1338, 1160, 1427, 1161, 1163, 1207, 899, 389, 517, 1159, 1253, 1440, 1036, 1267, 1719, 445, 1368, 1162, 1822, 1437, 1565, 1820, 667, 699, 1203, 1533].

Naturel [1337]. Naturens [1190]. Naturforschende [498]. Naturforscher [1059]. Naturforseherin [188]. Naturphilosophie [699]. Naturwissenschaft [146]. Naturwissenschaften [41]. Naumann [44, 44]. nautical [1478, 1475]. Nautics [446]. Naval [1753]. Navarrete [1768]. Navarro [1577, 1605]. Navicula [1002]. Navigational [433]. Nawbakht [868]. Nayrızı [896]. Nazi [1183]. Neal [1387]. Near [486]. nébuleuses [1523]. Needham [470, 446]. Negotiating [1741]. Neither [1499]. Nemorarius [346, 346]. Neptune [69, 1670]. Nerée [1463]. Netherlands [1021, 1524, 1600, 1637, 1539, 1688, 1521, 1291, 1587, 1217, 1686, 1665, 1284, 860, 1830, 1741, 864, 859, 1598, 447, 1333]. Network [1391, 1310, 1658]. Networking [1139, 1137]. Networks [1774, 1141]. Neuburger [90]. Neue [276]. Neuenschwander [1739, 1740, 1303, 1373]. Neues [78]. Neugebauer [549]. Neumann [1012]. Neumanns [1012]. Neuroscience [1372, 1792]. Neusis [537]. Neusis-Constructions [537]. neutrality [1613]. neutron [1512]. Neuzeit [1075]. neuzeitlicherer [254]. Never [1701, 1157]. Never-Ending [1701]. Newell [1603]. Newman [1125]. Newspaper [1146]. Newspapers [1418, 1144, 1412, 1143]. Newton [1745, 233, 471, 1269, 1193, 1192, 405, 1173, 1187, 1332, 1342, 1633, 761, 765, 769, 777, 802, 805, 836, 389, 856, 1047, 564, 199, 149, 283, 539, 588, 882, 520, 589]. Newtonian [1352, 1581, 1779, 1046, 411]. Newtonianism [1745]. Nexus [1082, 1608]. Niccolò [1192]. Niche [1309]. Nicholson [1528]. Nick [1380]. Nickelsen [1725]. Nicolaïdis [1640]. Nicolò [1493]. Nicomachus [694, 694]. Niels [338, 80, 338, 87, 80, 1169, 636]. Nielsen [1359, 1419, 461]. Nieto [1507, 1643, 1081]. Nieto-Galán [1507, 1643, 1081]. Nikolai [1366]. Nina [1082]. Nineteenth [1325, 144, 1768, 1355, 1414, 1372, 1327, 1544, 1602, 1784, 1520, 1651, 1828, 1769, 70, 1104, 1153, 1776, 787, 1792, 371, 894, 1091]. Nineteenth-Century [1355, 1414, 1327, 1768, 1602, 1784, 1520, 1828, 1769, 1776, 1792, 1091]. Ninth [505, 5]. Nirukti [180]. Nıshābūrı [1293]. Niskanen [1716]. NJ [1647]. No [320, 1230, 1388, 1656, 871, 1149]. Nobel [1331]. Nobili [760]. Nocks [1165]. nomenclature [46]. Nomenklatur [46]. Non [879, 915, 718, 1536]. Non-Hipparchian [879]. non-linear [1536]. Non-standard [915]. nor [1499]. Norden [1357, 1819]. Nordic [1341]. normalise [405]. Norman [1533]. North [1069, 1309]. Northern [1755]. Norton [1072, 1150, 1264]. Norway [1634]. Norwegian [1419]. Note [858, 1718, 199, 160, 538, 843, 550, 178, 197, 1465, 493, 550]. Notebaart [436]. Notebook [351]. Notes [184, 189, 203, 235, 783, 786, 788, 1116, 664, 278, 1408, 1128]. Notices [187, 191, 206, 214, 230, 296, 306, 316, 362, 373, 382, 416, 477, 492, 509]. Notion [1290]. nouvelle [663]. Nova [524]. Novel [1784, 1355]. November [1171, 1185, 1196, 1220]. Nuclear [1722, 1206, 1299, 860, 861, 1142, 1218, 1627, 1489, 1630, 1491, 1625, 1488, 1490, 1823, 1492]. nuclearity [1630]. Nullpunkt [125]. Number [217, 357, 717]. number-theoretic [357]. Numbers [1267, 572, 499, 505, 784, 1249]. Numeral [83]. Numerology [709]. Nunes [1478]. Nuremberg [1582]. Nursery [1251]. Nutrition [1419, 1735, 188]. NY [1664, 1645, 1668]. Nye [1281]. Nyhart [1207].

O [1545, 239, 1408, 1322, 457, 486]. oak [291]. Oakland [1586, 1601]. Obituaries [122, 19, 9, 110, 175, 90]. Obituary [183]. object [1577, 1629, 1628]. objective [1499]. Objectivity [1124]. Objects [1773, 1280, 1771, 1772, 1085, 1480]. Oblique [988, 198]. Obliquity [319]. obscure [893]. Observation [573, 326, 128]. Observational [1410]. Observations [1069, 734, 964, 609, 640, 1350, 768, 880, 704, 182]. Observatory [1088, 679]. Observe [1465]. Observers [1345, 1310, 1788]. Obstacles [1163]. occasion [52]. occidentalizare [1545]. Occurrence [436]. ocean [1616]. oceanic [1476]. Oceanographic [1274]. October [1229]. ocular [1688]. Oculum [449]. Oeders [27, 27]. Oersted [699]. Oertzen [1649]. of/from [1560]. Ofer [1343]. Office [1088, 1089]. Officers [1479, 1621]. Official [1063]. Ohm [705]. Oil [1676, 1164, 1474]. Oils [120]. Ökonomien [1279]. Olaf [576]. Old [1747, 1154, 1760, 1038, 254, 173, 1016, 78, 291, 1422]. Oldest [885]. Ole [1255, 1259]. Oliver [1507, 661]. Olivier [1325, 1055]. Olson [1074]. om [1190]. Omens [1267]. Omnia [360]. Omodeo [1587]. Oncology [1681]. ondes [718]. One [689, 666, 1416, 1028]. Only [1055]. Onnes [1078]. Oosterhoff [1749, 1541, 1542]. Open [1691]. openings [138]. Opera [360]. Operations [1321, 1006, 1006]. Ophthalmological [461]. Ophthalmology [1688]. Opitz [1369]. Oppenheimer [899, 1424]. Opponent [745]. opportunity [1486]. Optical [299, 73, 114, 405]. Optics [1325, 1173, 1187, 758, 534, 1339, 1457, 1818, 555, 1384]. Optikunterricht [555]. Optimism [1245]. optimist [1518]. Oral [1287, 1597]. Oran [1354]. Orbit [773, 267, 249]. Orbits [710, 442, 134]. Order [777, 1742, 1781, 1582, 1826, 1300, 1353]. Ordering [1324]. Ordre [1337]. Oreopithecus [1417]. Orfila [1081]. Organic [1585, 730, 1421, 741]. Organicism [1387]. organiques [741]. Organization [1721, 1382, 1790]. Oriental [60]. Orientation [571]. Orientations [1740]. Origin [93, 1769, 835, 283, 281, 426, 436, 463, 1486]. Original [646, 652, 105, 424, 397]. origine [1523]. origines [741, 244]. Origins [644, 1358, 1281, 1039, 744, 1216, 1181, 741, 1811, 1581, 1825, 244, 1721]. Orinoko. [1175]. Ornithologen [44]. Ornithologen-Briefe [44]. Ornithological [44]. O’Rourke [1519]. Orphaned [1627]. Orson [246]. Ørsted [387, 917, 699, 1190]. Oscillators [600]. Oskar [807, 807]. Osler [1253]. Osmotic [668]. Osteogenesis [95]. Osteopsathyrosis [95]. Ostkirchenkunst [14]. Ostwald [681]. Other [820, 752, 1666]. Otis [1155]. Otto [549]. Ottoman [1396, 1828]. oubliée [302]. our [1728, 227]. outer [46]. outlines [41]. Outsider [1798, 1140, 1354]. Outskirts [1140]. Overseas [1753]. overspecification [1535]. Overview [1172, 1452]. own [1728]. ownership [1686]. Oxford [1577, 1542, 1668, 1642, 1648, 1687]. Ozanam [675].

P [1711, 1657, 1444, 1401, 608, 1725, 485, 427, 1656, 309, 335]. P. [309, 1497, 1496]. pâna [1545]. PA [1607, 1602, 1541, 1603, 1585, 1578, 1582, 1649]. Padua [148, 148]. Paes [1227]. Pagan [1344, 1816]. page [1607]. Pageant [32]. painting [53]. Palaeontology [495, 1415]. Paleontological [1776]. Paleosyndemics [1703]. Palgrave [1519]. Palmerino [1266]. Palomares [1630]. Pañcasiddhāntikā [393, 466, 393]. Pandemias [1747]. Pandemic [1747, 1696, 1560, 1558]. Pandemics [1747, 1697, 1762, 1556]. Paolo [1665]. Papanelopoulou [1188]. Paper [1143, 1649]. paperback [1573]. Papers [424, 215, 1618]. Papiere [215]. Papyri [327]. Papyrus [738, 124, 663, 566]. Parabola [336]. Parabolas [710]. Paracelse [34]. Paracelsus [96, 277, 96, 277, 34, 489]. Paracelsus-Exzerpte [96]. paradigm [1461]. Paradoxes [915]. Parallel [747]. Parame’svara [1409]. Parameters [93, 679, 527, 720]. Parasitism [227]. Pardies [302, 244, 263, 375]. Paris [1543, 1606, 1523, 1579, 1335, 1279, 215, 1258, 1537, 92, 1735, 1731]. Pariser [215, 92]. Park [1415, 1422]. Parshall [1073]. Part [448, 1209, 1262, 1275, 556, 559, 1173, 1187, 1332, 1342]. Partial [801]. Particle [1134]. particles [518]. Particularism [1438]. Parts [744, 46]. Partus [23]. Pascal [1127, 587]. Paschalis [604]. Paschoud [1337]. Passage [1384, 893, 1818]. Passion [475, 1482]. Päßler [1175]. Passport [1742]. Past [1766, 1288, 1703, 1669, 1285, 1286, 1289, 1759, 1672, 1287, 1580]. Pasteur [848]. pasts [1549]. Patents [1393, 1801]. Path [988, 1744, 1238]. Pathogen [1704]. Pathogenic [1752, 1755]. Pathologie [48]. pathology [48]. Patricia [1236]. Patrick [1347]. Patriots [1396, 1828]. Pattern [499]. Patterned [1568]. Patterns [501, 489, 1212, 1680]. Paul [1021, 1543, 1370, 1164, 1539, 1567, 1217, 630]. Paula [1724]. Pauli´ ṡsasiddhānta [329]. Pay [339]. peace [1450]. Peculiar [744]. pedagogy [1513]. Peder [385, 404]. Pedersen [457, 576]. Pedra [1418]. Pedro [1478]. Peenemünde [1214]. Peirce [1510, 812]. Pelzbuches [248]. Peninsula [1492]. Penser [1337]. Peoples [1706]. Père [244]. Pérez [1506]. Perfect [505]. Perga [479]. Period [223, 317, 590, 825, 834, 846, 879, 1075, 430, 1038, 585, 1753]. Periodic [1130, 398, 1509, 1510, 1641, 1511, 1508, 1669, 1642, 1513, 1512, 1514]. Periodica [89]. periodicals [89]. periodicity [1515]. Periods [323, 81, 1663]. Periphery [1188, 1319, 1290]. Perks [163]. Perpetuum [1011, 527]. Persian [498, 289, 583, 582, 632, 693]. Persischen [289]. Persona [1716]. Personal [1466]. perspectiva [209]. Perspective [752, 1207, 1126, 758, 1170, 859, 209, 1510, 1456, 355, 1550, 208]. Perspectives [1072, 1163, 1312, 1690, 1608, 1446, 1538, 1084]. persuasive [1449]. Pertinens [74]. Pesetroig [487]. Pest [182]. Pestkranken [15]. Pestliteratur [159]. Pests [1362]. Peter [1728, 1219, 1426, 1124, 1205, 1178, 1267, 255, 255]. Peterborough [1584]. Petersburg [1484, 1361]. Petersen [1214]. Peterson [1421]. petit [532]. petroleum [1676]. Peurbach [1493]. Ph. [179]. Phaedra [1397]. pharmaceutical [291]. Pharmaco [461]. Pharmaco-Historical [461]. Pharmacographia [426]. Pharmacopoeia [376]. Pharmacopoeias [1607]. Pharmacy [1075, 1724, 1474]. pharmazeutischen [291]. Pharmazie [1075]. Phase [616]. Phenomena [760, 1564, 659]. Phenomenon [1308]. Philip [1402, 1216]. Philippinensibus} [1391]. Phillips [1654, 1653, 1338]. philology [1503]. Philomaths [172, 388]. Philoponus [238]. Philosopher [1474, 993]. Philosophers [568, 1668, 1521]. Philosophical [1518, 1604, 33, 1527]. Philosophische [33]. Philosophy [235, 1203, 1332, 1342, 473, 1037, 587, 1714, 411, 377, 1238, 1580, 1510, 1448, 699, 1584, 463]. Phiole [210]. Phlogiston [1157]. photonics [1461]. Photons [1524]. Phrenological [1307]. Physical [775, 761, 759, 1121, 515, 446, 1364, 1614, 841, 1139, 840, 1001, 1080, 668, 1058, 1179, 886, 1604, 1483]. Physician [1230]. Physicians [1059, 1719, 490, 54, 220, 1582]. Physicist [503, 1131, 867, 1520]. Physicists [1132, 1191]. Physico [1639]. Physico-theology [1639]. Physics [1430, 634, 292, 673, 787, 261, 369, 1402, 1314, 233, 446, 1614, 1024, 1241, 424, 593, 1252, 1140, 871, 1682, 895, 1823, 1299, 1083, 1334, 424]. physics-chemistry [895]. Physik [871]. Physikalische [1133]. physikalischen [1179]. Physiker [1131, 867, 1133]. physiological [89]. physiologischen [89]. Phytotrons [1444, 1809]. Pickering [1302]. Pickstone [1178]. Pico [1148]. Picture [1317, 8]. Pieces [1768]. Pierre [324]. Pieter [375]. Pietro [1587]. Pietsch [1328]. Piezoelectricity [1083]. Pighetti [472]. pile [1489]. Pilot [1757]. Pilot-area [1757]. Pimentel [1644]. Piñero [487]. pioneering [38]. Pionier [38]. Pisan [238]. Pittsburgh [1607, 1602, 1541, 1603, 1585, 1578, 1582, 1649]. Place [787, 1122]. Places [1305, 1592]. Plague [1695, 1696, 1694, 182, 1551, 15, 159, 1553]. plan [247]. Plancarte [1768]. Planck [293, 918, 919, 293, 424]. Plane [72]. planet [1816, 1058, 1344]. Planetary [126, 177, 710, 773, 775, 771, 224, 343, 680, 1034, 152, 325, 442, 679, 331, 846]. Planeten [134]. Planetenrechnung [354]. Planetis [1406]. Planets [381, 432]. Plant [546]. plants [1482, 1571, 1563]. Plato [112, 1122]. Platons [112]. Plausibility [611, 1231]. Playing [1411]. pleasant [1492]. Plimpton [572]. Pliny [688]. Pluralism [1322]. pneumatic [1564]. poem [618]. Pogrebysski [484]. point [125]. Points [988]. Poland [1660, 1662]. Polanyi [1281, 1811]. Polar [1357, 1819]. Polarity [601, 1125]. Polarographic [845]. Pole [1069]. Police [1101, 116]. Policies [1197, 1201, 1200, 1198, 1199, 1202]. Policy [1201, 1700, 1163, 1429, 1082, 1634]. Polish [1503, 1502]. Political [1197, 1341, 1557, 1486, 1462, 1609, 1663]. Politicians [1739]. Politics [1747, 1376, 1394, 1697, 1076, 1206, 1592, 1315, 1142, 1380, 1327, 1516, 1643, 1799, 1471, 1501, 1625, 1613, 1490, 1544, 1827]. Politik [1804, 1400]. Polyedergestalt [544]. Polygon [777]. Polygons [710]. polyhedral [544]. Pons [1644]. Pontecorvo [1299, 1823]. Popper [1003]. Popular [1755, 1415, 1451]. Popularisation [1251]. Popularization [1205].

Popularizing [1188]. Population [1711, 1090]. Porta [1564, 1565, 1562, 1563]. Portable [330]. Porter [1021]. Porto [1350]. portraits [1676]. Portugal [1671, 1153, 1224, 1226]. Portuguese [842, 1477, 1681, 1475, 894]. Portuondo [1189]. Posidonius [408]. Position [1004, 148, 829]. Positive [723]. Possibilities [525]. Possibility [628]. Post [1197, 1201, 1200, 1198, 1199, 1319, 1386, 1202, 1284, 1142]. Post-Colonial [1319]. Post-romantic [1386]. Post-War [1284, 1142]. Post-WW2 [1197, 1201, 1200, 1198, 1199, 1202]. postgenomics [1467]. Postulate [1050]. Postwar [1303]. Posy [992]. Potassium [601]. Potential [705, 1012]. Potentialtheorie [1012]. Pouch [1732]. pour [302, 263]. Pourquoi [718]. Power [761, 863, 860, 865, 861, 1186, 1241, 864, 1482, 1673, 1643, 1491, 1494, 1674, 1562, 540]. powers [1006, 1295]. Poxes [1701]. pp [1607, 1650, 1577, 1542, 1545, 1584, 1664, 1645, 1524, 1580, 1683, 1602, 1666, 1685, 1641, 1640, 1586, 1519, 1576, 1540, 1541, 1520, 1600, 1574, 1667, 1643, 1647, 1517, 1603, 1585, 1636, 1543, 1639, 1601, 1637, 1638, 1606, 1523, 1539, 1578, 1669, 1579, 1688, 1521, 1689, 1582, 1573, 1575, 1668, 1642, 1671, 1518, 1587, 1646, 1605, 1670, 1583, 1644, 1604, 1649, 1648, 1651, 1581, 1687, 1544, 1686, 1522, 1684, 1665]. pp. [1507, 1495, 1506]. Practica [1147]. Practical [776, 1306, 1731, 1505, 589]. Practice [1408, 1170, 1410, 1368, 1437, 84, 1820, 1539]. Practices [1365, 1405, 1407, 1209, 1737, 1476, 1568, 1182, 1592, 1649]. Practicing [1765]. Practitioners [1405]. Pragmatism [812]. Prague [1229, 845, 1615, 1059]. praktischen [84]. pratique [1689]. Pratt [246]. Pre [1021, 1775, 628, 1020]. Pre-Conquest [1775]. Pre-Euclidean [628]. Pre-Statistical [1021, 1020]. precession [298]. precursor [17]. Predecessor [279]. Prediction [1161, 1032, 596, 323]. Preface [1385]. Prehistoric [1039, 1412]. Prehistory [624]. Preliminary [144, 410, 1631]. Preliterate [60]. Prelude [881]. Premature [601]. premier [663, 650]. premières [263]. preparation [1733]. preparations [518]. preparedness [1558]. Prescription [1403]. Prescriptions [447]. Present [1766, 437, 1549, 1075, 484, 1580, 277, 1669, 1747]. preservation [254]. Preserved [343]. President [1208]. Presidential [1403]. presidents [1466]. Press [1607, 1650, 1577, 1542, 1584, 1602, 1666, 1685, 1586, 1576, 1541, 1574, 1667, 1643, 1647, 1517, 1603, 1585, 1639, 1416, 1497, 1601, 1495, 1578, 1145, 1582, 1573, 1755, 1415, 1414, 1668, 1642, 1518, 1646, 1670, 1649, 1648, 1651, 1581, 1687, 1544]. Presses [1579]. Preußen [1175]. Previously [105]. Price [406, 483]. Prigogine [608]. Princeton [1647]. Principal [426]. Principe [1667]. Principia [1633, 836, 856]. Principle [646, 758, 806, 331, 652, 778, 586]. principles [560, 240]. print [1542, 241]. printed [46]. Printing [851]. Prinzip [806]. Prinzipien [560]. Prinzips [778]. Prior [586]. Priority [567]. Private [1770, 1771, 1733]. Prize [1739, 1331, 1740]. Probabilism [1043]. Probability [1246, 369, 119]. Problem [761, 802, 1438, 433, 538, 493, 995, 1028, 663, 404, 1456, 391, 1679, 684, 1744, 500]. Problema [311]. Problème [493, 663, 1579, 684, 562].

Problems [710, 525, 131, 683, 1054, 209, 1448, 562, 391]. Procedure [773]. Procedures [991, 1114]. Process [1650, 1163]. Processes [502, 135, 1565]. processing [1474]. Produced [387]. Producing [1307]. Production [1351, 1600, 1677]. Productive [1662]. Professional [1419]. Professionalisation [983]. Professor [339, 342]. Prognosis [1759]. Prognostication [1300, 1826]. Program [673, 1596, 1610]. Programme [672]. Progress [67, 66, 1675]. Project [79, 1058, 751]. Projection [689, 551]. Projective [752]. Prolusione [1246]. promise [1451, 1450, 1449]. promises [1448, 1446]. Promising [1447]. promoting [1491]. Prony [803]. Proof [1403, 577, 1265]. proofs [113, 1665]. Propagation [79]. Proportion [32]. proportione [468, 529]. Proportionibus [346, 346]. Proportions [628]. Proposals [1363]. Propositions [836]. Prospects [1070]. Prosperity [1065]. Proto [782]. Proto-Ionic [782]. Proust [1683]. Prout [337]. Provenance [1769]. Provencal [629]. provencale [629]. Provincial [1773, 1771]. provisional [594]. Provisions [1065, 164]. Proxy [1228]. Prozac [1301, 1805]. Prussia [1713, 1065, 1090]. Psychiatry [202, 1601]. psychological [1449]. psychologie [1606]. psychology [1518, 1825, 1358]. Psychophysics [1712]. Ptolemaeus [134]. Ptolemaios [13]. Ptolemäus [128]. Ptolemy [689, 319, 128, 134, 1034, 1031, 640, 267, 328, 995, 880, 635, 798]. Public [1146, 1770, 1365, 1230, 1392, 1145, 1282, 1245, 1728, 1666, 1786, 1633, 1484]. publication [1532]. Publications [987, 245, 1055]. Publicity [1288]. Published [105]. Publishers [1540]. Publishing [1524, 1569]. Puerperal [279]. puissances [1006]. Pulte [1745]. Pun [505]. puppiforme [711]. puppiformes [711]. Purdue [1574]. Pursuing [1241]. Pursuit [1159, 1120, 1311]. Putscher [437]. Putting [1010, 1529]. Pyramids [571]. Pythagorean [207, 572].

Qūī [550]. Qūhī [550]. quadrants [429]. Quadratrix [324]. Quadratur [78]. Quadratures [1025]. quadrierbaren [1017]. Quaestiones [209]. Qualitative [177]. Qualities [283]. quand [650]. quanta [1524]. Quantenhypothese [264]. Quantenmechanik [1794, 1184]. Quantitative [89]. Quantities [818]. Quantum [1430, 918, 623, 1107, 600, 919, 630, 1334, 672, 424, 742, 512, 264, 895, 1794, 1234]. Quarantine [1555]. Queen [1142]. Quellen [376]. Querner [438]. Quest [1444, 1397, 1809, 1789, 1558]. Question [1769, 991, 1114, 1300, 1826]. Questions [1287, 208]. quina [1473]. Quinlan [1339]. Quinlan-McGrath [1339]. Quirino [831]. Qurra [224]. Qusayr [850].

R [987, 1177, 1125, 1354, 1269, 1440, 1130, 1247, 434, 500, 279, 1345, 1443]. Race [1712, 1438, 1576, 1801, 1393]. racial [1645, 1439]. racine [532]. Radio [1284, 1501, 1501, 1502]. radioactivity [1512]. radioisotopes [1629]. radium [1182]. Radius [685]. Raged [1695]. Raina [1336]. rainbow [555]. raison [650]. Rajender [1141]. Ramakers [1333]. Ramazzini [1570].

Ramée [309, 309]. Ramon [1749, 1541, 1081]. Raphael [1517]. Raphaële [1541, 1521]. rapprochement [650]. Rates [660]. rather [893]. Rational [700, 657, 71, 1429]. rationality [1791, 1370]. Ray [284, 722, 502]. Raynaud [1688]. Rayner [1669]. Rayner-Canham [1669]. Rays [251, 1593]. Raz [1343]. Re [880, 1548]. Re-analysis [880]. re-emerging [1548]. Reactions [660]. reactor [1627, 1489]. reactors [1491]. Read [1465, 1454]. Readers [1633, 1537]. Reading [1238, 1517]. real [1240]. Realism [1003, 1322]. Realität [704]. Reality [772, 704, 1134]. Realms [1045]. Rearranging [1080]. Reason [1370, 650, 1791]. Réaumur [606, 1368, 1820]. Rebecca [1370]. recalcitrance [1577]. Recasting [1349, 1817]. Received [11, 43, 141, 187, 191, 206, 214, 230, 243, 260, 296, 306, 316, 362, 477, 492, 509, 522, 542, 557, 613, 627, 643, 654, 670, 682, 696, 715, 725, 750, 797, 810, 824, 839, 866, 870, 891, 878, 898, 923, 949, 950, 951, 952, 953, 954, 955, 956, 957, 958, 959, 960, 961, 962, 963, 924, 925, 926, 927, 928, 929, 930, 931, 932, 933, 934, 935, 936, 937, 938, 939, 940, 941, 942, 943, 944, 945, 946, 947, 948, 980, 985, 1015, 1052]. received [22, 30, 58]. Recently [121]. Reception [1047, 1745, 1515, 460]. receptions [1604]. recette [5]. recipe [5]. Recipes [1306, 1635, 1564, 1565, 1566, 255]. Reciprocals [379]. Recognition [1742]. Reconceptualizing [1509]. reconciliation [650]. Reconfiguring [1253]. Reconsideration [337]. Reconsidering [1234]. Reconstructed [577]. Reconstruction [1410, 1715]. Records [16]. Recounting [1089]. Recovering [720]. Recruitment [1288]. Rectification [312, 671]. Redeeming [1766]. redemption [1447]. Reden [1583]. rediscovered [92]. Reeds [1038]. réel [1240]. Rees [313]. referees [1534]. reference [635]. Reflected [1413]. Reflections [1160, 829, 854, 1555, 1504, 475, 1690, 1466, 1468]. Reform [1295, 1813]. Reformation [1582, 1741]. Refraction [478, 406, 995, 589, 586]. Reframing [1309]. Refrigeration [1078]. Refutation [1263, 1679, 711, 540]. Regel [219]. regelmäßige [518]. regelmäßigen [98]. Regenbogens [555]. Regime [1367, 1785, 1377]. Regiomontanus [1718, 528]. Region [1357, 1819]. Register [1089]. Regius [1714]. Regular [726, 589, 98, 518]. regulations [1555]. Rehabilitated [1263]. Reich [1402, 1780, 1133]. Reichborn [9]. Reichborn-Kjennerud [9]. Reinhard [1183, 1084]. Reinhardt [1080]. Reinterpretation [1417]. Reis [1751]. Reise [27]. rejection [1459]. Related [581]. Relation [752, 546, 896, 1633]. Relations [223, 63, 645, 1610, 1611, 1609]. Relationship [587]. Relationships [1341, 1419]. Relativity [1181, 1522, 622, 672]. Relevance [1317]. relevant [74]. Religion [1639, 1074, 739, 1426]. religious [1447]. Remak [1155]. Remaking [827]. Remarks [820, 735, 617, 968, 86, 113, 383]. Remedies [1724]. Remote [1228]. Renacimiento [1605]. Renaissance [1148, 1376, 1586, 1076, 1587, 1323, 1339]. Rendering [1717]. René [530, 540]. René-Just [530, 540]. Renée [1517]. Renn [1647, 1334]. Rentetzi [1182]. Reorganization [1722]. Reply [1266]. Report [235, 1620, 717]. reported [704]. Reports [322]. Representation [1049, 53, 1688]. representations [1673, 1678]. reprinting [1531]. Reprints [205, 229, 242, 259, 266, 273, 315, 345, 353, 415, 422, 476, 510, 536, 547, 570, 186]. reproductive [247]. Republic [1229, 1569]. reputation [1569]. Research [1166, 1722, 831, 1634, 1121, 1743, 1362, 633, 1315, 1142, 438, 436, 1073, 1594, 1457, 1528, 1458, 1452, 1179, 545, 17, 1182, 1012, 33, 1274]. Researches [787, 303, 299, 445, 73]. resilience [1512]. resistance [712, 1492]. Resisted [805]. Resistive [761]. Resistor [731]. Resolution [771]. resources [1480]. Respect [436]. Response [1138, 1265, 1266, 1263, 1264]. Ressourcen [1804, 1400]. Results [481, 105]. Rethinking [1647, 1681, 1553]. Retrograde [1439]. rev [916]. Reveal [1405]. Revelations [439]. Reversal [601]. Review [1607, 1749, 1650, 1442, 819, 889, 890, 883, 923, 949, 950, 951, 952, 953, 954, 955, 956, 957, 958, 959, 960, 961, 962, 963, 924, 925, 926, 927, 928, 929, 930, 931, 932, 933, 934, 935, 936, 937, 938, 939, 940, 941, 942, 943, 944, 945, 946, 947, 948, 900, 922, 1711, 1577, 483, 1542, 1545, 1728, 462, 1584, 1664, 1645, 1747, 1524, 1580, 1683, 511, 1602, 1666, 700, 1685, 1444, 1507, 1641, 1712, 1640, 1586, 1519, 1576, 1540, 1541, 1746, 837, 1520, 455, 1600, 1574, 1748]. Review [1751, 456, 1667, 1643, 1647, 1290, 1517, 470, 1603, 1585, 461, 487, 1636, 1543, 472, 515, 473, 1639, 1497, 1601, 1637, 1745, 1638, 1606, 1743, 446, 435, 502, 459, 447, 1523, 1539, 1495, 471, 1578, 1496, 1669, 1579, 1688, 1521, 1440, 568, 501, 440, 489, 1724, 1567, 1713, 1689, 1725, 485, 458, 460, 1582, 1573, 484, 1714, 441, 426, 433, 424, 432, 434, 513, 503, 488, 498, 1506, 1575, 1555, 1715, 398, 491, 1668, 1642, 1441, 1671, 1518, 848, 425, 438, 427, 423, 1587, 457, 1646, 1605, 1726, 436, 1670, 1583, 499]. Review [514, 439, 1644, 1604, 1649, 1648, 1651, 490, 1750, 1581, 397, 486, 500, 1687, 512, 463, 475, 377, 560, 625, 1744, 1544, 1686, 520, 1522, 1684, 1665, 1443, 1716, 1727, 748, 1458, 247]. Reviews [1132, 1277, 1249, 1344, 10, 21, 29, 36, 42, 51, 57, 77, 82, 91, 97, 103, 111, 115, 123, 133, 140, 147, 153, 162, 176, 190, 204, 213, 216, 222, 228, 236, 258, 265, 280, 288, 295, 305, 314, 340, 344, 348, 352, 356, 361, 367, 378, 390, 399, 403, 407, 414, 431, 454, 469, 482, 497, 516, 521, 526, 531, 535, 541, 548, 553, 561, 565, 569, 597, 602, 607, 612, 621, 626, 631, 637, 642, 647, 653, 662, 669, 674, 678, 686, 687, 692, 695, 701, 707, 714, 719, 724]. Reviews [732, 737, 743, 749, 753, 757, 762, 766, 770, 781, 789, 793, 796, 800, 804, 809, 813, 823, 828, 832, 838, 849, 853, 857, 869, 874, 890, 877, 884, 922, 897, 901, 902, 903, 904, 905, 906, 907, 908, 909, 910, 911, 912, 913, 914, 969, 970, 971, 972, 973, 974, 975, 976, 977, 978, 979, 984, 1018, 1022, 1026, 1072, 1427, 1422, 1428, 1166, 1155, 1148, 1188, 1204, 1215, 1358, 1237, 1326, 1203, 1346, 1191, 1251, 1325, 1207, 1295, 1281, 1177, 1126, 1127, 1394, 1219, 1183, 1272, 1234, 1426, 1429, 1133, 1301]. Reviews [1381, 1302, 1214, 1377, 1179, 1397, 1124, 1402, 185, 1137, 1136, 1357, 1370, 1164, 1129, 1425, 1323, 357, 1268, 1128, 1086, 1240, 1206, 1280, 1253, 1352, 1356, 1081, 1184, 1401, 1176, 1205, 1299, 1061, 1154, 1125, 1383, 1135, 272, 1082, 1040, 1083, 1079, 1074, 1134, 1190, 1276, 1334, 1075, 1236, 1392, 1296, 1270, 1354, 1421, 1084, 1182, 1178, 1269, 1329, 1395, 1193, 1239, 1192, 1423, 1051, 1297, 1338, 1396, 1282, 1278, 1130, 1241, 1322, 1147, 1300, 1355, 1247, 1165, 1291, 1393, 1189, 1335, 1398, 1279, 1399, 1233, 1324, 398, 1348, 1372, 1131, 1371, 1216, 1353, 1267, 1382]. Reviews [1298, 195, 1347, 1380, 1336, 1175, 1217, 1384, 1343, 1328, 1149, 1085, 1080, 1212, 1327, 1378, 1048, 1073, 1194, 1368, 1150, 1180, 397, 1248, 1235, 1252, 1424, 1379, 1339, 377, 1271, 1340, 1218, 1400, 1250, 1333, 1238, 1232, 1181, 1174, 1345, 1349, 1359, 1195, 1369, 1337, 1213, 234]. Revision [1169]. Revisited [1034]. Revisiting [1246]. Revolution [1049, 844, 1127, 623, 1107, 1366, 672, 420, 1353, 1212, 66, 381, 1779, 1781, 1746, 1219, 1352, 1247, 1181]. Revolutionary [1177]. revolutionibus [585]. Revolutions [1252, 1468]. Reymond [1155, 1372, 1792, 1158]. RFE [1502]. rhetoric [1449]. Rhind [566]. Rhodes [1019]. Ribot [1449]. Ricardo [1645]. Ricci [192]. Rice [1073]. Richard [1542, 1541, 1794, 1184, 1074, 1398, 1181, 1749, 1794, 1637]. Richmann [755]. Richter [1788, 1345]. richtigen [53]. riddle [1732]. Riemann [760]. Ries [1359]. Rigaud [166]. Right [1005]. Rigor [835]. Ring [478]. Ringe [704]. Rings [760, 1001, 704]. Riordan [1399]. Rip [1291]. Riper [1135]. Rise [1711, 1207, 1746, 62, 1803, 1399, 843, 1728, 1664, 1486, 1582, 1814, 1441]. Rising [988, 118]. Risings [1014]. Risk [1439]. Risks [1756]. Risø [861]. Rita [1727]. Ritter [658, 658]. Ritual [966]. Ritz [967]. Rivalry [1218]. Rivière [1413]. Road [774, 822, 1615]. Robb [552]. Robert [1155, 1237, 451, 1401, 1300, 1327, 1424, 104, 472]. Roberto [1522]. Roberts [1600, 1264]. Roberval [300, 336]. Robin [1746]. Robison [591]. Robot [1165]. Robots [1440, 1395, 1822, 1824]. Robyn [1352]. Rochberg [1636]. Rock [1518]. Rockefeller [1201]. Rockets [1135]. Rodríguez [1726]. Rodríguez-Arribas [1726]. Rods [1664]. Roemer [1256]. Roger [1650, 591]. Role [1049, 600, 1416, 1309, 1731, 920, 1736, 1759, 1179]. Roles [1419]. Rollbild [137]. Rolle [1179]. Rolling [1744]. românesti [1545]. Roman [461, 825, 834, 846, 879, 435]. Romania [1545, 1431, 1501]. Romans [1298, 1806]. Romantic [1387, 1297, 1821, 1386, 699]. Romanticism [756, 1604, 1213, 1817, 1349]. romantischen [699]. Römer [1257, 1258, 1255, 1259]. Rommevaux [1240]. Ron [1249]. Ronald [1267]. Rond [1452, 1453]. Room [1765]. root [532]. Roots [1219, 1119, 1421, 708]. Roque [1645]. Rosen [1102, 1103]. Rosewater [1474]. Rossana [1715]. Rossiter [1371]. Rotation [412]. Rothmann [1009]. Rothschild [1733]. round [1588, 1070]. Rousseau [1213]. Rousset [56, 56]. route [1453]. Routes [1354, 1798]. Routine [1763]. Routledge [1580, 1507]. Rover [1395, 1824]. Royal [1348, 1455, 490, 1532, 1534, 1807, 1658]. royale [1667]. Rudolf [1155, 1075, 48]. ruin [1518]. Ruíz [175]. rule [219, 829]. Ruler [1687]. Ruse [1344]. Russell [196]. Russia [1360, 1361, 844, 1365, 1364]. Russian [484, 488, 1360, 1482, 1363, 1438, 1362]. Ruth [1195]. Rutten [601]. Ryan [1726]. Ryland [124].

S [812, 583, 459, 608, 1724, 1278, 1300, 1506, 1217, 1283, 1006, 1145, 617, 1332]. S-Matrix [617]. S. [1656]. S.J. [586]. Sūtra [703]. Sabine [1240]. Sacha [1684]. Sachse [1180]. Sacrobosco [1632]. Safāa’ih [558]. Safeguarding [1756]. sagas [290]. Sage [1481]. Sagen [138]. Sagredo [1380, 299, 1827]. saints [5]. Salary [339]. Salazar [1367]. Sally [1602]. Salomon [47]. Salvation [1753]. salves [194]. Salviati [1265]. Samaw’al [996]. Same [1259]. Sammelwerke [12]. Samuel [685, 699]. sana [1755]. Sánchez [1081]. Sanders [1510]. sanitary [1555]. Santa [1721]. Santo [1350]. Santos [1645]. Saros [1032]. SARS [1556, 1555]. SARS-CoV-2 [1555]. Sarton [110]. Satellites [1256]. Saturn [47, 641]. Savant [1327]. Savantes [1523]. Savants [1516, 1237]. Saxon [47]. Say [1695]. saying [893]. Scale [601]. Scaling [1266, 1264]. Scams [1412]. Scandinavia [151]. Scandinavian [1084, 1320]. Scerri [1668, 1130, 1642]. Schädeleröffnungen [138]. Schallgeschwindigkeit [610]. Schallortungsverfahren [610]. SCHCT [1171]. Scheffler [1746]. Scheme [198, 825]. Schemes [533, 1023]. Scheuchzer [132]. Schiemann [1203]. Schioler [435]. Schipperges [438]. Schism [440, 489]. Schlick [713]. Schlicks [713]. Schmidt [1419]. Schmidt-Nielsen [1419]. Schmieden [915]. Schmitz [1075]. Schoener [376]. Scholar [1430, 368, 1096]. Scholarly [1716]. Scholars [1229, 525, 1359, 1810, 1683]. Scholarship [1360, 1311, 408, 1683]. Scholia [992, 400]. Scholien [400]. Schöners [376]. School [1289, 845, 668]. schools [12]. Schot [1291]. Schreiben [15]. Schriften [38]. Schrödinger [616]. Schultze [1084, 1183]. Schwann [1155]. Schwarz [1349]. Sciability [1774]. SCIENCE [1338, 1197, 1201, 1769, 1063, 1146, 1344, 235, 889, 1070, 1138, 1171, 1185, 1196, 1220, 1330, 1360, 1431, 1445, 1427, 1422, 1428, 1042, 1728, 1361, 70, 1104, 1188, 1559, 1237, 1722, 1221, 1586, 1251, 1163, 60, 347, 812, 899, 1303, 1200, 261, 1434, 1365, 1033, 62, 1426, 61, 1795, 1643, 1290, 1739, 1144, 1433, 487, 1470, 515, 1312, 1432, 1425, 1729, 1206, 1401, 1205, 983, 1059, 1383, 1082, 1151, 1198, 1392, 1094, 1084, 1152, 1178, 1494, 1269, 67, 1734, 1145, 1297, 1396, 1282, 63, 221, 1278, 1714, 1245, 1199, 1355, 554, 1143, 1319, 172]. Science [433, 1189, 1373, 1142, 848, 1298, 1740, 1435, 1327, 71, 1819, 1750, 1283, 66, 463, 377, 1333, 1522, 1158, 1345, 1202, 1195, 1287, 1213, 1598, 1305, 1310, 1608, 1660, 1780, 146, 1451, 1469, 1666, 1784, 1507, 1786, 1612, 1787, 1547, 1788, 1486, 1661, 1468, 41, 1796, 1647, 1599, 1636, 1639, 1802, 1450, 1806, 65, 1808, 1597, 1448, 1449, 1446, 1481, 1811, 1573, 1471, 1671, 1616, 1816, 99, 1455, 1670, 1648, 1651, 1662, 1821, 35, 1466, 1828, 1663, 1588, 107, 1171, 1185, 1196, 1220, 1229, 1318, 1374, 1404, 1431, 1290, 1074, 1447, 1242, 1254]. Science [1273, 1292, 1490, 1208, 1260, 1283, 1460, 1580, 20, 1072, 1281, 1381, 473, 1423, 458, 491, 1267, 1343, 1212, 500, 1250, 1744, 1728, 511, 1251, 470, 1357, 1205, 446, 1074, 1236, 501, 1336, 1150, 1443]. Science-Industry [1082]. Sciences [218, 1683, 1358, 1221, 52, 1429, 515, 1405, 1743, 269, 1438, 1721, 1224, 282, 1266, 1335, 1139, 35, 1666, 1612, 1540, 1463, 1829, 1689, 1587, 1504, 1690, 1825, 1532, 1455, 1484, 1667, 1517, 1518, 1743, 1178, 488, 1369]. Scientific [1777, 1160, 1049, 1631, 1161, 1346, 1641, 1163, 790, 1177, 1127, 1200, 1219, 1748, 1774, 472, 1159, 630, 1079, 1276, 160, 1669, 1329, 1448, 1760, 677, 245, 1245, 372, 1506, 1616, 826, 438, 496, 645, 1117, 1162, 1218, 1305, 1778, 1468, 1526, 1629, 1530, 1480, 1613, 847, 1604, 1544, 1479, 1621, 1770, 1693, 1166, 434, 1398]. scientifique [1523]. Scientist [803, 1064, 188]. Scientists [1166, 1739, 1059, 1309, 1354, 1436, 1810, 1371, 1286, 1609, 1287, 1798, 1815, 1028, 1164, 1359]. SCK [1722]. Scoop [1418]. Scoops [1412]. Scott [1745]. Scribal [1517]. Script [525]. Scrutinized [1227]. Scuffles [1412]. Scull [1601]. Sculteti [1567]. Scurvy [3]. sea [1670, 433]. Seals [461]. Seamen [1753]. Search [1284, 406]. Searching [1224]. Seasonal [721]. Sebastian [1279]. Sechs [658]. secolului [1545]. Second [132, 1204, 37, 777, 539, 1247, 1594, 1070, 746, 886]. Secret [1189]. Secrets [1427]. section [1653, 4, 56]. Sections [479]. sector [4]. secular [1447]. Sedeño [1506]. Seduced [1779, 1352]. Seeber [544, 544]. Seed [143]. Seeds [1763]. Seeing [1824, 1221, 1395]. sein [699]. Seine [8, 847, 376, 540]. seiner [618, 92]. seismograph [1628]. Sektor [4]. Select [1277]. Selected [1641]. Selective [434]. Self [615, 1372, 1792, 1340]. Self-induction [615]. Sellers [1580]. Selling [1317, 1269]. selon [375]. Selskab [1520]. semiconductor [545]. Semmelweis [279]. Senenmut [968]. sensations [419]. sense [1650]. ‘Senses [1259]. Sensing [1228]. Sensory [1712]. sentinels [1556]. September [1021, 1070, 1229, 1330]. Series [740, 1317, 967, 650]. Serpent [1068, 575]. Service [685, 1447, 394]. Serving [1780, 1402]. Session [1375]. Seth [1576]. Sets [1128, 1664]. setting [1473]. Settlers [1756]. Seventeenth [1237, 1085, 634, 827]. Seville [1477]. Sex [1308, 1382, 1790]. sexo [1496]. sexologists [1471]. Sexology [1308]. Sexual [1390, 1471]. Shadows [1120]. Shank [1193, 1267, 1581]. shape [518]. Shaping [1137, 1752, 1648]. Shaul [1083, 1334]. Shazhou [1568]. sheet [78]. Shield [1274]. Shifting [1080]. Shinn [1310]. Ship [1375]. shipment [1629]. Ships [625]. Shock [1434, 1154]. Shores [1778, 1346]. Short [1442, 191, 206, 214, 230, 296, 306, 316, 362, 373, 382, 416, 477, 492, 509, 1783, 888, 1276, 256]. Short-Time [888]. Shorter [187]. Shows [1233]. Shuttleworth [1602]. Shyamala [1141]. si [1545]. Siam [1013]. Sichau [847]. Sicht [355]. Siddiqi [1296]. side [1657]. sided [1654]. Sidrak [282]. Siebenecks [98]. siècle [1237, 1144, 1689, 5]. Siegfried [1239]. Siegmund [1084, 1183]. Siegmund-Schultze [1084, 1183]. Sight [1384, 1818]. Sigmund [1065]. Significance [801, 872, 1642, 1130]. Sigval [1419]. Silent [1280]. Silke [1637, 1398, 1726]. Silvanus [1657, 1654, 1653]. Silver [554]. Simões [1671]. Simon [1600, 1149, 309, 309, 1659]. Simone [1299]. Simonin [1618]. Since [1154, 1237, 1371, 1815, 1662]. Sine [586]. Singapore [1641, 1669]. Single [1704]. Siphon [464]. Siraisi [1323]. Sisyphean [1739]. site [112]. situ} [1228]. Six [334, 658]. Sixes [888]. Sixth [20]. Size [274, 1630]. Sizes [343]. Sizing [773]. Sizing-Procedure [773]. SJ [1391]. skeleton [130]. Skepticism [1300, 1826]. Sketches [1302, 1812]. Skin [615, 1348, 338, 1807].

Skin-Effect [615]. skinned [7]. Skull [1416, 138]. Skulls [1480, 1483]. Sky [1314]. Slater [600]. Slavery [1278]. slaves [1656]. Small [321, 1312, 606, 291]. smallest [518]. Smallpox [1747, 1706, 1092]. Smaragdreise [87]. Smeaton [1658]. Smith [1650, 495, 1384]. Smoluchowski [780]. Snow [608, 1707]. Snowden [1747]. So-called [385]. sobre [1506]. Social [1281, 1219, 1246, 1144, 63, 1088, 1195, 1612, 1574, 1811, 68, 1666]. Socialism [1214, 1180]. socialist [1501, 867]. Sociedad [1747]. Societat [1171]. Société [20]. Societies [49, 161, 60, 20, 1540]. Society [1277, 1063, 1070, 1171, 1185, 1196, 1220, 1229, 1330, 1318, 1374, 1404, 1431, 70, 1104, 1747, 52, 1288, 1272, 1681, 1396, 1282, 221, 1242, 1254, 1273, 1292, 1372, 1217, 1087, 1091, 1180, 1208, 1260, 1283, 1460, 1792, 1550, 1828, 1534, 1614, 1807, 1658, 1348]. soil [1630]. Solaire [364, 450, 551, 506]. solaires [675, 302]. Solar [990, 323, 834, 879, 442, 1066, 374, 333]. Solid [493]. solida [581]. solide [493]. Solitary [1201]. Solomon [47]. Solovey [1666]. Solution [404, 391, 311]. Solutions [765]. Solvay [1234]. Solving [433, 1732]. Some [93, 820, 710, 1160, 574, 128, 449, 1680, 322, 617, 683, 624, 327, 1169, 836, 193, 1483, 1756, 215, 400, 254]. Somerville [1381, 1352, 1779, 1787]. Sommerfeld [264, 264]. Somnium [249]. Somos [1686]. Sonja [1540]. Sonnenbewegung [383]. Sonnenuhr [394]. Sophie [1247]. Sophus [519]. Soraya [1746]. Sørensen [1084]. Sörlin [1357]. Sorsby [462]. Soul [1358, 1402, 1780, 1782, 1825, 1347]. Sound [1173, 1187, 1469, 610]. Source [818, 217, 18, 458]. Sources [1632, 1246, 256, 376]. Sousa [1671]. South [1239, 1774, 921, 202]. Southern [1775]. sovereignty [1596]. Soviet [1296, 844, 1628, 1462]. soziale [1279]. Sozialist [867]. Space [552, 1296, 1321, 1725, 1176, 1574]. Spaceplane [1321]. spaces [1592, 1555]. Spain [1575, 1605, 1644, 1630, 1643, 1494, 1724, 684, 487]. Spanish [1326, 487, 1189, 1143, 1741, 1474]. Spary [1335, 1232]. Späte [96]. Spätmittelalters [159]. Special [1375, 672]. Specialists [508]. specialized [89]. species [129, 1646]. Specimens [1761]. Spectacle [1437, 1233]. Spectra [284]. Spectrography [1227]. Spectrum [389]. speculation [33]. Speculations [872, 351, 1252]. Speed [892, 1263, 1257, 994, 610]. Speed-Distance [1263]. Speeds [1266, 1264]. Spekulation [33]. speleology [1592]. Spencer [1129]. Sphaera [1632]. sphera [581]. Sphere [988, 1409, 1410, 1786, 1678, 1632, 1392]. spheres [1678]. Spherical [349]. Spiral [805, 537, 336]. Spirit [917, 86]. Splendid [455]. Splendor [1201]. Splitting [873]. Spotlight [1197, 1201, 1160, 1161, 1163, 1200, 1288, 1159, 1198, 1199, 1285, 1286, 1289, 1162, 1202, 1287, 1653]. Spots [1752]. Sprachen [594]. Spread [744, 173]. Springer [1664, 1524, 1683, 1522]. Spying [790]. squarable [1017]. Square [966, 1436, 873, 532, 805]. squaring [78, 1638]. Squelette [130]. St [1484]. St. [1361]. Stability [1122]. Stabius [1411]. Stage [1389]. Stahuis [1021]. Staley [1181]. Stalin [1366]. Stamhuis [1217]. Standard [825, 915]. Star [1720, 767, 656, 821, 1000, 729, 885]. Stark [264, 264]. Stars [1031, 480, 1678, 1603, 1493]. Start [601]. Stasis [1763]. State [1419, 1396, 1327, 1452, 1090, 1550, 1828]. state-level [1090]. States [1197, 1312, 843, 1629, 1706]. Stations [1363]. Statistical [1021, 1088, 1217, 1020, 1091, 720, 108]. Statistics [1090, 1092, 1087, 1091]. statistische [108]. Status [579, 1111, 1743, 436, 964]. Stavenhagen [439]. Stavros [1743]. stealing [1483]. Steam [1216]. Steele [1683]. Steigerwald [1585]. Stein [1065, 608]. Steinau [1416]. Stela [320]. Stellar [640]. Stelle [112]. Stellung [148]. Steno [1521]. Stensen [338, 338, 87, 276, 80, 636]. Stensenbriefe [276]. Stensens [87, 636]. Step [672, 1290]. Stephanus [993]. Stephen [1648, 1686, 1727]. stepping [1504]. stéréographique [551]. Stern [1684]. Steven [1148]. Stevin [309, 1659, 1149]. Stewardship [1663]. stiintei [1545]. Stillman [1265]. Stockholm [1615]. Stoic [408]. Stoics [634]. Stone [575]. stopped [1483]. storm [1131]. Story [1155, 1130, 1165, 1135, 1642]. Strange [1370, 1321, 1791]. Strano [1637]. Strasbourg [34]. Strategic [1197, 1202]. Strategy [475]. Stratigraphy [495]. Strazzoni [1714]. stresses [1611]. strings [1456]. Strömgren [1066]. structural [544]. Structure [722, 518, 530]. structures [1468]. Struggle [1402, 1780]. Strukturtheorie [530, 544]. Stubbornness [1020]. student [92]. Studie [1804, 1400]. Studien [386]. Studienzeit [92]. Studies [1540, 1636, 1010, 825, 834, 846, 879, 1693, 1807, 1521, 1186, 1319, 504, 1336, 1359, 386, 1810, 89, 302, 244, 375, 1660, 694, 263, 1215, 1348]. Study [1203, 1746, 831, 461, 507, 211, 1083, 1703, 865, 172, 546, 468, 529, 1631, 1615, 1733, 1494, 886, 1593, 1563, 165, 502]. Sturm [1370, 1131]. Stuttgart [1583]. Styles [1033, 1151]. suanzhi [1464]. Sub [1274]. Sub-Committee [1274]. Subatomic [1134]. Subjected [387]. subjective [1499]. Sublimation [85]. Sublime [1271]. substance [247]. substantiality [964]. subterraneus [1043]. Succession [799]. Successive [308]. Successors [1411]. Suff [270]. suggested [405]. suite [650]. suivie [5]. Súlba [703, 1025]. Śulba-sútras [1025]. Sulfides [445]. Sulfur [730]. Suman [1576]. Summaries [193]. Summary [527, 1652]. Summetria [1177]. Sums [801, 784]. Sun [988, 383, 28, 267, 332]. Sundial [365, 287, 364, 450, 506, 394]. Sundials [644, 330, 675, 302]. Super [1399, 1814, 209]. Superconducting [1399, 1814]. superheavy [1511]. Supplementary [113]. Supply [1066, 859]. Support [1725]. supposed [1558]. Surface [1719]. Surfaces [655]. Surgeons [1719]. Surgery [117]. Surman [1574, 1329]. Surprising [1305]. Surveilling [1757]. Survey [525, 239, 144, 894, 576, 434, 457]. Survival [1434, 1245, 684]. Survivance [684]. surviving [1731]. Susanne [1180]. Suspensions [852]. sútras [1025]. Svedberg [780]. Sven [1280, 1082, 1591]. Sverker [1357]. Swammerdam [16]. Swan [1539]. Swedish [1532, 1140]. Swenson [608]. Swift [1807]. Swindlers [1164]. Swine [1700]. Swinging [1744]. Switzerland [1683, 1519, 1638, 1522, 1676]. Sword [1274]. swords [290]. Sylla [1128]. Symbolism [739]. symmetric [386]. symmetrischen [386]. Symmetry [1177, 1037, 1159]. Symposium [474, 1615]. synthèse [99]. Synthesis [1347, 1782, 99]. Syphilis [1701]. System [318, 289, 744, 281, 865, 498, 1742, 398, 794, 417, 868, 693, 1509, 1511, 1508, 179, 1513, 717, 574, 1641]. systematic [4, 55]. systematische [4, 55]. Systems [863, 852, 1725, 1150]. Syzygy [1007]. Szpiech [1726].

T [435, 1296, 484, 1350, 1149, 475, 718]. Tūsī [632, 816]. Tūsı [1293]. Table [1070, 289, 666, 349, 641, 237, 401, 1007, 1000, 28, 1130, 392, 798, 566, 1510, 1669, 1642, 994, 1512]. tableau [1514]. Tables [379, 767, 1406, 722, 834, 846, 879, 152, 325, 1407, 268, 688, 178, 197, 720, 693, 717, 998, 527]. Tablet [207]. Tabulated [442]. Tachard [1013]. Tadhkirah [384]. Taesch [1543]. Tafeln [289]. Tagebuch [92, 80]. Tahdid [459]. Tail [1047]. Tailarach [1251]. Tailarach-Vielmas [1251]. take [1537, 1465]. Taking [1153]. Tale [735, 1730, 1206, 784, 1010, 1512]. talk [1599]. Tamil [100, 178, 197, 101]. Tang [1568]. Tangents [324, 300]. Tankelæser [1190]. Tarpan [1736]. task [119]. taste [1481]. Tau [691]. tautochronism [263, 375]. Tautochronisme [375, 263]. taxonomic [1733]. Tazzioli [1715]. teacher [24]. teachers [209, 148]. teaching [555, 1540]. teachings [139, 54, 181]. Teams [1395, 1824]. Technical [447, 1610]. Technique [1749, 759, 1282]. Technological [1748, 1186, 1450, 1461]. technologies [1517, 1565, 1488]. Technology [1138, 1330, 1428, 1188, 1697, 64, 892, 1290, 487, 1401, 446, 1154, 120, 145, 135, 1145, 1297, 63, 1241, 1199, 1143, 1830, 691, 1522, 1802, 1821, 1283, 1171, 1185, 1196, 1573, 1165, 1135, 1291]. Technoscience [1713]. Tècnica [1171]. Teeth [1068]. Teleology [1250]. Telephony [779, 1151]. Telescope [40]. Telesio [1587]. Television [1317]. Tell [1010]. Tell-tale [1010]. Temminck [44]. Temperature [660]. Template [1407]. Temporalities [1710, 1595]. Temporality [1408]. Temporis [23]. Tenements [462]. Tennis [1128]. Tension [705]. Tensions [1117]. Tenth [1204, 217]. Tercentenary [311]. Tere [1586]. Term [201, 174]. Terminology [1503, 1688]. Terms [430]. Terrall [1368]. Terre [1689]. Terrestrial [1045, 892, 633, 591]. Territories [1426]. Testament [439]. Testamentum [297]. tests [108]. Text [218, 817, 1019, 896, 200]. textbooks [1463]. Texte [713]. Texten [383]. Texts [317, 318, 574, 736, 702, 713, 383]. Textual [1410]. Teyler [132]. Thābit [224]. Thackray [501, 491]. Thales [323]. Theaetet [112]. Theaetetus [112]. Thebes [968]. Their [1422, 1155, 998, 787, 1158, 1769, 1641, 128, 562, 1530, 1456, 854, 550, 560, 540, 1609, 519]. Thell [1584]. Theodor [1155, 250, 780]. Theodore [1021, 1328]. Theodoros [112, 112]. Theodoros-Stelle [112]. Theological [1332]. Theology [1741, 1639]. Theorem [603, 505]. théorèmes [550]. Theorems [820, 550]. theoretic [357]. Theoretical [673, 787, 1421, 1520, 1457]. Theorie [386, 240, 711, 659, 244, 741, 263]. Theories [126, 443, 990, 1057, 67, 1226, 1646, 1252, 104]. Theory [573, 760, 644, 734, 559, 611, 1231, 1126, 1029, 1119, 1234, 406, 217, 224, 680, 1121, 1047, 517, 564, 600, 919, 211, 1293, 967, 630, 779, 965, 628, 554, 418, 546, 331, 417, 1008, 358, 1054, 333, 625, 119, 830, 386, 741, 1539, 240, 298, 1012, 1656, 1618, 711, 518, 530, 544, 393, 807, 244, 263, 659, 460, 512]. Thera [1039]. There [377]. Thermo [1474]. Thermo-chemical [1474].

Thermodynamics [1262, 1275, 507, 614, 350, 871, 746]. Thermodynamik [871, 746]. Theses [586]. Thesis [1473]. Thibaud [1606]. Thiel [1583]. thin [1764]. Things [1367]. Thinking [1752, 1590, 1085]. Third [1262, 1275, 1594]. Thirds [635]. Thomas [1370, 31, 1160, 1429, 142, 166, 232, 1017, 1741]. Thompson [1657, 1655, 1654, 1653, 1656]. Thomson [847, 1036, 847]. Thomson-Experimente [847]. Thoreau [1213]. Thought [347, 489, 434, 1259, 1276, 440]. Thoughts [63]. Thousand [1236]. Threat [1138]. Three [268, 328, 651, 494, 550, 1028, 255]. threefold [1493]. thrift [1635]. Throughout [81]. Throwing [1461, 1169]. Thule [688]. Thury [1689]. Thyrotoxicosis [278]. Thyself [1307]. Tidal [1231]. Tiere [210]. tigers [1556]. Tiggelen [1641, 1369]. Tilbury [657]. Tilling [471]. Tilt [1490]. Time [1225, 552, 601, 1758, 1497, 888, 1695, 697, 1761, 798, 746, 1555, 1131, 156, 1553, 1546, 475]. Timelines [1761]. Times [803, 1602, 1702, 118, 1081]. Timocharis [727, 640]. Tiroler [366]. Tissue [1282]. Titelbild [8]. Tod [136]. Together [1128]. Tokugawa [1497]. Toledo [998]. Tomb [968]. tomus [360]. tones [667]. Tongwen [1464]. Too [1201, 1316]. Tool [1222]. Tools [1411, 368, 1096, 1564]. tooth [255]. Tormented [1155]. Torquetum [465]. Tourniquets [1375]. Tower [785]. Track [1350]. Tractatus [876]. Tractional [698]. tracts [346, 291]. Trade [1383, 1808]. trading [1475]. Tradition [46, 831, 501, 1238, 1528, 1086, 1324]. Traditions [1334]. Trafficking [1182]. Training [1070]. traité [532, 646, 652, 1455]. Trajectories [1723, 1186, 1715]. Traktate [346]. Trans [1128]. Transactions [1527]. Transatlantic [1197, 1201, 1200, 1198, 1199, 1202]. Transcendental [698]. Transformation [1085, 1080, 544]. Transformations [1334, 1433]. Transforming [1722, 1469]. Transformism [1068]. Transhimalaya [1591]. Transit [1069]. Transition [1161, 1203]. Transits [322]. Translated [1687, 400]. Translating [1459]. translation [1459]. translations [88]. Transmission [1053, 1336, 1536]. Transmitting [1478]. transmutation [1646]. transmutations [1667]. Transnational [1201, 1137, 1573]. Trapezia [733]. Traum [54]. Traum-Lehren [54]. travel [27]. Travellers [1239]. Travelling [1223]. Travels [1351]. Treating [1571]. Treatise [218, 217, 343, 349, 479, 868, 150, 1536, 711]. Treatises [449, 632]. treatist [532]. treatment [84, 194, 212]. Trees [1328]. trends [1661]. Trepidation [1678, 224]. Tresch [1297]. trials [746]. Triangle [577, 207]. Triatomic [1244]. Tricks [784]. Trigonometry [1030, 392, 1174]. Triorchidie [6, 6]. Trios [1030]. Trisection [1168]. Trochu [1606]. trois [550]. Tropical [527]. tropicalism [1645]. Tropisms [425]. True [1007]. Truesdell [1098, 1099, 1095, 614, 1097, 1100]. Truitt [1440]. Truth [323, 1789, 1397]. tsai [665]. tube [1176]. Tuberculosis [1757, 1702]. tujing [1568]. Tumours [173]. Tunnel [1814, 1399]. Tunnels [1119]. Turchetti [1299]. turcs [429]. Turin [826]. Turkish [429]. turmoil [1462]. Turn [1432, 1308, 1505]. Turnhout [1640, 1540, 1689]. Turning [1528, 1531]. Tuscan [1776]. Tusı [384]. Twelfth [20]. Twentieth [1205, 1084, 1282, 1577, 1643, 842, 1119, 1139]. twentieth-century [1643, 1139]. Two [317, 1747, 1153, 896, 735, 1730, 1144, 1409, 722, 1408, 632, 697, 1186, 1266, 593, 1035, 302, 1517, 383, 1731, 1017, 1028, 5, 713, 429, 15, 241]. two-cornered [1017]. Twopenny [339]. Tycho [1720, 982, 729, 1008, 275, 358]. Tychonic [1679]. Tychonis [360]. type [544]. Types [1761]. Typhus [1708]. Typically [1756, 1020]. Tyrolean [366].

u.a [1179]. U.S. [1610]. Überlieferung [391, 248, 618]. übersetzt [400]. UK [1650, 1577, 1685, 1576, 1643, 1642, 1646, 1648, 1581, 1687]. Ukraine [1491]. Ulrich [1175]. Umayyade [364]. Umgestaltung [544]. unbekannten [96]. Unbound [1073]. Unbranding [1379, 1797]. Uncertainty [1397, 1789]. Underestimated [848]. Understanding [449, 1253, 1216, 1654]. underwater [1596]. Undescribed [7]. Undoing [1736]. Unearthing [1262, 1275]. Unexamined [1161]. Ungarnreise [636]. Unhealthy [1752]. Unification [559, 1270]. Unimpressed [1739]. Union [844]. uniqueness [1593]. United [1542, 1637, 1197, 1629, 1706]. Unity [917]. Universe [690, 246, 1493]. universels [675]. Universitaires [1579]. Universitat [1605]. Universities [1229, 1574]. University [1607, 1021, 1229, 1577, 1542, 1602, 1685, 1586, 1576, 1541, 1574, 1667, 1643, 1647, 1517, 1603, 1585, 1639, 1497, 1601, 1578, 1582, 1573, 1668, 1642, 1646, 1670, 1649, 1648, 1651, 1581, 1687, 1544, 339, 1341, 840, 339]. Unknown [1731, 720, 96]. unmaking [1645]. Unmoved [1739]. Unnoticed [1706]. Unorthodox [1744]. Unpublished [358, 658, 1618]. Unreasonable [339]. Unresponsive [1739]. unseeing [1675]. Unsichtbare [1179]. Unstable [1141]. Untersuchungen [338, 89, 108, 694]. until [256]. Untranslated [23]. Unusual [1004, 580]. Unveiling [1744]. Unvereinbarkeit [494]. unveröffentlichte [658]. Unversitat [1575]. Up-and-down [1593]. Upheaval [844]. Upon [459]. Upright [1131]. Uranium [1316, 1218]. Uranus [704]. Urban [1507, 1247]. Urge [1750]. Urinary [106]. Urologie [1804, 1400]. Ursachen [128]. Ursula [1232, 1713]. Urteil [293]. Uruk [1683]. USA [863]. Use [442, 480, 1564, 856]. Used [998, 461, 836]. Using [1307, 1288, 1760, 1285, 1286, 1563, 1289, 1287]. utopia [1462]. Utopian [1297, 1821].

V [1214, 1178, 432, 372, 593, 88, 88]. V-2 [1214]. V.8 [814]. Vaccination [1092]. vaccine [1558]. València [1605]. Valentini [1234]. Valleriani [1248]. Valley [525]. Valson [423]. Value [580, 331]. Values [703, 501, 1515, 1508]. Vanden [1148, 1149]. Vanpaemel [1664]. van’t [668]. Varāha [393, 393]. Varia [50, 160, 18, 35, 35]. Variant [1678]. Variation [275]. Variations [603]. Varieté [1389]. Variety [177]. Various [1020]. Vasistha [393, 418, 393]. Vast [1169]. Vat. [876]. Vavilov [1283, 1366]. Vegetable [426]. vegetation [1571]. Velocities [1050]. Velocity [676, 1050, 1256, 794].

Venetian [1631]. Venetiis [1002]. Ventilation [1753]. Ventura [1645]. Venus [134, 1069, 754, 134, 322, 418, 393, 466, 702]. Venustheorie [393]. Verbreitungsart [182]. Verbrennungsexperiments [252]. vergessene [713]. Vergleich [335, 231]. Verité [1389]. Verlagsgesellschaft [1583]. vermeintliche [6]. Versa [1770]. Versammlung [1059]. verschollene [711]. Verschwinden [712]. Version [816, 358, 1537]. Versions [1409]. Versuch [338, 847]. Versuche [746]. versus [996, 149, 166, 1554, 668, 408, 1610]. Vertesi [1395]. Vertical [1594, 1596, 1589, 1595, 1590]. Verticalities [1597, 1588]. verticality [1592]. Very [1442, 1783, 479]. Verzeichnis [594]. Vesalius [1686, 8]. Vesals [8]. Vesical [106]. Vestiges [1535]. via [1445]. vial [210]. vibrating [1456]. Vibration [387, 263]. Vibrations [244]. Vice [1770]. victims [15]. Victor [1353, 1605]. Victorian [1388, 983, 965, 1089]. Vidal [1358]. Videnskabernes [1520]. Vie [1543, 5]. Vielmas [1251]. Vienna [1182]. Viète [98]. Viètes [215, 98, 215]. View [614, 412, 875]. Views [1560, 592, 1210, 540]. vii [1602]. Viii [131, 1638]. VIII.31 [1631]. Villaseñor [1586]. Vinca [1626]. Vincent [1125, 1515]. Vinci [75]. Vinciana [32]. Vindicated [408]. vipers [1554]. Viral [1762]. Virchow [1155, 48, 48]. Virtual [600, 1620]. virtue [1569, 1127]. Virtuoso [472, 1194]. Viruses [1746]. visibilities [1677]. Visibility [152, 237]. Visible [1677, 1717, 1673, 1326]. Vision [1424]. Visions [1422, 1383, 860, 1057, 1399, 1808, 1814]. Visual [1049, 1326, 1383, 1328, 1535, 1676, 1578, 1808, 1428]. vital [226]. vitality [1585]. vitesse [994]. Vizinho [1011]. Vleuten [1137]. Vögelin [563]. Vogt [818]. Vol [1075, 1583]. Volcanic [554]. Volny [1523]. Voltaire [1581]. Volterra [1246]. Volume [1093, 1062, 1178, 1688, 219]. Volumenbestimmung [219]. vom [618, 540, 309]. Vorarbeiten [518]. Voraussetzungen [8]. Vorläufer [17]. vorläufiges [594]. Vormärz [1059]. Vortex [965]. Vos [1724]. Voskuhl [1340]. Voyage [87, 636, 1013]. voyages [1476]. vs [1003, 1037, 865]. Vuillemin [1337]. Vulcan [1748]. vulgarisation [1689].

W [1442, 1422, 456, 447, 426, 1393, 1399, 1371, 514, 1328, 704, 1804]. W. [704]. Waage [660]. Waal [1747]. Waddell [1427]. Waerden [511, 498]. Waetzmann [667, 667]. Wahrscheinlichkeitsrechnung [119]. wa’l [380]. wa’l-muqābalah [380]. Wäldern [1175]. Walker [1133, 1180]. walking [1651]. Wallis [1602, 827, 225]. Walter [1733]. Wappen [6]. War [1166, 1747, 844, 1370, 1314, 1401, 1299, 1715, 1754, 1202, 1201, 1802, 1450, 1823, 1502, 1420, 1284, 1791, 1794, 1500, 1312, 1615, 1450, 1317, 1809, 1316, 1245, 1315, 1313, 1613, 1522, 1142, 1166, 1444]. Warden [426]. Wars [1193, 1218, 1024, 841]. wartime [1610]. Warwick [1645]. Was [339, 1220, 829, 650, 1417, 1745, 1020, 1483]. wasn’t [1558, 677]. Wässer [366]. Watches [486]. Water [435, 886, 763, 1322, 1023]. Water-Lifting [435]. Waterclock [464]. waterclocks [887]. waters [366]. Watson [1272]. Watt [1241, 1216]. Watts [1195]. Wave [616]. Waves [837, 718, 1501].

Waxes [120]. Way [1262, 1275, 1364]. Wayenborgh [1688]. Ways [598, 1020]. Weal [1230]. Weather [1397, 1685, 1789, 1595]. weeks [1549]. Wegener [508]. Weib [618]. weiblichen [247]. Weimar [630]. Weiß [699, 699]. Weiterwirken [544]. Well [461]. Weltraum [1176]. Weltsystem [179]. wenig [38]. Went [761]. Wenzel [1007]. Were [998, 400, 1464]. werk [86]. Werke [88]. Werrett [1600]. West [1574, 221, 1677, 479]. Western [1272, 1549, 594, 968, 1464]. westlichen [594]. Westman [1300]. wetting [247]. Wheels [1321, 435]. Whewell [496]. which [405, 400]. Whiggish [1157]. Whitrow [513]. wichtigsten [560]. wide [829]. Widerlegung [711, 540]. Widerstandes [712]. Widmalm [1082]. wieder [92]. Wien [1050]. Wigelsworth [1269]. Wild [1736]. Wilding [1380]. Wilhelm [1667, 658, 1176, 519, 1180, 1050, 681, 546, 430, 586]. Will [1490]. William [169, 1429, 1125, 451, 1528, 1606, 495, 496]. Williams [1750]. Willy [342]. Wilson [1687, 1282]. Winchester [997]. Wind [1119, 865, 1186, 864, 1677, 863]. Windmills [436]. winds [1677]. Wings [1321, 334]. winning [1486]. Winslowiana [92, 92]. Winter [1206]. Winterl [917]. Wirkung [806]. Wise [1072, 1150]. Wissen [1279]. Wissenschaft [107]. Wissenschaftler [1028]. Wissenschaftliche [1583]. wissenschaftlichen [847]. Wissenschaftsakademien [1583]. Wissenschaftsbeziehungen [645]. Within [1140]. Without [1795, 1068, 1050, 495, 1150, 1443]. Witt [1747]. Wittich [1567]. Wittichi [1567]. Wolfe [1746]. Wolfenbüttel [594, 594]. Wolfgang [1748]. Woman [1307, 618]. Women [1288, 1141, 1471, 1311, 1286, 1815, 1289, 26, 1287, 1641, 1506, 1641, 1309, 1289, 1371, 1195]. wonderful [1149]. Wooden [7]. word [1541]. worden [400]. words [1656]. Work [1042, 455, 1010, 993, 350, 882, 1255, 1349, 639, 552, 1534, 1179, 1817, 335]. Working [1649]. Works [102, 160, 423, 86, 1456, 52, 88]. Workshop [368, 1096]. World [1229, 1166, 1376, 1210, 1683, 1346, 1641, 829, 1795, 1381, 1775, 1425, 1253, 1802, 1450, 690, 1752, 1669, 1371, 1043, 1443, 1778, 1787, 1240, 1815, 1505, 1149, 1314, 1024, 841, 1607, 1189, 1136, 1401]. World-views [1210]. World-wide [829]. worldmaking [1806, 1298]. worlds [1472]. worldviews [1448]. worldwide [1686]. Worm [1817, 1349]. Wormbs [1082]. Worse [1369, 1829]. Wrestling [1267]. Writer [1268]. Writing [23, 1573, 1311]. Writings [1405, 38]. written [1464]. Wrong [830, 817]. Wrong. [1417]. WW2 [1197, 1201, 1200, 1198, 1199, 1202].

x [1683, 1666, 1582, 1665, 284]. X-Ray [284]. xii [1517, 1585, 1578, 1521]. xiii [1524, 1580]. xiii.10 [873]. xiv [1542, 1600, 1637, 1522, 360]. XIVe [1006]. XIVth [736]. XIX [1545, 212, 1686]. XVI [84, 1539, 1688, 1587]. XVIIe [1237]. XVIII [194, 1541]. XVth [736]. XX [1545]. xxii [1664, 1684]. xxiv [1641]. xxvii [1519]. xxxii [1636].

Yūnus [580, 401]. Yalcınkaya [1396]. yao [665]. Yazid [439]. Ydea [1575]. Year [1013, 1236, 527, 636]. Years [1747, 507, 1682, 1366, 298, 89, 44, 1460, 1466, 1514, 485]. Yemen [1000]. York [1664, 1645, 1668, 1490]. Young [1430]. Yugoslavia [38]. Yukawa [596, 742]. Yulia [1497]. Yves [1618].

Zīij [558]. Zīj [580, 237]. Zījes [644]. Zach [166]. Zacher [456]. Zacut [1011]. Zacuto [527]. Zahlensystem [717]. zahlentheoretische [357]. Zahnrezepte [255]. Zarqāllu [224]. Zarqālluh [728]. Zarqal [333]. Zaslavsky [499]. Zbyszewski [1226]. Zealand [1060, 1595]. Zeichern [1179]. Zeit [1131]. Zeitgenossen [182]. zeitgenössisches [293]. Zeitrichtung [746]. Zeitvorstellung [746]. Zentren [107]. zero [125]. Zeuthen [818, 815]. Zeytinburnu [1757]. Zij [409]. Zilsel [1473]. Zittel [1539]. Zodiac [872, 480]. zone [1475]. Zoo [1729, 1733, 1735, 1737, 1735]. Zoologe [645]. Zoological [1484]. Zoologiques [768]. Zoologist [645]. Zoos [1731, 1736]. Zum [555, 694, 1065, 113, 31]. Zur [1179, 8, 41, 112, 545, 386, 1075, 254, 1012, 159, 291, 108, 219, 807, 53, 194, 12, 391, 452, 778, 419, 26]. Zürcher [86, 86]. Zurich [498]. Zusammenfassungen [193]. zwei [383, 1028, 713, 15, 241]. Zweiten [746]. Zwischen [44, 871, 1133, 264, 1175, 293]. Zyklische [746].

References

Anonymous:1950:I

[1] Anonymous. Introductory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(1):1, September 1950. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hjelmslev:1950:EAA

[2] Johannes Hjelmslev. Eudoxus’ axiom and Archimedes’ lemma. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(1):2–11, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Guerra:1950:HAC

[3] Francisco Guerra. Hispanic–American contribution to the history of scurvy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(1):12–23, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hofmann:1950:GSN

[4] Jos. E. Hofmann. Über Gregorys systematische Näherungen für den Sektor eines Mittelpunktkegelschnittes. (German) [On Gregory’s systematic approximations for the sector of a midpoint conic section]. Centaurus:

An International Journal of the History of Science and its Cultural Aspects, 1(1):24–37, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wickersheimer:1950:VSC

[5] Ernest Wickersheimer. Une vie des saints Côme et Damien dans un manuscrit médical du IXe siècle, suivie d’une recette de collyre attribuée a la mère des deux saints (French) [A life of the saints Como and Damien in a medical manuscript of the Ninth Century, followed by a recipe of drops attributed to the mother of the two saints]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1 (1):38–42, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Belloni:1950:WVT

[6] Luigi Belloni. Über das Wappen und die vermeintliche Triorchidie von Bartolomeo Colleoni. (German) [On the crest and the alleged Triorchidie of Bartolomeo Colleoni]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(1):43–61, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Russell:1950:UWE

[7] K. F. Russell. An undescribed wooden Écorché [skinned] figure. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(1):62–65, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Artlelt:1950:TFV

[8] Walter Artlelt. Das Titelbild zur “Fabrica” Vesals und Seine kunstgeschichtlichen Voraussetzungen. (German) [The cover picture for “Fabrica Vesalius” and its art-historical conditions]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1 (1):66–77, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1950:OIR

[9] Anonymous. Obituaries: Ingjald Reichborn-Kjennerud (27 June 1865–7 August 1949). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(1):78–86, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1950:Ra

[10] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(1):88–94, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1950:BRBa

[11] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(1): 95–96, September 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1950:AAG

[12] Reinhold F. G. Müller. Zur Aufnahme der Altmedizin in die grossen Sammelwerke der indischen Ärzteschulen. (German) [accommodating age medicine in the great collections of Indian medical schools]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):97–116, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachman:1950:HP

[13] A. G. Drachman. Heron and Ptolemaios. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):117–131, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wendt:1950:HAO

[14] C. H. W. Wendt. Die heiligen Ärzte in der Ostkirchenkunst. (German) [Holy doctors in Eastern Church art]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):132–138, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schadelbauer:1950:ZSK

[15] Karl Schadelbauer. Zwei Schreiben König Maximilians über die Isolierung von Pestkranken. (German) [Two letters of King Maximilian on the isolation of plague victims]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):139–142, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Engell:1950:RJS

[16] H. Engell. Records on Jan Swammerdam in the Amsterdam archives. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):143–155, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lesky:1950:GVH

[17] Erna Lesky. Galen als Vorläufer der Hormonforschung. (German) [Galen as a precursor of hormone research]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):156–162, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ranganathan:1950:VSM

[18] S. R. Ranganathan. Varia: Source materials for the history of medicine. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):163–165, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1950:OCC

[19] Anonymous. Obituaries: Cecil Clifford Dobeil (1886–1949). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):166–173, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1950:SMS

[20] A. G. Drachmann. Societies and meetings: Sixth international congress for the history of science and Twelfth congress of the Société internationale d’histoire de la médecine. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):174–175, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1950:Rb

[21] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):176–190, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1950:BRBb

[22] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(2):

191–192, December 1950. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Farrington:1951:TPM

[23] Benjamin Farrington. Temporis Partus Masculus: An untranslated writing of Francis Bacon. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(3):193–205, March 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VonDiepgen:1951:KAA

[24] Paul Von Diepgen. Der Kirchenlehrere Augustin und Die Anatomie Im Mittelalter. (German) [The church teacher Augustine and anatomy of the Middle Ages]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(3):206–211, March 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jonckheere:1951:CAE

[25] Frans Jonckheere. La circonsion des anciens Égyptiens. (French) [Circumcision of the Ancient Egyptians]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(3):212–234, March 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VonMarzell:1951:GFC

[26] Heinrich Von Marzell and Günther Schmid. Zur Geschichte Des Frauenblattes (Chrysanthemum Balsamital). (German) [history of the journal for women (Chrysanthemum Balsamital)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(3):235–241, March 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VonAnker:1951:GCO

[27] Jean Von Anker. Georg Christian Oeders Botanische Reise in Europa Um Die Mitte Des Achtzehnten Jahrhunderts. (German) [Georg Christian Oeders botanical travel in Europe in the middle of the Eighteenth Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(3):242–265, March 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Neugebauer:1951:GTM

[28] O. Neugebauer. A Greek table for the motion of the Sun. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(3):266–270, March 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1951:Ra

[29] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(3):271–289, March 1951. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1951:BRBa

[30] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1 (3):290–292, March 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VonHofmann:1951:GTB

[31] J. E. Von Hofmann. Zum Gedenken An Thomas Bradwardine. (German) [Commemorating Thomas Bradwardine]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(4):293– 308, June 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Steinitz:1951:PPI

[32] Kate Trauman Steinitz. A pageant of proportion in illustrated books of the 15th and 16th Century in the Elmer Belt Library of Vinciana. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(4):309–333, June 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VonSchlechta:1951:HPS

[33] Karl Von Schlechta. Hirnforschung und Philosophische Spekulation Im Griechischen Altertum. (German) [Brain research and philosophical speculation in Greek Antiquity]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(4):334–355, June 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wickersheimer:1951:PSF

[34] Ernest Wickersheimer. Paracelse à Strasbourg. (French) [Paracelsus in Strasbourg]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(4):356–365, June 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VandeVelde:1951:VMH

[35] A. J. J. Van de Velde. Varia: Un musée de l’histoire des sciences en Belgique. (French) [Varia: a museum of the history of science in Belgium]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(4):366–377, June 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1951:Rb

[36] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 1(4):378–392, June 1951. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1951:ASK

[37] A. G. Drachmann. On the alleged second ktesibios. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(1):1–10, September 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schmid:1951:ESJ

[38] Alfred Schmid. Die Elektrotherapeutischen Schriften von Jos. Franciscus Domin. Einem wenig bekannten Pionier der Elektrotherapie in Jugoslavien. (German) [The electrotherapeutic writings of Jos. Franciscus dominguez. An acquaintance of pioneering electrotherapy in Yugoslavia]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(1):11–38, September 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VonBrozovic:1951:BJF

[39] Leander Von Brozović. Biographisches über Joseph Franz Domin. (German) [Biographical data about Joseph Franz Domin]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(1):38–43, September 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosen:1951:WDG

[40] Edward Rosen. When did Galileo make his first telescope? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(1):44–51, September 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Feremutsch:1951:GHC

[41] Kurt Feremutsch. Die Grundzuege der Hirnanatomie bei Carl Gustav Carus (1789–1869): Ein Beitrag zur Geschichte der Medizin und der Naturwissenschaften des beginnenden 19. Jahrhunderts. (German) [The broad outlines of the brain anatomy of Carl Gustav Carus (1789–1869): A contribution to the history of medicine and science of the early 19th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(1):52–85, September 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1951:Rc

[42] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(1):86–94, September 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1951:BRBb

[43] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(1): 95–96, September 1951. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Stresemann:1952:OBJ

[44] Erwin Stresemann and Peter Thomsen. Ornithologen-Briefe Aus Den Jahren 1816 bis 1820 Gewechselt Zwischen J. F. Naumann und C. J. Temminck. (German) [Ornithological letters from the years 1816 to 1820 exchanged between J. F. Naumann and C. J. Temminck]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2 (2):97–139, September 1952. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schmidt:1952:CA

[45] Olaf Schmidt. On the computation of the Ahargana. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(2):140–180, September 1952. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Belloni:1952:FNA

[46] Luigi Belloni. Die figurierte Nomenklatur der ausseren Körperteile des Menschen in der Tradition der griechischen Manuskripte und in den ersten gedruckten anatomischen Büchern. (German) [The illustrated nomenclature of the outer parts of the human body in the tradition of Greek manuscripts and the first printed anatomy books]. Centaurus:

An International Journal of the History of Science and its Cultural Aspects, 2(2):181–200, September 1952. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ers:1952:AMA

[47] Gerhard Ers. Der anatomische Merkspruch des Angelsächsischen Salomon und Saturn. (German) [The anatomical mnemonic of the Anglo-Saxon Solomon and Saturn]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(2):201–204, September 1952. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hasche-Klunder:1952:RVI

[48] Irmingard Hasche-Klünder. Rudolf Virchow, Infektion und Infektionskrankheit, Bakteriologie und Pathologie. (German) [Rudolf Virchow, infection and infectious disease, bacteriology and pathology]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(2):205–250, September 1952. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1952:SM

[49] Anonymous. Societies and meetings. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(2):251, September 1952. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fueter:1952:V

[50] Ed. Fueter. Varia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(2):252, September 1952. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1952:R

[51] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(2):253–264, September 1952. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dijksterhuis:1953:CHA

[52] E. J. Dijksterhuis. Christiaan Huygens: An address delivered at the annual meeting of the Holland Society of Sciences at Haarlem, May 13th, 1950, on the occasion of the completion of Huygens’s Collected Works. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(3):265–282, March 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Herrlinger:1953:FEA

[53] Robert Herrlinger. Zur Frage der ersten, anatomisch richtigen Darstellung des menschlichen Körpers in der Malerei. (German) [On the first anatomically correct representation of the human body in the art of painting]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(3):283–288, March 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1953:TLA

[54] Reinhold F. G. Müller. Traum-Lehren altindischer Ärzte. (German) [Dream teachings of Ancient Indian physicians]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2 (3):289–313, March 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schwarz:1953:SAA

[55] Georg Theodor Schwarz. Die systematische Arbeitsweise Albrecht von Hallers 1708–1777. (German) [The systematic approach of Albrecht von Haller 1708–1777]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(3):314–348, March 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Quecke:1953:FRE

[56] Kurt Quecke. François Rousset und die erste Monographie über den Kaiserschnitt. (German) [François Rousset and the first monograph on the Caesarean section]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(3):349–363, March 1953. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1953:Ra

[57] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(3):364–373, March 1953. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1953:BRB

[58] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 2(3): 373, March 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosenfeld:1953:I

[59] L. Rosenfeld. Introduction. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):7–11, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Childe:1953:SPS

[60] V. Gordon Childe. Science in preliterate societies and the ancient oriental civilisations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):12–23, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Forbes:1953:MES

[61] R. J. Forbes. Metals and early science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):24–31, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Farrington:1953:RAS

[62] B. Farrington. The rise of abstract science among the Greeks. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):32–39, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Needham:1953:TSR

[63] Joseph Needham. Thoughts on the social relations of science and technology in China. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):40–48, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Forbes:1953:MTM

[64] R. J. Forbes. Metallurgy and technology in the Middle Ages. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):49–57, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lilley:1953:CEH

[65] S. Lilley. Cause & effect in the history of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3 (1):58–72, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Taton:1953:FRP

[66] R. Taton. The French revolution and the progress of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):73–89, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mason:1953:IPT

[67] S. F. Mason. The idea of progress and theories of evolution in science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):90–106, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Shryock:1953:ISI

[68] Richard H. Shryock. The interplay of social and internal factors in modern medicine: an historical analysis. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):107–125, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pannekoek:1953:DN

[69] A. Pannekoek. The discovery of Neptune. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):126–137, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bernal:1953:SIS

[70] J. D. Bernal. Science, industry and society in the Nineteenth Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):138–165, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Singer:1953:SCR

[71] Dorothea Waley Singer. Science and confidence in the rational mind. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):166–182, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1953:PAA

[72] A. G. Drachmann. The plane astrolabe and the anaphoric clock. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):183–189, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Winter:1953:ORI

[73] H. J. J. Winter. The optical researches of Ibn Al-Haitham. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):190–210, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hintzsche:1953:MAH

[74] Erich Hintzsche. Manuscriptum Alberti Halleri ad Historiam Medicinae Pertinens: (Deutung des Manuscriptes A 96 der Burgerbibliothek Bern). (Latin/German) [Albert Halleri’ manuscript of the relevant history of medicine: Interpretation of the manuscript A 96 of the Bern Municipal Library]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):211–221, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bottger:1953:ELV

[75] Herbert Böttger. Die Embryologie Leonardos Da Vinci. (German) [The embryology of Leonardo da Vinci]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):222–235, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ronchese:1953:D

[76] F. Ronchese. Les dartres. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):236–243, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1953:Rb

[77] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(1):244–278, September 1953. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hofmann:1954:ANQ

[78] J. E. Hofmann. Altes und Neues von der Quadratur des Descartesschen Blattes. (German) [Old and new on the squaring of the Cartesian sheet]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(2):279–295, March 1954. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anker:1954:ECP

[79] Jean Anker. An Eighteenth Century project for the propagation of botanical knowledge. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(2):296–304, March 1954. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schlichting:1954:TNS

[80] Th. H. Schlichting. Das Tagebuch von Niels Stensen. (German) [The journal of Niels Stensen]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(2):305–310, March 1954. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nice:1954:IPT

[81] Margaret Morse Nice. Incubation periods throughout the ages. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(2):311–359, March 1954. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1954:R

[82] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 3(2):360–374, March 1954. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Irani:1955:ANF

[83] Rida A. K. Irani. Arabic numeral forms. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(1):1–12, March 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Korbler:1955:KPA

[84] Juraj Körbler. Über die Krebsbehandlung des praktischen Arztes im XVI. Jahrhundert. (German) [On cancer treatment in medical practice of the 16th Centuries]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(1):13–22, March 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Levey:1955:EAD

[85] Martin Levey. Evidences of ancient distillation, sublimation and extraction in Mesopotamia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(1):23–33, March 1955. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heller:1955:KBK

[86] Siegfried Heller. Kritische Bemerkungen zu dem 4. Kapitel des Buches von Josef Zürcher ≫Aristoteles’ werk und geist≪. (German) [Critical remarks on the 4th chapter of the book by Joseph Zürcher Aristotle’s works and spirit]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(1):34–50, March 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Scherz:1955:NSS

[87] Gustav Scherz. Niels Stensens Smaragdreise. (German) [Niels Stensen’s emerald voyage]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(1):51–57, March 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lesky:1955:DGB

[88] Erna Lesky. Deutsche Galenübersetzungen zu Beintker-Kahlenberg, Werke des Galenos, V. (German) [German translations of Galen in Beintker-Kahlenberg, Works of Galen, V]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(1):58–62, March 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rothschuh:1955:QUE

[89] K. E. Rothschuh and A. Schäfer. Quantitative Untersuchungen über die Entwicklung des physiologischen Fachschrifttums (Periodica) in den letzten 150 Jahren. (German) [Quantitative studies on the development of specialized physiological literature (periodicals) in the last 150 years]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(1):63–66, March 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Underwood:1955:OMN

[90] E. Ashworth Underwood. Obituaries: Max Neuburger (8 December 1868–15 March 1955). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(1):67–79, March 1955. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1955:Ra

[91] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(1):80–96, March 1955. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hintzsche:1955:AHB

[92] Erich Hintzsche. Albrecht Hallers “Manuscripta Winslowiana”, ein wieder aufgefundenes Tagebuch aus seiner Pariser Studienzeit. (German)

[Albrecht Haller’s “Manuscripta Winslowiana”, a rediscovered diary from his Paris student days]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(2):97–121, September 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aaboe:1955:BOS

[93] Asger Aaboe. On the Babylonian origin of some Hipparchian parameters. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(2):122–125, September 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Levey:1955:DAA

[94] Martin Levey. Dyeing auxiliaries in ancient Mesopotamia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(2):126–131, September 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Valentin:1955:GOO

[95] Bruno Valentin. Die Geschichte der Osteopsathyrosis (Osteogenesis Imperfecta). (German) [The history of osteopsathyrosis (Osteogenesis Imperfecta)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(2):132–147, September 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eis:1955:SPE

[96] Gerhard Eis. Späte Paracelsus-Exzerpte aus unbekannten Handschriften. (German) [Late Paracelsus excerpts from unknown manuscripts]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(2):148–162, September 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1955:Rb

[97] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(2):163–176, September 1955. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hofmann:1956:VKR

[98] Jos. E. Hofmann. Über Viètes Konstruktion des regelmäßigen Siebenecks. (German) [On Viète’s construction of the regular heptagon]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(3):177–184, March 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sarton:1956:HBS

[99] George Sarton. Henri Berr (1863–1954): La synthèse de l’histoire et l’histoire de la science. (French) [Henri Berr (1863–1954): the synthesis of history and the history of science]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(3):185–197, March 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rav:1956:MMT

[100] I. V. M. Krishna Rav. The motion of the Moon in Tamil astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(3):198–220, March 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VanDerWaerden:1956:TA

[101] B. L. van der Waerden. Tamil astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(3):221–234, March 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Haas:1956:MAJ

[102] Karlheinz Haas. Die Mathematischen Arbeiten von Johann Hudde (1628–1704) Burgermeister von Amsterdam. (German) [The mathematical works of Johann Hudde (1628–1704), Mayor of Amsterdam]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(3):235–284, March 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1956:Ra

[103] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(3):285–296, March 1956. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Turner:1956:RHT

[104] H. D. Turner. Robert Hooke and theories of combustion. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(4):297–310, September 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hatton:1956:AJE

[105] A. P. Hatton and L. Rosenfeld. An analysis of Joule’s experiments on the expansion of air: (including original results not previously published). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(4):311–318, September 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Williamson:1956:AUV

[106] Raymond Williamson. An ancient urinary vesical calculus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(4):319–324, September 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonSchipperges:1956:AZA

[107] Heinrich von Schipperges. Assimilations-Zentren Arabischer Wissenschaft Im 12. Jahrhundert. (German) [The Arabic science assimilation center in the 12th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(4):325–350, September 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tutzke:1956:SUB

[108] Dietrich Tutzke. Über statistische Untersuchungen als Beitrag zur Geschichte des Hebammenwesens im ausgehenden 18. Jahrhundert. (German) [On statistical tests as a contribution to the history of midwifery in the late 18th Centuries]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(4):351–359, September 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Crommelin:1956:FCM

[109] C. A. Crommelin. La force d’un corps en mouvement. (French) [The force of a body in movement]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(4):360–368, September 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1956:OGS

[110] Anonymous. Obituaries: George Sarton (31 August 1884–22 March 1956). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(4):369–381, September 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1956:Rb

[111] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 4(4):382–384, September 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heller:1956:BDT

[112] Siegfried Heller. Ein Beitrag zur Deutung der Theodoros-Stelle in Platons Dialog ≪Theaetet≫. (German) [A contribution to the interpretation of Theodoros’ site in Plato’s Dialogue ≪Theaetetus≫]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(1):1–58, December 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hofmann:1956:EBG

[113] Jos. E. Hofmann. Ergänzende Bemerkungen zum ≪ geometrischen≫ Irrationalitätsbeweis der alten Griechen. (German) [Supplementary remarks on ≪ geometric≫ irrationality proofs of the Ancient Greeks]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(1):59–72, December 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Winter:1956:AOM

[114] H. J. J. Winter. The Arabic optical mss. in the British Isles. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(1):73–88, December 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1956:Rc

[115] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(1):89–96, December 1956. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosen:1957:FCM

[116] George Rosen. The fate of the concept of medical police 1780–1890. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(2):97–113, June 1957. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kock:1957:CLS

[117] Wolfram Kock. Carolus Linnaeus and surgery. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5 (2):114–120, June 1957. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lesley:1957:BRT

[118] Mark Lesley. Bīrünī on rising times and daylight lengths. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(2):121–141, June 1957. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Biermann:1957:AAW

[119] Kurt-R. Biermann. Eine Aufgabe aus den Anfängen der Wahrscheinlichkeitsrechnung. (German) [A task from the beginnings of probability theory]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(2):142–150, June 1957. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Levey:1957:TOF

[120] Martin Levey. Technology of oils, fats and waxes in ancient Mesopotamia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(2):151–163, June 1957. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pazzini:1957:RDN

[121] ?. Pazzini. A recently discovered new Morgagni. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(2):164–166, June 1957. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anker:1957:O

[122] Jean Anker and A. G. Drachmann. Obituaries. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5 (2):167–169, June 1957. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1957:R

[123] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(2):170–175, June 1957. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1958:APR

[124] B. L. van der Waerden. The astronomical papyrus Ryland 27. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(3–4):177–191, June 1958. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Huber:1958:NBE

[125] Peter Huber. Über Den Nullpunkt Der Babylonischen Ekliptik. (German) [On the zero point of the Babylonian ecliptic]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(3–4):192–208, June 1958. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aaboe:1958:BPT

[126] Asger Aaboe. On Babylonian planetary theories. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5 (3–4):209–277, June 1958. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1958:HAE

[127] A. G. Drachmann. How Archimedes expected to move the Earth. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(3–4):278–282, June 1958. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Czwalina:1958:EBP

[128] Arthur Czwalina. Über einige Beobachtungsfehler des Ptolemäus und die Deutung ihrer Ursachen. (German) [some observation errors of Ptolemy and the importance of their cause]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(3–4):283–306, June 1958. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hooykaas:1958:CIS

[129] R. Hooykaas. The concepts of “individual” and “species” in chemistry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(3–4):307–322, June 1958. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jonckheere:1958:CED

[130] Frans Jonckheere. La conception Égyptienne du squelette. (French) [The Egyptian concept of the skeleton]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(3–4):323–338, June 1958. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brinch:1958:MPH

[131] Ove Brinch. The medical problems of Henry viii. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5 (3–4):339–369, June 1958. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Altena:1958:SFS

[132] C. O. van Regteren Altena. A second fossil from Scheuchzer’s collection in the Teyler Museum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(3–4):370–372, June 1958. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1958:R

[133] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 5(3–4):373–376, June 1958. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Czwalina:1959:PBP

[134] Arthur Czwalina. Ptolemaeus, Die Bahnen Der Planeten Venus und Merkur. (German) [Ptolemy, the orbits of Venus and Mercury]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(1):1–35, March 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Levey:1959:FTA

[135] Martin Levey. Food and its technology in ancient mesopotamia: the earliest chemical processes and chemicals. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(1):36–51, March 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Korbler:1959:KTK

[136] Juraj Körbler. Krankheit und Tod Kaiser Diokletians. (German) [The illness and death of Emperor Diocletian]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(1):52–54, March 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hintzsche:1959:RAA

[137] E. Hintzsche. Ein rollbild altchinesischer anatomie. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(1):55–67, March 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Miiller:1959:SNI

[138] Reinhold F. G. Miiller. Schädeleröffnungen nach indischen Sagen. (German) [Skull openings in Indian legends]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(1):68–81, March 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lienhard:1959:KCN

[139] Siegfried Lienhard. Konstitution und Charakter Nach Den Lehren Der Altindischen Medizin. (German) [constitution and character of the teachings of Ancient Indian medicine]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(1):82–93, March 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1959:Ra

[140] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(1):94–108, March 1959. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1959:BRB

[141] Anonymous. Book received: Book received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(1):109–112, March 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lohne:1959:TH

[142] Johs. Lohne. Thomas Harriott (1560–1621). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(2): 113–121, June 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lorch:1959:GNS

[143] Jacob W. Lorch. Gleanings on the naked seed controversy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(2):122–128, June 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Deas:1959:CCE

[144] Herbert D. Deas. Crystallography and crystallographers in England in the Early Nineteenth Century: a preliminary survey. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(2):129–148, June 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Levey:1959:CTA

[145] Martin Levey. Clay and its technology in ancient Mesopotamia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(2):149–156, June 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Berger:1959:JHL

[146] Peter Berger. Johann Heinrich Lamberts Bedeutung in Der Naturwissenschaft Des 18. Jahrhunderts. (German) [Johann Heinrich Lambert’s importance in the science of the 18th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6 (2):157–254, June 1959. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1959:Rb

[147] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 6(2):255–256, June 1959. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Konig:1960:SAM

[148] Klaus G. König. Die Stellung der Anatomen unter den medizinischen Lehrern in Padua im 16. Jahrhundert. (German) [The position of anatomists among medical teachers in Padua in the 16th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(1):1–5, September 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lohne:1960:HVN

[149] Johs Lohne. Hooke versus Newton. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(1):6–52, September 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Keil:1960:DBM

[150] Gundolf Keil. Der deutsche Branntweintraktat des Mittelalters. (German) [The German brandy treatise of the Middle Ages]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(1):53–100, September 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Richards:1960:LS

[151] Peter Richards. Leprosy in Scandinavia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(1):101–133, September 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kennedy:1960:PVT

[152] E. S. Kennedy and Muhammad Agha. Planetary visibility tables in Islamic astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(1):134–140, September 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1960:Ra

[153] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(1):141–143, September 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosen:1960:CABa

[154] Edward Rosen. Copernicus and Al-Bitruji. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(2):145–151, December 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosen:1960:CABb

[155] Edward Rosen. Copernicus and Al-Bitruji. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(2):152–156, December 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Scherz:1960:BBG

[156] Gustav Scherz. Briefe aus der Bartholinerzeit. (German) [Letters from the Bartholin time]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(2):157–196, December 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sokol:1960:BBJ

[157] Stanisaw Sokól. Die Bibliothek eines Barbiers aus dem Jahre 1550. (German) [The library of a barber from 1550]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(2):197–206, December 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Seide:1960:BGM

[158] Jacob Seide. The barnacle goose myth in the Hebrew literature of the Middle Ages. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(2):207–212, December 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schultheiss:1960:BPS

[159] Emil Schultheiss. Beitrag zur Pestliteratur des Spätmittelalters. (German) [Contribution to the plague literature of the Late Middle Ages]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(2):213–219, December 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lohne:1960:VNH

[160] Johns Lohne. Varia: a note on Harriott’s scientific works. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7 (2):220–221, December 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1960:SMS

[161] Anonymous. Societies and meetings: Societies and meetings. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(2):222, December 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1960:Rb

[162] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 7(2):223–235, December 1960. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1963:MJP

[163] Olaf Pedersen. Master John Perks and his mechanical curves. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):1–18, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonOrth:1963:GGA

[164] Hermann von Orth. ΔIAIT A ΓEP ONT ΩN: die Geriatrie der griechischen Antike. (Greek/German) [Elder provisions: geriatrics of Ancient Greece]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):19–47, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonWatermann:1963:AHC

[165] Rembert von Watermann. Alexander von Humboldt und die chemische Erforschung der “Gesundheit der Luft”. (German) [Alexander von Humboldt and the chemical study of the “health of the air”]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):48–68, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lohne:1963:FFT

[166] Johs Lohne. The fair fame of Thomas Harriott: Rigaud versus Baron von Zach. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):69–84, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonSchonfeldt:1963:LAA

[167] Klaus von Schönfeldt. Lucidarius — Auszüge in astronomischen Handschriften des 15. Jahrhunderts. (German) [Excerpts in astronomical manuscripts of the 15th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):85–90, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachman:1963:FAH

[168] A. G. Drachman. Fragments from Archimedes in Heron’s mechanics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):91–146, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dewhurst:1963:DWC

[169] Kenneth Dewhurst. Dr. William Cole’s (1635–1716) letters to Locke. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):147–173, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dahl:1963:LCC

[170] Per F. Dahl. Ludvig A. Colding and the conservation of energy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):174–188, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonThiele:1963:EMB

[171] Joachim von Thiele. Ernst Mach — Bibliographie. (German) [Ernst Mach — bibliography]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):189–237, March 1963. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1963:PCE

[172] Olaf Pedersen. The “philomaths” of 18th Century England: a study in amateur science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):238–262, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Onuigbo:1963:AOD

[173] Wilson I. B. Onuigbo. The age-old dictum on the spread of tumours. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):263–268, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lorch:1963:CCH

[174] J. Lorch. The carpel — a case-history of an idea and a term. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):269–291, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1963:OAR

[175] Anonymous. Obituaries: Aníbal Ruíz Moreno (1907–1960). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):292–296, March 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1963:Ra

[176] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 8(1):297–300, March 1963. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aaboe:1963:GQP

[177] Asger Aaboe. On a Greek qualitative planetary model of the epicyclic variety. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(1):1–10, September 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Toomer:1963:NTA

[178] G. J. Toomer. A note on Tamil astronomical tables. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(1):11–15, September 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1963:PMK

[179] Konrad Müller. Ph. Melanchthon und das kopernikanische Weltsystem. (German) [Ph. Melanchthon and the Copernican system]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(1):16–28, September 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1963:NIA

[180] Reinhold F. G. Müller. Nirukti indischer Ärzte. (German) [Nirukti Indian doctors]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(1):29–34, September 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1963:ALA

[181] Reinhold F. G. Müller. Amputations-Lehren in altindischer Medicin?. (German) [Amputation teachings in ancient Indian medicine?]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(1):35–37, September 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Diosi:1963:BZV

[182] Peter Diosi. Bemerkungen eines Zeitgenossen über die Verbreitungsart der Pest im Burzenlande im Jahre 1786. (German) [Observations of contemporaries on the method of dissemination of the plague in Burzenlande in 1786]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(1):38–44, September 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Snorrason:1963:OEG

[183] E. Snorrason. Obituary: Edvard Gotfredsen 1899–1963. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(1):45–46, September 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1963:NI

[184] Anonymous. Notes and informations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(1):47–50, September 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Herivel:1963:R

[185] J. W. Herivel, Johannes Lohne, Asger Aaboe, Olaf Schmidt, Axel V. Nielsen, Else Plejl, Olaf Pedersen, Olaf Pedersen, E. Bastholm, Eyvind Bastholm, E. Snorrason, and Vilh. Maller-Christensen. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(1):51–66, September 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Scavenius:1963:RNE

[186] I. Scavenius. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(2):67–70, December 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1963:SNBa

[187] Anonymous. Shorter notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9 (2):71–72, December 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Strubing:1963:NEH

[188] E. Strubing. Nahrung und Ernährung bei Hildegard von Bingen, Äbtissin, Ärztin und Naturforseherin (1098–1179). (German) [Food and nutrition in Hildegard of Bingen, abbess and medical natural scientist (1098–1179)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(2):73–124, December 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1963:NIN

[189] Anonymous. Notes and informations: Notes and informations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(2):125–128, December 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1963:Rb

[190] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(2):129–136, December 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1963:SNBb

[191] Anonymous. Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(2): 137–138, December 1963. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hofmann:1964:BGM

[192] Jos. E. Hofmann. Über die Exercitatio geometrica des M. A. Ricci. (German) [On the Exercitatio geometrica of M. A. Ricci]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(3):139–193, March 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1964:ZED

[193] Reinhold F. G. Muller. Zusammenfassungen einiger derzeitiger Beurteilungen indischer Medizin. (German) [Summaries of some current assessments of Indian medicine]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(3):194–211, March 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Korbler:1964:BKB

[194] Juraj Körbler. Zur Behandlung der Krebskrankheit mit Bleisalben im XVIII. Jahrhundert. (German) [On the treatment of cancer with lead salves in the 18th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(3):212–215, March 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosenfeld:1964:R

[195] L. Rosenfeld, Else Plejl, and Wilh. Møller-Christensen. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(3):216–218, March 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nielsen:1964:CHH

[196] Axel V. Nielsen. Contributions to the history of the Hertzsprung–Russell diagram. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):219–253, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Toomer:1964:FNT

[197] G. J. Toomer. A further note on Tamil astronomical tables. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):254–256, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hamadanizadeh:1964:MIS

[198] Javad Hamadanizadeh. A Medieval interpolation scheme for oblique ascensions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):257–265, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knudsen:1964:NNC

[199] Ole Knudsen. A note on Newton’s concept of force. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):266–271, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Werlin:1964:FJG

[200] Josef Werlin. Ein Feuerwerkstext aus dem 16. Jahrhundert. (German) [A fireworks text of the 16th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):272– 287, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Diosi:1964:LTC

[201] Peter Diosi. Long term changes in the natural history of infectious diseases. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):288–292, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pettersson:1964:MMS

[202] Olof Pettersson. Magic and medicine in South African Bantu psychiatry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):293–316, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1964:NI

[203] Anonymous. Notes and information. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):317–318, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1964:Ra

[204] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):319–323, June 1964. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1964:RNE

[205] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):324– 327, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1964:SNBa

[206] Anonymous. Short notices of books received: Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 9(4):328–329, June 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DeSollaPrice:1964:BPT

[207] Derek J. De Solla Price. The Babylonian “Pythagorean Triangle” tablet. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(1):1–13, September 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vescovini:1964:QPD

[208] Graziella Federici Vescovini. Les questions de ‘perspective’ de Dominicus de Clivaxo. (French) [The questions of ‘perspective’ of Dominicus de Clivaxo]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(1):14–18, September 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1964:QMD

[209] Anonymous. Quaestiones magistri dominici de clivaxo super perspectiva. (Latin) [The problems of the church teachers of clivaxo [????] on perspective]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(1):18–28, September 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eis:1964:TPG

[210] Gerhard Eis. Tiere aus der Phiole. (German) [Animals from the vial]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(1):29–34, September 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kargon:1964:DCT

[211] Robert Kargon. The decline of the caloric theory of heat: a case study. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(1):35–39, September 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Korbler:1964:KBX

[212] Juraj Korbler. Eine Krebsbehandlung zu Beginn des XIX. Jahrhunderts. (German) [A cancer treatment at the beginning of the 19th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(1):40–47, September 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1964:Rb

[213] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(1):48–61, September 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1964:SNBb

[214] Anonymous. Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(1): 62–64, September 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Busard:1964:EPV

[215] H. L. L. Busard. Über einige Papiere aus Viètes Nachlaß in der Pariser Bibliothèque Nationale. (German) [Some papers from Viètes estate in the National Library in Paris]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(2):65–126, December 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1964:Rc

[216] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(2):127–128, December 1964. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1965:TNT

[217] Bernard R. Goldstein. A treatise on number theory from a Tenth Century Arabic source. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(3):129–134, March 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1965:TFT

[218] Anonymous. Text: The first treatise of the division of abstract sciences. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(3):135–160, March 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Werlin:1965:RVF

[219] Josef Werlin. Eine Regel zur Volumenbestimmung von Fässern. (German) [A rule for determining the volume of barrels]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(3):161–164, March 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1965:KAA

[220] Reinhold F. G. Müller. Können die altindischen Ärzte als Materialisten gelten?. (German) [Can the Ancient Indian physicians be considered materialists?]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(3):165–173, March 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Needham:1965:SSE

[221] Joseph Needham. Science and society in East and west. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(3):174–197, March 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1965:Ra

[222] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(3):198–212, March 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aaboe:1965:PRB

[223] Asger Aaboe. On period relations in Babylonian astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(4):213–231, June 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1965:TTA

[224] Bernard R. Goldstein. On the theory of trepidation: According to Thābit b. Qurra and al-Zarqāllu and its implications for homocentric planetary theory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(4):232–247, June 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonScriba:1965:WHG

[225] Christoph J. von Scriba. Wallis und Harriot. (German) [Wallis and Harriot]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(4):248–257, June 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonJorgensen:1965:BLG

[226] Bent Søren von Jørgensen. Berzelius und die Lebenskraft. (German) [Berzelius and the vital force]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(4):258–281, June 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Friedmann:1965:HOK

[227] Herbert Friedmann. The history of our knowledge of avian brood parasitism. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(4):282–304, June 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1965:Rb

[228] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(4):305–315, June 1965. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1965:RNE

[229] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(4):316– 317, June 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1965:SNB

[230] Anonymous. Short notices of books received: Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 10(4):318–320, June 1965. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1966:VMB

[231] B. L. van der Waerden. Vergleich der mittleren Bewegungen in der babylonischen, griechischen und indischen Astronomie. (German) [comparison of the mean motions in the Babylonian, Greek and Indian astronomy]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(1):1–18, March 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lohne:1966:THM

[232] Vot J. A. Lohne. Thomas Harriot als Mathematiker. (German) [Thomas Harriot as mathematician]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(1):19–45, March 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kargon:1966:NBH

[233] Robert Kargon. Newton, Barrow and the hypothetical physics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(1):46–56, March 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1966:IBB

[234] Reinhold F. G. Müller. Indische Blut-Bewertungen. (German) [Indian blood reviews]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(1):57–62, March 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1966:NIR

[235] Anonymous. Notes and information: Report on the Danish national committee for the history and philosophy of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(1):63–66, March 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1966:Ra

[236] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(1):67–72, March 1966. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kennedy:1966:CVT

[237] E. S. Kennedy and Mardiros Janjanian. The crescent visibility table in Al-Khwārizmī’s Zīj. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(2):73–78, June 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Grant:1966:APA

[238] Edward Grant. Aristotle, Philoponus, Avempace, and Galileo’s Pisan dynamics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(2):79–93, June 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cartwright:1966:HS

[239] Brian S. Cartwright. N_{2}O_{4}: a historical survey. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(2):96–110, June 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lauer:1966:GMT

[240] Hans H. Lauer. Grundzüge einer medizinischen Theorie in der Kulturmorphologie Ibn Chalduns. (German) [Principles of medical theory in the cultural morphology of Ibn Khaldun]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(2):111– 127, June 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sommer:1966:ZDL

[241] H. Sommer. Zwei Druckfragmente von Lorenz Fries (1483?–1531). (German) [Two print fragments of Lorenz Fries (1483?–1531)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(2):128–140, June 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1966:RNEa

[242] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(2):141– 142, June 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1966:BRBa

[243] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11 (2):143–144, June 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

ZiggelaarSJ:1966:OTV

[244] A. Ziggelaar, S.J. Aux origines de la théorie des vibrations harmoniques: Le Père Ignace Gaston Pardies (1636–1673): Études sur Ignace Gaston Pardies I. (French) [On the origins of the theory of harmonic vibrations: Father Ignace Gaston Pardies (1636–1673): studies on Ignace Gaston Pardies I]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):145–151, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ones:1966:SPH

[245] Gordon Ones. The scientific publications of Henry Kater. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):152–189, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Skabelund:1966:CAF

[246] Donald Skabelund. Cosmology on the American frontier: Orson Pratt’s key to the universe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):190–204, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1966:IBW

[247] Reinhold F. G. Müller. Die indische Bewertung des weiblichen Fortplanzungsstoffes (drtava). (German) [The Indian review of the female reproductive plan wetting substance [????] (drtava) ]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11 (3):205–209, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Werlin:1966:HUP

[248] Josef Werlin. Eine Heidelberger Überlieferung des Pelzbuches von Gottfried von Franken. (German) [A Heidelberg lore of the fur book by Gottfried of Franconia]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):210–214, December 1966. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosen:1966:MOK

[249] Edward Rosen. The Moon’s orbit in Kepler’s Somnium. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):217–221, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thiele:1966:BGT

[250] Joachim Thiele. Briefe von Gustav Theodor Fechner und Ludwig Boltzmann an Ernst Mach. (German) [Letters of Gustav Theodor Fechner and Ludwig Boltzmann to Ernst Mach]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):222–235, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Badash:1966:EGF

[251] Lawrence Badash. An Elster and Geitel failure: Magnetic deflection of beta rays. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):236–240, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bohm:1966:JMG

[252] Walter Böhm. John Mayow und die Geschichte des Verbrennungsexperiments. (German) [John Mayow and the history of combustion experiments]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):241–258, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1966:IDM

[253] Reinhold F. G. Müller. Indisch: Dravya (medizingeschichtlich). (German) [Indian: Dravya (medical history)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):259–260, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1966:EBB

[254] Reinhold F. G. Muller. Einige Belege zur Bewahrung alter Lehrsätze in neuzeitlicherer indischer Medizin. (German) [Some evidence for preservation of old dogmas in modern Indian medicine]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3): 270–274, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Werun:1966:DZP

[255] Josef Werun. Drei Zahnrezepte von Peter Ernst. (German) [Three tooth recipes from Peter Ernst]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):275–278, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schultheiss:1966:SHE

[256] Emil Schultheiss and Louis Tardy. Short history of epidemics in Hungary until the great cholera epidemic of 1831 (on the basis of historical sources). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):279–301, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1966:BCB

[257] Anonymous. Brief communications: Brief communications. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):302–309, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1966:Rb

[258] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):310–319, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1966:RNEb

[259] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3):320–323, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1966:BRBb

[260] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 11(3): 324–327, December 1966. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1968:ASP

[261] A. G. Drachmann. Archimedes and the science of physics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(1):1–11, March 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Boutelle:1968:AA

[262] Marion Boutelle. The almanac of Azarquiel. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(1): 12–19, March 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

ZiggelaarSJ:1968:PDD

[263] August Ziggelaar, S.J. Les premières démonstrations du tautochronisme de la cycloïde, et une conséquence pour la théorie de la vibration harmonique. Études sur Ignace Gaston Pardies, II. (French) [The first demonstrations of the cycloid’s tautochronism, and a consequence for the theory of harmonic vibration. Studies on Ignace Gaston Pardies]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(1):21–37, March 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hermann:1968:FDZ

[264] Armin Hermann. Die frühe Diskussion zwischen Stark und Sommerfeld über die Quantenhypothese (1). (German) [Early discussions between Stark and Sommerfeld on the quantum hypothesis (1)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(1):38–59, March 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1968:Ra

[265] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(1):60–70, March 1968. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1968:RNEa

[266] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(1):71–72, March 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Petersen:1968:DLA

[267] Viggo M. Petersen and Olaf Schmidt. The determination of the longitude of the apogee of the orbit of the Sun according to Hipparchus and Ptolemy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(2):73–96, June 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Neugebauer:1968:TCT

[268] O. Neugebauer. Three Copernican tables. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(2):97– 106, June 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kennedy:1968:DCH

[269] E. S. Kennedy. The digital computer and the history of the exact sciences. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(2):107–113, June 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Broszinski:1968:HHS

[270] Hartmut Broszinski. Das Heilmittelglossar des Hans Suff von Göppingen. (German) [The drunkenness-cure glossary of Hans Suff of Göppingen]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(2):114–131, June 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1968:BCB

[271] Anonymous. Brief communications: Brief communications. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(2):132–137, June 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knudsen:1968:R

[272] Ole Knudsen. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(2):138–141, June 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1968:RNEb

[273] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(2):142– 144, June 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Toomer:1968:SLE

[274] G. J. Toomer. The size of the lunar epicycle according to Hipparchus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(3):145–150, September 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thoren:1968:TBD

[275] Victor E. Thoren. Tycho Brahe’s discovery of the variation. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(3):151–166, September 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Scherz:1968:NSG

[276] Gustav Scherz. Neue Stensenbriefe. (German) [New correspondence of Stensen]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(3):167–181, September 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Korbler:1968:KPG

[277] Juraj Körbler. Die Krebslehre des Paracelsus und die Gegenwart. (German) [The cancer doctrine of Paracelsus and the present]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(3):182–191, September 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Greenwald:1968:NHG

[278] Isidor Greenwald. Notes on the history of goitre and of thyrotoxicosis in Denmark. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(3):192–196, September 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Turunen:1968:CRE

[279] Aarno Turunen. C. R. Ehrström, the Finnish predecessor of Ignaz Semmelweis, the defeater of puerperal fever. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(3):197–201, September 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1968:Rb

[280] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(3):202–208, September 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Neugebauer:1968:OBB

[281] O. Neugebauer. The origin of ≫System B≪ of Babylonian astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(4):209–214, December 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). To W. K. Feller as a token of lifelong friendship.

Nichols:1968:SBA

[282] Robert E. Nichols, Jr. Sidrak and Bokkus on the atmospheric and Earth sciences. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(4):215–232, December 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

McGuire:1968:OND

[283] I. E. McGuire. The origin of Newton’s doctrine of essential qualities. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(4):233–260, December 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heimann:1968:MIX

[284] P. M. Heimann. Moseley’s interpretation of X-ray spectra. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(4):261–274, December 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1968:ME

[285] B. L. van der Waerden. Mendel’s experiments. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(4): 275–288, December 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tchorbadjiev:1968:EDC

[286] S. Tchorbadjiev. The early days of chemistry in Bulgaria. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(4):289–302, December 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nielsen:1968:SAD

[287] Axel V. Nielsen. A sundial from about 1200 A.D. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(4): 303–307, December 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1968:Rc

[288] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 12(4):308–320, December 1968. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Burckhardt:1969:ASP

[289] J. J. Burckhardt and B. L. van der Waerden. Das astronomische System der Persischen Tafeln I. (German) [The astronomical system of the Persian Table I]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(1):1–28, March 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1969:NSE

[290] A. G. Drachmann. On the named swords, especially in the Icelandic sagas. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(1):29–36, March 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Telle:1969:AEM

[291] Joachim Telle. Altdeutsche Eichentraktate aus medizinischen Handschriften (Beiträge zur pharmazeutischen Kleinliteratur im ausgehenden Mittelalter). (German) [Old German oak tracts of medical manuscripts (contributions to the small pharmaceutical literature of the late Middle Ages)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(1):37–61, March 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bluh:1969:EMH

[292] Otto Blüh. Ernst Mach as an historian of physics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(1):62–84, March 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thiele:1969:ZUK

[293] Joachim Thiele. Ein zeitgenössisches Urteil über die Kontroverse zwischen Max Planck und Ernst Mach. (German) [A contemporary assessment of the controversy between Max Planck and Ernst Mach]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(1):85–90, March 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1969:BCBa

[294] Anonymous. Brief communications: Brief communications. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(1):91–98, March 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1969:Ra

[295] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(1):99–110, March 1969. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1969:SNBa

[296] Anonymous. Short notices of books received: Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(1):111–112, March 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jorgensen:1969:TG

[297] Bent Søren Jørgensen. Testamentum gebri. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):113– 119, June 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moesgaard:1969:EYC

[298] Kristian Peder Moesgaard. The 1717 Egyptian years and the Copernican theory of precession. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):120–138, June 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1969:SOR

[299] Olaf Pedersen. Sagredo’s optical researches. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):139– 150, June 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1969:RMT

[300] Kirsti Møller Pedersen. Roberval’s method of tangents. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):151–182, June 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knudsen:1969:LBD

[301] Ole Knudsen and Kurt Møller Pedersen. The link between ‘determination’ and conservation of motion in Descartes’ dynamics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):183–186, June 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ziggelaar:1969:IOD

[302] A. Ziggelaar. Une invention oubliée: Deux instruments pour la construction des cadrans solaires: Études sur Ignace Gaston Pardies III. (French) [A forgotten invention: Two instruments for the construction of sundials: studies on Ignace Gaston Pardies III]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):187– 197, June 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:1969:JER

[303] Gordon Jones. Joule’s early researches. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):198– 219, June 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1969:BCBb

[304] Anonymous. Brief communications: Brief communications. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):220–227, June 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1969:Rb

[305] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):228–238, June 1969. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1969:SNBb

[306] Anonymous. Short notices of books received: Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(2):238–240, June 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1969:DHD

[307] A. G. Drachmann. A detail of Heron’s dioptra. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3): 241–247, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kennedy:1969:EMS

[308] E. S. Kennedy. An early method of successive approximations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3):248–250, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Verdonk:1969:EPR

[309] J. J. Verdonk. Vom Einfluß des P. de la Ramée auf Simon Stevin. (German) [On the influence of P. de la Ramée on Simon Stevin]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3):251–262, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wilson:1969:EKA

[310] Curtis Wilson. The error in Kepler’s acronychal data for Mars. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3):263–268, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bruins:1969:PAT

[311] Evert M. Bruins. Problema Alhaseni at the tercentenary of Huygens’ Solution. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3):269–277, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sloth:1969:CHR

[312] Flemming Sloth. Chr. Huygens’ rectification of the cycloid. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3):278–284, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eis:1969:HRG

[313] Gerhard Eis. Zu Heinrich von Rees. (German) [To Heinrich von Rees]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3):285–290, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1969:Rc

[314] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3):291–312, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1969:RNE

[315] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3):313–315, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1969:SNBc

[316] Anonymous. Short notices of books received: Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(3):316–321, September 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aaboe:1969:TLT

[317] Asger Aaboe and Abraham Sachs. Two lunar texts of the Achaemenid period from Babylon. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):1–22, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bernsen:1969:CCB

[318] Lis Bernsen. On the construction of column B in system A of the astronomical cuneiform texts. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):23–28, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Britton:1969:PDO

[319] John P. Britton. Ptolemy’s determination of the obliquity of the ecliptic. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):29–41, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Caminos:1969:AED

[320] Ricardo A. Caminos. An ancient Egyptian donation stela in the archaeological museum of Florence (inv. no. 7207). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):42–46, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1969:CLS

[321] A. G. Drachmann. Capital letters and small letters in mathematics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):47–48, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1969:SMR

[322] Bernard R. Goldstein. Some Medieval reports of Venus and Mercury transits. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):49–59, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hartner:1969:EPT

[323] Willy Hartner. Eclipse periods and Thales’ prediction of a solar eclipse: Historic truth and modern myth. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):60–71, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jensen:1969:PFM

[324] Claus Jensen. Pierre Fermat’s method of determining tangents of curves and its application to the conchoid and the quadratrix. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):72–85, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kennedy:1969:AKP

[325] A. S. Kennedy and Walid Ukashah. Al-Khwarizmi’s planetary latitude tables. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):86–96, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Oppenheim:1969:DCO

[326] A. Leo Oppenheim. Divination and celestial observation in the last Assyrian empire. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):97–135, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Parker:1969:SDM

[327] Richard A. Parker. Some Demotic mathematical papyri. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):136–141, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Petersen:1969:TLM

[328] Viggo M. Petersen. The three lunar models of Ptolemy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):142–171, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pingree:1969:LPS

[329] David Pingree. The later Pauli´ ṡsasiddhānta. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1): 172–241, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DeSollaPrice:1969:PSA

[330] Derek J. De Solla Price. Portable sundials in antiquity, including an account of a new example from aphrodisias. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):242–266, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schmidt:1969:MVP

[331] Olaf Schmidt. A mean value principle in Babylonian planetary theory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):267–286, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Swerdlow:1969:HDS

[332] Noel Swerdlow. Hipparchus on the distance of the Sun. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):287–305, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Toomer:1969:STZ

[333] G. J. Toomer. The solar theory of az-Zarqal: a history of errors. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 14(1):306–336, December 1969. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Solon:1971:SWI

[334] Peter Solon. The six wings of Immanuel Bonfils and Michael Chrysokokkes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(1):1–20, March 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1971:BMA

[335] B. L. van der Waerden. Berichtigung zu meiner Arbeit ≪Vergleich der mittleren Bewegungen in der babylonischen, griechischen und indischen Astronomie≫ (Centaurus 11, p. 1–18). (German) [Correction on my work ≪Comparison of mean motions in Babylonian, Greek and Indian astronomy≫ ]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(1):21–24, March 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1971:RCA

[336] Kirsti Moller Pedersen. Roberval’s comparison of the arclength of a spiral and a parabola. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(1):26–43, March 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Larder:1971:PHR

[337] David F. Larder. Prout’s hypothesis: a reconsideration. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(1):44–50, March 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kohlsaat:1971:UNS

[338] Heike Kohlsaat and Carl Schirren. Die Untersuchungen von Niels Stensen über die Haut. Versuch einer kritischen Betrachtung. (German) [The investigations of Niels Stensen on the skin. Attempt at a critical examination]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(1):51–71, March 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Warburg:1971:WUC

[339] Erik Warburg. Was the University of Christian III a twopenny university, or was the salary demanded by Leonhart Fuchs unreasonable? The professors’ pay in the 16th Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(1):72–106, March 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1971:Ra

[340] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(1):107–112, March 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bechler:1971:ACE

[341] Zeev Bechler. Aristotle corrects Eudoxus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(2):113– 123, June 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1971:NIG

[342] A. G. Drachmann. A new interpretation of the golden horns. A comment on Professor Willy Hartner’s book. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(2):124–134, June 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1971:PDS

[343] Bernard R. Goldstein and Noel Swerdlow. Planetary distances and sizes in an anonymous Arabic treatise preserved in Bodleian Ms. Marsh 621. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(2):135–170, June 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1971:Rb

[344] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(2):171–189, June 1971. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1971:RNE

[345] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(2):190– 191, June 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Busard:1971:TBP

[346] H. L. L. Busard. Die Traktate De Proportionibus von Jordanus Nemorarius und Campanus. (German) [The tracts De proportionibus of Jordanus Nemorarius and Campanus]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(3):193–227, December 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Coleman:1971:BCC

[347] William Coleman. Bateson and chromosomes: Conservative thought in science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(3):228–314, December 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1971:Rc

[348] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 15(3):315–328, December 1971. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jensen:1972:ANM

[349] Claus Jensen. Abū nasr Mansūr’s approach to spherical astronomy as developed in his treatise “the table of minutes”. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(1): 1–19, March 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Talbot:1972:ATW

[350] G. R. Talbot and A. J. Pacey. Antecedents of thermodynamics in the work of Guillaume Amontons. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(1):20–40, March 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knudsen:1972:LKN

[351] Ole Knudsen. From Lord Kelvin’s notebook: Ether speculations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(1):41–53, March 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1972:Ra

[352] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(1):54–62, March 1972. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1972:RNE

[353] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(1):63–64, March 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1972:AP

[354] B. L. van der Waerden. Aegyptische planetenrechnung. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(2):65–91, June 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Krieger:1972:KBA

[355] Helmut Krieger. Konstruktion bogengleicher algebraischer Kurven in historischer Sicht. (German) [Construction of algebraic curves of equal arc in historical perspective]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(2):92–162, June 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1972:Rb

[356] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(2):163–168, June 1972. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hofmann:1972:RBZ

[357] J. E. Hofmann. Reviews: Über eine zahlentheoretische Aufgabe Fermats. (German) [On a number-theoretic exercise of Fermat]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(3):169–202, September 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thoren:1972:UVT

[358] Victor E. Thoren. An ‘unpublished’ version of Tycho Brahe’s lunar theory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(3):203–206, September 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Iostelii:1972:LD

[359] M. Melchioris Iöstelii. Lvnæ Deltqvivm. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(3):207–230, September 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Christianson:1972:ABB

[360] John Robert Christianson. Addenda to Tychonis Brahe Opera Omnia tomus XIV. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(3):231–247, September 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1972:Rc

[361] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(3):248, September 1972. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1972:SNB

[362] Anonymous. Short notices of books received: Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(3):253–256, September 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1972:LBG

[363] Bernard R. Goldstein. Levi ben Gerson’s lunar model. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16 (4):257–284, December 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Janin:1972:CSM

[364] Louis Janin. Le cadran solaire de la Mosquée Umayyade à Damas. (French) [The sundial of the Umayyade mosque in Damascus]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(4):285–298, December 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Livingston:1972:MIC

[365] John Livingston. The Mukhula, an Islamic conical sundial. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(4):299–308, December 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eis:1972:BCT

[366] Gerhard Eis. Brief communications: Ein Tiroler Fragment der “Gebrannten Wässer” Gabriel von Lebensteins. (German) [A fragment of the Tyrolean “Burnt waters of life” of Gabriel von Lebenstein]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(4):309–319, December 1972. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1972:Rd

[367] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 16(4):320, December 1972. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Truesdell:1973:SWT

[368] C. Truesdell. The Scholar’s workshop and tools. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(1):1–10, March 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Garber:1973:AIP

[369] Elizabeth Garber. Aspects of the introduction of probability into physics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(1):11–40, March 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Herivel:1973:IFB

[370] John Herivel. The influence of Fourier on British mathematics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(1):40–57, March 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Klein:1973:MEE

[371] Martin J. Klein. Mechanical explanation at the end of the Nineteenth Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(1):58–82, March 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pihl:1973:SAV

[372] Mogens Pihl. The scientific achievements of L. V. Lorenz. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(1):83–94, March 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1973:SNa

[373] Anonymous. Short notices. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(1):95–96, March 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Swerdlow:1973:ABD

[374] Noel Swerdlow. Al-Battānī’s determination of the solar distance. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(2):97–105, June 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

ZiggelaarSJ:1973:TCS

[375] August Ziggelaar, S.J. Le tautochronisme de la cycloïde selon Pieter Baert: Études sur Ignace Gaston Pardies, IV. (French) [The tautochronism of the cycloide according to Pieter Baert: studies on Ignace Gaston Pardies, IV]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(2):106–118, June 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Telle:1973:AJS

[376] Joachim Telle. Das Arzneibuch Johannes Schöners und seine mittelhochdeutschen Quellen. (German) [The pharmacopoeia of John Schoener and his Middle High German sources]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(2):119– 141, June 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Truesdell:1973:RBT

[377] C. Truesdell. Reviews: Is There a Philosophy of Science? An essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(2):142–172, June 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1973:Ra

[378] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(2):173–174, June 1973. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bruins:1973:TRI

[379] Evert M. Bruins. Tables of reciprocals with irregular entries. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(3):177–188, September 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Saliba:1973:MAJ

[380] George A. Saliba. The meaning of al-jabr wa’l-muqābalah. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(3):189–204, September 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wilson:1973:IPK

[381] Curtis Wilson. The inner planets and the Keplerian revolution. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(3):205–248, September 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1973:SNb

[382] Anonymous. Short notices. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(3):249, September 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hubner:1973:KBZ

[383] Wolfgang Hübner. Kritische Bemerkungen zu zwei Texten über die Sonnenbewegung. (German) [Critical remarks on two texts on the movement of the Sun]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(4):253–259, December 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Livingston:1973:NAD

[384] John W. Livingston. Nas.ir al-Din al-T .usī’s al-Tadhkirah: a category of Islamic astronomical literature. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(4):260–275, December 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jorgensen:1973:PHF

[385] Niels Therkel Jørgensen and Kristian Peder Moesgaard. Peder Horrebow’s formulation of the so-called Mayer formula. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(4): 276–279, December 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knobloch:1973:LST

[386] Eberhard Knobloch. Leibnizens Studien zur Theorie der symmetrischen Funktionen. (German) [Leibniz’ studies in the theory of symmetric functions]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(4):280–294, December 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tunbridge:1973:LOE

[387] Paul A. Tunbridge. A letter from ørsted on the effects produced in bodies subjected to vibration. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(4):295–300, December 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wallis:1973:BPM

[388] P. J. Wallis. British philomaths — Mid-Eighteenth Century and earlier. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(4):301–314, December 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gruner:1973:DFL

[389] S. M. Gruner. Defending Father Lucas: a consideration of the Newton– Lucas dispute on the nature of the spectrum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(4): 315–329, December 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1973:Rb

[390] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 17(4):330–334, December 1973. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Neuenschwander:1974:UAL

[391] E. Neuenschwander. Zur Überlieferung der Archytas — Lösung des delischen Problems. (German) [On the lore of Archytas — solution of the Delian problem]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(1):1–5, March 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Toomer:1974:CTH

[392] G. J. Toomer. The chord table of Hipparchus and the early history of Greek trigonometry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(1):6–28, March 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wirth:1974:VVB

[393] Peter B. Wirth. Die Venustheorie von Vasistha in der Pañcasiddhāntikā des Varāha Mihira. (German) [The Venus theory of Vasistha in the Pañcasiddhāntikā des Varāha Mihira]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(1):29– 43, March 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rohr:1974:SAI

[394] René R. J. Rohr. Sonnenuhr und Astrolabium im Dienste der Moschee. (German) [Sundial and astrolabe in the service of the mosque]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(1):44–56, March 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Broszinski:1974:FMG

[395] Hartmut Broszinski. Eine Freiburger medizinische Glossensammlung aus dem 15. Jahrhundert. (German) [A Freiburg medical gloss collection from the 15th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(1):57–63, March 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1974:BCBa

[396] Anonymous. Brief communications: Brief communications. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(1):64–69, March 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Stromholm:1974:RBG

[397] Per Strømholm. Reviews: Original Golden Hit. An essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(1):70–75, March 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rancke-Madsen:1974:RBP

[398] E. Rancke-Madsen. Reviews: The Periodic System of Chemical Elements. An essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(1):76–80, March 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1974:Ra

[399] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(1):81–96, March 1974. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Busard:1974:EES

[400] H. L. L. Busard. Über einige Euklid-Scholien, die den Elementen von Euklid, übersetzt von Gerard von Cremona, angehängt worden sind. (German) [Some Euclidean scholia, which were the Elements of Euclid, translated and extended by Gerard of Cremona]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(2):97–128, June 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

King:1974:DAT

[401] David A. King. A double-argument table for the lunar equation attributed to ibn Yūnus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(2):129–146, June 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heimann:1974:CFC

[402] P. M. Heimann. Conversion of forces and the conservation of energy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(2):147–161, June 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1974:Rb

[403] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(2):162–172, June 1974. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jorgensen:1974:PHS

[404] Niels Therkel Jørgensen. Peder Horrebow’s solution of the Keplerproblem. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(3):173–183, September 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bechler:1974:NLF

[405] Z. Bechler. Newton’s law of forces which are inversely as the mass: a suggested interpretation of his later efforts to normalise a mechanistic model of optical dispersion. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(3):184–222, September 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Frankel:1974:SCT

[406] Eugene Frankel. The search for a corpuscular theory of double refraction: Malus, Laplace and the price competition of 1808. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(3):223–245, September 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1974:Rc

[407] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(3):246–252, September 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Taisbak:1974:PVA

[408] C. M. Taisbak. Posidonius vindicated at all costs? Modern scholarship versus the stoic Earth measurer. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(4):253–269, December 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kunitzsch:1974:NLA

[409] P. Kunitzsch. New light on al-Battani’s Zij. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(4):270–274, December 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1974:LBG

[410] Bernard R. Goldstein. Levi ben Gerson’s preliminary lunar model. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(4):275–288, December 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pampusch:1974:EMH

[411] Anita M. Pampusch. ≫experimental≪, ≫metaphysical≪, and ≫hypothetical≪ philosophy in Newtonian methodology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(4):289–300, December 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thoren:1974:EVG

[412] Victor Thoren, Charles Gow, and Kent Honeycutt. An early view of galactic rotation. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(4):301–314, December 1974. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1974:BCBb

[413] Anonymous. Brief communications: Brief communications. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(4):315–318, December 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1974:Rd

[414] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(4):319–324, December 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1974:RNE

[415] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(4):325, December 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1974:SN

[416] Anonymous. Short notices. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 18(4):326–328, December 1974. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schmidt:1975:SAA

[417] Olaf Schmidt. A system of axioms for the Archimedean theory of equilibrium and centre of gravity. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):1–35, March 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pingree:1975:VTV

[418] David Pingree. Vasistha’s theory of Venus: the misinterpretation of an emendation. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):36–39, March 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thiele:1975:ABB

[419] J. Thiele. Zur Analyse der Bewegungsempfindungen. Briefe von Josef Breuer an Ernst Mach. (German) [On the analysis of the sensations of movement. Letters of Josef Breuer to Ernst Mach]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):40–53, March 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Oldroyd:1975:MCR

[420] D. R. Oldroyd. Mineralogy and the ‘chemical revolution’. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):54–71, March 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eis:1975:EMB

[421] Gerhard Eis. Der Einfluss der mittelalterlichen Botanik auf die Gartenliteratur des 16.–18. Jahrhunderts. (German) [The influence of Medieval botany on the garden literature of the 16th–18th Centuries]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):71–80, March 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1975:RNE

[422] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):81–84, March 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosenfeld:1975:RBW

[423] L. Rosenfeld. Review: Life and Works of Baron Cauchy — French — by C. A. Valson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):81, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1975:RBO

[424] K. M. Pedersen. Review: Planck’s Original Papers in Quantum Physics — Classic Papers in Physics — by H. Kangro. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1): 81–82, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosenfeld:1975:RBM

[425] L. Rosenfeld. Review: Forced Movements, Tropisms, and Animal Conduct — by J. Loeb. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):82, ???? 1975. CODEN CENTA4. ISSN 0008-8994.

Pedersen:1975:RBI

[426] O. Pedersen. Review: Pharmacographia Indica — History of Principal Drugs of Vegetable Origin Met with in British India — by W. Dymock, C. J. K. Warden, and D. Hooper. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):83–84, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosenfeld:1975:RBT

[427] L. Rosenfeld. Review: Climate Through Ages — by C. E. P. Brooks. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):84, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lorch:1975:AJI

[428] R. P. Lorch. The astronomy of Jābir ibn Aflah. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):85–107, June 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rohr:1975:DAQ

[429] René R. J. Rohr and Louis Janin. Deux astrolabes — quadrants turcs. (French) [Two astrolabes — Turkish quadrants]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19 (2):108–124, June 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wanscher:1975:HWJ

[430] Johan Henrik Wanscher. The history of Wilhelm Johannsen’s genetical terms and concepts from the period 1903 to 1926. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):125–147, June 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1975:Ra

[431] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):148–156, June 1975. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1975:RBP

[432] O. Pedersen. Review: Atlas of Planets — by V. D. Callatay and A. Dollfus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):148–149, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1975:RBN

[433] K. M. Pedersen. Review: Birth of Navigational Science — Solving in 18th-Century of Problem of Finding Longitude at Sea — by E. G. Forbes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):149, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1975:RBT

[434] O. Pedersen. Review: Scientific Thought 1900–1960 — Selective Survey — by R. Harre. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):150, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Keller:1975:RBI

[435] A. Keller. Review: Roman and Islamic Water-Lifting Wheels — by T. Schioler. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):151–152, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schioler:1975:RBR

[436] T. Schioler. Review: Windmills — Research Status with Respect to Origin and Occurrence — German — by J. C. Notebaart. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):152–153, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bastholm:1975:BMI

[437] E. Bastholm. History of Medical Illustrating. 2. 1600 to Present Day — German — by M. Putscher. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):153, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosenfeld:1975:RBS

[438] L. Rosenfeld. Review: Toward Scientific Research, 1822-1972 — German — by H. Querner and H. Schipperges. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):154, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sheppard:1975:RBA

[439] H. Sheppard. Review: Testament of Alchemy, Being Revelations of Morienus to Khalid Ibn Yazid — by L. Stavenhagen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):154–155, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moller-Christensen:1975:RBL

[440] V. Møller-Christensen. Review: Divided Legacy — History of Schism in Medical Thought — by H. L. Coulter. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):155, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1975:RBG

[441] K. M. Pedersen. Review: Letters — German — by L. Euler. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(2):155–156, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moesgaard:1975:EPL

[442] K. P. Moesgaard. Elements of planetary, lunar, and solar orbits, 1900 BC to AD 1900, tabulated for historical use. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(3):157– 181, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bikerman:1975:TCA

[443] J. J. Bikerman. Theories of capillary attraction. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19 (3):182–206, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoyer:1975:HDC

[444] U. Hoyer. How did Carnot calculate mechanical equivalent of heat? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(3):207–219, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Snelders:1975:RDC

[445] H. A. M. Snelders. Researches of Dutch chemists about nature of sulfides. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(3):220–237, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Keller:1975:RBC

[446] A. Keller. Review: Science and Civilization in China. 4. Physics and Physical Technology. 3. Civil Engineering and Nautics — by Joseph Needham. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(3):238–239, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Koning:1975:RBT

[447] D. A. W. Koning. Review: Medical and Technical Prescriptions of Central Netherlands — Dutch — by W. L. Braekman. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(3):239–240, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bruins:1975:PAE

[448] Evert M. Bruins. The part in Ancient Egyptian mathematics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):241–251, December 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Evans:1975:DOA

[449] Gillian R. Evans. Duc Oculum. Aids to understanding in some mediaeval treatises on the abacus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):252–263, December 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Janin:1975:CSJ

[450] Louis Janin. Un cadran solaire juif. (French) [A Jewish sundial]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):264–272, December 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cantor:1975:WRG

[451] G. N. Cantor. William Robert Grove, the correlation of forces, and the conservation of energy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):273–290, December 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schutt:1975:EAG

[452] Hans-Werner Schütt. Zur Entdeckungsgeschichte des Apatits. (German) [History of the discovery of Apatite]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):291–298, December 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rancke-Madsen:1975:DE

[453] E. Rancke-Madsen. The discovery of an element. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4): 299–313, December 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1975:Rb

[454] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):314–328, December 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Doldsamplonius:1975:RBW

[455] Y. Doldsamplonius. Review: Splendid Work on Arithmetic — French — by ?. Assamawal. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):314–315, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fleckenstein:1975:RBC

[456] J. O. Fleckenstein. Review: Important Commentaries on Dyadik by G. W. Leibniz — German — by H. Zacher. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):315– 317, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Saliba:1975:RBA

[457] G. A. Saliba. Review: Survey of Almagest — by O. Pedersen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):317–318, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

North:1975:RBB

[458] J. D. North. Review: Source Book in Medieval Science — by E. Grant. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):319–320, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

King:1975:RBU

[459] D. A. King. Review: Commentary Upon Birunis Kitab-Tahdid-Al-Amakin — by E. S. Kennedy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):320–323, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

North:1975:RBC

[460] J. D. North. Review: Reception of Copernicus Heliocentric Theory — by J. Dobrzycki. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):323–325, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Guerra:1975:RBO

[461] F. Guerra. Review: Ancient Ophthalmological Agents — Pharmaco-Historical Study of Collyria and Seals for Collyria Used During Roman Antiquity, as Well as of Most Frequent Components of Collyria — by H. Nielsen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):325, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bastholm:1975:RBC

[462] E. Bastholm. Review: Tenements of Clay — by A. Sorsby. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):326, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Trenn:1975:RBA

[463] T. J. Trenn. Review: Modern Age and Enlightenment — Origin of Modern Science and Philosophy — German — by J. Mittelstrass. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(4):326–328, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Drachmann:1976:KWH

[464] A. G. Drachmann. Ktesibios’s waterclock and Heron’s adjustable siphon. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):1–10, March 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lorch:1976:AIJ

[465] R. P. Lorch. The astronomical instruments of Jābir ibn Aflah and the Torquetum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):11–35, March 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1976:MVP

[466] B. L. van der Waerden. On the motion of Venus in the Pañcasiddhāntikā. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):35–44, March 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosen:1976:CB

[467] Edward Rosen. Copernicus’ Axioms. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):44–49, March 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sorensen:1976:SBP

[468] Knud Erik Sørensen. A study of the De proportione motus by Marcus Marci de Kronland. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):50–76, March 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1976:Ra

[469] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):77–78, March 1976. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Glahn:1976:RBC

[470] E. Glahn. Review: Science and Civilization in China — by Joseph Needham. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):77–78, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lohne:1976:RBN

[471] J. A. Lohne. Review: Correspondence of Newton, I — by Hall and Tilling. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):78–80, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hall:1976:RBB

[472] M. B. Hall. Review: Robert Boyle and Scientific Virtuoso Movement — Italian — by C. Pighetti. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):80–81, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heimann:1976:RBI

[473] P. M. Heimann. Review: Historical Introduction to Philosophy of Science — by J. Losee. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):81, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moesgaard:1976:SC

[474] K. P. Moesgaard. Symposium on Copernicus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1): 81–82, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Trenn:1976:RBP

[475] T. J. Trenn. Review: Time, Passion, and Knowledge — Reflections on Strategy of Existence — by J. T. Fraser. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):82– 83, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1976:RNEa

[476] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):84, March 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1976:SNBa

[477] Anonymous. Short notices of books received: Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(1):85–88, March 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bruin:1976:ERE

[478] Frans Bruin and Margaret Bruin. The equator ring, equinoxes and atmospheric refraction. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):89–111, June 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Unguru:1976:VEA

[479] Sabetai Unguru. A very early acquaintance with Apollonius of Perga’s treatise on conic sections in the Latin West. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):112– 128, June 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moesgaard:1976:BSZ

[480] Kristtan Peder Moesgaard and Leif Kahl Kristensen. The bright stars of the zodiac, a catalogue for historical use. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):129– 158, June 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Campbell:1976:EMR

[481] Margaret Campbell. Explanations of Mendel’s results. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):159–174, June 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1976:Rb

[482] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):175–184, June 1976. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ashfaque:1976:RBG

[483] S. M. Ashfaque. Review: Gears from Greeks — Antikythera Mechanism — Calendar Computer from ca 80 BC — by D. D. Price. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):175–176, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Olszewski:1976:RBA

[484] E. Olszewski. Review: Mechanics from Antiquity to Present Day — Russian — by A. T. Grigorian and I. B. Pogrebysski. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):176–179, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Neugebauer:1976:RBY

[485] O. Neugebauer. Review: Kepler, 400 Years — by A. Beer and P. Beer. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):179, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tekeli:1976:RBC

[486] S. Tekeli. Review: European Clocks and Watches in the Near East — by O. Kyrz. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):179–181, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Guerra:1976:RBB

[487] F. Guerra. Review: Historical Bibliography on Science and Technology in Spain — Spanish — by J. M. López Piñero, M. Pesetroig, and L. García Ballester. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):181, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Petri:1976:RBA

[488] W. Petri. Review: Development of 1st Academy of Natural Sciences — Russian — by J. C. Kopelevic. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):181–182, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moller-Christensen:1976:RBL

[489] V. Møller-Christensen. Review: Divided Legacy — History of Schism in Medical Thought. 1. Patterns Emerge — Hippocrates to Paracelsus — by H. L. Coulter. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):182–183, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Snorrason:1976:RBR

[490] E. Snorrason. Review: History of Royal College of Physicians of London — by A. M. Cooke. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):183–184, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ranckemadsen:1976:RBD

[491] E. Ranckemadsen. Review: John Dalton — Critical-Assessment of His Life and Science — by A. Thackray. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):184, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1976:SNBb

[492] Anonymous. Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2): 185–188, June 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sesiano:1976:MIA

[493] Jacques Sesiano. Un mémoire d’Ibn al-Hay\underline{th}am sur un problème arithmétique solide. (French) [A note of Ibn Al-Hay\underline{th}am on a solid arithmetic problem]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3):189–195, September 1976. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoyer:1976:UDK

[494] Ulrich Hoyer. Über die Unvereinbarkeit der drei Keplerschen Gesetze mit der Aristotelischen Mechanik. (German) [On the incompatibility of Kepler’s three laws with Aristotelian mechanics]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3): 196–209, September 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Laudan:1976:WSS

[495] Rachel Laudan. William Smith. Stratigraphy without palaeontology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3):210–226, September 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ruse:1976:SMW

[496] Michael Ruse. The scientific methodology of William Whewell. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3):227–257, September 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1976:Rc

[497] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3):258–267, September 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pingree:1976:RBS

[498] D. Pingree. Review: Heliocentric System in Greek, Persian and Indican Astronomy — German — Naturforschende Gesellschaft in Zurich — by B. L. van der Waerden. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3):258–260, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Scriba:1976:RBC

[499] C. J. Scriba. Review: Africa Counts — Numbers and Pattern in African Culture — by C. Zaslavsky. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3):260–261, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Telle:1976:RBM

[500] J. Telle. Review: Problem of Metals in Antique Science — French — by R. Halleux. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3):261–263, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mehrtens:1976:RBV

[501] H. Mehrtens. Review: Science and Values — Patterns of Tradition and Change — by A. Thackray and E. Mendelsohn. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20 (3):263–264, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Keller:1976:RBN

[502] A. G. Keller. Review: Diffusion of New Industrial Processes — International Study — by L. Nabseth and G. F. Ray. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20 (3):264–265, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1976:RBG

[503] K. M. Pedersen. Review: Harold G. J. Moseley: Life and Letters of an English Physicist, 1887–1915 — by J. L. Heilbron. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3):265–266, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Petri:1976:HMS

[504] W. Petri. Historical-mathematical studies. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(3):266– 267, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Taisbak:1976:PNM

[505] C. M. Taisbak. Perfect numbers: a mathematical pun? An analysis of the last theorem in the ninth book of Euclid’s Elements. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20 (4):269–275, December 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rohr:1976:CSA

[506] Richard Rohr. Le cadran solaire analemmatique. (French) [The analemmatic sundial]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):276–278, December 1976. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hutchison:1976:MHS

[507] Keith Hutchison. Mayer’s hypothesis: a study of the early years of thermodynamics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):279–304, December 1976. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Frankel:1976:AWS

[508] Henry Frankel. Alfred Wegener and the specialists. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4): 305–324, December 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1976:SNBc

[509] Anonymous. Short notices of books received: Short notices of books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):325–326, December 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1976:RNEb

[510] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):327–329, December 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brackbernsen:1976:RBA

[511] L. Brackbernsen. Review: Science Awakening. 2. — by B. L. van der Waerden. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):327–328, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thorsen:1976:RBQ

[512] J. Thorsen. Review: History of Quantum Theory — German — by F. Hund. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):328–329, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1976:RBC

[513] O. Pedersen. Review: Isis Cumulative Bibliography — by M. Whitrow. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):330–331, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Scriba:1976:RBI

[514] C. J. Scriba. Review: Dialog on Introduction to Arithmetic and Algebra — German — by G. W. Leibniz. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):331–332, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heimann:1976:RBA

[515] P. M. Heimann. Review: Method and Appraisal in Physical Sciences — Critical Background to Modern Science, 1800–1905 — by C. Howson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):332, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1976:Rd

[516] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(4):330–332, December 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heimann:1977:GNL

[517] P. M. Heimann. ≫geometry and nature≪: Leibniz and Johann Bernoulli’s theory of motion. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(1):1–26, March 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wiederkehr:1977:FIR

[518] Karl Heinrich Wiederkehr. Von frühen Ideen über eine regelmäßige Gestalt kleinster Materieteilchen bis zu Delisles und Bergmans Vorarbeiten für Haüys Kristallstrukturtheorie. (German) [Early ideas of a regular shape smallest particles of matter up to Delisle and Bergmans preparations for haüys’ crystal structure theory. (German)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(1):27–43, March 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kauffman:1977:CWB

[519] George B. Kauffman. Christian Wilhelm Blomstrand (1826–1897) and Sophus Mads Jørgensen (1837–1914). Their correspondence from 1870 to 1897. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(1):44–63, March 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Waff:1977:NMM

[520] Crajg B. Waff. Newton and the motion of the Moon an essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(1):64–75, March 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1977:Ra

[521] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(1):76–84, March 1977. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1977:BRB

[522] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(1): 85–87, March 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sesiano:1977:MAI

[523] Jacques Sesiano. Les méthodes d’analyse indéterminée chez Abu Kasmil. (French) [the indeterminate-analysis methods of Abu Kasmil]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(2):89–105, June 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lohne:1977:NEC

[524] J. A. Lohne. Nova Experimenta Crystalli Islandici Disdiaclastici. (Latin) [new experiments on Icelandic birefringent crystals]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(2):106–148, June 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ashfaque:1977:AIV

[525] Syed Mohammad Ashfaque. Astronomy in the Indus valley civilization: a survey of the problems and possibilities of the ancient Indian astronomy and cosmology in the light of Indus script decipherment by the Finnish scholars. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(2):149–193, June 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1977:Rb

[526] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(2):194–200, June 1977. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Swerdlow:1977:SDP

[527] Noel M. Swerdlow. A summary of the derivation of the parameters in the Commentariolus from the Alfonsine Tables with an appendix on the length of the tropical year in Abraham Zacuto’s Almanach Perpetuum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(3–4):201–213, December 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Folkerts:1977:RMG

[528] Menso Folkerts. Regiomontanus als Mathematiker. (German) [Regiomontanus as mathematician]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(3–4):214–245, December 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sorensen:1977:SBP

[529] Knud Erik Sørensen. A study of the De proportione motus by Marcus Marci de Kronland. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(3–4):246–277, December 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wiederkehr:1977:RJH

[530] Karl Heinrich Wiederkehr. René-Just Haüys Strukturtheorie der Kristalle. (German) [René-Just Haüy’s crystal-structure theory]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(3–4):278–299, December 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1977:Rc

[531] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 21(3–4):300–328, December 1977. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Allard:1978:PTI

[532] André Allard. Le petit traité d’Isaac Argyre sur la racine carrée. (French) [The little treatist of Isaac Argyre on the square root]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(1):1–43, September 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hamadanizadeh:1978:ISD

[533] Javad Hamadanizadeh. Interpolation schemes in Dastūr al-Munajjūn. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(1):44–52, September 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knudsen:1978:EDD

[534] Ole Knudsen. Electric displacement and the development of optics after Maxwell. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(1):53–60, September 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1978:Ra

[535] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(1):61–71, September 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1978:RNE

[536] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(1):72–76, September 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knorr:1978:ANC

[537] Wilbur R. Knorr. Archimedes’ Neusis-constructions in spiral lines. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(2):77–98, December 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1978:NBP

[538] Kirsti Møller Pedersen. A note on Bartholin and the problem of Debeaune. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(2):99–107, December 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nicholas:1978:NES

[539] John M. Nicholas. Newton’s extremal second law. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(2): 108–130, December 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wiederkehr:1978:RJH

[540] Karl Heinrich Wiederkehr. René-Just Haüys Konzeption vom individuellen integrierenden Molekül, ihre Widerlegung, und seine Ansichten über kristallbildende Kräfte. (German) [René-Just Haüy’s concept of individually-integrated molecules, their refutation, and his views on crystal-forming power]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(2):131–156, December 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1978:Rb

[541] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(2):157–161, December 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1978:BRB

[542] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(2): 162–164, December 1978. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sharma:1979:LDJ

[543] S. D. Sharma and S. S. Lishk. Length of the day in Jaina astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(3):165–176, March 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wiederkehr:1979:WHI

[544] Karl Heinrich Wiederkehr. Das Weiterwirken der Haüyschen Idee von der Polyedergestalt der Moleküle in der Chemie, die Umgestaltung der Haüyschen Strukturtheorie durch Seeber und Delafosse, und Bravais’ Entdeckung der Gittertypen. (German) [The continuing influence of the idea of the Haüys polyhedral theory of molecules in chemistry, the transformation of Haüys’ structural theory by Seeber and Delafosse, and Bravais’ discovery of lattice type]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(3):177–186, March 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kaiser:1979:KBB

[545] Walter Kaiser. Karl Bädekers Beitrag zur Halbleiterforschung. (German) [Karl Baedeker’s contribution to semiconductor research]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(3):187–200, March 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Roll-Hansen:1979:GTW

[546] Nils Roll-Hansen. The genotype theory of Wilhelm Johannsen and its relation to plant breeding and the study of evolution. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(3):201–235, March 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1979:RNE

[547] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(3):236– 240, March 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1979:Ra

[548] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(3):241–256, March 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sachs:1979:ONB

[549] J. Sachs and G. J. Toomer. Otto Neugebauer, bibliography, 1925–1979. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(4):257–280, June 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sesiano:1979:NTT

[550] J. Sesiano. Note sur trois théorèmes de mécanique d’al-Qūhī et leur conséquence. (French) [Note on three theorems in mechanics by d’al-Qūī and their consequences]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(4):281–297, June 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Janin:1979:ACS

[551] Louis Janin. Astrolabe et cadran solaire en projection stéréographique horizontale. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(4):298–314, June 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Briginshaw:1979:AGS

[552] A. J. Briginshaw. The axiomatic geometry of space–time: An assessment of the work of A. A. Robb. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(4):315–323, June 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1979:Rb

[553] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 22(4):324–332, June 1979. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Paisley:1979:SSA

[554] P. B. Paisley and D. R. Oldroyd. Science in the Silver Age: Aetna, a classical theory of volcanic activity. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(1):1–20, September 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ziggelaar:1979:ERD

[555] August Ziggelaar. Die Erklärung des Regenbogens durch Marcantonio de Dominis, 1611. Zum Optikunterricht am Ende des 16. Jahrhunderts. (German) [The explanation of the rainbow by Marc Antonio de Dominis, 1611. On teaching optics at the end of the 16th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(1):21–50, September 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Buchwald:1979:HEMa

[556] J. Z. Buchwald. The Hall effect and Maxwellian electrodynamics in the 1880’s. Part I: the discovery of a new electric field. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23 (1):51–99, September 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1979:BRB

[557] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(1): 100–104, September 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

King:1979:NLB

[558] David A. King. New light on the Zīij al-Safāa’ih of Abū Ja^{c} far al-Khāzin. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(2):105–117, December 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Buchwald:1979:HEMb

[559] J. Z. Buchwald. The Hall effect and Maxwellian electrodynamics in the 1880’s. Part II: the unification of theory, 1881–1893. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23 (2):118–162, December 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Truesdell:1979:ERB

[560] C. Truesdell. An essay review of Geschichte der mechanischen Prinzipien und ihrer wichtigsten Anwendungen. (German) [History of mechanical principles and their main applications]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(2):163–175, December 1979. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1979:Rc

[561] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(2):176, December 1979. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Folkerts:1980:PEI

[562] Menso Folkerts. Probleme der Euklidinterpretation und ihre Bedeutung für die Entwicklung der Mathematik. (German) [Problems of Euclid interpretation and their importance for the development of mathematics]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(3):185–215, March 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jervis:1980:VCE

[563] Jane L. Jervis. Vögelin on the comet of 1532: Error analysis in the 16th Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(3):216–229, March 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hendry:1980:NTC

[564] John Hendry. Newton’s theory of colour. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(3):230– 251, March 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1980:Ra

[565] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(3):252–255, March 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1980:TRP

[566] B. L. van der Waerden. The (2: n ) table in the Rhind Papyrus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(4):259–274, June 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:1980:APC

[567] Helge Kragh. Anatomy of a priority conflict: the case of element 72. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(4):275–301, June 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

McMullin:1980:GAP

[568] Ernan McMullin. Galileo against the philosophers an essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(4):302–311, June 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1980:Rb

[569] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(4):312–313, June 1980. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1980:RNE

[570] Anonymous. Reprints and new editions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 23(4):329– 332, June 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Neugebauer:1980:OP

[571] O. Neugebauer. On the orientation of pyramids. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24 (1):1–3, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schmidt:1980:PPN

[572] Olaf Schmidt. On Plimpton 322. Pythagorean numbers in Babylonian mathematics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):4–13, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aaboe:1980:OTB

[573] Asger Aaboe. Observation and theory in Babylonian astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):14–35, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brack-Bernsen:1980:SIE

[574] Lis Brack-Bernsen. Some investigations on the ephemerides of the Babylonian Moon texts, System A. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):36–50, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moesgaard:1980:FMS

[575] Kristian Peder Moesgaard. The full Moon serpent. A foundation stone of ancient astronomy?: the full Moon serpent. A foundation stone of ancient astronomy? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):51–96, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Toomer:1980:HEB

[576] G. J. Toomer. Hipparchus’ empirical basis for his lunar mean motions: a historical footnote to Olaf Pedersen, a survey of the Almagest, 161– 64. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):97–109, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Taisbak:1980:APH

[577] Chr. M. Taisbak. An Archimedean proof of Heron’s formula for the area of a triangle; reconstructed. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):110–116, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1980:CBC

[578] B. L. van der Waerden. The conjunction of 3102 B.C. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):117–131, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1980:SMA

[579] Bernard R. Goldstein. The status of models in Ancient and Medieval astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):132–147, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hartner:1980:UVL

[580] Willy Hartner. An unusual value for the length of the meridian degree: 66 1/2 miles, in Ibn Yūnus’ H .ākimitic Zīj. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):148–152, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lorch:1980:SSR

[581] Richard Lorch. The sphera solida and related instruments. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):153–161, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kennedy:1980:AEP

[582] E. S. Kennedy. Astronomical events from a Persian astrological manuscript. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):162–177, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gingerich:1980:AKA

[583] Owen Gingerich. Appendix to E. S. Kennedy “astronomical events from a Persian astrological manuscript”. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):178–180, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

North:1980:AFC

[584] J. D. North. Astrology and the fortunes of churches. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):181–211, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Swerdlow:1980:LPM

[585] Noel M. Swerdlow. Long-period motions of the Earth in De revolutionibus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):212–245, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ziggelaar:1980:SLR

[586] August Ziggelaar. The Sine Law of Refraction derived from the Principle of Fermat — prior to Fermat? The theses of Wilhelm Boelmans S.J. in 1634. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):246–262, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lutzen:1980:RBP

[587] Jesper Lützen. The relationship between Pascal’s mathematics and his philosophy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):263–272, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Shapiro:1980:NHE

[588] Alan E. Shapiro. Newton and Huygens’ explanation of the 22^{∘} halo. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):273–287, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Whiteside:1980:KNF

[589] D. T. Whiteside. Kepler, Newton and Flamsteed on refraction through a ‘regular aire’: the mathematical and the practical. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):288–315, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Andersen:1980:IMD

[590] Kirsti Andersen. An impression of mathematics in Denmark in the period 1600–1800. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):316–334, September 1980. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:1980:RJB

[591] Kurt Møller Pedersen. Roger Joseph Boscovich and John Robison on terrestrial aberration. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):335–345, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knudsen:1980:CVI

[592] Ole Knudsen. 19th Century views on induction in moving conductors. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):346–360, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pihl:1980:TCV

[593] Mogens Pihl. Two contributions of L. V. Lorenz to mathematical physics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 24(1):361–368, September 1980. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Folkerts:1981:MMH

[594] Menso Folkerts. Mittelalterliche mathematische Handschriften in westlichen Sprachen in der Herzog August Bibliothek Wolfenbüttel Ein vorläufiges Verzeichnis. (German) [Mathematical Medieval manuscripts in Western languages in the Herzog August Library in Wolfenbüttel: A provisional list]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(1):1–49, April 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Smith:1981:MLA

[595] G. C. Smith. De Morgan and the laws of algebra. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(1):50–70, April 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brown:1981:YPM

[596] Laurie M. Brown. Yukawa’s prediction of the meson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(1):71–132, April 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1981:Ra

[597] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(1):133–154, April 1981. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Evans:1981:BGF

[598] Gillian R. Evans. Boethius’ geometry and the four ways. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(2):161–165, July 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Siesser:1981:CGE

[599] William G. Siesser. Christian Gottfried Ehrenberg: Founder of micropaleontology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(2):166–188, July 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hendry:1981:BKS

[600] John Hendry. Bohr–Kramers–Slater: a virtual theory of virtual oscillators and its role in the history of quantum mechanics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(2):189–221, July 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Glen:1981:FPA

[601] William Glen. The first potassium–argon geomagnetic polarity reversal time scale; a premature start by Martin G. Rutten. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(2):222–238, July 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1981:Rb

[602] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(2):239–252, July 1981. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knorr:1981:HCC

[603] Wilbur Richard Knorr. The hyperbola-construction in the conics, book II: Ancient variations on a theorem of Apollonius. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25 (3):253–291, October 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Neugebauer:1981:CPC

[604] O. Neugebauer. On the Computus Paschalis of “Cassiodorus”. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(3):292–302, October 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1981:ACC

[605] B. R. Goldstein and D. Pingree. Astronomical computations for 1299 from the Cairo Geniza. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(3):303–318, October 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller:1981:MRD

[606] Gerhard H. Müller and Marshall R. Wheeler. Misinterpretations of Réaumur’s description of small flies. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(3):319–327, October 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1981:Rc

[607] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(3):328–348, October 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Klein:1982:AEF

[608] M. J. Klein. Albert Einstein — four commemorative lectures — Loyd S. Swenson, Jr., C. P. Snow, Howard Stein, and Ilya Prigogine. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(3):344–345, October 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knorr:1982:OEH

[609] Wilbur R. Knorr. Observations on the early history of the conies. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(1):1–24, April 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ullmann:1982:EMS

[610] Dieter Ullmann. Die ersten Messungen der Schallgeschwindigkeit in Luft und das Schallortungsverfahren von Jonas Meldercreutz. (German) [The first measurements of the speed of sound in air and the sound localization method of Jonas Meldercreutz]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(1):25–37, April 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Campbell:1982:MTC

[611] Margaret Campbell. Mendel’s theory: Its context and plausibility. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(1):38–69, April 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1982:Ra

[612] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(1):70–81, April 1982. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1982:BRBa

[613] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(1): 82–84, April 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lervig:1982:WHC

[614] Philip Lervig. What is heat? C. Truesdell’s view of thermodynamics. A critical discussion. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(2):85–122, July 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jordan:1982:DHS

[615] D. W. Jordan. D. E. Hughes self-induction and the skin-effect. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(2):123–153, July 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:1982:ESW

[616] Helge Kragh. Erwin Schrödinger and the wave equation: the crucial phase. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(2):154–197, July 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Grythe:1982:SRE

[617] Inge Grythe. Some remarks on the early S-matrix. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(2):198–203, July 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Haage:1982:ABN

[618] Bernhard D. Haage. Das alchemistische Bildgedicht vom ‘Nackten Weib’ in seiner bisher ältesten Überlieferung. (German) [The alchemical image from the poem ‘Naked woman’ in its earliest archaeological lore]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(2):204–214, July 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Evans:1982:BCB

[619] Gillian R. Evans. Brief communications: Brief communications. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(2):215–217, July 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

King:1982:BCB

[620] David A. King. Brief communications: Brief communications. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(2):217–218, July 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1982:Rb

[621] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(2):219–240, July 1982. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goodstein:1982:IMR

[622] Judith R. Goodstein. The Italian mathematicians of Relativity. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(3):241–261, October 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Galison:1982:DMF

[623] Peter Galison. The discovery of the muon and the failed revolution against quantum electrodynamics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(3):262–316, October 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Melnikov:1982:SDP

[624] V. P. Mel’nikov. Some details in the prehistory of the discovery of element 72. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(3):317–322, October 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Truesdell:1982:ECT

[625] C. Truesdell. Euler’s contribution to the theory of ships and mechanics: An essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(3):323–335, October 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1982:Rc

[626] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(3):336–340, October 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1982:BRBb

[627] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 26(3): 341–344, October 1982. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Larsen:1984:PPE

[628] Mogens Esrom Larsen. On the possibility of a pre-Euclidean theory of proportions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(1):1–25, April 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sesiano:1984:AML

[629] Jacques Sesiano. Une arithmétique médiévale en langue provençale. (French) [A Medieval arithmetic in the Provençal language]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(1):26–75, April 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kraft:1984:ASK

[630] P. Kraft and P. Kroes. Adaptation of scientific knowledge to an intellectual environment. Paul Forman’s “Weimar Culture, Causality, and Quantum Theory, 1918–1927”: Analysis and criticism. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(1):76–99, April 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1984:Ra

[631] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(1):100–108, April 1984. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kennedy:1984:TPA

[632] E. S. Kennedy. Two Persian astronomical treatises by Nasīr al-Dīn al Tūsī. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(2):109–120, July 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

OHara:1984:GMM

[633] James G. O’Hara. Gauß‘ method for measuring the terrestrial magnetic force in absolute measure: Its invention and introduction in geomagnetic research. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(2):121–147, July 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Barker:1984:SCP

[634] Peter Barker and Bernard R. Goldstein. Is Seventeenth Century physics indebted to the Stoics? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(2):148–164, July 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Taisbak:1984:BCE

[635] C. M. Taisbak. Brief communications: Eleven eighty-thirds. Ptolemy’s reference to Eratosthenes in Almagest 1.12. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(2):165– 167, July 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vida:1984:BCN

[636] Tivadar Vida. Brief communications: Niels Stensens Ungarnreise im Jahre 1669. (German) [Niels Stensen’s Hungarian voyage in the year 1669]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(2):167–172, July 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1984:Rb

[637] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(2):173–188, July 1984. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Boker:1984:NIA

[638] R. Böker. Über Namen und Identifizierung der ägyptischen Dekane. (German) [The name and identification of the Egyptian Deans]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(3):189–217, October 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanMaanen:1984:HVH

[639] Jan A. van Maanen. Hendrick van Heuraet (1634–1660?): His life and mathematical work. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(3):218–279, October 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Maeyama:1984:ASO

[640] Y. Maeyama. Ancient stellar observations: Timocharis, Aristyllus, Hipparchus, Ptolemy — the dates and accuracies. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27 (3):280–310, October 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:1984:GST

[641] Alexander Jones. A Greek Saturn table. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(3):311–317, October 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1984:Rc

[642] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(3):318–336, October 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1984:BRB

[643] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 27(3): 337–339, October 1984. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Berggren:1985:OAB

[644] J. L. Berggren. The origins of al-Bīrūnī’s “Method of the Zījes” in the theory of sundials. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(1):1–16, April 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Scheele:1985:ZHK

[645] Irmtraut Scheele. Der Zoologe Henrik Krøyer (1799–1870) und die deutsch-dänischen Wissenschaftsbeziehungen. (German) [The zoologist Henrik Krøyer (1799–1870) and German–Danish scientific relations]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(1):17–30, April 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fraser:1985:DPOa

[646] Craig Fraser. D’Alembert’s principle: the original formulation and application in Jean d’Alembert’s Traité de Dynamique (1743). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(1):31–61, April 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1985:Ra

[647] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(1):62–93, April 1985. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hulden:1985:AGD

[648] Bjarne Huldén, Kate Larsen, and Olaf Pedersen. Aage Gerhard Drachmann (1891–1980): a bibliography. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(2):102–107, July 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Langermann:1985:BBD

[649] Y. Tzvi Langermann. The book of bodies and distances of Habash al-Hāsīb. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(2):108–128, July 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Curchin:1985:QDP

[650] Leonard Curchin and Roger Herz-Fischler. De quand date le premier rapprochement entre la suite de Fibonacci et la division en extreme et moyenne raison?. (French) [When was the first reconciliation between the Fibonacci series and division in extreme and mean reason?]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(2):129–138, July 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Andersen:1985:GMT

[651] Kirsti Andersen and Henrik Meyer. Georg Mohr’s three books and the Gegenübung auf Compendium Euclidis Curiosi. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28 (2):139–144, July 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fraser:1985:DPOb

[652] Craig Fraser. D’Alembert’s Principle: the original formulation and application in Jean d’Alembert’s Traité de Dynamique (1743). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(2):145–159, July 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1985:Rb

[653] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(2):160–162, July 1985. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1985:BRB

[654] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(2): 163–167, July 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hogenduk:1985:DGC

[655] Jan P. Hogenduk. Diocles and the geometry of curved surfaces. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(3):169–184, October 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1985:SLH

[656] Bernard R. Goldstein. Star lists in Hebrew. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(3):185–208, October 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chekin:1985:ERM

[657] L. S. Chekin. Elements of the rational method in Gervase of Tilbury’s cosmology and geography. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(3):209–217, October 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Guiot:1985:SUB

[658] Jean-Paul Guiot. Sechs unveröffentlichte Briefe von Johann Wilhelm Ritter (1776–1810). (German) [Six unpublished letters of Johann Wilhelm Ritter (1776–1810)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(3):218–243, October 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonRuger:1985:MTK

[659] Alexander von Rüger. Die Molekularhypothese in der Theorie der Kapillarerscheinungen (1805–1873). (German) [The molecular hypothesis in the theory of capillary phenomena (1805–1873)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28 (3):244–276, October 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ohrstrom:1985:GWI

[660] Peter Øhrstrøm. Guldberg and Waage on the influence of temperature on the rates of chemical reactions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(3):277–287, October 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Buchwald:1985:OHM

[661] Jed Z. Buchwald. Oliver Heaviside, Maxwell’s apostle and Maxwellian apostate. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(3):288–330, October 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1985:Rc

[662] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(3):331–348, October 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Couchoud:1986:ENI

[663] Sylvia Couchoud. Essai d’une nouvelle interprétation du premier problème du Papyrus Mathématique Démotique 10520 du British Museum. (French) [Essay on a new interpretation of the first problem of the Demotic Mathematical Papyrus 10520 in the British Museum]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(1):1–4, March 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bromley:1986:NAM

[664] Allan G. Bromley. Notes on the Antikythera mechanism. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(1):5–27, March 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Yano:1986:CYJ

[665] Michio Yano. The ch’i-yao jang-tsai-chüeh and its ephemerides. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(1):28–35, March 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dobrzycki:1986:OCT

[666] Jerzy Dobrzycki. One Copernican table. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(1):36–39, March 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ullmann:1986:HKW

[667] Dieter Ullmann. Helmholtz–Koenig–Waetzmann und die Natur der Kombinationstöne. (German) [Helmholtz–Koenig–Waetzmann and the nature of the combination tones]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(1):40–52, March 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Snelders:1986:DPC

[668] H. A. M. Snelders. The Dutch physical chemist J. J. van Laar (1860– 1938) versus J. H. van’t Hoff’s ‘Osmotic School’. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(1):53–71, March 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1986:Ra

[669] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(1):72–83, March 1986. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1986:BRBa

[670] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(1): 84–87, March 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hofmann:1986:JBK

[671] J. E. Hofmann. Johann Bernoullis Kreisrektifikation durch Evolventenbildung. (German) [Johann Bernoulli’s circle rectification by formation of the involute]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(2):89–99, June 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nugayev:1986:SRS

[672] R. M. Nugayev. Special Relativity as a step in the development of the quantum programme: Revolution in a revolution. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(2):100–109, June 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cushing:1986:IHM

[673] James T. Cushing. The importance of Heisenberg’s S-matrix program for the theoretical high-energy physics of the 1950’s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(2):110–149, June 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1986:Rb

[674] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(2):150–164, June 1986. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rohr:1986:CSU

[675] René R. J. Rohr. Les cadrans solaires universels de Jacques Ozanam. (French) [The sundials of Jacques Ozanam]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(3):165–177, September 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DAgostino:1986:MDA

[676] S. D’Agostino. Maxwell’s dimensional approach to the velocity of light. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(3):178–204, September 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nersessian:1986:WWL

[677] Nancy J. Nersessian. “Why wasn’t Lorentz Einstein?” An examination of the scientific method of H. A. Lorentz. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(3): 205–242, September 1986. CODEN CENTA4. ISSN 0008-8994 (print),

1600-0498 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1111/j.1600-0498.1986.tb00857.x/abstract.

Anonymous:1986:Rc

[678] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(3):243–248, September 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Saliba:1986:DNP

[679] George Saliba. The determination of new planetary parameters at the Maragha Observatory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(4):249–271, December 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1986:LBG

[680] Bernard R. Goldstein. Levi ben Gerson’s theory of planetary distances. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(4):272–313, December 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Leegwater:1986:DWO

[681] Arie Leegwater. The development of Wilhelm Ostwald’s chemical energetics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(4):314–337, December 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1986:BRBb

[682] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(4): 338–339, December 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:1987:SBM

[683] Alexander Jones. On some borrowed and misunderstood problems in Greek catoptrics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(1):1–17, April 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sesiano:1987:SMH

[684] Jacques Sesiano. Survivance médiévale en hispanie d’un problème né en Mésopotamie. (French) [Medieval survival in New Spain of a problem born in Mesopotamia]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(1):18–61, April 1987. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kurola:1987:CGS

[685] Osmo Kurola. Classical geometry in the service of geodesy: Samuel Klingenstierna’s calculation of the radius of curvature of the ellipse. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(1):62–85, April 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1987:Ra

[686] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(1):86–89, April 1987. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1987:Rb

[687] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(1):90–92, April 1987. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Roseman:1987:HTT

[688] Christina Horst Roseman. Hour tables and Thule in Pliny’s Natural History. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(2):93–105, July 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Andersen:1987:CPO

[689] Kirsti Andersen. The central projection in one of Ptolemy’s map constructions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(2):106–113, July 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:1987:BWG

[690] Helge Kragh. The beginning of the world: Georges Lemaître and the expanding universe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(2):114–139, July 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Treitel:1987:CTD

[691] Jonathan Treitel. Confirmation with technology: the discovery of the Tau lepton. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(2):140–180, July 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1987:Rc

[692] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(2):181–195, July 1987. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1987:ASP

[693] B. L. van der Waerden. The astronomical system of the Persian tables II. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(3):197–211, October 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brentjes:1987:UNA

[694] Sonja Brentjes. Untersuchungen zum Nicomachus Arabus. (German) [Studies on Nicomachus Arabus]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(3):212–292, October 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1987:Rd

[695] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(3):293–309, October 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1987:BRB

[696] Anonymous. Book received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 30(3): 310–314, October 1987. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kennedy:1988:TMA

[697] E. S. Kennedy. Two Medieval approaches to the equation of time. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(1):1–8, April 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bos:1988:TML

[698] H. J. M. Bos. Tractional motion and the legitimation of transcendental curves. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(1):9–62, April 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wiederkehr:1988:CSW

[699] K. H. Wiederkehr. Christian Samuel Weiß (1780–1856) unter dem Einfluss der romantischen Naturphilosophie und sein Briefwechsel mit H. Chr. Oersted über Atomistik und Dynamik. (German) [Christian Samuel Weiss (1780–1856) under the influence of Romantic philosophy of nature and his correspondence with H. Chr. Ørsted about atomism and dynamism]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(1):63–85, April 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Buchwald:1988:RHE

[700] Jed Buchwald. The rational and the historical: An essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(1):86–92, April 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1988:Ra

[701] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(1):93–102, April 1988. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderWaerden:1988:MVG

[702] B. L. van der Waerden. The motion of Venus in Greek, Egyptian and Indian texts. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(2):105–113, July 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gupta:1988:NIV

[703] R. C. Gupta. New Indian values of π from the Mānava Súlba Sūtra. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(2):114–125, July 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonSchmeidler:1988:RWH

[704] Felix von Schmeidler. Über die Realität der von W. Herschel berichteten Beobachtungen der Ringe des Uranus. (German) [On the reality of the observations of the rings of Uranus reported by W. Herschel]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(2):126–140, July 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Archibald:1988:TPO

[705] Thomas Archibald. Tension and potential from Ohm to Kirchhoff. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(2):141–163, July 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fischer:1988:BC

[706] K. A. F. Fischer. Brief communication. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(2):164– 167, July 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1988:Rb

[707] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(2):168–180, July 1988. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muroi:1988:ECR

[708] Kazuo Muroi. Extraction of cube roots in Babylonian mathematics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(3):181–188, October 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Neugebauer:1988:GN

[709] O. Neugebauer and G. Saliba. On Greek numerology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31 (3):189–206, October 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aiton:1988:PPS

[710] E. J. Aiton. Polygons and parabolas: Some problems concerning the dynamics of planetary orbits. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(3):207–221, October 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Weiss:1988:WTM

[711] Burghard Weiss. Die Widerlegung der Theorie der molécules puppiformes. Eine verschollene Abhandlung von Augustin Jean Fresnel. (German) [The refutation of the theory of puppiforme molecules. A lost treatise of Augustin Jean Fresnel]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(3):222–258, October 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schubert:1988:VEW

[712] Helmut Schübert. Das Verschwinden des elektrischen Widerstandes. (German) [hTe disappearance of electrical resistance]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31 (3):259–299, October 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hentschel:1988:ZVT

[713] Klaus Hentschel. Zwei vergessene Texte Moritz Schlicks. (German) [Two forgotten texts of Moritz Schlick]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(3):300–311, October 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1988:Rc

[714] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(3):312–337, October 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1988:BRB

[715] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(3): 338–340, October 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Britton:1989:EFE

[716] John P. Britton. An early function for eclipse magnitudes in Babylonian astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(1):1–52, April 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sesiano:1989:KZG

[717] Jacques Sesiano. Koptisches Zahlensystem und (griechisch-)koptische Multiplikationstafeln nach einem arabischen Bericht. (German) [Coptic number system and (Greek-) Coptic multiplication tables according to an Arab report]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(1):53–65, April 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DAgostino:1989:PHN

[718] S. D’Agostino. Pourquoi Hertz et non pas Maxwell, a-t-il découvert les ondes électriques?. (French) [Why Hertz, and not Maxwell, discovered electric waves]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(1):66–76, April 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1989:Ra

[719] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(1):77–83, April 1989. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanDalen:1989:SMR

[720] Benno van Dalen. A statistical method for recovering unknown parameters from Medieval astronomical tables. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(2):85–145, July 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rochberg-Halton:1989:BSH

[721] Francesca Rochberg-Halton. Babylonian seasonal hours. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(2):146–170, July 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hogendijk:1989:MST

[722] Jan P. Hogendijk. The mathematical structure of two Islamic astrological tables for ‘casting the rays’. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(2):171–202, July 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:1989:CCJ

[723] Helge Kragh. Concept and controversy: Jean Becquerel and the positive electron. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(2):203–240, July 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1989:Rb

[724] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(2):241–248, July 1989. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1989:BRB

[725] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(2): 249–256, July 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knorr:1989:ACR

[726] Wilbur R. Knorr. On Archimedes’ construction of the regular heptagon. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(3):257–271, October 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1989:EHA

[727] Bernard R. Goldstein and Alan C. Bowen. On early Hellenistic astronomy: Timocharis and the first Callippic calendar. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32 (3):272–293, October 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Puig:1989:AZG

[728] Roser Puig. Al-Zarqālluh’s graphical method for finding lunar distances. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(3):294–309, October 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moesgaard:1989:TBD

[729] Kristian Peder Moesgaard. Tycho Brahe’s discovery of changes in star latitudes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(3):310–323, October 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jensen:1989:EHO

[730] K. A. Jensen. The early history of organic sulfur chemistry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(3):324–335, October 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kryzhanovsky:1989:BCI

[731] L. N. Kryzhanovsky. Brief communication: the invention of the resistor. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(3):336–339, October 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1989:Rc

[732] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 32(3):340–351, October 1989. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brack-Bernsen:1990:BTB

[733] Lis Brack-Bernsen and Olaf Schmidt. Bisectable trapezia in Babylonian mathematics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(1):1–38, April 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brack-Bernsen:1990:BLT

[734] Lis Brack-Bernsen. On the Babylonian lunar theory: a construction of column Φ from horizontal observations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(1):39– 56, April 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Britton:1990:TTC

[735] John P. Britton. A tale of two cycles: Remarks on column Φ. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(1):57–69, April 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mancha:1990:IAH

[736] J. L. Mancha. Ibn al-Haytham’s homocentric epicycles in Latin astronomical texts of the XIVth and XVth centuries. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(1): 70–89, April 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1990:Ra

[737] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(1):90–91, April 1990. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:1990:BGA

[738] Alexander Jones. Babylonian and Greek astronomy in a papyrus concerning Mars. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(2):97–114, July 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

North:1990:ASM

[739] J. D. North. Astronomical symbolism in the Mithraic religion. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(2):115–148, July 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hayashi:1990:CMS

[740] T. Hayashi, T. Kusuba, and M. Yano. The correction of the Mādhava series for the circumference of a circle. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(2):149– 174, July 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kubbinga:1990:OTC

[741] H. H. Kubbinga. Les origines de la théorie cellulaire; les ≪ molécules organiques≫ de Buffon. (French) [The origins of cellular theory: Buffon’s ‘organic molecules’]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(2):175–213, July 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rechenberg:1990:YHQ

[742] Helmut Rechenberg and Laurie M. Brown. Yukawa’s heavy quantum and the mesotron (1935–1937). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(2):214–252, July 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1990:Rb

[743] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(2):253–292, July 1990. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoyrup:1990:PPA

[744] Jens Høyrup. On parts of parts and ascending continued fractions an investigation of the origins and spread of a peculiar system. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(3):293–324, October 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). Otto Neugebauer in memoriam.

Mancosu:1990:DCE

[745] Paolo Mancosu and Ezio Vailati. Detleff Clüver: An early opponent of the Leibnizian differential calculus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(3):325–344, October 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bierhalter:1990:ZZZ

[746] Günter Bierhalter. Zyklische Zeitvorstellung, Zeitrichtung und die frühen Versuche einer Deduktion des Zweiten Hauptsatzes der Thermodynamik. (German) [cyclical conception of time, the direction of time, and the early trials of a deduction of the Second Law of Thermodynamics]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(3):345–367, October 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Britz:1990:CFH

[747] Dieter Britz. Cold fusion: An historical parallel. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33 (3):368–372, October 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

ZiggelaarSJ:1990:DJD

[748] August Ziggelaar, S.J. Did the Jesuits denounce Galileo?: An essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(3):373–377, October 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1990:Rc

[749] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(3):378–380, October 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1990:BRB

[750] Anonymous. Books received: Book received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 33(3): 381–388, October 1990. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hogendijk:1991:DBP

[751] Jan P. Hogendijk. Desargues’ Brouillon Project and the conics of Apollonius. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(1):1–43, March 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Andersen:1991:DMP

[752] Kirsti Andersen. Desargues’ method of perspective its mathematical content, its connection to other perspective methods and its relation to Desargues’ ideas on projective geometry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(1):44–91, March 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1991:Ra

[753] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(1):92–96, March 1991. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Britton:1991:CBM

[754] John P. Britton and Christopher B. F. Walker. A 4th Century Babylonian model for Venus: B.M. 33552. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(2):97–118, June 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kryzhanovsky:1991:REE

[755] Leonid N. Kryzhanovsky. Richmann’s experiment and the electrophorus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(2):119–124, June 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hutchison:1991:IAL

[756] Keith Hutchison. Idiosyncrasy, achromatic lenses, and early romanticism. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(2):125–171, June 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1991:Rb

[757] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(2):172–192, June 1991. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knorr:1991:PLP

[758] Wilbur R. Knorr. On the principle of linear perspective in Euclid’s optics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(3):193–210, September 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fraser:1991:MTP

[759] Craig G. Fraser. Mathematical technique and physical conception in Euler’s investigation of the elastica. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(3):211–246, September 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Archibald:1991:RTE

[760] Thomas Archibald. Riemann and the theory of electrical phenomena: Nobili’s rings. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(3):247–271, September 1991. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Erlichson:1991:HNW

[761] Herman Erlichson. How Newton went from a mathematical model to a physical model for the problem of a first power resistive force. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(3):272–283, September 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1991:Rc

[762] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(3):284–288, September 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Montserrat:1991:WCL

[763] Jesús M. Montserrat and Luis Navarro. The water cycle in Lucretius. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(4):289–308, December 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Berggren:1991:MIM

[764] J. L. Berggren. Medieval Islamic methods for drawing azimuth circles on the astrolabe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(4):309–344, December 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Erlichson:1991:NFI

[765] Herman Erlichson. Newton’s first inverse solutions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34 (4):345–366, December 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1991:Rd

[766] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 34(4):367–376, December 1991. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Christiansen:1992:DST

[767] H. Dalgas Christiansen. Decanal star tables for lunar houses in Egypt? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(1):1–27, April 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Provencal:1992:OZD

[768] Philippe Provençal. Observations zoologiques de’Abd al-Latîf al-Bag̀dâdî. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(1):28–45, April 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Erlichson:1992:NHC

[769] Herman Erlichson. Newton and Hooke on centripetal force motion. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(1):46–63, April 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1992:Ra

[770] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(1):64–65, April 1992. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Davis:1992:KRI

[771] A. E. L. Davis. Kepler’s resolution of individual planetary motion. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(2):97–102, July 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Davis:1992:KDL

[772] A. E. L. Davis. Kepler’s ‘distance law’ — myth not reality. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(2):103–120, July 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Davis:1992:GEK

[773] A. E. L. Davis. Grading the eggs (Kepler’s sizing-procedure for the planetary orbit). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(2):121–142, July 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Davis:1992:KRD

[774] A. E. L. Davis. Kepler’s road to Damascus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(2):143– 164, July 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Davis:1992:KPF

[775] A. E. L. Davis. Kepler’s physical framework for planetary motion. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(2):165–190, July 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dold-Samplonius:1992:PAM

[776] Yvonne Dold-Samplonius. Practical Arabic mathematics: Measuring the Muqarnas by al-Kāshī. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(3):193–242, October 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Erlichson:1992:NPM

[777] Herman Erlichson. Newton’s polygon model and the second order fallacy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(3):243–258, October 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Seidler:1992:EDP

[778] Jan Seidler and Irena Seidlerová. Zur Entstehungsgeschichte des Dopplerschen Prinzips. (German) [The genesis of Doppler Principle]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(3):259–304, October 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:1992:LLE

[779] Helge Kragh. Ludvig Lorenz and the early theory of long-distance telephony. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(3):305–324, October 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sredniawa:1992:CMS

[780] Bronisław Średniawa. The collaboration of Marian Smoluchowski and Theodor Svedberg on Brownian motion and density fluctuations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(3):325–355, October 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1992:Rb

[781] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 35(3):356–375, October 1992. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lorenzen:1993:PIC

[782] Eivind Lorenzen. The Proto-Ionic capital from Larisa (a brief history of ancient metrology). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(1):1–21, April 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1993:Na

[783] Anonymous. Notes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(1):20–21, April 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Taisbak:1993:THS

[784] Christian Marinus Taisbak. A tale of half sums and differences ancient tricks with numbers. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(1):22–32, April 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Erlichson:1993:GHT

[785] Herman Erlichson. Galileo and high tower experiments. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(1):33–45, April 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1993:Nb

[786] Anonymous. Notes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(1):45, April 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DAgostino:1993:HRT

[787] Salvo D’Agostino. Hertz’s researches and their place in Nineteenth Century theoretical physics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(1):46–77, April 1993. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1993:Nc

[788] Anonymous. Notes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(1):78–82, April 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1993:Ra

[789] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(1):83–96, April 1993. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Christensen:1993:SSI

[790] Dan Ch. Christensen. Spying on scientific instruments. The career of Jesper Bidstrup. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(2):209–244, July 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Darrigol:1993:EMB

[791] Olivier Darrigol. The electrodynamics of moving bodies from Faraday to Hertz. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(3):245–360, October 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1993:C

[792] Anonymous. Corrigendum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(3):361, October 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1993:Rb

[793] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 36(3):362–381, October 1993. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Riley:1994:LVF

[794] Leigh Riley. The lunar velocity function in system B first-crescent ephemerides. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(1):1–51, January 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Behboud:1994:GGA

[795] Ali Behboud. Greek geometrical analysis. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(1):52–86, January 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1994:Ra

[796] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(1):87–92, January 1994. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1994:BRBa

[797] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37 (1):93–96, January 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanDalen:1994:PTE

[798] Benno van Dalen. On Ptolemy’s table for the equation of time. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(2):97–153, April 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Litten:1994:CNM

[799] Freddy Litten. Die Carathéodory-Nachfolge in München 1938–1944. (German) [The Carathéodory-succession in Munich 1938–1944]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(2):154–172, April 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1994:Rb

[800] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(2):173–182, April 1994. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brack-Bernsen:1994:FBC

[801] Lis Brack-Bernsen and Olaf Schmidt. On the foundations of the Babylonian column Φ: Astronomical significance of partial sums of the lunar four. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(3):183–209, July 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Erlichson:1994:HNP

[802] Herman Erlichson. Huygens and Newton on the problem of circular motion. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(3):210–229, July 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bradley:1994:PBB

[803] Margaret Bradley. Prony the bridge-builder: the life and times of Gaspard de Prony, educator and scientist. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(3):230– 268, July 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1994:Rc

[804] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(3):269–278, July 1994. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Erlichson:1994:RIS

[805] Herman Erlichson. Resisted inverse-square centripetal force motion along Newton’s great ‘look-alike’, the equiangular spiral. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(4): 279–303, December 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mayerhofer:1994:PKW

[806] Boris Mayerhofer. Das Prinzip der kleinsten Wirkung bei Hermann von Helmholtz. (German) [The Principle of Least Action by Hermann von Helmholtz]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(4):304–320, December 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wolff:1994:HKG

[807] Stefan L. Wolff. Von der Hydrodynamik zur kinetischen Gastheorie — Oskar Emil Meyer. (German) [On hydrodynamics in the kinetic theory of gases — Oskar Emil Meyer]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(4):321–348, December 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:1994:LE

[808] Helge Kragh. Letters to the Editor. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(4):349–354, December 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1994:Rd

[809] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(4):355–367, December 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1994:BRBb

[810] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 37(4): 368–371, December 1994. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sesiano:1996:KAM

[811] J. Sesiano. Le Kitāb al-Misāh.a d’Abū Kāmil. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38 (1):1–21, January 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dauben:1996:CPE

[812] Joseph W. Dauben. Charles S. Peirce, evolutionary pragmatism and the history of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(1):22–82, January 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1996:Ra

[813] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(1):83–95, January 1996. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vitrac:1996:DVB

[814] B. Vitrac. La définition V.8 des Eléments d’Euclide. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(2-3):97–121, July 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Taisbak:1996:ZEB

[815] Christian Marinus Taisbak. Zeuthen and Euclid’s Data 86: Algebra – or a lemma about intersecting hyperbolas? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(2-3):122– 139, July 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Berggren:1996:AKF

[816] J. L. Berggren. Al-Kūhī’s “filling a lacuna in Book II of Archimedes” in the version of nas.īr al-dīn-T .ūsī. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(2-3):140–207, July 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knorr:1996:WTE

[817] Wilbur R. Knorr. The wrong text of Euclid: on Heiberg’s text and its alternatives. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(2-3):208–276, July 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Caving:1996:DBH

[818] Maurice Caving. The debate between H. G. Zeuthen and H. Vogt (1909– 1915) on the historical source of the knowledge of irrational quantities. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(2-3):277–292, July 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1996:Rb

[819] Anonymous. Review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(2-3):293–294, July 1996. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aaboe:1996:DOR

[820] A. Aaboe and J. L. Berggren. Didactical and other remarks on some theorems of Archimedes and infinitesimals. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(4):295–316, December 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:1996:IAK

[821] Bernard R. Goldstein and José ChabáS. Ibn al-Kammâd’s star list. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(4):317–334, December 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:1996:GGR

[822] Helge Kragh. Gamow’s game: The road to the Hot Big Bang. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(4):335–361, December 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL

http://adsabs.harvard.edu/abs/1996Cent...38..335K;   http:/
/onlinelibrary.wiley.com/doi/10.1111/j.1600-0498.1996.tb00020.
x/abstract.

Anonymous:1996:Rc

[823] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(4):362–376, December 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1996:BRB

[824] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(4): 377–381, December 1996. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:1997:SARa

[825] Alexander Jones. Studies in the astronomy of the Roman period: I. The standard lunar scheme. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(1):1–36, January 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Roero:1997:GBB

[826] Clara Silvia Roero. Giovanni Battista Benedetti and the scientific environment of Turin in the 16th Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(1):37–66, January 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Malet:1997:BWR

[827] Antoni Malet. Barrow, Wallis, and the remaking of Seventeenth Century indivisibles. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(1):67–92, January 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1997:Ra

[828] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(1):93–96, January 1997. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chemla:1997:RWW

[829] Karine Chemla. Reflections on the world-wide history of the Rule of False Double Position, or: How a loop was closed. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(2):97– 120, April 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eckert:1997:MAF

[830] Michael Eckert. Mathematik auf Abwegen: Ferdinand Lindemann und die Elektronentheorie. (German) [Mathematics gone wrong: Ferdinand Lindemann and electron theory]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(2):121–140, April 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dragoni:1997:QMR

[831] Giorgio Dragoni and Giulio Maltese. Quirino Majorana’s research on gravitational absorption: a case study in the misinterpreted experiment tradition. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(2):141–187, April 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1997:Rb

[832] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(2):188–192, April 1997. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anjing:1997:HDA

[833] Qu Anjing. On hypotenuse diagrams in ancient China. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(3):193–210, July 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:1997:SARb

[834] Alexander Jones. Studies in the astronomy of the Roman period: II. tables for solar longitude. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(3):211–229, July 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kouremenos:1997:MRO

[835] Theokritos Kouremenos. Mathematical rigor and the origin of the exhaustion method. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(3):230–252, July 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Erlichson:1997:ENU

[836] Herman Erlichson. Evidence that Newton used the calculus to discover some of the propositions in his Principia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(3):253– 266, July 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DAgostino:1997:HHE

[837] Salvo D’Agostino. Heinrich Hertz and electric waves: An essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(3):267–272, July 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1997:Rc

[838] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(3):273–284, July 1997. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1997:BRB

[839] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(3): 285–289, July 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Roberts:1997:PCI

[840] Gerrylynn K. Roberts. Physical chemists for industry: the making of the chemist at University College London, 1914–1939. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(4): 291–310, December 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nielsen:1997:IDP

[841] Anita Kildebaek Nielsen and Helge Kragh. An institute for dollars: Physical chemistry in Copenhagen between the World Wars. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(4):311–331, December 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

AmorimDaCosta:1997:MPC

[842] A. M. Amorim Da Costa. The mirror of the Portuguese chemical laboratories in the first decades of the Twentieth Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39 (4):332–342, December 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Salkauskas:1997:NRC

[843] Mudis Šalkauskas. A note on the rise of chemistry in the Baltic states. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(4):343–348, December 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brooks:1997:CWR

[844] Nathan M. Brooks. Chemistry in war, revolution, and upheaval: Russia and the Soviet Union, 1900–1929. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(4):349–367, December 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Calascibetta:1997:CCB

[845] Franco Calascibetta. Chemistry in Czechoslovakia between 1919 and 1939: J. Heyrovskýa and the Prague Polarographic School. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 39(4):368–381, December 1997. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:1998:SAR

[846] Alexander Jones. Studies in the astronomy of the Roman period III. Planetary epoch tables. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(1):1–41, January 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sichau:1998:NES

[847] C. Sichau. Ein nationales Experiment und seine Auswirkungen auf einen wissenschaftlichen Versuch: Die Einführung des Government Grant und die Joule–Thomson-Experimente von C. Sichau. (German) [A national experiment and its impact on a scientific experiment: The introduction of the Grant Government and the Joule–Thomson experiments of c. sichau]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(1):42–80, January 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Roll-hansen:1998:PUH

[848] Nils Roll-hansen. Pasteur: an underestimated hero of science: An essay review. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(1):81–93, January 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1998:Ra

[849] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(1):94–96, January 1998. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brunet:1998:FCQ

[850] J.-P. Brunet, R. Nadal, and Cl. Vibert-Guigue. The fresco of the cupola of Qusayr^{c} Amra. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(2):97–123, April 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chabaas:1998:EPA

[851] Joséa Chabáas and Antoni Roca. Early printing of astronomy: the Lunari of Bernat de Granollachs. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(2):124–134, April 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lauridsen:1998:HAM

[852] Emil Kring Lauridsen and Niels Abrahamsen. The history of astatic magnet systems and suspensions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(2):135–169, April 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1998:Rb

[853] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(2):170–176, April 1998. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mendell:1998:REC

[854] Henry Mendell. Reflections on Eudoxus, Callippus and their curves: Hippopedes and Callippopedes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(3–4):177–275, October 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sesiano:1998:EFG

[855] J. Sesiano. An early form of Greek algebra. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(3–4):276– 302, October 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Guicciardini:1998:DNU

[856] Niccoló Guicciardini. Did Newton use his calculus in the Principia? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(3–4):303–344, October 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1998:Rc

[857] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 40(3–4):345–384, October 1998. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1999:ENE

[858] Anonymous. Editorial note: Editorial note. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(1–2):1– 2, April 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vleuten:1999:CCE

[859] Erik V. D. Vleuten. Constructing centralized electricity supply in Denmark and The Netherlands: An actor group perspective. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(1–2):3–36, April 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lagaaij:1999:DVP

[860] J. A. C. Lagaaij and G. P. J. Verbong. Different visions of power the introduction of nuclear power in The Netherlands 1955–1970. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(1–2):37–63, April 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nielsen:1999:RAI

[861] Henry Nielsen, Keld Nielsen, Flemming Petersen, and Hans Siggaard. Risø and the attempts to introduce nuclear power into Denmark. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(1–2):64–92, April 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rudiger:1999:NGC

[862] Mogens Rüdiger. The natural gas controversy the introduction of natural gas in Denmark, 1972–84. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(1–2):93–111, April 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heymann:1999:FSW

[863] Matthias Heymann. A fight of systems? Wind power and electric power systems in Denmark, Germany, and the USA. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41 (1–2):112–136, April 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Verbong:1999:WPN

[864] G. P. J. Verbong. Wind power in The Netherlands, 1970–1995. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(1–2):137–160, April 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nielsen:1999:IWP

[865] Kristian Hvidtfelt Nielsen. Interpreting wind power vs. the electric power system: a Danish case-study. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(1–2):161–177, April 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1999:BRBa

[866] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41 (1–2):178–181, April 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wolff:1999:LAP

[867] Stefan L. Wolff. Leo Arons — Physiker und Sozialist. (German) [Leo Arons — physicist and socialist]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(3):183–212, July 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VanBrummelen:1999:ASM

[868] Glen Van Brummelen. The astronomical system in Mūsā ibn Nawbakht’s astrological treatise, the Kitāb al-Kāmil. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(3):213– 243, July 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1999:Ra

[869] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(3):244–250, July 1999. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1999:BRBb

[870] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(3): 251, July 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Berger:1999:GZP

[871] Jutta Berger. Grenzgänge zwischen Physik und Chemie: Thermodynamik und chemische Kinetik — kein Happy-End im 19. Jahrhundert. (German) [Boundaries between physics and chemistry: thermodynamics and chemical kinetics — no happy ending in the 19th Century]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(4):253–279, October 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brack-Bernsen:1999:BZS

[872] Lis Brack-Bernsen and Hermann Hunger. The Babylonian zodiac: Speculations on its invention and significance. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(4):280–292, October 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Taisbak:1999:SSA

[873] Christian Marinus Taisbak. Splitting a square analysis of Euclid’s elements xiii.10. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(4):293–295, October 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:1999:Rb

[874] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 41(4):296–318, October 1999. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Montserrat:2000:AVH

[875] Jesús M. Montserrat and Luis Navarro. The atomistic view of heat in Lucretius. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(1):1–20, January 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoyrup:2000:JFB

[876] Jens Høyrup. Jacobus de Florentia, Tractatus algorismi (1307), the chapter on algebra (Vat. Lat. 4826, fols 36^{v} –45^{v}). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(1): 21–69, January 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2000:Ra

[877] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(1):70–74, January 2000. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2000:BRB

[878] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(1): 75, January 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:2000:SAR

[879] Alexander Jones. Studies in the astronomy of the Roman Period IV solar tables based on a non-Hipparchian model. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(2):77– 88, April 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Steele:2000:RAE

[880] John M. Steele. A re-analysis of the eclipse observations in Ptolemy’s Almagest. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(2):89–108, April 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mertens:2000:DDB

[881] Joost Mertens. The development of the dry battery: Prelude to a mass consumption article (1882–1908). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(2):109–134, April 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tonietti:2000:DNM

[882] Tito M. Tonietti. Does Newton’s musical model of gravitation work?: a mistake and its meaning. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(2):135–149, April 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2000:RER

[883] Anonymous. Review essay: Review essay. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(2):150– 153, April 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2000:Rb

[884] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(2):154–157, April 2000. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VonSpaeth:2000:DOE

[885] Ove Von Spaeth. Dating the oldest Egyptian star map. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(3):159–179, July 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Michel-Nozieres:2000:SMB

[886] C. Michel-Nozières. Second Millennium Babylonian water clocks: a physical study. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(3):180–209, July 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fermor:2000:DBW

[887] J. Fermor and J. M. Steele. The design of Babylonian waterclocks: Astronomical and experimental evidence. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(3):210–222, July 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Huber:2000:BST

[888] Peter J. Huber. Babylonian short-time measurements: Lunar sixes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(3):223–234, July 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2000:RAN

[889] Anonymous. Review article: National histories of science, the Dutch case. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(3):235–239, July 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2000:RAR

[890] Anonymous. Review article: Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(3):240– 247, July 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2000:BR

[891] Anonymous. Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(3):248, July 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Frercks:2000:CTE

[892] Jan Frercks. Creativity and technology in experimentation: Fizeau’s terrestrial determination of the speed of light. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(4):249–287, December 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Topper:2000:KWS

[893] David Topper. “I know that what I am saying is rather obscure … ”: on clarifying a passage in Galileo’s dialogue. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(4):288–296, December 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Saraiva:2000:SPM

[894] Luis M. Ribeiro Saraiva. A survey of Portuguese mathematics in the Nineteenth Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 42(4):297–318, December 2000. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2001:BQC

[895] Helge Kragh. Before quantum chemistry: Erich Hückel and the physicschemistry interface. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(1):1–16, January 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brentjes:2001:TCE

[896] Sonja Brentjes. Two comments on Euclid’s Elements? On the relation between the Arabic text attributed to al-Nayrızı and the Latin text ascribed to Anaritius. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(1):17–55, January 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Ra

[897] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(1):56–67, January 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BR

[898] Anonymous. Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(1):68–69, January 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Day:2001:ONS

[899] Michael A. Day. Oppenheimer on the nature of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):73–112, April 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRRa

[900] Anonymous. Book review: Review essay: Chinese mathematics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):113–121, April 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rb

[901] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rc

[902] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rd

[903] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Re

[904] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rf

[905] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rg

[906] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rh

[907] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Ri

[908] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rj

[909] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rk

[910] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rl

[911] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rm

[912] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Rn

[913] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:Ro

[914] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):122–135, April 2001. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Spalt:2001:CSN

[915] Detlef D. Spalt. Curt Schmieden’s non-standard analysis — a method of dissolving the standard paradoxes of analysis. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):137–175, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VonKoch:2001:MBR

[916] Johannes Von Koch. Die Mondeklipsen von BM 41004 rev. 18. (German) [The lunar eclipse of BM 41004 rev. 18]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):176– 183, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jacobsen:2001:SUO

[917] Anja Skaar Jacobsen. Spirit and unity: Ørsted’s fascination by Winterl’s chemistry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):184–218, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Darrigol:2001:HDM

[918] Olivier Darrigol. The historians’ disagreements over the meaning of Planck’s quantum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):219–239, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoffmann:2001:ECP

[919] Dieter Hoffmann. On the experimental context of Planck’s foundation of quantum theory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):240–259, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Singh:2001:MBR

[920] Rajinder Singh. Max Born’s role in the lattice dynamic controversy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):260–277, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gislen:2001:SEA

[921] Lars Gislén and J. C. Eade. South East Asian eclipse calculations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):278–307, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRRb

[922] Anonymous. Book review: Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):308–310, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBa

[923] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBb

[924] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBc

[925] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBd

[926] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBe

[927] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBf

[928] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBg

[929] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBh

[930] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBi

[931] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBj

[932] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBk

[933] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBl

[934] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBm

[935] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBn

[936] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBo

[937] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBp

[938] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBq

[939] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBr

[940] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBs

[941] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBt

[942] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBu

[943] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBv

[944] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBw

[945] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBx

[946] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBy

[947] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBz

[948] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBaa

[949] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBab

[950] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBac

[951] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBad

[952] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBae

[953] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBaf

[954] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBag

[955] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBah

[956] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBai

[957] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBaj

[958] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBak

[959] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBal

[960] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBam

[961] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBan

[962] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2001:BRBao

[963] Anonymous. Book review: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):311– 313, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoyrup:2002:ESC

[964] Jens Høyrup. Existence, substantiality, and counterfactuality observations on the status of mathematics according to Aristotle, Euclid, and others. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):1–31, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2002:VAV

[965] Helge Kragh. The vortex atom: a Victorian theory of everything. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):32–114, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knudsen:2002:ACS

[966] Toke Lindegaard Knudsen. On altar constructions with square bricks in ancient Indian ritual. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):115–126, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Konno:2002:RDL

[967] Hiroyuki Konno. Ritz’s discovery of the Lyman series before 1913 and Lyman’s indifference to the Bohr theory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):127– 139, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Leitz:2002:RAA

[968] Christian Leitz. Remarks about the appearance of Mars in the Tomb of Senenmut in Western Thebes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):140–142, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Ra

[969] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):143–144, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Rb

[970] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):144–145, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Rc

[971] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):145–146, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Rd

[972] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):146–147, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Re

[973] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):147–148, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Rf

[974] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):148–149, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Rg

[975] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):149–150, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Rh

[976] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):150–151, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Ri

[977] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):151–153, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Rj

[978] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):153–154, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Rk

[979] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):154–155, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:BR

[980] Anonymous. Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):156–157, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nakane:2002:EHH

[981] Michiyo Nakane and Craig G. Fraser. The early history of Hamilton– Jacobi dynamics 1834–1837. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(3–4):161–227, December 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Christianson:2002:LTB

[982] J. R. Christianson. The legacy of Tycho Brahe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(3–4): 228–247, December 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kjaergaard:2002:CAP

[983] Peter C. Kjærgaard. Competing allies: Professionalisation and the hierarchy of science in Victorian Britain. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(3–4):248–288, December 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:Rl

[984] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(3–4):289–310, December 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2002:BRB

[985] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44 (3–4):311, December 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2003:EEI

[986] Anonymous. Editorial: Editors’ introduction. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45 (1-4):1–3, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2003:PBR

[987] Anonymous. Publications of Bernard R. Goldstein. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45 (1-4):4–15, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brack-Bernsen:2003:PMR

[988] Lis Brack-Bernsen. The path of the moon, the rising points of the sun, and the oblique great circle on the celestial sphere. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45 (1-4):16–31, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rochberg:2003:LDB

[989] Francesca Rochberg. Lunar data in Babylonian horoscopes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):32–45, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Britton:2003:CSA

[990] John P. Britton. On corrections for solar anomaly in Babylonian lunar theories. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):46–58, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bowen:2003:CME

[991] Alan C. Bowen. Cleomedes and the measurement of the Earth: a question of procedures. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):59–68, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:2003:PBS

[992] Alexander Jones. A posy of Almagest scholia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45 (1-4):69–78, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pingree:2003:GEW

[993] David Pingree. A Greek ephemeris for 796: the work of Stephanus the Philosopher? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):79–82, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tihon:2003:TGV

[994] Anne Tihon. Une table grecque de vitesse lunaire (Laurentianus 28/26). (French) [A Greek table of lunar speed (Laurentianus 28/26)]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):83–99, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Smith:2003:PAI

[995] A. Mark Smith. Ptolemy, Alhacen, and Ibn Mu’adh and the problem of atmospheric refraction. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):100–115, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Berggren:2003:ASV

[996] J. L. Berggren and Glen Van Brummelen. Al-Samaw’al versus al-Kūh? on the depression of the horizon. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):116–129, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

North:2003:W

[997] John North. Winchester 1067. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):130–141, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chabas:2003:WAT

[998] José Chabás. Were the Alfonsine Tables of Toledo first used by their authors? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):142–150, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jacquart:2003:BGA

[999] Danielle Jacquart. Bernard de Gordon et l’astrologie. (French) [Bernard de Gordon and astrology]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):151–158, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kunitzsch:2003:STM

[1000] Paul Kunitzsch and Y. Tzvi Langermann. A star table from Medieval Yemen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):159–174, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Shank:2003:RFH

[1001] Michael H. Shank. Rings in a fluid heaven: the equatorium-driven physical astronomy of Guido de Marchia (fl. 1292–1310). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):175–203, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

King:2003:CEC

[1002] David A. King. 14th-Century England or 9th-Century Baghdad? New insights on the elusive astronomical instrument called Navicula de Venetiis. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):204–226, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Freudenthal:2003:IRC

[1003] Gad Freudenthal. ‘Instrumentalism’ and ‘realism’ as categories in the history of astronomy: Duhem vs. Popper, Maimonides vs. Gersonides. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):227–248, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Glasner:2003:GUP

[1004] Ruth Glasner. Gersonides unusual position on “position”. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):249–263, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mancha:2003:RAH

[1005] J. L. Mancha. Right ascensions and Hippopedes: Homocentric models in Levi ben Gerson’s astronomy I. first anomaly. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4): 264–283, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Levy:2003:IBX

[1006] Tony Lévy. Immanuel Bonfils (XIVe s.): fractions décimales, puissances de 10 et opérations arithmétiques. (French) [Immanuel Bonfils (14th Century): decimal fractions, powers of 10, and arithmetic operations]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):284–304, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kremer:2003:WFT

[1007] Richard L. Kremer. Wenzel Faber’s table for finding true syzygy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 45(1-4):305–329, December 2003. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Swerdlow:2004:TBE

[1008] N. M. Swerdlow. Tycho Brahe’s early lunar theory and the lunar eclipse of 31 January 1599. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(1):1–40, February 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Barker:2004:HRC

[1009] Peter Barker. How Rothmann changed his mind. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46 (1):41–57, February 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hon:2004:PEH

[1010] Giora Hon. Putting error to (historical) work: Error as a tell-tale in the studies of Kepler and Galileo. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(1):58–81, February 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Samso:2004:AZJ

[1011] Julio Samsó. Abraham Zacut and José Vizinho’s Almanach Perpetuum in Arabic (16th–19th C.). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(1):82–97, February 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schlote:2004:CNF

[1012] Karl-Heinz Schlote. Carl Neumanns Forschungen zur Potentialtheorie. (German) [Carl Neumann’s research on potential theory]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(2):99–132, April 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gislen:2004:AAI

[1013] Lars Gislén. Analysis of the astronomical information in Tachard: “Voyage to Siam in the Year 1685”. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(2):133–144, April 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hollywood:2004:ARB

[1014] Louise Hollywood and John M. Steele. Acronycal risings in Babylonian astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(2):145–162, April 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2004:BRB

[1015] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46 (2):163–188, April 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Young:2004:AIH

[1016] C. K. Young. Archimedes’s iron hand or claw – a new interpretation of an old mystery. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(3):189–207, August 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schonbeck:2004:TCQ

[1017] J. Schönbeck. Thomas Clausen und die quadrierbaren Kreisbogenzweiecke. (German) [Thomas Clausen and the squarable two-cornered arc]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(3):208–229, August 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2004:Ra

[1018] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(3):230–241, August 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Netz:2004:ERH

[1019] Reviel Netz. Eudemus of Rhodes, Hippocrates of Chios and the earliest form of a Greek mathematical text. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(4):243–286, December 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Stamhuis:2004:SVW

[1020] Ida H. Stamhuis and Paul M. M. Klep. The stubbornness of various ways of knowledge was not typically Dutch; the statistical mind in a pre-statistical era. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(4):287–317, December 2004. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2004:CSM

[1021] Anonymous. Comment on: The Statistical Mind in a Pre-Statistical Era: The Netherlands 1750–1850, edited by Paul M. M. Ckep and Ida H. Stahuis delivered at The Free University of Amsterdam on September 4, 2002 by Theodore M. Porter, University of California in Los Angeles. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(4):318–330, December 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2004:Rb

[1022] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 46(4):331–332, December 2004. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brack-Bernsen:2005:DEM

[1023] Lis Brack-Bernsen. The “days in excess” from MUL.APIN on the “first intercalation” and “water clock” schemes from MUL.APIN. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(1):1–29, February 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Maas:2005:IIA

[1024] Ad Maas. Institutionalised individualism Amsterdam physics between the World Wars. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(1):30–59, February 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Delire:2005:QCA

[1025] Jean Michel Delire. Quadratures, circulature and the approximation of \sqrt{2} in the Indian Śulba-sútras. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(1):60–71, February 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2005:R

[1026] Anonymous. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(1):72–85, February 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hogendijk:2005:AME

[1027] Jan P. Hogendijk. Applied mathematics in Eleventh Century Al-Andalus: Ibn Mu^{c}ādh al-Jayyān? and his computation of astrological houses and aspects. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(2):87–114, May 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VonHeering:2005:PZW

[1028] Peter Von Heering and Daniel Osewold. Ein Problem, zwei Wissenschaftler, drei Instrumente. (German) [One problem, two scientists, three instruments]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(2):115–139, May 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Deltete:2005:BLE

[1029] Robert J. Deltete. Die Lehre von der Energie. (German) [The Theory of Energy]: Georg Helm’s energetic manifesto. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(2):140– 162, May 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Duke:2005:HET

[1030] Dennis W. Duke. Hipparchus’ eclipse trios and early trigonometry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(2):163–177, May 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Macfarlane:2005:BCS

[1031] Roger T. Macfarlane and Paul S. Mills. Bright and conspicuous stars in Ptolemy and Hipparchus: on the mistranslation of ϵκϕανης. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(2):178–180, May 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brack-Bernsen:2005:EPL

[1032] Lis Brack-Bernsen and John M. Steele. Eclipse prediction and the length of the Saros in Babylonian astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(3):181–206, September 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fangerau:2005:NSJ

[1033] Heiner Fangerau and Irmgard Müller. National styles? Jacques Loeb’s analysis of German and American science around 1900 in his correspondence with Ernst Mach. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(3):207–225, September 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:2005:PPM

[1034] Alexander Jones and Dennis Duke. Ptolemy’s planetary mean motions revisited. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(3):226–235, September 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sidoli:2005:HDA

[1035] Nathan Sidoli. Heron’s dioptra 35 and analemma methods: An astronomical determination of the distance between two cities. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(3):236–258, September 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Navarro:2005:JJT

[1036] Jaume Navarro. J. J. Thomson on the nature of matter: Corpuscles and the continuum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(4):259–282, October 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hon:2005:KVL

[1037] Giora Hon. Kant vs. Legendre on symmetry: Mirror images in philosophy and mathematics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(4):283–297, October 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Yuste:2005:AGO

[1038] Piedad Yuste. Algebra and geometry in the old Babylonian period: Matters concerning reeds. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(4):298–315, October 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fragoulis:2005:OAG

[1039] D. Fragoulis, A. Skembris, C. Papaodysseus, P. Rousopoulos, Th. Panagopoulos, M. Panagopoulos, C. Triantafyllou, A. Vlachopoulos, and C. Doumas. Origins and application of geometry in the Thera prehistoric civilization ca. 1650 BC. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(4):316–340, October 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2005:R

[1040] Helge Kragh. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(4):341–343, October 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2005:I

[1041] Anonymous. Index. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 47(4):344–346, October 2005. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Barker:2006:NWE

[1042] Peter Barker. New work in early modern science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(1):1–2, January 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Waddell:2006:WIM

[1043] Mark A. Waddell. The world, as it might be: Iconography and probabilism in the Mundus subterraneus of Athanasius Kircher. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(1):3–22, January 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tredwell:2006:MCE

[1044] Katherine A. Tredwell. The Melanchthon Circle’s English epicycle. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(1):23–31, January 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Boner:2006:KLC

[1045] Patrick J. Boner. Kepler’s living cosmology: Bridging the celestial and terrestrial realms. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(1):32–39, January 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Friesen:2006:HNE

[1046] John Friesen. Hutchinsonianism and the Newtonian Enlightenment. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(1):40–49, January 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heidarzadeh:2006:RNT

[1047] Tofigh Heidarzadeh. The reception of Newton’s theory of cometary tail formation. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(1):50–65, January 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sorensen:2006:R

[1048] Henrik Kragh Sørensen and Charlotte Methuen. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(1):66–68, January 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Baldasso:2006:RVR

[1049] Renzo Baldasso. The role of visual representation in the scientific revolution: a historiographic inquiry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(2):69–88, April 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hon:2006:AVE

[1050] Giora Hon and Bernard R. Goldstein. Adding velocities without exceeding the velocity of light: Wilhelm Wien’s algorithm (1904) and Albert Einstein’s light postulate (1905). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(2):89–113, April 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moesgaard:2006:R

[1051] Kristian Peder Moesgaard, Peder Anker, Mark Clark, Marta Gómez Martínez, Anja Skaar Jacobsen, Jole Shackelford, Michael Eckert, Jacqueline Stedall, Kristian Hvidtfelt Nielsen, and Helge Kragh. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(2):114–129, April 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2006:BRB

[1052] Anonymous. Books received: Books received. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48 (2):130–131, April 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Forsdyke:2006:HTI

[1053] Donald R. Forsdyke. Heredity as transmission of information: Butlerian ‘intelligent design’. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(3):133–148, July 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tonietti:2006:MCF

[1054] Tito M. Tonietti. The mathematical contributions of Francesco Maurolico to the theory of music of the 16th Century (the problems of a manuscript). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(3):149–200, July 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sorensen:2006:LOM

[1055] Henrik Kragh Sørensen. Louis Olivier: a mathematician only known through his publications in Crelle’s Journal during the 1820s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(3):201–231, July 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Andersen:2006:E

[1056] Hanne Andersen. Editorial. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(4):233, October 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Magruder:2006:GVE

[1057] Kerry V. Magruder. Global visions and the establishment of theories of the Earth. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(4):234–257, October 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Taylor:2006:MGP

[1058] Kenneth L. Taylor. Marivetz, Goussier, and Planet Earth: a Late Enlightenment geo-physical project. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(4):258–283, October 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Klemun:2006:NSG

[1059] Marianne Klemun. Natural science and geology as a medium of integration: the Versammlung deutscher Naturforscher und Ärzte in Prague in 1837 and the meetings of German natural scientists and physicians during the ‘Vormärz’ (1822–1848). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(4):284–297, October 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Grapes:2006:AMD

[1060] Rodney Grapes. Alexander McKay and the discovery of lateral displacement on faults in New Zealand. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(4):298–313, October 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kjaergaard:2006:R

[1061] Peter C. Kjærgaard, Rajinder Singh, Fokko Jan Dijksterhuis, Micah Ross, Donald B. Wagner, Paolo Palmieri, Helge Kragh, Bruce J. Hunt, Christian Forstner, Marie Louise Thomsen, Adrian Rice, Jens Høyrup, Henrik Knudsen, Peter C. Kjærgaard, Henry Nielsen, Giora Hon, and Helge Kragh. Reviews. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48(4):314–336, October 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2006:AII

[1062] Anonymous. Author index: Index for volume 48. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 48 (4):337–339, October 2006. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Andersen:2007:ECO

[1063] Hanne Andersen. Editorial: Centaurus — the official journal of the European society for the history of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(1):1–2, February 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knobloch:2007:AHE

[1064] Eberhard Knobloch. Alexander von Humboldt — the explorer and the scientist. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(1):3–14, February 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brusch:2007:PMP

[1065] Björn Brüsch. Provisions made for prosperity and affluence: Karl Sigmund Franz Freiherr von Stein zum Altenstein and the establishment of the Gärtnerlehranstalt in Prussia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(1):15–55, February 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rebsdorf:2007:BSI

[1066] Simon O. Rebsdorf. Bengt Strömgren: Interstellar glow, helium content, and solar life supply, 1932–1940. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(1):56–79, February 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Frercks:2007:ID

[1067] Jan Frercks. Immaterial devices. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(2):81–113, May 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DeBont:2007:STC

[1068] Raf De Bont. A serpent without teeth. The conservative transformism of Jean-Baptiste d’Omalius d’Halloy (1783–1875). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(2):114–137, May 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aspaas:2007:GNP

[1069] Per Pippin Aspaas and Truls Lynne Hansen. Geomagnetism by the North Pole, anno 1769: the magnetic observations of Maximilian Hell during his Venus Transit Expedition. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(2):138–164, May 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2007:EHS

[1070] Anonymous. ESHS: History of Science in Education and Training in Europe: What new prospects? Round table discussion at the Second European Society for the History of Science Conference in Cracow on 9 September 2006. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(2):165–171, May 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Berggren:2007:AHA

[1071] Len Berggren. Asger Hartvig Aaboe: 26 April 1922 to 19 January 2007. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(2):172–177, May 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Arabatzis:2007:RBE

[1072] Theodore Arabatzis. Reviews: Growing Explanations. Historical Perspectives on Recent Science — Edited by M. Norton Wise. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(2):178–179, May 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sorensen:2007:RBU

[1073] Henrik Kragh Sørensen. Reviews: Mathematics Unbound: the Evolution of an International Mathematical Research Community, 1800–1945 — Edited by Karen H. Parshall and Adrian C. Rice. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(2):179–181, May 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2007:RBR

[1074] Helge Kragh. Reviews: Science and Religion, 400 BC–AD 1550: From Aristotle to Copernicus — by Edward Grant and Science and Religion, 1450–1900: From Copernicus to Darwin — by Richard G. Olson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(2):181–182, May 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kruse:2007:RBP

[1075] Poul R. Kruse. Reviews: Geschichte der Pharmazie. Vol. II: Von der Frühen Neuzeit bis zur Gegenwart. (German) [History of Pharmacy. Vol II: From the Early Modern Period to the Present] — by Rudolf Schmitz. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(2):182–183, May 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hayton:2007:MBC

[1076] Darin Hayton. Martin Bylica at the Court of Matthias Corvinus: Astrology and politics in Renaissance Hungary. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):185–198, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nielsen:2007:FDD

[1077] Anita Kildebæk Nielsen. Fashioning and demarcation of the Danish chemical community in the 19th century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):199–226, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VanDelft:2007:FLC

[1078] Dirk VanDelft. Facilitating Leiden’s cold: the International Association of Refrigeration and the internationalisation of Heike Kamerlingh Onnes’s Cryogenic Laboratory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):227–245, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2007:RBH

[1079] Helge Kragh. Reviews: From Here to Eternity. Ernst Haeckel and Scientific Faith — by Mario A. Di Gregorio. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):246–247, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Simoes:2007:RBR

[1080] Ana Simões. Reviews: Shifting and Rearranging: Physical Methods and the Transformation of Modern Chemistry — by Carsten Reinhardt. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):247–248, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jacobsen:2007:RBM

[1081] Anja Skaar Jacobsen. Reviews: Chemistry, Medicine, and Crime: Mateu J. B. Orfila (1787–1853) and His Times — Edited by José Ramón Bertomeu–Sánchez; Agustí Nieto-Galan. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):249, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knudsen:2007:RBS

[1082] Henrik Knudsen. Reviews: The Science-Industry Nexus: History, Policy, Implications — Edited by Karl Grandinl; Nina Wormbs; Sven Widmalm. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):250–251, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2007:RBB

[1083] Helge Kragh. Reviews: The Beginnings of Piezoelectricity. A Study in Mundane Physics — by Shaul Katzir. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):251–252, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lykknes:2007:RBS

[1084] Annette Lykknes. Reviews: Perspectives on Scandinavian Science in the Early Twentieth Century — Edited by Reinhard Siegmund-Schultze; H. Kragh Sørensen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):253–254, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Shapiro:2007:RBO

[1085] Alan E. Shapiro. Reviews: Thinking with Objects: the Transformation of Mechanics in the Seventeenth Century — by Domenico Bertoloni Meli.

Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):254–255, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoyrup:2007:RBE

[1086] Jens Høyrup. Reviews: Essays on Early Medieval Mathematics: the Latin Tradition — by Menso Folkerts. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(3):255–256, August 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Stamhuis:2007:ECH

[1087] Ida H. Stamhuis and Hanne. Andersen. Editorial: the control of a healthy society: Institutionalizing statistics in the 19^{th} century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(4):257, November 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Porter:2007:ISO

[1088] Theodore M. Porter. Introduction: the statistical office as a social observatory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(4):258–260, November 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wolfenstein:2007:RNG

[1089] Gabriel K. Wolfenstein. Recounting the nation: the general register office and Victorian bureaucracies. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(4):261–288, November 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Labbe:2007:ISN

[1090] Morgane Labbé. Institutionalizing the statistics of nationality in Prussia in the 19th Century (from local bureaucracy to state-level census of population). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(4):289–306, November 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Stamhuis:2007:NCS

[1091] Ida H. Stamhuis. A Nineteenth-Century statistical society that abandoned statistics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(4):307–336, November 2007. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rusnock:2007:MSH

[1092] Andrea Rusnock. Medical statistics and hospital medicine: the case of the smallpox vaccination. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 49(4):337–359, November 2007. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Andersen:2008:ECV

[1093] Hanne Andersen. Editorial: Centaurus at volume 50: Looking into the archives. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):1–3, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lilley:2008:ACE

[1094] S. Lilley. From the archives: Cause & effect in the history of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):4–18, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2008:CLT

[1095] Helge Kragh. Commentary 01 on Lilley 1953 and Truesdell 1970. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):19–20, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Truesdell:2008:ASW

[1096] C Truesdell. From the archives: the Scholar’s workshop and tools. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):21–30, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:2008:CCL

[1097] Kurt Møller Pedersen. Commentary: Commentary 02 on Lilley 1953 and Truesdell 1970. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):31, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Arabatzis:2008:CCL

[1098] Theodore Arabatzis. Commentary: Commentary 03 on Lilley 1953 and Truesdell 1973. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):32–36, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chang:2008:CCL

[1099] Hasok Chang. Commentary: Commentary 04 on Lilley 1953 and Truesdell 1973. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):37–42, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schickore:2008:CCL

[1100] Jutta Schickore. Commentary: Commentary 05 on Lilley 1953 and Truesdell 1973. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):43–45, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rosen:2008:AFC

[1101] George Rosen. From the archives: the fate of the concept of medical police 1780–1890. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):46–62, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hamlin:2008:CCR

[1102] Christopher Hamlin. Commentary: Commentary 01 on Rosen 1957. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):63–69, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knipper:2008:CCR

[1103] Michael Knipper. Commentary: Commentary 02 on Rosen 1957. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):70–72, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bernal:2008:ASI

[1104] J. D. Bernal. From the archives: Science, industry and society in the Nineteenth Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):73–100, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cowan:2008:CCB

[1105] Ruth Schwartz Cowan. Commentary: Commentary 01 on Bernal 1953. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):101–102, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Haard:2008:CCB

[1106] Mikael Hård. Commentary: Commentary 02 on Bernal 1953. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):103–104, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Galison:2008:ADM

[1107] Peter Galison. From the archives: the discovery of the muon and the failed revolution against quantum electrodynamics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):105–159, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Galison:2008:CCG

[1108] Peter Galison. Commentary: Commentary 01 on Galison 1982. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):160–161, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Franklin:2008:CCG

[1109] Allan Franklin. Commentary: Commentary 02 on Galison 1982. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):162–165, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kaiser:2008:CCG

[1110] David Kaiser. Commentary: Commentary 03 on Galison 1983. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):166–167, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:2008:ASM

[1111] Bernard R. Goldstein. From the archives: the status of models in Ancient and Medieval astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):168–183, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Goldstein:2008:CCG

[1112] Bernard R. Goldstein. Commentary: Commentary 01 on Goldstein 1980. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):184–188, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Barker:2008:CCG

[1113] Peter Barker. Commentary: Commentary 02 on Goldstein 1980. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):189–194, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bowen:2008:ACM

[1114] Alan C. Bowen. From the archives: Cleomedes and the measurement of the Earth: a question of procedures. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):195– 204, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knobloch:2008:CB

[1115] Eberhard Knobloch. Commentary on Bowen 2003. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50 (1–2):205, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2008:NCN

[1116] Anonymous. Notes on contributors: Notes on contributors. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(1–2):206–209, February 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Steinle:2008:IEF

[1117] Friedrich Steinle and Gregor Schiemann. Introduction: the empirical and the formal — tensions in scientific knowledge. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50 (3):211–213, August 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Darrigol:2008:ECC

[1118] Olivier Darrigol. Empirical challenges and concept formation in the history of hydrodynamics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(3):214–232, August 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eckert:2008:TWT

[1119] Michael Eckert. Theory from wind tunnels: Empirical roots of Twentieth Century fluid dynamics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(3):233–253, August 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jonkers:2008:PMS

[1120] Art. R. T. Jonkers. The pursuit of magnetic shadows: the formalempirical dipole field of Early-Modern geomagnetism. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(3):254–289, August 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Good:2008:BDM

[1121] Gregory A. Good. Between data, mathematical analysis and physical theory: Research on Earth’s magnetism in the 19th century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(3):290–304, August 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hon:2008:WKE

[1122] Giora Hon and Bernard R. Goldstein. What keeps the Earth in its place? The concept of stability in Plato and Aristotle. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):305–323, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Boner:2008:KEA

[1123] Patrick J. Boner. Kepler’s early astrological calendars: Matter, methodology and multidisciplinarity. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):324–328, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hentschel:2008:RBL

[1124] Klaus Hentschel. Reviews: Objectivity — by Lorraine Daston and Peter Galison. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):329–330, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Klein:2008:RBA

[1125] Ursula Klein. Reviews: The artificial and the natural: an evolving polarity — Edited by Bernadette Bensaude-Vincent and William R. Newman. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):330–332, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Davis:2008:RBG

[1126] Philip J. Davis. Reviews: The Geometry of An Art: the History of the Mathematical Theory of Perspective from Alberti to Monge — by Kirsti Andersen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):332–334, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dijksterhuis:2008:RBG

[1127] Fokko Jan Dijksterhuis. Reviews: The Good Life in the Scientific Revolution. Descartes, Pascal, Leibniz, and the Cultivation of Virtue — by Matthew Jones. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):334–335, November 2008. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hon:2008:RBA

[1128] Giora Hon. Reviews: The Art of Conjecturing, Together with Letter to a Friend on Sets in Court Tennis — by Jacob Bernoulli (Trans. with an introduction and notes by Edith Dudley Sylla). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4): 335–337, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hjermitslev:2008:RBS

[1129] Hans Henrik Hjermitslev. Reviews: Herbert Spencer and the invention of modern life — by Mark Francis. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):338–339, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nielsen:2008:RBP

[1130] Anita Kildebæk Nielsen. Reviews: The Periodic Table: Its Story and Its Significance — by Eric R. Scerri. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):339–341, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Reiter:2008:RBI

[1131] Wolfgang L. Reiter. Reviews: Aufrecht im Sturm der Zeit: Der Physiker James Franck 1882–1964. (German) [Upright in the storm of time: the physicist James Franck 1882–1964] — by Jost Lemmerich. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):341–342, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aaserud:2008:RBM

[1132] Finn Aaserud. Reviews: The Mental Aftermath: the Mentality of German Physicists 1945–1949 — by Klaus Hentschel. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4): 342–343, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Forstner:2008:RBZ

[1133] Christian Forstner. Reviews: Physiker zwischen Autonomie und Anpassung. Die Deutsche Physikalische Gesellschaft im Dritten Reich — Edited by Dieter Hoffmann and Mark Walker. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4): 344–345, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2008:RBM

[1134] Helge Kragh. Reviews: Particle Metaphysics: a Critical Account of Subatomic Reality — by Brigitte Falkenburg. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):345, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knopp:2008:RBM

[1135] Matthias Knopp. Reviews: Rockets and missiles: the life story of a technology — by A. Bodoing van Riper. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):346–347, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heymann:2008:RBF

[1136] Matthias Heymann. Reviews: Ingenium: five machines that changed the world — by Mark Denny. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):347, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heymann:2008:RBE

[1137] Matthias Heymann. Reviews: Networking Europe. Transnational Infrastructures and the Shaping of Europe, 1850–2000 — Edited by Erik van der Vleuten and Arne Kaijser. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 50(4):348–349, November 2008. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2009:EJU

[1138] Anonymous. Editorial: Journals under threat: a joint response from history of science, technology and medicine Editors. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(1):1–4, February 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rentetzi:2009:IGN

[1139] Maria Rentetzi and Sally G. Kohlstedt. Introduction: Gender and networking in Twentieth-Century physical sciences. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51 (1):5–11, February 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wennerholm:2009:OPE

[1140] Staffan Wennerholm. On the outskirts of physics: Eva von Bahr as an outsider within in early 20th Century Swedish experimental physics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(1):12–36, February 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kohlstedt:2009:UNA

[1141] Sally G. Kohlstedt and Suzanne M. Fischer. Unstable networks among women in academe: the legal case of Shyamala Rajender. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(1):37–62, February 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rentetzi:2009:GSP

[1142] Maria Rentetzi. Gender, science and politics: Queen Frederika and nuclear research in Post-war Greece. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(1):63–87, February 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Papanelopoulou:2009:MPS

[1143] Faidra Papanelopoulou and Peter C. Kjærgaard. Making the paper: Science and technology in Spanish, Greek and Danish newspapers around 1900. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(2):89–96, May 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gonzalez-Silva:2009:IES

[1144] Matiana González-Silva and Néstor Herran. Ideology, elitism and social commitment: Alternative images of science in two fin de siècle Barcelona newspapers. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(2):97–115, May 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mergoupi-Savaidou:2009:PIS

[1145] Eirini Mergoupi-Savaidou, Faidra Papanelopoulou, and Spyros Tzokas. The public image(s) of science and technology in the Greek daily press, 1908–1910. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(2):116–142, May 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Andersen:2009:DPI

[1146] Casper Andersen and Hans H. Hjermitslev. Directing public interest: Danish newspaper science 1900–1903. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(2):143–167, May 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Oaks:2009:RBB

[1147] Jeffrey A. Oaks. Reviews: Fibonacci’s De Practica Geometrie — by Barnabas Hughes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(2):168–169, May 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Boner:2009:RBL

[1148] Patrick J. Boner. Reviews: The Limits of Influence: Pico, Louvain, and the Crisis of Renaissance Astrology — by Steven Vanden Broecke. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(2):169–170, May 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Shackelford:2009:RBN

[1149] Jole Shackelford. Reviews: ‘Magic is no magic’, the wonderful world of Simon Stevin — by Jozef T. Devreese and Guido Vanden Berghe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(2):170–171, May 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Steigerwald:2009:RBL

[1150] Joan Steigerwald. Reviews: Science without laws: model systems, cases, exemplary narratives — Edited by Angela N. H. Creager, Elizabeth Lunbeck and N. Norton Wise. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(2):172–173, May 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2009:SSE

[1151] Helge Kragh. Styles of science and engineering: the case of early longdistance telephony. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(3):175–188, August 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Magruder:2009:JSA

[1152] Kerry V. Magruder. Jesuit science after Galileo: the cosmology of Gabriele Beati. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(3):189–212, August 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bonifacio:2009:TIA

[1153] Vitor Bonifácio, Isabel Malaquias, and João Fernandes. Taking internal advantage of external events — two astronomical examples from Nineteenth Century Portugal. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(3):213–234, August 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kjaergaard:2009:RBS

[1154] Peter C. Kjærgaard. Reviews: The Shock of the Old: Technology and Global History Since 1900 — by David Edgerton. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51 (3):235–236, August 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Becker:2009:RBL

[1155] Daniel Becker. Reviews: Müller’s Lab. The Story of Jakob Henle, Theodor Schwann, Emil du Bois-Reymond, Hermann von Helmholtz, Rudolf Virchow, Robert Remak, Ernst Haeckel, and Their Brilliant, Tormented Advisor — by Laura Otis. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(3):236–237, August 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2009:E

[1156] Anonymous. Erratum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(3):238, August 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chang:2009:WNB

[1157] Hasok Chang. We have never been whiggish (about phlogiston). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):239–264, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wegener:2009:SIG

[1158] Daan Wegener. Science and internationalism in Germany: Helmholtz, Du Bois-Reymond and their critics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):265–287, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hon:2009:SNS

[1159] Giora Hon and Bernard R. Goldstein. Spotlight on: the nature of scientific change: in pursuit of conceptual change: the case of Legendre and symmetry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):288–293, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ariew:2009:SNS

[1160] Roger Ariew. Spotlight on: The nature of scientific change: Some reflections on Thomas Kuhn’s account of scientific change. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):294–298, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bowen:2009:SNS

[1161] Alan C. Bowen. Spotlight on: the nature of scientific change: From description to prediction: an unexamined transition in Hellenistic astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):299–304, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Steinle:2009:SNS

[1162] Friedrich Steinle. Spotlight on: the nature of scientific change: Scientific change and empirical concepts. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):305–313, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chen:2009:SNS

[1163] Xiang Chen and Peter Barker. Spotlight on: the nature of scientific change: Process concepts and cognitive obstacles to change: Perspectives on the history of science and science policy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):314–320, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hinton:2009:RBS

[1164] Diana Davids Hinton. Reviews: Scientists and Swindlers: Consulting on Coal and Oil in America, 1820–1890 — by Paul Lucier. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):321–322, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pedersen:2009:RBR

[1165] Marco Nørskov Pedersen. Reviews: The Robot: the Life Story of a Technology — by Lisa Nocks. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):323–324, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Barker:2009:RBR

[1166] Peter Barker. Reviews: Scientific Research in World War II: What Scientists Did in the War — Edited by Ad Maas and Hans Hooijmaijers. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(4):324–326, November 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lutzen:2010:IHK

[1167] Jesper Lützen and Henrik Kragh Sørensen. Introduction: in honour of Kirsti Andersen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(1):1–3, February 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lutzen:2010:AGI

[1168] Jesper Lützen. The algebra of geometric impossibility: Descartes and Montucla on the impossibility of the duplication of the cube and the trisection of the angle. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(1):4–37, February 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sorensen:2010:TSL

[1169] Henrik Kragh Sørensen. Throwing some light on the vast darkness that is analysis: Niels Henrik Abel’s critical revision and the concept of absolute convergence. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(1):38–72, February 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Marcussen:2010:KAH

[1170] Marianne Marcussen. To Kirsti: ‘From Art History with Love… ’ perspective: Then and now, and artistic practice. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52 (1):73–88, February 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2010:ESHa

[1171] Anonymous. European Society for the History of Science —- announcement: 4th International Conference of the European Society for the History of Science (Barcelona, 18–20 November 2010) Hosted by: Societat Catalana d’Història de la Ciència i de la Tècnica (SCHCT) the circulation of science and technology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(1):89–90, February 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Blanchard:2010:MFO

[1172] Elodie Vieille Blanchard. Modelling the future: an overview of the ‘limits to growth’ debate. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):91–116, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Darrigol:2010:ABLa

[1173] Olivier Darrigol. The analogy between light and sound in the history of optics from the Ancient Greeks to Isaac Newton. Part 1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):117–155, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wepster:2010:RBM

[1174] Steven Wepster. Reviews: The Mathematics of the Heavens and the Earth: the Early History of Trigonometry — by Glen van Brummelen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):156–157, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rupke:2010:RBD

[1175] Nicolaas Rupke. Reviews: Ein ‘Diplomat aus den Wäldern des Orinoko.’ Alexander von Humboldt als Mittler zwischen Preußen und Frankreich — by Ulrich Päßler. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):157–158, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Katzir:2010:RBK

[1176] Shaul Katzir. Reviews: Der Komet in der Entladungsröhre Eugen Goldstein, Wilhelm Foerster und die Elektrizität im Weltraum. (German) [The comet in the discharge tube: Eugen Goldstein, Wilhelm Foerster and electricity in space] — by Michael Hedenus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):158–160, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Darvas:2010:RBS

[1177] György Darvas. Reviews: From Summetria to Symmetry: the Making of a Revolutionary Scientific Concept — by Giora Hon and Bernard R. Goldstein. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):160–162, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Maienschein:2010:RBC

[1178] Jane Maienschein. Reviews: The Cambridge History of Science. Volume 6. The Modern Biological and Earth Sciences — edited by Peter J. Bowler and John V. Pickstone. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):162–163, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heering:2010:RBH

[1179] Peter Heering. Reviews: Unsichtbare Hände. Zur Rolle von Laborassistenten, Mechanikern, Zeichern u.a. Amanuenses in der physikalischen Forschungs- und Entwicklungsarbeit. (German) [Invisible hands. On the role of lab assistants, mechanics, Zeichern [??], etc. Amanuenses in physical research and development work] — edited by Klaus Hentschel. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):163–164, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Steinhauser:2010:RBK

[1180] Thomas Steinhauser. Reviews: The Kaiser Wilhelm Society under National Socialism — edited by Susanne Heim, Carola Sachse and Mark Walker. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):164–166, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wazeck:2010:RBG

[1181] Milena Wazeck. Reviews: Einstein’s Generation: the Origins of the Relativity Revolution — by Richard Staley. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):166– 168, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lykknes:2010:RBM

[1182] Annette Lykknes. Reviews: Trafficking materials and gendered experimental practices: radium research in early 20th century Vienna — by Maria Rentetzi. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):168–170, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eckert:2010:RBF

[1183] Michael Eckert. Reviews: Mathematicians Fleeing from Nazi Germany. Individual Fates and Global Impact — by Reinhard Siegmund-Schultze. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):170–171, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jacobsen:2010:RBI

[1184] Anja Skaar Jacobsen. Reviews: Quantenmechanik im Kalten Krieg: David Bohm und Richard Feynman — by Christian Forstner. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):171–172, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2010:ESHb

[1185] Anonymous. European Society for the History of Science — announcement: 4th International Conference of the European Society for the History of Science, Barcelona, 18–20 November 2010: the circulation of science and technology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(2):173–174, May 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nielsen:2010:TTM

[1186] Kristian H. Nielsen. Technological trajectories in the making: Two case studies from the contemporary history of wind power. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(3):175–205, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Darrigol:2010:ABLb

[1187] Olivier Darrigol. The analogy between light and sound in the history of optics from the Ancient Greeks to Isaac Newton. Part 2. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(3):206–257, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bowler:2010:RBS

[1188] Peter J. Bowler. Reviews: Popularizing Science and Technology in the European Periphery, 1800–2000 — edited by Faidra Papanelopoulou. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(3):258–259, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Perez:2010:RBS

[1189] Nuria Valverde Pérez. Reviews: Secret Science: Spanish Cosmography and the New World — by María M. Portuondo. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52 (3):259–260, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2010:RBT

[1190] Helge Kragh. Reviews: Naturens Tankelæser: En Biografi om Hans Christian Ørsted — by Dan Ch. Christensen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52 (3):260–262, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cantor:2010:RBI

[1191] Geoffrey Cantor. Reviews: Driven to Innovate: a Century of Jewish Mathematicians and Physicists — by Ioan James. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52 (3):262–264, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Meli:2010:RBN

[1192] Domenico Bertoloni Meli. Reviews: Isaac Newton on Mathematical Certainty and Method — by Niccolò Guicciardini. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52 (3):264–265, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mazzotti:2010:RBN

[1193] Massimo Mazzotti. Reviews: The Newton Wars and the Beginning of the French Enlightenment — by John Bennett Shank. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(3):265–267, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Spary:2010:RBE

[1194] Emma Spary. Reviews: The English Virtuoso: Art, Medicine and Antiquarianism in the Age of Empiricism — by Craig Ashley Hanson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(3):267–268, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wils:2010:RBS

[1195] Kaat Wils. Reviews: Women in Science. A Social and Cultural History — by Ruth Watts. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(3):268–270, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2010:ESHc

[1196] Anonymous. European Society for the History of Science — announcement: 4th International Conference of the European Society for the History of Science, Barcelona, 18–20 November 2010 the circulation of science and technology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(3):271–272, August 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Abir-Am:2010:SPWa

[1197] Pnina G. Abir-Am. Spotlight on post-WW2 transatlantic science policies: Comparing strategic, political, and economic agendas in the United States of America, Europe, and East Asia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):273–279, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Krige:2010:SPW

[1198] John Krige. Spotlight on: Post-WW2 transatlantic science policies: Building the arsenal of knowledge. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):280–296, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Oreskes:2010:SPW

[1199] Naomi Oreskes. Spotlight on: Post-WW2 transatlantic science policies: Science, technology and free enterprise. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):297–310, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Doel:2010:SPW

[1200] Ronald E. Doel. Spotlight on: Post-WW2 transatlantic science policies: Does scientific intelligence matter? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):311–322, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Abir-Am:2010:SPWb

[1201] Pnina G. Abir-Am. Spotlight on: Post-WW2 transatlantic science policies: the Rockefeller Foundation and the post-WW2 transnational ecology of science policy: from solitary splendor in the inter-war era to a ‘me too’ agenda in the 1950s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):323–337, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Westwick:2010:SPW

[1202] Peter J. Westwick. Spotlight on: Post-WW2 transatlantic science policies: the international history of the strategic defense initiative: American influence and economic competition in the late cold war. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):338–351, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cahan:2010:RBH

[1203] David Cahan. Reviews: Hermann von Helmholtz’s Mechanism: the Loss of Certainty. A Study on the Transition from Classical to Modern Philosophy of Nature — by Gregor Schiemann. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):352–353, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brain:2010:RBT

[1204] Robert Brain. Reviews: A Tenth of a Second: a History — by Jimena Canales. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):353–355, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Keene:2010:RBA

[1205] Melanie Keene. Reviews: Science for All: the Popularization of Science in Early Twentieth–Century Britain — by Peter J. Bowler. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):355–356, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hughes:2010:RBN

[1206] Jeff Hughes. Reviews: A Nuclear Winter’s Tale: Science and Politics in the 1980s — by Lawrence Badash. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):356–358, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cittadino:2010:RBN

[1207] Eugene Cittadino. Reviews: Modern Nature: the Rise of the Biological Perspective in Germany — by Lynn K. Nyhart. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4): 358–360, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Strbanova:2010:ESH

[1208] Soňa Štrbáňová. European Society for the History of Science — letter from the Future President. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 52(4):361–362, November 2010. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VanBosstraeten:2011:DCC

[1209] Truus Van Bosstraeten. Dogs and Coca-Cola: Commemorative practices as part of laboratory culture at the Heymans Institute Ghent, 1902–1970. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):1–30, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bordoni:2011:BEM

[1210] Stefano Bordoni. Beyond electromagnetic and mechanical world-views: J. Larmor’s models of matter and energy in the early 1890s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):31–54, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sidoli:2011:HAD

[1211] Nathan Sidoli. Heron of Alexandria’s date. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):55–61, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Simon:2011:RBH

[1212] Jonathan Simon. Reviews: the Historiography of the Chemical Revolution: Patterns of Interpretation in the History of Science — by John McEvoy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):62–63, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Zemplen:2011:RBN

[1213] Gábor Á. Zemplén. Reviews: Autobiography and Natural Science in the Age of Romanticism: Rousseau, Goethe, Thoreau — by Kuhn Bernhard. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):63–65, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Giffard:2011:RBF

[1214] Hermione Giffard. Reviews: Missiles for the Fatherland: Peenemünde, National Socialism and the V-2 Missile — by Michael B. Petersen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):65–67, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brentjes:2011:RBA

[1215] Sonja Brentjes. Reviews: Studies in al-Kimya′: Critical Issues in Latin and Arabic Alchemy and Chemistry — Ahmad Y. al-Hassan. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):67, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Roberts:2011:RBW

[1216] Lissa Roberts. Reviews: James Watt, Chemist: Understanding the Origins of the Steam Age — by David Philip Miller. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53 (1):68–69, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Saetnan:2011:RBS

[1217] Ann Saetnan. Reviews: The Statistical Mind in Modern Society. The Netherlands 1850–1940 — edited by Jacques G. S. J. van Maarseveen, Paul M. M. Klep & Ida H. Stamhuis. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):69–71, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Turchetti:2011:RBW

[1218] Simone Turchetti. Reviews: Uranium Wars. The Scientific Rivalry That Created the Nuclear Age — by Amir D. Aczel. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53 (1):71–73, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dupre:2011:RBS

[1219] Sven Dupré. Reviews: the Social and Economic Roots of the Scientific Revolution — by Gideon Freudenthal & Peter McLaughlin. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):73–74, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2011:ESHa

[1220] Anonymous. European Society for the History of Science — announcement: the 2010 General Assembly was held at the conclusion of the 4th International Conference of the European Society for the History of Science, in the Institut d’Estudis Catalans, Barcelona, on 20 November 2010. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(1):75–76, February 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Carneiro:2011:ISI

[1221] Ana Carneiro and Marianne Klemun. Instruments of science — instruments of geology; introduction to Seeing and measuring, constructing and judging: Instruments in the history of the Earth sciences. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(2):77–85, May 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Klemun:2011:GHF

[1222] Marianne Klemun. The Geologist’s hammer — ‘fossil’ tool, equipment, instrument and/or badge? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(2):86–101, May 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vaccari:2011:TII

[1223] Ezio Vaccari. Travelling with instruments: Italian geologists in the field in the 18th and 19th centuries. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(2):102–115, May 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Malaquias:2011:SMI

[1224] Isabel Malaquias and Manuel S. Pinto. Searching for modernization — instruments in the development of Earth sciences in Portugal (18th Century). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(2):116–134, May 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brandstetter:2011:TMM

[1225] Thomas Brandstetter. Time machines: Model experiments in geology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(2):135–145, May 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mota:2011:BLM

[1226] Teresa Salomé Mota. A bursting landscape in the middle of Portugal: Theories and experiments by Georges Zbyszewski. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(2):146–163, May 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DeMFigueiroa:2011:MSA

[1227] Silvia F. De M. Figueirôa. Minerals scrutinized: Alberto Betim Paes Leme (1883–1938) and the application of spectrography. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(2):164–175, May 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Good:2011:MIE

[1228] Gregory A. Good. Measuring the inaccessible Earth: Geomagnetism, in situ measurements, remote sensing, and proxy data. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(2):176–189, May 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2011:ESHb

[1229] Anonymous. European Society for the History of Science — announcement: Universities in Central Europe — Crossroads of Scholars from all over the World 29 September 2011–1 October, 2011, Charles University in Prague, Czech Republic, first circular. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(2):190– 191, May 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Henry:2011:MMN

[1230] John Henry. ‘Mathematics made no contribution to the public weal’: Why Jean Fernel (1497–1558) became a physician. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53 (3):193–220, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Clutton-Brock:2011:PGT

[1231] Martin Clutton-Brock and David Topper. The plausibility of Galileo’s tidal theory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):221–235, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Walton:2011:RBE

[1232] Steven A. Walton. Reviews: Materials and Expertise in Early Modern Europe: Between Market and Laboratory — edited by Ursula Klein and Emma C. Spary. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):236–237, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Qureshi:2011:RBD

[1233] Sadiah Qureshi. Reviews: Spectacle of Deformity: Freak Shows and Modern British Culture — by Nadja Durbach. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53 (3):237–238, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Faye:2011:RBT

[1234] Jan Faye. Reviews: Quantum Theory at the Crossroads. Reconsidering the 1927 Solvay Conference — by Guido Bacciagaluppi and Antony Valentini. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):238–240, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thomas:2011:RBF

[1235] Marion Thomas. Reviews: Looking for a Few Good Males. Female Choice in Evolutionary Biology — by Erika Lorraine Milam. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):240–242, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kwa:2011:RBF

[1236] Chunglin Kwa. Reviews: Science: a Four Thousand Year History — by Patricia Fara. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):242–244, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brockliss:2011:RBF

[1237] Laurence Brockliss. Reviews: Échanges franco-britanniques entre savants depuis le XVIIe siècle. Franco–British Interactions in Science since the Seventeenth Century — edited by Robert Fox and Bernard Joly. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):244–245, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermij:2011:RBA

[1238] Rienk Vermij. Reviews: Copernicus and the Aristotelian Tradition. Education, Reading, and Philosophy in Copernicus’ Path to Heliocentrism — by André Goddu. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):245–247, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

McClellan:2011:RBC

[1239] James E. McClellan III. Reviews: Knowledge and Colonialism: Eighteenth-Century Travellers in South Africa — by Siegfried Huigen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):247–248, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoyrup:2011:RBC

[1240] Jens Høyrup. Reviews: Mathématiques et connaissance du monde réel avant Galilée. (French) [Mathematics and knowledge of the real world before Galileo] — edited by Rommevaux, Sabine. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53 (3):248–251, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nuvolari:2011:RBP

[1241] Alessandro Nuvolari. Reviews: Pursuing Power and Light. Technology and Physics from James Watt and Albert Einstein — by Bruce J. Hunt. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):251–252, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pisano:2011:ESHa

[1242] Raffaele Pisano and M. Rosa Massa-Esteve. European Society for the History of Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):253–254, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2011:C

[1243] Anonymous. Corrigendum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(3):255, August 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2011:CME

[1244] Helge Kragh. A controversial molecule: the early history of triatomic hydrogen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):257–279, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Onghena:2011:SCS

[1245] Sofie Onghena. The survival of 19th-Century scientific optimism: the public discourse on science in Belgium in the aftermath of the Great War (ca. 1919–1930). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):280–305, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Durand:2011:RSB

[1246] Antonin Durand and Laurent Mazliak. Revisiting the sources of Borel’s interest in probability: Continued fractions, social involvement, Volterra’s prolusione. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):306–332, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Peckham:2011:RBM

[1247] Robert Peckham. Reviews: Urban Modernity: Cultural Innovation in the Second Industrial Revolution — by Miriam R. Levin, Sophie Forgan, Martina Hessler, Robert H. Kargon and Morris Low. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):333–334, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tarrant:2011:RBE

[1248] Neil Tarrant. Reviews: Galileo Engineer — by Matteo Valleriani. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):334–336, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Allen:2011:RBI

[1249] Garland E. Allen. Reviews: Biology and Ideology: from Descartes to Dawkins — Edited by Denis Alexander and Ron Numbers. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):336–338, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Valleriani:2011:RBT

[1250] Matteo Valleriani. Reviews: Explanation and Teleology in Aristotle’s Science — by Mariska Leunissen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):338–340, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chappey:2011:RBN

[1251] Jean-Luc Chappey. Reviews: Science in the Nursery. The Popularisation of Science in Britain and France, 1761–1901 — Edited by Laurence Tailarach-Vielmas. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):340–342, November 2011. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thomas:2011:RBS

[1252] William Thomas. Reviews: Higher Speculations: Grand Theories and Failed Revolutions in Physics and Cosmology — by Helge Kragh. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):342–343, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hutton:2011:RBW

[1253] Sarah Hutton. Reviews: Reconfiguring the World: Nature, God, and Human Understanding from the Middle Ages to Early Modern Europe — by Margaret J. Osler. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):343–345, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pisano:2011:ESHb

[1254] Raffaele Pisano. European Society for the History of Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 53(4):346–347, November 2011. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tybjerg:2012:ICC

[1255] Karin Tybjerg. Introduction: Cosmos and cogwheels — the work of Ole Rømer (1644–1710). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(1):1–3, February 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kristensen:2012:RJS

[1256] Leif Kahl Kristensen and Kurt Møller Pedersen. Roemer, Jupiter’s satellites and the velocity of light. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(1):4–38, February 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Willmoth:2012:RFC

[1257] Frances Willmoth. Römer, Flamsteed, Cassini and the speed of light. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(1):39–57, February 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dijksterhuis:2012:CPI

[1258] Fokko Jan Dijksterhuis. Conjunctions in Paris: Interactions between Rømer and Huygens. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(1):58–76, February 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tybjerg:2012:SHS

[1259] Karin Tybjerg. ‘Senses and Hands to the Same Degree as Thought’ — Ole Rømer’s mechanical astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(1):77–102, February 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Strbanova:2012:ESHa

[1260] Soňa Štrbáňová and Milada Sekyrková. European Society for the History of Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(1):103–104, February 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VanWyhe:2012:AHD

[1261] John Van Wyhe and Mark J. Pallen. The ‘Annie Hypothesis’: Did the death of his daughter cause Darwin to ‘give up Christianity’? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):105–123, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bordoni:2012:UBMa

[1262] Stefano Bordoni. Unearthing a buried memory: Duhem’s third way to thermodynamics. Part 1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):124–147, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Norton:2012:ARA

[1263] John D. Norton and Bryan W. Roberts. Article with response: Article: Galileo’s refutation of the speed-distance law of fall rehabilitated. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):148–164, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Palmerino:2012:ARR

[1264] Carla Rita Palmerino. Article with response: Response: Aggregating speeds and scaling motions: a response to Norton and Roberts. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):165–176, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Laird:2012:ARR

[1265] W. R. Laird. Article with response: Response: Stillman Drake on Salviati’s Proof. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):177–181, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Norton:2012:ARR

[1266] John D. Norton and Bryan W. Roberts. Article with response: Reply to response: the scaling of speeds and distances in Galileo’s two new sciences: a reply to Palmerino and Laird. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):182– 191, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Robson:2012:RBN

[1267] Eleanor Robson. Reviews: Wrestling with Nature: From Omens to Science — edited by Peter Harrison, Ronald L. Numbers and Michael H. Shank. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):192–193, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Holmes:2012:RBW

[1268] John Holmes. Reviews: Darwin the Writer — by George Levine. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):193–194, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mandelbrote:2012:RBS

[1269] Scott Mandelbrote. Reviews: Selling Science in the Age of Newton. Advertising and the Commoditization of Knowledge — by Jeffrey R. Wigelsworth. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):194–196, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lehner:2012:RBU

[1270] Christoph Lehner. Reviews: Einstein’s Unification — by Jeroen van Dongen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):196–199, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Truitt:2012:RBD

[1271] E. R. Truitt. Reviews: Sublime Dreams of Living Machines: the Automaton in the European Imagination — by Minsoo Kang. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):199–200, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fangerau:2012:RBM

[1272] Heiner Fangerau. Reviews: Forensic Medicine in Western Society: a History — by Katherine D. Watson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):200–201, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pisano:2012:ESHa

[1273] Raffaele Pisano. European Society for the History of Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(2):202–203, May 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Turchetti:2012:SSB

[1274] Simone Turchetti. Sword, shield and buoys: a history of the NATO Sub-Committee on Oceanographic Research, 1959–1973. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):205–231, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bordoni:2012:UBMb

[1275] Stefano Bordoni. Unearthing a buried memory: Duhem’s third way to thermodynamics. Part 2. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):232–249, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2012:RBS

[1276] Helge Kragh. Reviews: A Short History of Scientific Thought — by John Henry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):250–251, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ahnert:2012:RBE

[1277] Thomas Ahnert. Reviews: Considerations, Encouragements, Improvements. Die Select Society in Edinburgh, 1754–1764 — by Iris Fleßenkämper. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):251–252, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Niekerk:2012:RBA

[1278] Carl Niekerk. Reviews: The Anatomy of Blackness: Science and Slavery in an Age of Enlightenment — by Andrew S. Curran. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):252–254, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Phillips:2012:RBA

[1279] Denise Phillips. Reviews: Wissen, Arbeit, Freundschaft. Ökonomien und soziale Beziehungen an den Akademien in London, Paris und Berlin um 1700 — by Sebastian Kühn. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):254–255, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hunter:2012:RBM

[1280] Matthew C. Hunter. Reviews: Silent Messengers: the Circulation of Material Objects of Knowledge in the Early Modern Low Countries — edited by Sven Dupré and Christoph Lüthy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):255–257, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dahms:2012:RBP

[1281] Hans-Joachim Dahms. Reviews: Michael Polanyi and his Generation. Origins of the Social Construction of Science — by Mary Jo Nye. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):257–259, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nathoo:2012:RBC

[1282] Ayesha Nathoo. Reviews: Tissue Culture in Science and Society: the Public Life of a Biological Technique in Twentieth Century Britain — by Duncan Wilson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):259–261, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Strbanova:2012:ESHb

[1283] Soňa Štrbáňová. European Society for the History of Science: the Moscow S. I. Vavilov Institute for the History of Science and Technology celebrates its eightieth anniversary. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(3):262–264, August 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Elbers:2012:ENF

[1284] Astrid Elbers. The establishment of the new field of radio astronomy in the Post-War Netherlands: a search for Allies and funding. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(4):265–285, November 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rossiter:2012:SUPa

[1285] Margaret W. Rossiter. Spotlight on: Using the past to enhance the future: Introduction to spotlight on using the past to enhance the future. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(4):286–287, November 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rossiter:2012:SUPb

[1286] Margaret W. Rossiter. Spotlight on: Using the past to enhance the future: the more you look, the more you find: Archives of recent American women scientists. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(4):288–291, November 2012. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Zanish-Belcher:2012:SUP

[1287] Tanya Zanish-Belcher. Spotlight on: Using the past to enhance the future: the archives of women in science and engineering and future directions for oral history: Questions for women scientists. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(4):292–298, November 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eller:2012:SUP

[1288] Troy Eller. Spotlight on: Using the past to enhance the future: Publicity, recruitment, and history: Society of women engineers. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(4):299–304, November 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sedgwick:2012:SUP

[1289] Jessica Sedgwick. Spotlight on: Using the past to enhance the future: the archives for women in medicine: Documenting Women’s experiences and contributions at Harvard medical school. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(4):305–310, November 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gavroglu:2012:ERS

[1290] Kostas Gavroğlu. Essay review: the STEP (Science and Technology in the European Periphery) initiative: Attempting to historicize the notion of European science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(4):311–327, November 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Peeters:2012:RBM

[1291] Evert Peeters. Reviews: Technology and the Making of The Netherlands. The Age of Contested Modernization, 1890–1970 — Edited by Schot H. Lintsen and A. Rip. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(4):328–329, November 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pisano:2012:ESHb

[1292] Raffaele Pisano. European Society for the History of Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 54(4):330–331, November 2012. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kirchner:2013:CTM

[1293] Eric Kirchner and Mohammad Bagheri. Color theory in Medieval Islamic lapidaries: Nıshābūrı, Tūsı and Kāshānı. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):1–19, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Blumenthal:2013:LAC

[1294] Geoffrey Blumenthal. On Lavoisier’s achievement in chemistry. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):20–47, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cook:2013:RBC

[1295] Harold J. Cook. Reviews: Inventing Chemistry: Herman Boerhaave and the Reform of the Chemical Arts — by John C. Powers. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):48–49, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lamy:2013:RBC

[1296] Jérôme Lamy. Reviews: Into the Cosmo. Space Exploration and Soviet Culture — edited by James T. Andrew and Asif A. Siddiqi. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):49–50, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller-Sievers:2013:RBR

[1297] Helmut Muller-Sievers. Reviews: The Romantic Machine. Utopian Science and Technology after Napoleon — by John Tresch. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):50–52, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Romano:2013:RBD

[1298] Elisa Romano. Reviews: What Did the Romans Know? An Inquiry into Science and Worldmaking — by Daryn Lehoux. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):52–54, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kemper:2013:RBP

[1299] Claudia Kemper. Reviews: The Pontecorvo Affair: a Cold War Defection and Nuclear Physics — by Simone Turchetti. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):54–56, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Omodeo:2013:RBC

[1300] Pietro D. Omodeo. Reviews: The Copernican Question: Prognostication, Skepticism, and Celestial Order — by Robert S. Westman. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):56–58, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Freis:2013:RBM

[1301] David Freis. Reviews: From Melancholia to Prozac: a History of Depression — Clark Lawlor. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):58–59, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Garcia-Sancho:2013:RBC

[1302] Miguel Garcia-Sancho. Reviews: The Cybernetic Brain: Sketches of Another Future — by Andrew Pickering. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):59–61, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Debru:2013:NLE

[1303] Claude Debru. Neuenschwander Lecture, ESHS 2012: Postwar science in divided Europe: a continuing cooperation. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):62–69, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2013:FIC

[1304] Anonymous. Fifth International Conference of the ESHS. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(1):70–71, February 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonOertzen:2013:FSS

[1305] Christine von Oertzen, Maria Rentetzi, and Elizabeth S. Watkins. Finding science in surprising places: Gender and the geography of scientific knowledge. Introduction to ‘beyond the academy: Histories of gender and knowledge’. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(2):73–80, May 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Leong:2013:CKF

[1306] Elaine Leong. Collecting knowledge for the family: Recipes, gender and practical knowledge in the early modern English household. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(2):81–103, May 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bittel:2013:WKT

[1307] Carla Bittel. Woman, know thyself: Producing and using phrenological knowledge in 19th-Century America. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(2):104– 130, May 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Leng:2013:EPF

[1308] Kirsten Leng. An ‘elusive’ phenomenon: Feminism, sexology and the female sex drive in Germany at the turn of the 20th century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(2):131–152, May 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kohlstedt:2013:INS

[1309] Sally Gregory Kohlstedt. Innovative niche scientists: Women’s role in reframing North American museums, 1880–1930. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55 (2):153–174, May 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonOertzen:2013:SCM

[1310] Christine von Oertzen. Science in the cradle: Milicent Shinn and her home-based network of baby observers, 1890–1910. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(2):175–195, May 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pomata:2013:ACW

[1311] Gianna Pomata. Amateurs by choice: Women and the pursuit of independent scholarship in 20th Century historical writing. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(2):196–219, May 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heymann:2013:IPC

[1312] Matthias Heymann and Janet Martin-Nielsen. Introduction: Perspectives on Cold War science in small European states. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(3):221–242, August 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Roberts:2013:IIC

[1313] Peder Roberts. Intelligence and internationalism: The Cold War career of Anton Bruun. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(3):243–263, August 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoeneveld:2013:CBS

[1314] Friso Hoeneveld and Jeroen van Dongen. Out of a clear blue sky? FOM, the bomb, and the boost in Dutch physics funding after World War II. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(3):264–293, August 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Petersen:2013:PUM

[1315] Nikolaj Petersen. The politics of US military research in Greenland in the Early Cold War. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(3):294–318, August 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nielsen:2013:THH

[1316] Henry Nielsen and Henrik Knudsen. Too hot to handle: The controversial hunt for uranium in Greenland in the Early Cold War. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(3):319–343, August 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kinney:2013:SGB

[1317] D. J. Kinney. Selling Greenland: The big picture television series and the Army’s bid for relevance during the Early Cold War. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(3):344–357, August 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2013:ESHa

[1318] Anonymous. The European Society for the History of Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(3):358–359, August 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Patiniotis:2013:BLG

[1319] Manolis Patiniotis. Between the local and the global: History of science in the European periphery meets post-colonial studies. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):361–384, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Turner:2013:CHG

[1320] Laura E. Turner and Henrik Kragh Sørensen. Cultivating the herb garden of Scandinavian mathematics: The congresses of Scandinavian mathematicians, 1909–1925. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):385–411, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Launius:2013:SCA

[1321] Roger D. Launius. The strange career of the American spaceplane: The long history of wings and wheels in human space operations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):412–432, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nye:2013:RBW

[1322] Mary Jo Nye. Reviews: Is Water H_{2}O? Evidence, Realism and Pluralism — by Hasok Chang. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):433–434, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hodson:2013:RBL

[1323] Niall Hodson. Reviews: Communities of Learned Experience: Epistolary Medicine in the Renaissance — by Nancy G. Siraisi. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):435–436, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Raby:2013:RBL

[1324] Megan Raby. Reviews: Ordering Life: Karl Jordan and the Naturalist Tradition — by Kristin Johnson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):437–438, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chen-Morris:2013:RBH

[1325] Raz Chen-Morris. Reviews: A History of Optics from Greek Antiquity to the Nineteenth Century — by Olivier Darrigol. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4): 438–439, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brockmann:2013:RBE

[1326] Sophie Brockmann. Reviews: Visible Empire: Botanical Expeditions and Visual Culture in the Spanish Enlightenment — by Daniela Bleichmar. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):439–441, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Simon:2013:RBS

[1327] Jonathan Simon. Reviews: The Savant and the State: Science and Cultural Politics in Nineteenth-Century France — Robert Fox. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):441–443, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sepkoski:2013:RBL

[1328] David Sepkoski. Reviews: Trees of Life: A Visual History of Evolution — Theodore W. Pietsch. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):443–444, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Marchetti:2013:RBN

[1329] Christian Marchetti. Reviews: The Nationalization of Scientific Knowledge in the Habsburg Empire 1848–1918 — edited by Mitchell G. Ash and Jan Surman. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4):444–446, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2013:ESH

[1330] Anonymous. European Society for the History of Science: 6th International Conference of the ESHS “Communicating Science, Technology And Medicine” Lisbon, 4–6 September 2014. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 55(4): 447–448, November 2013. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Helsvig:2014:APM

[1331] Kim G. Helsvig. The Abel prize: The missing Nobel in mathematics? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1):1–30, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ducheyne:2014:GANa

[1332] Steffen Ducheyne. ‘s Gravesande’s appropriation of Newton’s natural philosophy, Part I: Epistemological and theological issues. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1):31–55, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2014:RBS

[1333] Koen Vermeir. Reviews: Art and Science in the Early Modern Netherlands — edited by Eric Jorink and Bart Ramakers. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56 (1):56–57, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2014:RBT

[1334] Helge Kragh. Reviews: Traditions and Transformations in the History of Quantum Physics — edited by Shaul Katzir, Christoph Lehner, and Jürgen Renn. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1):58–59, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Perez:2014:RBE

[1335] Nuria Valverde Pérez. Reviews: Eating the Enlightenment. Food and the Sciences in Paris, 1670-1760 — by Emma C. Spary. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1):59–61, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Roy:2014:RBB

[1336] Tirthankar Roy. Reviews: Science between Europe and Asia: Historical Studies on the Transmission, Adoption and Adaptation of Knowledge — edited by Feza Günergun and Dhruv Raina. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1):61–62, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wolfe:2014:RBO

[1337] Charles T. Wolfe. Reviews: Penser L’Ordre Naturel, 1680–1810 — edited by Adrien Paschoud and Nathalie Vuillemin. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1): 62–65, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muller-Wille:2014:RBN

[1338] Staffan Müller-Wille. Reviews: Acolytes of NATURE: Defining natural SCIENCE in Germany, 1770–1850 — by Denise Phillips. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1):65–67, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tramelli:2014:RBA

[1339] Barbara Tramelli. Reviews: Influences: Art, Optics and Astrology in the Italian Renaissance — by Mary Quinlan-McGrath. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1):67–68, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Truitt:2014:RBE

[1340] E. R. Truitt. Reviews: Androids in the Enlightenment: Mechanics, Artisans, and Cultures of the Self — by Adelheid Voskuhl. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(1):68–70, February 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dhondt:2014:HDD

[1341] Pieter Dhondt. Honorary doctoral degrees as expressions of political and cultural relationships at Nordic University jubilees (1840–1911). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(2):71–96, May 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ducheyne:2014:GANb

[1342] Steffen Ducheyne. ’s Gravesande’s appropriation of Newton’s natural philosophy, Part II: Methodological issues. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(2):97– 120, May 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schneider:2014:RBS

[1343] Ivo Schneider. Reviews: Baroque Science — by Ofer Gal and Raz Chen-Morris. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(2):121–123, May 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anker:2014:RGH

[1344] Peder Anker. Reviews: The Gaia Hypothesis: Science on a Pagan Planet — by Michael Ruse. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(2):123–124, May 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Westermann:2014:REO

[1345] Andrea Westermann. Reviews: The Earthquake Observers. Disaster Science from Lisbon to Richter — by Deborah R. Coen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(2):124–126, May 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Canizares-Esguerra:2014:RSK

[1346] Jorge Canizares-Esguerra. Reviews: Shores of Knowledge: New World Discoveries and the Scientific Imagination — by Joyce Appleby. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(2):126–128, May 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rothman:2014:RKC

[1347] Aviva Rothman. Reviews: Kepler’s Cosmological Synthesis: Astrology, Mechanism and the Soul — by Patrick J. Boner. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(2):128–129, May 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Reinarz:2014:RSS

[1348] Jonathan Reinarz. Reviews: Studies of Skin Color in the Early Royal Society — by Cristina Malcolmson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(2):129–131, May 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

White:2014:RWW

[1349] Deborah Elise White. Reviews: Worm Work: Recasting Romanticism — by Janelle A. Schwarz. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(2):131–133, May 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Orr:2014:NOG

[1350] Mary Orr. New observations on a geological hotspot track: Excursions in Madeira and Porto Santo (1825) by Mrs T. Edward Bowdich. Centaurus:

An International Journal of the History of Science and its Cultural Aspects, 56(3):135–166, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Raposo:2014:MLC

[1351] Pedro M. P. Raposo, Ana Simões, Manolis Patiniotis, and José R. Bertomeu-Sánchez. Moving localities and creative circulation: Travels as knowledge production in 18th-century Europe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56 (3):167–188, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hutton:2014:RSL

[1352] Sarah Hutton. Reviews: Seduced by Logic: Émilie du Châtelet, Mary Somerville and the Newtonian Revolution — by Robyn Arianrhod. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(3):189–190, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Roberts:2014:RMM

[1353] Lissa Roberts. Reviews: Matter and Method in the Long Chemical Revolution: Laws of Another Order — by Victor Boantza. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(3):190–192, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Loison:2014:ROS

[1354] Laurent Loison. Reviews: Outsider Scientists: Routes to Innovation in Biology — Edited by Oran Harman and Michael R. Dietrich. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(3):192–194, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Page:2014:RNS

[1355] Michael Page. Reviews: Novel Science: Fiction and the Invention of Nineteenth-Century Geology — by Adelene Buckland. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(3):194–195, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Illy:2014:REB

[1356] József Illy. Reviews: Einstein’s Berlin: In the Footsteps of a Genius — by Dieter Hoffmann. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(3):195–197, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heymann:2014:RSG

[1357] Matthias Heymann. Reviews: Science, Geopolitics and Culture in the Polar Region. Norden Beyond Borders — Edited by Sverker Sörlin. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(3):197–198, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brock:2014:RSS

[1358] Adrian C. Brock. Reviews: The Sciences of the Soul: The Early Modern Origins of Psychology — by Fernando Vidal. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56 (3):198–200, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wille:2014:RSS

[1359] Robert-Jan Wille. Reviews: Scientists and Scholars in the Field. Studies in the History of Fieldwork and Expeditions — by Kristian H. Nielsen, Michael Harbsmeier and Christopher J. Ries. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(3): 200–201, August 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2014:FCR

[1360] Anonymous. Foreword: Contemporary Russian scholarship on the history of science in Russia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(4):203–204, November 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bayuk:2014:HSR

[1361] Dimitri A. Bayuk. History of science in Russia: The IIET in Moscow and St. Petersburg. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(4):205–228, November 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Loskutova:2014:ERI

[1362] Marina Loskutova. Early research on insect pests in the Russian empire: Bureaucracy, academic community and local knowledge in the 1830s–1840s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(4):229–253, November 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fedotova:2014:FES

[1363] Anastasia Fedotova. Forestry experimental stations: Russian proposals of the 1870s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(4):254–274, November 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Krivosheina:2014:LWA

[1364] Galina Krivosheina. Long way to the anthropological exhibition: The institutionalization of physical anthropology in Russia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(4):275–304, November 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Elina:2014:BLP

[1365] Olga Elina. Between local practices and global knowledge: Public initiatives in the development of agricultural science in Russia in the 19th Century and early 20th Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(4):305–329, November 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kolchinsky:2014:NVY

[1366] Eduard I. Kolchinsky. Nikolai Vavilov in the years of Stalin’s ‘revolution from above’ (1929–1932). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 56(4):330–358, November 2014. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DoMarGago:2015:TDG

[1367] Maria Do Mar Gago. Things of darkness: Genetics, melanins and the regime of Salazar (1936–1952). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(1):1–27, February 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Spary:2015:RCN

[1368] E. C. Spary. Reviews: Catching Nature in the Act: Réaumur and the Practice of Natural History in the Eighteenth Century — by Mary Terrall. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(1):28–29, February 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wils:2015:RBW

[1369] Kaat Wils. Reviews: For Better or For Worse? Collaborative Couples in the Sciences — edited by Annette Lykknes, Donald L. Opitz and Brigitte Van Tiggelen. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(1):29–31, February 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heymann:2015:RHR

[1370] Matthias Heymann. Reviews: How Reason almost Lost its Mind: The Strange Career of Cold War Rationality — by Paul Erickson, Judy L. Klein, Lorraine Daston, Rebecca Lemov, Thomas Sturm and Michael D. Gordin. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(1):31–33, February 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rentetzi:2015:RWS

[1371] Maria Rentetzi. Reviews: Women Scientists in America: Forging a New World since 1972 — by Margaret W. Rossiter. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57 (1):33–35, February 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Reiss:2015:RED

[1372] Christian Reiß. Reviews: Emil du Bois-Reymond: Neuroscience, Self and Society in Nineteenth Century Germany — by Gabriel Finkelstein. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(1):35–36, February 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Renn:2015:NLE

[1373] Jürgen Renn. Neuenschwander Lecture, ESHS 2014: From the history of science to the history of knowledge — and back. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(1):37–53, February 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2015:ESHa

[1374] Anonymous. European Society for the History of Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(1):54–60, February 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chilvers:2015:FTS

[1375] Christopher A. J. Chilvers. Five tourniquets and a ship’s bell: The special session at the 1931 Congress. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):61–95, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Blumenthal:2015:ACE

[1376] Geoffrey Blumenthal. Astrology and Copernicus’s early experiences in the world of Renaissance politics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):96–115, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Giffard:2015:REM

[1377] Hermione Giffard. Reviews: Engineers and the Making of the Francoist Regime — by Lino Camprubí. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):116–117, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Simon:2015:RLM

[1378] Jonathan Simon. Reviews: The Limits of Matter: Chemistry, Mining and Enlightenment — by Hjalmar Fors. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):117– 119, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Timmermann:2015:RGU

[1379] Carsten Timmermann. Reviews: Generic: The Unbranding of Modern Medicine — by Jeremy A. Greene. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):119–120, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rothman:2015:RGI

[1380] Aviva Rothman. Reviews: Galileo’s Idol: Gianfrancesco Sagredo & the Politics of Knowledge — by Nick Wilding. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):120– 122, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fulford:2015:RMS

[1381] Tim Fulford. Reviews: Mary Somerville and the World of Science — by Allan Chapman. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):122–123, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rocha:2015:RCS

[1382] Leon Antonio Rocha. Reviews: The Classification of Sex: Alfred Kinsey and the Organization of Knowledge — by Donna J. Drucker. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):123–125, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Knoeff:2015:RCV

[1383] Rina Knoeff. Reviews: Commercial Visions: Science, Trade and Visual Culture in the Dutch Golden Age — by Dániel Margócsy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):125–126, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schemmel:2015:RSL

[1384] Matthias Schemmel. Reviews: From Sight to Light: The Passage from Ancient to Modern Optics — by A. Mark Smith. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(2):126–128, May 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2015:IPE

[1385] Koen Vermeir. Introduction: Preface: Electricity and imagination. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(3):129–130, August 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2015:EIP

[1386] Koen Vermeir. Electricity and imagination: Post-romantic electrified experience and the gendered body. An introduction. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(3):131–155, August 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gilmore:2015:JNL

[1387] Paul Gilmore. John Neal’s lightning imagination: Electricity against romantic organicism. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(3):156–172, August 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Morus:2015:NMD

[1388] Iwan Rhys Morus. No mere dream: Material culture and electrical imagination in Late Victorian Britain. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(3):173–191, August 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Otto:2015:EEA

[1389] Ulf Otto. Enter electricity: An allegory’s stage appearance between Verité and Varieté. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(3):192–211, August 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Halliday:2015:EHC

[1390] Sam Halliday. Electricity and homosexuality: from 19th-century American sexual health literature to D. H. Lawrence. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57 (3):212–228, August 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kroupa:2015:EEP

[1391] Sebestian Kroupa. Ex epistulis Philippinensibus: Georg Joseph Kamel SJ (1661–1706) and his correspondence network. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4): 229–259, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lalli:2015:RHA

[1392] Roberto Lalli. Reviews: Heisenberg in the Atomic Age: Science and the Public Sphere — by Cathryn Carson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):260–261, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pemberton:2015:RGG

[1393] Stephen Pemberton. Reviews: The Genealogy of a Gene: Patents, HIV/ AIDS, and Race — by Myles W. Jackson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):261–263, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dooley:2015:RCC

[1394] Brendan Dooley. Reviews: The Crown and the Cosmos. Astrology and the Politics of Maximilian I — by Darin Hayton. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4): 263–265, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Marktanner:2015:RSL

[1395] Alina Marktanner. Reviews: Seeing Like a Rover. How Robots, Teams and Images Craft Knowledge of Mars — by Janet Vertesi. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):265–266, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Murphy:2015:RLP

[1396] Jane Murphy. Reviews: Learned Patriots: Debating Science, State, and Society in the 19th-Century Ottoman Empire — by Alper Yalçınkaya. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):266–268, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Henderson:2015:RMU

[1397] Gabriel Henderson. Reviews: Masters of Uncertainty: Weather Forecasters and the Quest for Ground Truth — by Phaedra Daipha. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):268–269, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Perkins:2015:RSI

[1398] Emma Perkins. Reviews: Scientific Instruments on Display — by Silke Ackermann, Richard L. Kremer and Mara Miniati. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4): 269–271, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pisano:2015:RTV

[1399] Raffaele Pisano. Reviews: Tunnel Visions: The Rise and Fall of the Superconducting Super Collider — by Michael Riordan, Lillian Hoddeson and Adrienne W. Kolb. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):271–273, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Turda:2015:RUN

[1400] Marius Turda. Reviews: Urologie und Nationalsozialismus. Eine Studie zu Medizin und Politik als Ressourcen füreinander — by Matthis Krischel. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):273–274, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kangaslahti:2015:RWF

[1401] Kate Kangaslahti. Reviews: World’s Fairs on the Eve of War: Science, Technology and Modernity, 1937–1942 — by Robert H. Kargon, Karen Fiss, Morris Low and Arthur P. Molella. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):274–276, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hentschel:2015:RSR

[1402] Klaus Hentschel. Reviews: Serving the Reich: The Struggle for the Soul of Physics under Hitler — by Philip Ball. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):276–277, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chemla:2015:PAE

[1403] Karine Chemla. Presidential address ESHS 2014: Proof, generality and the prescription of mathematical action: A nanohistorical approach to communication. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):278–300, November 2015. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2015:ESHb

[1404] Anonymous. European Society for the History of Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):301–306, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Husson:2016:IHD

[1405] Matthieu Husson and Richard L. Kremer. Introduction: How do writings in the early astral sciences reveal mathematical practices and practitioners? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(1–2):1–5, February/May 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Falk:2016:LMA

[1406] Seb Falk. Learning medieval astronomy through tables: The case of the equatorie of the planetis. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(1–2):6–25, February/May 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Li:2016:TTC

[1407] Liang Li. Template tables and computational practices in early modern Chinese calendric astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(1–2):26–45, February/May 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Husson:2016:ETC

[1408] Matthieu Husson. Exploring the temporality of complex computational practice: Two eclipse notes by John of Murs in the ms Escorial O II 10. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(1–2):46–65, February/May 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hirose:2016:TVD

[1409] Sho Hirose. Two versions of a description of the armillary sphere in Parame’svara’s Goladipika. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(1–2):66–86, February/ May 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Morgan:2016:SCT

[1410] Daniel Patrick Morgan. Sphere confusion: A textual reconstruction of astronomical instruments and observational practice in first-millennium CE China. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(1–2):87–103, February/May 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kremer:2016:PGT

[1411] Richard L. Kremer. Playing with geometrical tools: Johannes Stabius’s Astrolabium imperatorium (1515) and its successors. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(1–2):104–134, February/May 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hochadel:2016:ISS

[1412] Oliver Hochadel, Miquel Carandell Baruzzi, and Clara Florensa. Introduction: Scuffles, scoops and scams: The construction of prehistoric knowledge in newspapers. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(3):135–147, August 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cataldi:2016:IMM

[1413] Maddalena Cataldi. Inventing the Menton Man. Rivière’s discovery as reflected in the French media. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(3):148–165, August 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Podgorny:2016:DPF

[1414] Irina Podgorny. The daily press fashions a heroic intellectual: The making of Florentino Ameghino in Late Nineteenth-Century Argentina. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(3):166–184, August 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Plana:2016:MPP

[1415] Laura Valls Plana. A mammoth in the park: Palaeontology, press and popular culture in Barcelona (1870–1910). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(3):185–202, August 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hochadel:2016:OSM

[1416] Oliver Hochadel. One skull and many headlines: The role of the press in the Steinau Hoax of 1911. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(3):203–218, August 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Florensa:2016:DWW

[1417] Clara Florensa. ‘Darwin was wrong.’ The international media coverage of the Oreopithecus ’ reinterpretation (1956–1959). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58 (3):219–238, August 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Baruzzi:2016:FAS

[1418] Miquel Carandell Baruzzi. The first American scoop: The Pedra Furada controversy in newspapers (1978–2015). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(3):239– 256, August 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Elvbakken:2016:RBA

[1419] Kari Tove Elvbakken and Annette Lykknes. Relationships between academia, state and industry in the field of food and nutrition: The Norwegian chemist Sigval Schmidt-Nielsen (1877–1956) and his professional roles, 1900–1950. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):257–280, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cozzoli:2016:EBD

[1420] Daniele Cozzoli. Ethno-biology during the Cold War: Biocca’s expedition to Amazonia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):281–309, November 2016. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Loison:2016:RLO

[1421] Laurent Loison. Reviews: The Life Organic, The Theoretical Biology Club and the Roots of Epigenetics — by Erik L. Peterson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):310–311, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Au:2016:ROA

[1422] Sokhieng Au. Reviews: Old Age, New Science: Gerontologists and Their Biosocial Visions, 1900–1960 — by Hyung W. Park. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):311–313, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mody:2016:RCH

[1423] Cyrus C. M. Mody. Reviews: A Companion to the History of American Science — by Georgina M. Montgomery and Mark A. Largent. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):313–315, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thorpe:2016:RHV

[1424] Charles Thorpe. Reviews: The Hope and Vision of J. Robert Oppenheimer — by Michael A. Day. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):315–316, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hochadel:2016:RCH

[1425] Oliver Hochadel. Reviews: A Companion to the History of Science, Blackwell Companions to World History— by Bernard Lightman. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):316–318, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Flipse:2016:RTS

[1426] Abraham C. Flipse. Reviews: The Territories of Science and Religion— by Peter Harrison. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):318–320, November 2016. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Aspaas:2016:RJS

[1427] Per Pippin Aspaas. Reviews: Jesuit Science and the End of Nature’s Secrets— by Mark A. Waddell. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):320–322, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Baldasso:2016:RVC

[1428] Renzo Baldasso. Reviews: Visual Cultures in Science and Technology. A Comparative History— by Klaus Hentschel. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):322–324, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fontaine:2016:RRA

[1429] Philippe Fontaine. Reviews: Rational Action: The Sciences of Policy in Britain and America, 1940–1960— by William Thomas. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):324–326, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Badino:2016:YSL

[1430] Massimiliano Badino. Young Scholar Lecture ESHS 2016: What have the historians of quantum physics ever done for us? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):327–346, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2016:ESH

[1431] Anonymous. European Society for the History of Science: New feature: Now history of science in Romania. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):347–352, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heymann:2017:TEH

[1432] Matthias Heymann. 1970s: Turn of an era in the history of science? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):1–9, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12146.

Graf:2017:DST

[1433] Rüdiger Graf. Détente science? Transformations of knowledge and expertise in the 1970s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):10–25, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1600-0498.12148.

Egan:2017:SSC

[1434] Michael Egan. Survival science: Crisis disciplines and the shock of the environment in the 1970s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):26–39, February– May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1600-0498.12149.

Seefried:2017:GSF

[1435] Elke Seefried. Globalized science. The 1970s futures field. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):40–57, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1111/1600-0498.12150.

Mody:2017:SSE

[1436] Cyrus C. M. Mody. Square scientists and the excluded middle. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):58–71, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12147.

Young:2017:NSE

[1437] Mark Thomas Young. Nature as spectacle; experience and empiricism in early modern experimental practice. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):72– 96, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print),

1600-0498 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/1600-0498.12155.

Knight:2017:GRM

[1438] Nathaniel Knight. Geography, race and the malleability of man: Karl von Baer and the problem of academic particularism in the Russian human sciences. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):97–121, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12154.

Couttenier:2017:RBC

[1439] Maarten Couttenier. ‘With the risk of being called retrograde’. Racial classifications and the attack on the Aryan myth by Jean-Baptiste d’Omalius d’Halloy (1783–1875). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):122–151, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-

0498 (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1111/1600-0498.12144.

Margocsy:2017:BRM

[1440] Dániel Margócsy. Book review: Medieval Robots: Mechanism, Magic, Nature, and Art — by E. R. Truitt. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):152–153, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-

0498 (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1111/1600-0498.12139.

Reinarz:2017:BRR

[1441] Jonathan Reinarz. Book review: Rise of the Modern Hospital: An Architectural History of Health and Healing, 1870–1940 — by J. Kisacky. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):153–155, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12151.

Anderson:2017:HCV

[1442] Robert G. W. Anderson. Book review: The History of Chemistry. A Very Short Introduction — by W. H. Brock. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):155– 156, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print),

1600-0498 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/1600-0498.12152.

Widmalm:2017:SFC

[1443] Sven Widmalm. Book review: Science without Frontiers: Cosmopolitanism and National Interests in the World of Learning, 1870–1940 — by R. Fox. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):156–157, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12153.

Camprubi:2017:BRE

[1444] Lino Camprubi. Book review: Engineering the Environment. Phytotrons and the Quest for Climate Control in the Cold War — by D. P. Munns. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):157–159, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12156.

Anonymous:2017:IEB

[1445] Anonymous. Israel and Europe. Building bridges via the history of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(1–2):160–165, February–May 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12145.

Mulberger:2017:PSH

[1446] Annette Mülberger and Jaume Navarro. The promises of science. Historical perspectives. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(3):167–172, August 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12157.

Navarro:2017:PRS

[1447] Jaume Navarro. Promising redemption. Science at the service of secular and religious agendas. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(3):173–188, August 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1600-0498.12158.

Mormann:2017:SWP

[1448] Thomas Mormann. Scientific worldviews as promises of science and problems of philosophy of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(3):189–203, August 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1600-0498.12159.

Mulberger:2017:PRM

[1449] Annette Mülberger. The persuasive rhetoric of a manifesto (1870): Ribot’s promise of an “independent” psychological science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(3):204–222, August 2017. CODEN CENTA4. ISSN 0008-8994 (print),

1600-0498 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/1600-0498.12160.

Katzir:2017:SWP

[1450] Shaul Katzir. “In war or in peace:” the technological promise of science following the First World War. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(3):223–237, August 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1600-0498.12161.

Bowler:2017:PSE

[1451] Peter J. Bowler. The promise of science in early 20th-century popular literature. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(3):238–250, August 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12162.

Guilbaud:2017:HOC

[1452] Alexandre Guilbaud and Christophe Schmit. A historiographical overview of the current state of research into Jean Le Rond D’Alembert (1717–1783). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(4):251–262, November 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12168.

Launay:2017:JRA

[1453] Françoise Launay. From Jean Le Rond to D’Alembert: a route enlightened by new archival data. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(4):263–273, November 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1600-0498.12170.

Passeron:2017:HRA

[1454] Irène Passeron. How to read D’Alembert’s correspondence? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(4):274–284, November 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1111/1600-0498.12165.

Schmit:2017:BNS

[1455] Christophe Schmit. Beginnings of a new science. D’Alembert’s Traité de dynamique and the French Royal Academy of Sciences around 1740. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(4):285–299, November 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12167.

Jouve:2017:FWA

[1456] Guillame Jouve. The first works of D’Alembert and Euler about the problem of vibrating strings from the perspective of their correspondence. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(4):300–307, November 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12171.

Ferlin:2017:NIM

[1457] Fabrice Ferlin. New insights into major theoretical research in optics in the Age of Enlightenment. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(4):308–319, November 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1600-0498.12163.

Guilbaud:2017:DCE

[1458] Alexandre Guilbaud. D’Alembert’s contribution to the Encyclopédie: A review of research on the inventory of his articles. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(4):320–328, November 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1111/1600-0498.12166.

Comtesse:2017:BAR

[1459] Dagmar Comtesse and Moritz Epple. Between appropriation and rejection: Translating D’Alembert into German, and D’Alembert on translation. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 59(4):329–341, November 2017. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1600-0498.12164.

Vermeir:2018:CYE

[1460] Koen Vermeir. Celebrating 15 years of the European Society for the History of Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):1–2, February–May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Krasnodebski:2018:TLP

[1461] Marcin Krasnodebski. Throwing light on photonics: The genealogy of a technological paradigm. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):3–24, February–May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mazliak:2018:BSE

[1462] Laurent Mazliak. The beginnings of the Soviet encyclopedia. The utopia and misery of mathematics in the political turmoil of the 1920s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):25–51, February–May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Figueiroa:2018:FAB

[1463] Silvia Fernanda de M. Figueirôa. A French author in a Brazilian library: Nerée Boubée (1806–1862) and his textbooks on geological sciences. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):52–68, February–May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Zhu:2018:HWW

[1464] Yiwen Zhu. How were Western written calculations introduced into China? — an analysis of the Tongwen suanzhi (arithmetic guidance in the common language, 1613). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):69–86, February–May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Leong:2018:RDO

[1465] Elaine Leong. Read. Do. Observe. Take note! Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60 (1–2):87–103, February–May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2018:FYH

[1466] Koen Vermeir, Claude Debru, Robert Fox, Eberhard Knobloch, Helge Kragh, Sona Strbánová, Fabio Bevilacqua, Karine Chemla, and Antoni Malet. Fifteen years of the history of science in Europe: Personal reflections by the ESHS presidents. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):104–123, February–May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Loison:2018:GGP

[1467] Laurent Loison. The gene: From genetics to postgenomics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):124–125, February–May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dyck:2018:KSS

[1468] Maarten Van Dyck. Kuhn’s structures of scientific revolutions at fifty. Reflections on a science classic. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):126–128, February– May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bruyninckx:2018:AET

[1469] Joeri Bruyninckx. The age of electroacoustics: Transforming science and sound. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):128–130, February–May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heggie:2018:SEE

[1470] Vanessa Heggie. Science in an extreme environment: The 1963 American Mount Everest expedition. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):130–131, February– May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Oosterhuis:2018:SPF

[1471] Harry Oosterhuis. Sexual politics and feminist science: Women sexologists in Germany 1900–1933. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(1–2):131–133, February– May 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sanchez:2018:ACE

[1472] Antonio Sánchez and Henrique Leitão. Artisanal culture in early modern Iberian and Atlantic worlds. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(3):135–140, August 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eamon:2018:CCQ

[1473] William Eamon. Corn, cochineal, and quina: The “Zilsel Thesis” in a colonial Iberian setting. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(3):141–158, August 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vos:2018:RPO

[1474] Paula De Vos. Rosewater and philosophers’ oil: Thermo-chemical processing in medieval and early modern Spanish pharmacy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(3):159–172, August 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sanchez:2018:CEG

[1475] Antonio Sánchez. Charts for an empire. A global trading zone in early modern Portuguese nautical cartography. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(3):173–188, August 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Leitao:2018:IAP

[1476] Henrique Leitão. Instruments and artisanal practices in long distance oceanic voyages. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(3):189–202, August 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Collins:2018:IPA

[1477] Edward Collins. Interactions of Portuguese artisanal culture in the maritime enterprise of 16th-century Seville. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(3):203– 215, August 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Almeida:2018:TNC

[1478] Bruno Almeida. Transmitting nautical and cosmographical knowledge in the 16th and 17th centuries: The case of Pedro Nunes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(3):216–229, August 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2018:EOS

[1479] Anonymous. The ESHS officers and the Scientific Council. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(3):230, August 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Klemun:2018:SBC

[1480] Marianne Klemun, Marina Loskutova, and Anastasia Fedotova. Skulls and blossoms: Collecting and the meaning of scientific objects as resources from the 18th to the 20th century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(4):231–237, November 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Napolitani:2018:SBT

[1481] Maddalena Napolitani. “Born with the taste for science and the arts”: The science and the aesthetics of Balthazar-Georges Sage’s mineralogy collections, 1783–1825. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(4):238–256, November 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Elina:2018:PPC

[1482] Olga Elina. A passion for plants: Collections and power games in botany in the Russian Empire from the 18th to the early 19th century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(4):257–275, November 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tammiksaar:2018:SWS

[1483] Erki Tammiksaar and Ken Kalling. “I was stealing some skulls from the bone chamber when a bigamist cleric stopped me.” Karl Ernst von Baer and the development of physical anthropology in Europe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(4):276–293, November 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Slepkova:2018:ZMS

[1484] Nadezhda V. Slepkova and Tatiana I. Yusupova. The Zoological Museum of the St. Petersburg Academy of Sciences, 1860s–1910: From an academic institute to a public museum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(4):294–314, November 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fedotova:2018:KME

[1485] Anastasia Fedotova, Tomasz Samojlik, and Piotr Daszkiewicz. Killing for museums: European bison as a museum exhibit. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(4):315–332, November 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Diagre-Vanderpelen:2018:RFB

[1486] Denis Diagre-Vanderpelen. The rise and fall of the Belgian Forestry Museum and Geographic Arboretum (1900–1980): A political origin and a winning opportunity for science? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 60(4):333–349, November 2018. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2019:E

[1487] Koen Vermeir. Editorial. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):1, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sastre-Juan:2019:FFB

[1488] Jaume Sastre-Juan and Jaume Valentines-Álvarez. Fun and fear: The banalization of nuclear technologies through display. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61 (1–2):2–13, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Boyle:2019:SBA

[1489] Alison Boyle. “Banishing the atom pile bogy”: Exhibiting Britain’s first nuclear reactor. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):14–32, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sastre-Juan:2019:SIY

[1490] Jaume Sastre-Juan. “If You Tilt This game, will it explode?”: The politics of nuclear display at the New York Hall of Science (1966–1973). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):33–50, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kasperski:2019:CNR

[1491] Tatiana Kasperski. Children, nation and reactors: Imagining and promoting nuclear power in contemporary Ukraine. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61 (1–2):51–69, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Valentines-Alvarez:2019:MFA

[1492] Jaume Valentines-Álvarez and Ana Macaya-Andrés. Making fun of the atom: Humor and pleasant forms of anti-nuclear resistance in the Iberian Peninsula, 1974–1984. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):70–90, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nothaft:2019:AUN

[1493] C. Philipp E. Nothaft. An Alfonsine universe: Nicolò Conti and Georg Peurbach on the threefold motion of the fixed stars. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):91–110, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Malet:2019:SPF

[1494] Antoni Malet. Science and power: Francoist Spain (1939–1975) as a case study. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):111–132, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lecuyer:2019:BRM

[1495] Christophe Lécuyer. Book review: Models of innovation: The history of an idea, by Benoît Godin, Cambridge, MA: MIT Press, 2017, 344 pp. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):133–134, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Loredo-Narciandi:2019:BRD

[1496] José Carlos Loredo-Narciandi. Book review: Detrás de la cortina. El sexo en España (1790–1950), by Jean Louis Guereña, Madrid: Cátedra 2018, 630 p. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):135–136, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Horiuchi:2019:BRM

[1497] Annick Horiuchi. Book review: Making Time: Astronomical Time Measurement in Tokugawa Japan, by Yulia Frumer, Chicago, IL: University of Chicago Press, 2018, 272 p. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):137–139, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Regier:2019:IHE

[1498] Jonathan Regier and Koen Vermeir. Introduction: History of early astronomy in Centaurus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(1–2):139–142, February 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rev:2019:NON

[1499] István Rév. Neither objective nor subjective. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61 (3):143–152, August 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Georgiev:2019:CWA

[1500] Georgi Georgiev. Cold War atmosphere: Distorted information and facts in the case of Free Europe balloons. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(3):153–177, August 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Petrinca:2019:RWM

[1501] Ruxandra Petrinca. Radio waves, memories, and the politics of everyday life in socialist Romania: The case of Radio Free Europe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(3):178–199, August 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Walewska-Choptiany:2019:LTI

[1502] Joanna Walewska-Choptiany. Listening through the Iron Curtain: RFE and Polish Radio in the “fog of war”. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(3):200– 231, August 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Surman:2019:TBC

[1503] Jan Surman. Terminology between chemistry and philology: A Polish interdisciplinary debate in 1900? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(3):232–253, August 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Simoes:2019:LBS

[1504] Ana Simões. Looking back, stepping forward: Reflections on the sciences in Europe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(3):254–267, August 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sanchez:2019:PKE

[1505] Antonio Sánchez. Practical knowledge and empire in the early modern Iberian world. Towards an artisanal turn. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(3):268–281, August 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pla:2019:BRG

[1506] Mònica Balltondre Pla. Book review: García, Dauder, S., Sedeño, Eulalia Pérez. Las “mentiras” científicas sobre las mujeres [Scientific “lies” about women]. Editorial Catarata, Madrid, 2017, 256 pp. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(3):282–284, August 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Canedo-Rodriguez:2019:BRE

[1507] Montserrat Cañedo-Rodríguez. Book review: Edited by Oliver Hochadel, Agustín Nieto-Galán. Urban histories of science. Making knowledge in the city, 1820–1940. Abingdon: Routledge, 2019, 252 pp. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(3):284–285, August 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lykknes:2019:PSM

[1508] Annette Lykknes and Brigitte Van Tiggelen. The periodic system: The (multiple) values of an icon. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):287–298, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bensaude-Vincent:2019:RCE

[1509] Bernadette Bensaude-Vincent. Reconceptualizing chemical elements through the construction of the periodic system. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4): 299–310, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Campbell:2019:PTI

[1510] Chris Campbell. The periodic table as an icon: A perspective from the philosophy of Charles Sanders Peirce. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):311–328, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2019:PSI

[1511] Helge Kragh. The periodic system and the idea of a chemical element: From Mendeleev to superheavy elements. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):329–344, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VanTiggelen:2019:TRP

[1512] Brigitte Van Tiggelen and Annette Lykknes. A tale of resilience: The periodic table after radioactivity and the discovery of the neutron. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):345–359, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Robinson:2019:CPP

[1513] Ann E. Robinson. Chemical pedagogy and the periodic system. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):360–378, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonWulfingen:2019:PTF

[1514] Bettina Bock von Wülfingen. The periodic tableau: Form and colours in the first 100 years. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):379–404, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pulkkinen:2019:VPM

[1515] Karoliina Pulkkinen. Values and periodicity: Mendeleev’s reception of the equations of Mills, Chicherin, and Vincent. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4): 405–423, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fox:2019:SDP

[1516] Robert Fox. Savants and diplomats: The politics of commemoration at the Berthelot centenary, 1927. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):424–442, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Giannetto:2019:BRR

[1517] Enrico Giannetto. Book review: Raphael,Renée [Jennifer]. Reading Galileo: Scribal technologies and the two new sciences. Baltimore, MD: Johns Hopkins University Press, 2017, xii + 265 pp. ISBN: 978-1-4214-2177-3. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):443–444, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rieppel:2019:BRA

[1518] Lukas Rieppel. Book review: Adrian Currie. Rock, bone, and ruin: An optimist’s guide to the historical sciences. Cambridge, MA: MIT Press. Life and mind: Philosophical issues in biology and psychology, 2018, 376 pp. ISBN: 978-0-262-03726-6. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):444–446, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cellamare:2019:BRM

[1519] Davide Cellamare. Book review: Marjorie O’Rourke Boyle. Cultural anatomies of the heart in Aristotle, Augustine, Aquinas, Calvin, and Harvey. Cham, Switzerland: Palgrave Macmillan, 2018, xxvii + 171 pp. ISBN: 978-3-319-93652-9. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):446–448, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Darrigol:2019:BRH

[1520] Olivier Darrigol. Book review: Helge Kragh. Ludvig Lorenz: A nineteenth-century theoretical physicist. Copenhagen, Denmark: Det kongelige Danske Videnskabernes Selskab, 2018, 280 pp. ISBN 978-87-7304-417-9. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):448–449, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mare:2019:BRE

[1521] Lucio Mare. Book review: Edited by Andrault, Raphaële, Lærke, Mogens. Steno and the philosophers. Leiden, The Netherlands: Brill. Brill’s Studies in Intellectual History, 2018, xii + 291 pp. ISBN: 978-90-04-36064-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):449–451, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Walter:2019:BRR

[1522] Scott A. Walter. Book review: Roberto Lalli. Building the general relativity and gravitation community during the Cold War. Cham, Switzerland: Springer. Springer Briefs in History of Science and Technology, 2017, xiv + 168 pp. ISBN: 978-3-319-54654-4. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):451–453, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kragh:2019:BRV

[1523] Helge Kragh. Book review: Volny Fages. Savantes nébuleuses: L’origine du monde entre marginalité et autorité scientifique (1860–1920). Paris, France: Editions EHESS, 2018, 362 pp. ISBN: 978-2-7132-2726-4. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):453–455, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Blum:2019:BRH

[1524] Alexander S. Blum. Book review: Hentschel, Klaus. Photons: The history and mental models of light quanta. Dordrecht, The Netherlands: Springer International Publishing, 2018, xiii + 231 pp. ISBN: 978-3-319-95251-2. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 61(4):455–456, November 2019. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2020:WAE

[1525] Koen Vermeir. What about editors? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):1–4, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fyfe:2020:IEE

[1526] Aileen Fyfe and Anna Gielas. Introduction: Editorship and the editing of scientific journals, 1750–1950. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):5–20, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Moxham:2020:ALJ

[1527] Noah Moxham. “Accoucheur of literature”: Joseph Banks and the Philosophical Transactions, 1778–1820. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):21–37, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gielas:2020:TTI

[1528] Anna Gielas. Turning tradition into an instrument of research: The editorship of William Nicholson (1753–1815). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):38–53, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Stoeger:2020:PAM

[1529] Alexander Stoeger. Putting astronomy on the map: The launch of the first geographical-astronomical journal. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):54–68, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jenkins:2020:CSJ

[1530] Bill Jenkins. Commercial scientific journals and their editors in Edinburgh, 1819–1832. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):69–81, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vandendriessche:2020:TJE

[1531] Joris Vandendriessche. Turning journals into encyclopaedias: Medical editorship and reprinting in the Low Countries (1815–1860). Centaurus:

An International Journal of the History of Science and its Cultural Aspects, 62(1):82–97, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Beckman:2020:ELP

[1532] Jenny Beckman. Editors, librarians, and publication exchange: The Royal Swedish Academy of Sciences in the long 19th century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):98–110, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Baldwin:2020:BBE

[1533] Melinda Baldwin. The business of being an editor: Norman Lockyer, Macmillan and Company, and the editorship of Nature, 1869–1919. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):111–124, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fyfe:2020:ERC

[1534] Aileen Fyfe. Editors, referees, and committees: Distributing editorial work at the Royal Society journals in the late 19th and 20th centuries. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):125–140, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Carman:2020:VEO

[1535] Christián C. Carman. Vestiges of the emergence of overspecification and indifference to visual accuracy in the mathematical diagrams of medieval manuscripts. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):141–157, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wartenberg:2020:NLT

[1536] Ilana Wartenberg. A non-linear transmission of Euclid’s Elements in a medieval Hebrew calendrical treatise. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):158–180, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Elior:2020:WDM

[1537] Ofer Elior. What did medieval readers take to be “Al-H .ajjaj’s version” of Euclid’s Elements? The evidence of MS Paris, BnF, héb. 1011. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):181–197, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2020:CEP

[1538] Koen Vermeir. Centaurus at 70: Editors’ perspectives. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):198–205, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kusukawa:2020:BRE

[1539] Sachiko Kusukawa. Book review: Edited by Christoph Lüthy, Claudia Swan, Paul J. J. M. Bakker, Claus Zittel. Image, imagination, and cognition: Medieval and Early Modern theory and practice. Leiden, The Netherlands: Brill. Intersections 55, 2018. xvi + 324 pp. ISBN: 978-90-04-36574-2. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):206–208, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chaney:2020:BRE

[1540] Eric Chaney. Book review: Edited by Sonja Brentjes. Teaching and learning the sciences in Islamicate societies (800–1700). Turnhout, Belgium: Brepols Publishers, Studies on the Faculty of Arts, History and Influence 3, 2019, 334 pp. ISBN 978-2-503-57445-5. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62 (1):208–209, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Corneanu:2020:BRM

[1541] Sorana Corneanu. Book review: Marr, Alexander, Garrod, Raphaële, Ramón Marcaida, José, Oosterhoff, Richard J.. Logodaedalus: word histories of ingenuity in Early Modern Europe. Pittsburgh, PA: University of Pittsburgh Press, 2018, xviii + 358 pp. ISBN: 978-0-8229-4541-3. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):209–211, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Axworthy:2020:BRO

[1542] Angela Axworthy. Book review: Oosterhoff, Richard J.. Making mathematical culture: University and print in the circle of Lefèvre d’Étaples. Oxford, United Kingdom: Oxford University Press, 2018, xiv + 276 pp. ISBN: 978-0-19-882352-0. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):211–213, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hajek:2020:BRB

[1543] Kim M. Hajek. Book review: Le Bras, Anatole. Un enfant à l’asile. Vie de Paul Taesch (1874–1914). Paris, France: CNRS Editions, 2018, 298 pp. ISBN: 978-2-271-11500-3. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):213–215, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vandendriessche:2020:BRC

[1544] Joris Vandendriessche. Book review: Csiszar, Alex. The scientific journal: Authorship and the politics of knowledge in the Nineteenth Century. Chicago, IL: Chicago University Press, 2018, 368 pp. ISBN: 978-0-226-55337-5. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(1):216–217, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Babes:2020:BRG

[1545] Ovidiu Babes. Book review: George Andrei Iavorenciuc. Cunoastere si occidentalizare: O istorie a stiintei românesti de la jumatatea secolului XIX pâna la începutul secolului XX. Cluj-Napoca, Romania: Editura Mega, 2018, 494 pp. ISBN: 978-606-543-951-1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62 (1):218, February 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2020:EDH

[1546] Koen Vermeir. Editorial: Doing history in the time of COVID-19. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):219–222, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Charters:2020:HSM

[1547] Erica Charters and Richard A. McKay. The history of science and medicine in the context of COVID-19. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):223– 233, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Green:2020:EDR

[1548] Monica H. Green. Emerging diseases, re-emerging histories. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):234–247, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Geissler:2020:LEP

[1549] P. Wenzel Geissler and Ruth J. Prince. Layers of epidemy: Present pasts during the first weeks of COVID-19 in western Kenya. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):248–256, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Leung:2020:CSS

[1550] Angela Ki Che Leung. Chinese state and society in epidemic governance: a historical perspective. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):257–262, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Henderson:2020:IEF

[1551] John Henderson. The invisible enemy: Fighting the plague in early modern Italy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):263–274, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Heitman:2020:AAF

[1552] Kristin Heitman. Authority, autonomy and the first London Bills of Mortality. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):275–284, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Varlik:2020:RHP

[1553] Nükhet Varlik. Rethinking the history of plague in the time of COVID-19. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):285–293, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pelling:2020:BVE

[1554] Margaret Pelling. “Bosom vipers”: Endemic versus epidemic disease. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):294–301, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pouget:2020:BRQ

[1555] Benoît Pouget. Book review: Quarantine, cholera, and international health spaces: Reflections on 19th-century European sanitary regulations in the time of SARS-CoV-2. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):302–310, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Keck:2020:ATC

[1556] Frédéric Keck. Asian tigers and the Chinese dragon: Competition and collaboration between sentinels of pandemics from SARS to COVID-19. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):311–320, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

DAbramo:2020:HPE

[1557] Flavio D’Abramo and Sybille Neumeyer. A historical and political epistemology of microbes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):321–330, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dolan:2020:IWS

[1558] Brian Dolan. It wasn’t supposed to be a coronavirus: The quest for an influenza A(H5N1)-derived vaccine and the limits of pandemic preparedness. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):331–343, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bretelle-Establet:2020:SDG

[1559] Florence Bretelle-Establet. Science, demons, and gods in the battle against the COVID-19 epidemic. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):344–353, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Birn:2020:HNF

[1560] Anne-Emanuelle Birn. How to have narrative-flipping history in a pandemic: Views of/from Latin America. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):354–369, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:2020:CHH

[1561] David S. Jones. COVID-19, history, and humility. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(2):370–380, May 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rusu:2020:GBD

[1562] Doina-Cristina Rusu and Dana Jalobeanu. Giovan Battista Della Porta and Francis Bacon on the creative power of experimentation. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):381–392, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rusu:2020:UIS

[1563] Doina-Cristina Rusu. Using instruments in the study of animate beings: Della Porta’s and Bacon’s experiments with plants. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62 (3):393–405, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Borrelli:2020:GBD

[1564] Arianna Borrelli. Giovan Battista Della Porta’s construction of pneumatic phenomena and his use of recipes as heuristic tools. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):406–424, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jalobeanu:2020:ERG

[1565] Dana Jalobeanu. Enacting recipes: Giovan Battista Della Porta and Francis Bacon on technologies, experiments, and processes of nature. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):425–446, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pastorino:2020:BRB

[1566] Cesare Pastorino. Beyond recipes: The Baconian natural and experimental histories as an epistemic genre. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):447–464, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Morawiec:2020:BRI

[1567] Adam Morawiec. Book review: Itinerarium Wittichi ex Calendarium Sculteti: New biographical evidence on the Breslau mathematician Paul Wittich (ca. 1546–ca. 1587). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):465–478, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lycas:2020:PGS

[1568] Alexis Lycas. The patterned guidelines of Shazhou (Shazhou tujing) and geographical practices in Tang China. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):479–497, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Spary:2020:PVM

[1569] E. C. Spary. Publishing virtue: Medical entrepreneurship and reputation in the Republic of Letters. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):498–521, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Eric:2020:AAA

[1570] Cindy Hodoba Eric. Artificial apertures: The archaeology of Ramazzini’s De fontium in 17th-century Earth historiography. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):522–541, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jalobeanu:2020:TPL

[1571] Dana Jalobeanu and Oana Matei. Treating plants as laboratories: A chemical natural history of vegetation in 17th-century England. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):542–561, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Manias:2020:DH

[1572] Chris Manias. Dinosaur histories. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):562–565, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Olsakova:2020:BRJ

[1573] Doubravka Olsáková. Book review: John Krige, How knowledge moves. Writing the transnational history of science and technology. Chicago, IL: The University of Chicago Press, 2019, 408 pp., paperback. ISBN: 978-0-226-60599-9. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):566–567, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Duque:2020:BRJ

[1574] Carl Antonius Lemke Duque. Book review: Jan Jakub, Surman Universities in Imperial Austria 1848–1918. A social history of a multilingual space. West Lafayette, IN: Purdue University Press, 2019, 460 pp., $49.95, ISBN: 978-1-55753-837-6. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):567–570, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Portuondo:2020:BRY

[1575] María M. Portuondo. Book review: Ydea astronomica de la fabrica del mundo y movimiento de los cuerpos celestiales. Cedillo Díaz, Juan Miguel Angel Granada Félix Gómez Crespo Barcelona, Spain: Unversitat de Barcelona, 2019, 484 pp. ISBN: 978-84-9168-164-9. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62 (3):570–571, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chakrabarti:2020:BRS

[1576] Pratik Chakrabarti. Book review: Suman, Seth, Difference and disease: Medicine, race, and the eighteenth-century British Empire. Cambridge, UK: Cambridge University Press, 2018, 336 pp. ISBN: 978-1-108-41830-0. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):572–573, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Arabatzis:2020:BRJ

[1577] Theodore Arabatzis. Book review: Jaume Navarro, Ether and modernity: The recalcitrance of an epistemic object in the early twentieth century. Oxford, UK: Oxford University Press, 2018, 272 pp., Hardcover, $85.00. ISBN: 978-0-19-879725-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):573–575, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lopez:2020:BRD

[1578] Susana Gómez López. Book review: Domenico Bertoloni, Meli Mechanism. A visual, lexical, and conceptual history. Pittsburgh, PA: University of Pittsburgh Press, 2019, xii+188 pp. ISBN: 978-0-8229-4547-5; 0-8229-4547-9. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):576–577, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Malara:2020:BRM

[1579] Ivan Malara. Book review: Mehl, Édouard. Descartes et la fabrique du monde. Le problème cosmologique, de Copernic à Descartes. Paris, France: Presses Universitaires de France. Épiméthée, 2019, 424 pp. ISBN: 978-2-13-080065-1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):578–579, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Borrelli:2020:BRE

[1580] Arianna Borrelli. Book review: Emily, Herring, Kevin Matthew Jones, Konstantin Kiprijanov, Laura Sellers Chilton. The past, present, and future of integrated history and philosophy of science. London, England: Routledge, 2019, xiii+256 pp. ISBN: 978-0-8153-7985-0; 978-1-351-21482-7. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):579–581, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Storni:2020:BRS

[1581] Marco Storni. Book review: Shank, John Bennett. Before Voltaire: The French origins of “Newtonian mechanics”, 1680–1715. Chicago, IL and London, UK: The University of Chicago Press, 2018, ix + 444 pp. ISBN: 978-0-226-50929-7. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):581–583, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nutton:2020:BRM

[1582] Vivian Nutton. Book review: Murphy, Hannah. A new order of medicine. The rise of physicians in Reformation Nuremberg. Pittsburgh, PA: University of Pittsburgh Press, 2019, x + 262 pp. ISBN: 978-0-8229-4560-4. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):583–584, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schwarz:2020:BRB

[1583] Angela Schwarz. Book review: Berg Matthias, Thiel Jens. Europäische Wissenschaftsakademien im “Krieg der Geister”. Reden und Dokumente 1914 bis 1920 (Acta Historica Leopoldina, Vol. 72). Stuttgart, Germany: Wissenschaftliche Verlagsgesellschaft, 2018, 319 pp. ISBN: 978-3-8047-3797-6. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):584–586, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Begley:2020:BRC

[1584] Justin Begley. Book review: Cavendish, Margaret. Edited by Anne M. Thell. Grounds of natural philosophy. Peterborough, ON: Broadview Press, 2020, 264 pp. ISBN: 978-1-55481-387-2. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62 (3):586–588, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gregory:2020:BRS

[1585] Frederick Gregory. Book review: Steigerwald, Joan. Experimenting at the boundaries of life: Organic vitality in Germany around 1800. Pittsburgh, PA: University of Pittsburgh Press, 2019, xii + 460 pp. ISBN: 978-0-8229-4553-6. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):588–589, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Celani:2020:BRA

[1586] Ruben Celani. Book review: Álvarez Mari-Tere, Villaseñor Black Charlene. Renaissance futurities: Science, art, invention. Oakland, CA: University of California Press, 2019, 238 pp. ISBN: 978-0-520-29698-5; 978-0-520-96951-3. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):590–592, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rusu:2020:BRO

[1587] Doina-Cristina Rusu. Book review: Omodeo Pietro Daniel. Bernardino Telesio and the natural sciences in the Renaissance. Leiden, The Netherlands: Brill, 2019, xvi + 285 pp. ISBN: 978-90-04-35263-6; 978-90-04-35264-3. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(3):592–594, August 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vonHardenberg:2020:IRR

[1588] Wilko Graf von Hardenberg and Martin Mahony. Introduction — up, down, round and round: Verticalities in the history of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):595–611, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anthony:2020:MMM

[1589] Patrick Anthony. Mines, mountains, and the making of a vertical consciousness in Germany ca. 1800. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):612–630, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Reidy:2020:DVT

[1590] Michael S. Reidy. Darwin’s vertical thinking: Mountains, mobility, and the imagination in 19th-century geology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):631–646, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bergwik:2020:EES

[1591] Staffan Bergwik. Elevation and emotion: Sven Hedin’s mountain expedition to Transhimalaya, 1906–1908. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):647–669, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mattes:2020:NSD

[1592] Johannes Mattes. National spaces and deepest places: Politics and practices of verticality in speleology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):670–696, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Minor:2020:JML

[1593] Adriana Minor. Up-and-down journeys: The making of Latin America’s uniqueness for the study of cosmic rays. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):697–719, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Achermann:2020:VGS

[1594] Dania Achermann. Vertical glaciology: The second discovery of the third dimension in climate research. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):720–743, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Henry:2020:MWV

[1595] Matthew Henry. Making weather vertical: Meteorology and the temporalities of infrastructural atmospheres in New Zealand, ca. 1920–1950. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):744–762, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Adler:2020:DHC

[1596] Antony Adler. Deep horizons: Canada’s underwater habitat program and vertical dimensions of marine sovereignty. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):763–782, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Merchant:2020:VOH

[1597] Paul Merchant. Verticalities in oral histories of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):783–796, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanWingerden:2020:SEE

[1598] Pieter van Wingerden. Science on the edge of empire: E. A. Forsten (1811–1843) and the Natural History Committee (1820–1850) in the Netherlands Indies. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):797–821, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gordin:2020:WDW

[1599] Michael D. Gordin. What do we talk about when we talk about language and science? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):822–825, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Dupre:2020:BRL

[1600] Sven Dupré. Book review: Lissa L. Roberts, Simon Werrett. Compound histories: Materials, governance and production, 1760–1840. Leiden, The Netherlands and Boston, MA: Brill, 2018, xiv — 388 pp. ISBN: 978-90-04-32549-4. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):826–828, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Huertas:2020:BRA

[1601] Rafael Huertas. Book review: Andrew, Scull. Psychiatry and its discontents. Oakland, CA: University of California Press, 2019, 345 pp. ISBN: 978-0-520-97357-2; 978-0-520-30549-6. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4): 828–830, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brain:2020:BRB

[1602] Robert Brain. Book review: Bonea, Amelia, Dickson, Melissa, Shuttleworth, Sally, Wallis, Jennifer. Anxious Times: Medicine & modernity in nineteenth-century Britain. Pittsburgh, PA: University of Pittsburgh Press, 2019, vii + 312 pp. ISBN: 978-0-8229-4551-2. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):830–832, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gonzalez:2020:BRN

[1603] Daniel López González. Book review: Newell Catherine L.. Destined for the stars: Faith, the future and America’s final frontier. Pittsburgh, PA: University of Pittsburgh Press, 2019, 292 pp. ISBN: 978-0-8229-4556-7; 0-8229-4556-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):832–833, November 2020. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Smith:2020:BRD

[1604] Roger Smith. Book review: Dezső Gurka, Changes in the image of man from the Enlightenment to the Age of Romanticism: Philosophical and scientific receptions of (physical) anthropology in the 18–19th centuries. Budapest, Hungary: Gondolat, 2019, 280 pp. ISBN: 978-963-693-300-5. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):834–835, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sanchez:2020:BRN

[1605] Mariana Sanchez. Book review: Navarro Brotons Víctor. Jerónimo Muñoz: Matemáticas, cosmología y humanismo en la época del Renacimiento. València, Spain: Universitat de València, 2019, 324 pp. ISBN: 978-84-9134-439-1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):835–837, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jesus:2020:BRT

[1606] Paulo Jesus. Book review: Thibaud, Trochu. William James: Une autre histoire de la psychologie. Paris, France: CNRS Éditions, 2018, 375 pp. ISBN: 978-2-271-12300-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):837–839, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Achim:2020:BRM

[1607] Miruna Achim. Book review: Matthew James Crawford, Joseph M. Gabriel. Drugs on the page: Pharmacopoeias and healing knowledge in the early modern Atlantic world. Pittsburgh, PA: University of Pittsburgh Press, 2019. ix + 384 pp. ISBN: 978-0-8229-4562-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 62(4):839–841, November 2020. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Adamson:2021:GPS

[1608] Matthew Adamson and Roberto Lalli. Global perspectives on science diplomacy: Exploring the diplomacy-knowledge nexus in contemporary histories of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):1–16, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Zaidi:2021:SPE

[1609] S. Waqar H. Zaidi. Scientists as political experts: Atomic scientists and their claims for expertise on international relations, 1945–1947. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):17–31, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Zhang:2021:TAV

[1610] Li Zhang and Yanmei Zhu. Technical assistance versus cultural export: George Cressey and the U.S. Cultural Relations Program in wartime China, 1942–1946. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):32–50, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Adamson:2021:FFU

[1611] Matthew Adamson and Simone Turchetti. Friends in fission: US–Brazil relations and the global stresses of atomic energy, 1945–1955. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):51–66, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Castelo:2021:IAC

[1612] Cláudia Castelo and Frederico Ágoas. Inter-African cooperation in the social sciences in the era of decolonization: a case of science diplomacy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):67–83, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Robinson:2021:AND

[1613] Ann E. Robinson. Attempting neutrality: Disciplinary and national politics in a Cold War scientific controversy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):84–102, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lalli:2021:CEC

[1614] Roberto Lalli. Crafting Europe from CERN to Dubna: Physics as diplomacy in the foundation of the European Physical Society. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):103–131, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Janac:2021:RSC

[1615] Jirí Janác and Doubravka Olsáková. On the road to Stockholm: A case study of the failure of Cold War international environmental initiatives (Prague Symposium, 1971). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):132–149, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Robinson:2021:SIS

[1616] Sam Robinson. Scientific imaginaries and science diplomacy: The case of ocean exploitation. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):150–170, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rothman:2021:KEC

[1617] Aviva Rothman. Kepler’s epitome of Copernican astronomy in context. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):171–191, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Storni:2021:CGH

[1618] Marco Storni. Cartography, geodesy, and the heliocentric theory: Yves Simonin’s unpublished papers. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):192–209, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Charters:2021:HEE

[1619] Erica Charters and Kristin Heitman. How epidemics end. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):210–224, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Savoia:2021:VBE

[1620] Paolo Savoia. Virtual Bologna ESHS 2020 conference report. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):225–227, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2021:EOS

[1621] Anonymous. The ESHS Officers and the Scientific Council. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):228, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2021:Ca

[1622] Anonymous. Corrigendum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):229, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2021:Cb

[1623] Anonymous. Corrigendum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):230, February 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Anonymous:2021:E

[1624] Anonymous. Erratum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(1):231, February 2021. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rentetzi:2021:MCP

[1625] Maria Rentetzi and Kenji Ito. The material culture and politics of artifacts in nuclear diplomacy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):233–243, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Higuchi:2021:MIH

[1626] Toshihiro Higuchi and Jacques E. C. Hymans. Materialized internationalism: How the IAEA made the Vinca Dosimetry Experiment, and how the experiment made the IAEA. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):244–261, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Adamson:2021:OAF

[1627] Matthew Adamson. Orphaned atoms: The first Moroccan reactor and the frameworks of nuclear diplomacy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):262–276, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jacobsen:2021:SDO

[1628] Lif Lund Jacobsen, Irina Fedorova, and Julia Lajus. The seismograph as a diplomatic object: The Soviet–American exchange of instruments, 1958– 1964. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):277–295, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ito:2021:SOM

[1629] Kenji Ito. The scientific object and material diplomacy: The shipment of radioisotopes from the United States to Japan in 1950. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):296–319, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Florensa:2021:NMS

[1630] Clara Florensa. A nuclear monument the size of a football field: The diplomatic construction of soil nuclearity in the Palomares accident (Spain, 1966). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):320–338, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bardi:2021:SIC

[1631] Alberto Bardi. Scientific interactions in colonial, multilinguistic, and interreligious contexts: Venetian Crete and the manuscript Marcianus latinus VIII.31 (2614). A preliminary study. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):339– 352, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brannon:2021:SAA

[1632] James Brannon. The Sphere of Anthony Ascham: Sources for the earliest-known English-language cosmography based on Sacrobosco’s De Sphaera. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):353–380, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ducheyne:2021:RFE

[1633] Steffen Ducheyne and Jip van Besouw. Readers of the first edition of Newton’s Principia on the relation between gravity, matter, and divine and natural causation: British public debates, 1687–1713. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):381–395, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Elvbakken:2021:RKP

[1634] Kari Tove Elvbakken and Helle Margrete Meltzer. Research, knowledge, and policy on goitre and iodine in Norway (1850–2016). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):396–415, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Matei:2021:RTE

[1635] Oana Matei. Recipes and thrift in early modern and modern knowledge. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):416–420, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gysembergh:2021:BRA

[1636] Victor Gysembergh. Book review: Alan C. Bowen and Francesca Rochberg. Hellenistic astronomy: The science in its contexts, Brill’s Companions to Classical Studies, Leiden: Brill, 2020, xxxii + 751 pp. ISBN: 978-90-04-40056-6; 978-90-04-24336-1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2): 421–422, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Husson:2021:BRR

[1637] Matthieu Husson. Book review: Richard Dunn, Silke Ackermann, Giorgio Strano. Heaven and earth united: Instruments in astrological contexts. Leiden, The Netherlands: Brill, 2018, xiv + 286 pp. ISBN: 978-90-04-38143-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):423–424, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jesseph:2021:BRC

[1638] Douglas Jesseph. Book review: Crippa, Davide. The impossibility of squaring the circle in the 17th century: A debate among Gregory, Huygens and Leibniz. Basel, Switzerland: Birkhäuser, 2019, viii + 184 pp. ISBN: 978-3-030-01637-1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):424–427, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Henry:2021:BRA

[1639] John Henry. Book review: Ann Blair and Kaspar Greyerz. Physicotheology: Religion and science in Europe, 1650–1750. Baltimore, MD: Johns Hopkins University Press, 2020, ix + 274 pp. ISBN: 978-1-4214-3846-7. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):427–429, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Caudano:2021:BRE

[1640] Anne-Laurence Caudano. Book review: Edited by Efthymios Nicolaïdis, Greek alchemy from late antiquity to early modernity, Turnhout, Belgium: Brepols, 2018, 198 pp. + 21 col. ill. ISBN: 978-2-503-58191-0. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):430–431, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Carneiro:2021:BRE

[1641] Ana Carneiro. Book review: Edited by Annette Lykknes, Brigitte Tiggelen. Women in their element: Selected women’s contributions to the periodic system. Singapore, Singapore: World Scientific, 2019, xxiv + 531 pp. ISBN: 978-981-12-0628-3. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):431–433, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rayner-Canham:2021:BRS

[1642] Geoff Rayner-Canham. Book review: Scerri, Eric. The periodic table: Its story and its significance (2nd ed.). Oxford, UK: Oxford University Press, 2020, 504 pp. ISBN: 978-0-19-091436-3. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):433– 434, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gamez:2021:BRN

[1643] Santiago Guzmán Gámez. Book review: Nieto-Galan, Agustí. The politics of chemistry: Science and power in twentieth-century Spain. Cambridge, UK: Cambridge University Press, 2019, 284 pp. ISBN: 978-1-108-48243-1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):435–436, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Slater:2021:BRP

[1644] John Slater. Book review: Pimentel, Juan. Fantasmas de la ciencia española. Madrid, Spain: Marcial Pons Historia/Fundación Jorge Juan, 2020, 418 pp. ISBN: 978-84-17945-01-5. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):436– 438, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Berlivet:2021:BRE

[1645] Luc Berlivet. Book review: Edited by Anderson, Warwick, Roque, Ricardo, Santos, Ricardo Ventura. Luso-tropicalism and its discontents: The making and unmaking of racial exceptionalism. New York, NY: Berghahn, 2019, 316 pp. ISBN: 978-1-78920-113-0. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):438–440, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sanchez-Gonzalez:2021:BRJ

[1646] José Carlos Sánchez-González. Book review: Jenkins, Bill. Evolution before Darwin: Theories of the transmutation of species in Edinburgh, 1804–1834. Edinburgh, UK: Edinburgh University Press, 2019, 232 pp. ISBN: 978-1-4744-4578-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):440–442, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gaukroger:2021:BRJ

[1647] Stephen Gaukroger. Book review: Jürgen Renn. evolution of knowledge: Rethinking science for the Anthropocene. Princeton, NJ: Princeton University Press, 2020, 584 pp. ISBN: 978-0-691-17198-2. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):442–444, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Smith:2021:BRG

[1648] Roger Smith. Book review: Gaukroger, Stephen. Civilization and the culture of science: Science and the shaping of modernity, 1795–1935. Oxford, UK: Oxford University Press, 2020, 519 pp. ISBN: 978-0-19-884907-0. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):444–446, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Smith:2021:BRE

[1649] Bonnie G. Smith. Book review: Edited by Bittel Carla, Leong Elaine, von Oertzen Christine. Working with paper: Gendered practices in the history of knowledge. Pittsburgh, PA: University of Pittsburgh Press, 2019, 310 pp. ISBN: 978-0-8229-4559-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):446– 447, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Alvarez:2021:BRS

[1650] Teresa Álvarez. Book review: Smith, Roger. The sense of movement: An intellectual history. London, UK: Process Press, 2019, 444 pp. ISBN: 978-1-899209-24-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):448–449, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Smith:2021:BRM

[1651] Roger Smith. Book review: Mayer, Andreas. The science of walking: Investigations into locomotion in the long nineteenth century. Chicago, IL: University of Chicago Press, 2020, 232 pp. ISBN: 139780226328355; 139780226352480. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):449–451, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Simoes:2021:MEC

[1652] Ana Simões and Maria Paula Diogo. Mapping the ESHS community: A brief summary. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(2):452, May 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gooday:2021:SPT

[1653] Graeme Gooday. Silvanus Phillips Thompson (1851–1916): an introduction to the spotlight section. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):453–458, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gooday:2021:SMS

[1654] Graeme Gooday. “A many-sided crystal”: Understanding the manifold legacy of Silvanus Phillips Thompson (1851-1916). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63 (3):459–474, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cantor:2021:TB

[1655] Geoffrey Cantor. Thompson, biographer. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):475–488, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Stanley:2021:NSW

[1656] Matthew Stanley. No slaves to words: S. P. Thompson’s theory of history. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):489–498, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Arapostathis:2021:LBC

[1657] Stathis Arapostathis and Anna Guagnini. Living in between: the commercial side of Silvanus P. Thompson’s engineering. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63 (3):499–512, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Morris:2021:SJL

[1658] Andrew M. A. Morris. “The joint labours of ingenious men”: John Smeaton’s Royal Society network and the Eddystone Lighthouse. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):513–531, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermij:2021:SST

[1659] Rienk Vermij. Simon Stevin through the lens of his dedications. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):532–545, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ash:2021:HSC

[1660] Mitchell G. Ash. History of science in Central and Eastern Europe: Studies from Poland, Hungary, and Croatia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3): 546–552, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Duancic:2021:RTH

[1661] Vedran Duancić. Recent trends in the history of science in Croatia. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):553–568, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Surman:2021:PMH

[1662] Jan Surman. Productive marginalities: the history of science in/about Poland since 1989. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):569–584, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Zemplen:2021:HSH

[1663] Gábor Á. Zemplén. History of science in Hungary: Stewardship and audience in periods of institutional and political change. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):585–602, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bendegem:2021:BRB

[1664] Jean Paul Van Bendegem. Book review: De Bock, Dirk & Vanpaemel, Geert. Rods, sets and arrows: the rise and fall of modern mathematics in Belgium. New York, NY: Springer, 2019, xxii + 293 pp. ISBN: 978-3-030-20598-0; 978-3-030-20599-7. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):603–604, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Watson:2021:BRF

[1665] Katherine D. Watson. Book review: Francesco Paolo Ceglia. The body of evidence: Corpses and proofs in Early Modern European medicine. Leiden, The Netherlands: Brill, 2020, x + 355 pp. ISBN: 978-90-04-28481-4. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):604–606, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bryson:2021:BRS

[1666] Dennis Bryson. Book review: Solovey, Mark. Social science for what? Battles over public funding for the “other sciences” at the National Science Foundation. Cambridge, MA: MIT Press, 2020, 398 + x pp. ISBN: 978-0-262-53905-0. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):606–608, August 2021. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Franckowiak:2021:BRL

[1667] Rémi Franckowiak. Book review: Lawrence M. Principe. The transmutations of chymistry: Wilhelm Homberg and the Académie royale des sciences. Chicago, IL: The University of Chicago Press, 2020, 464 pp. ISBN: 978-0-226-70078-6. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):609–611, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rayner-Canham:2021:BRE

[1668] Geoff Rayner-Canham. Book review: Eric Scerri, Ghibaudi, Elena. What is a chemical element? A collection of essays by chemists, philosophers, historians, and educators. New York, NY: Oxford University Press, 2020, 312 pp. ISBN: 978-0-19-093378-4. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):611–613, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lykknes:2021:BRR

[1669] Annette Lykknes. Book review: Rayner-Canham, Geoff. The periodic table: Past, present, and future. Singapore: World Scientific, 2020, 312 pp. ISBN: 978-981-12-1848-4. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):613–614, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schotte:2021:BRA

[1670] Margaret Schotte. Book review: Adler, Antony. Neptune’s laboratory: Fantasy, fear, and science at sea. Cambridge, MA: Harvard University Press, 2019, 256 pp. ISBN: 978-0-674-97201-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63 (3):614–616, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Renn:2021:BRS

[1671] Jürgen Renn. Book review: Simões, Ana, Sousa, Ana Matilde. The global adventure of science: Einstein, Eddington and the eclipse. Lisbon, Portugal: Chili com Carne, 2019, 245 pp. ISBN: 978-989-8363-41-1. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(3):616–617, August 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2021:CPF

[1672] Koen Vermeir. Centaurus: Past and future. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):619–620, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Frey:2021:MPV

[1673] Felix Frey and Jonas Schädler. Making power visible: Codifications, infrastructures, and representations of energy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4): 621–630, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Russ:2021:DEF

[1674] Daniela Russ. Deciphering economic futures: Electricity, calculation, and the power economy, 1880–1930. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):631–650, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

vanderStraeten:2021:MPM

[1675] Jonas van der Straeten. Measuring progress in megawatt: Colonialism, development, and the “unseeing” electricity grid in East Africa. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):651–674, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hindelang:2021:OMC

[1676] Laura Hindelang. Oil media: Changing portraits of petroleum in visual culture between the US, Kuwait, and Switzerland. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4): 675–694, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hesse:2021:VWP

[1677] Nicole Hesse. Visible winds: The production of new visibilities of wind energy in West Germany, 1973–1991. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):695–713, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gessner:2021:TSV

[1678] Samuel Gessner. Trepidation spheres: Variant representations of the eighth sphere and the debate about the movement of the apogees and the fixed stars in Alfonsine astronomy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):714–754, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Recio:2021:EEK

[1679] Gonzalo Luis Recio. The Earth’s eccentricity in Kepler’s refutation of the Tychonic approach to the problem of Mars. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):755–761, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gantet:2021:DMG

[1680] Claire Gantet. The dissemination of mesmerism in Germany (1784– 1815): Some patterns of the circulation of knowledge. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):762–778, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Medori:2021:ASC

[1681] Beatriz Medori. The American Society for the Control of Cancer in the Portuguese Institute of Oncology’s Bulletin: Rethinking nationalism. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):779–803, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bevilacqua:2021:YHP

[1682] Fabio Bevilacqua and Salvatore Esposito. 40 years of history of physics in Italy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):804–806, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brack-Bernsen:2021:BRC

[1683] Lis Brack-Bernsen. Book review: Christine Proust and John Steele. Scholars and scholarship in Late Babylonian Uruk. Cham, Switzerland: Springer, Why the Sciences of the Ancient World Matter 2, 2019, x + 274 pp. ISBN: 978-3-030-04175-5. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):807–809, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wartenberg:2021:BRS

[1684] Ilana Wartenberg. Book review: Stern, Sacha. The Jewish calendar controversy of 921/2 CE. Leiden, Boston: Brill, 2019, xxii, 576 pp. ISBN: 978-90-04-38866-6. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):810–811, November 2021. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Burnett:2021:BRL

[1685] Charles Burnett. Book review: Lawrence-Mathers, Anne. Medieval meteorology: Forecasting the weather from Aristotle to the Almanac. Cambridge, UK: Cambridge University Press, 2020, 224 pp., ISBN: 978-1-108-40600-0. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):811–814, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Verwaal:2021:BRM

[1686] Ruben E. Verwaal. Book review: Margócsy, Dániel Somos, Mark and Joffe, Stephen N.. The Fabrica of Andreas Vesalius: A worldwide descriptive census, ownership, and annotations of the 1543 and 1555 editions. Leiden, The Netherlands: Brill, 2018, xix + 517 pp. ISBN: 978-90-04-33629-2. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):814–816, November 2021. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tho:2021:BRM

[1687] Tzuchien Tho. Book review: Mugnai, Massimo, Van Ruler, Han, Wilson, Martin. Leibniz: Dissertation on combinatorial art, Translated with introduction and commentary. Oxford, UK: Oxford University Press, 2020, 320 pp. ISBN: 978-0-19-883795-4. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):816–818, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Mantovani:2021:BRR

[1688] Mattia Mantovani. Book review: Raynaud, Dominique. Eye representation and ocular terminology from antiquity to Helmholtz. Amsterdam, The Netherlands: Wayenborgh. Hirschberg History of Ophthalmology — The Monographs. Volume 16, 2020, xvi + 636 pp. ISBN: 978-90-6299-468-7. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):819–822, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Naze:2021:BRA

[1689] Yaël Nazé. Book review: Aubin, David. Femmes, vulgarisation et pratique des sciences au siècle des Lumières: Les Dialogues sur l’astronomie et la Lettre sur la figure de la Terre de César–François Cassini de Thury. Turnhout: Brepols, 310 pp. ISBN: 978-2-503-58603-8. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4):822–823, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Simoes:2021:HRS

[1690] Ana Simões and Maria Paula Diogo. Historiographical reflections on sciences in Europe: Perspectives from Centaurus. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 63(4): 824–829, November 2021. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2022:CCO

[1691] Koen Vermeir and Theodore Arabatzis. Centaurus: Continuing as an open access journal. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Charters:2022:IEA

[1692] Erica Charters. Information, expertise, and authority: The many ends of epidemics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Linton:2022:WDE

[1693] Natalie M. Linton, Francesca Anne Lovell-Read, Emma Southall, Hyojung Lee, Andrei R. Akhmetzhanov, Robin N. Thompson, and Hiroshi Nishiura. When do epidemics end? Scientific insights from mathematical modelling studies. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Stenseth:2022:EPE

[1694] Nils Christian Stenseth, Katharine Rose Dean, and Barbara Bramanti. The end of plague in Europe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Jones:2022:LTW

[1695] Lori Jones. The last time that we can say the plague raged: Historicizing epidemics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Slack:2022:EPC

[1696] Paul Slack. End of a pandemic? Contemporary explanations for the end of plague in Eighteenth-Century England. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fang:2022:EPM

[1697] Xiaoping Fang. Ending pandemics in Mao’s China: Politics, technology, and epidemiology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Naono:2022:LME

[1698] Atsuko Naono. The last mile of elimination of malaria in Myanmar? Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Segata:2022:CBE

[1699] Jean Segata. Chikungunya in Brazil, an endless epidemic. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Berridge:2022:MER

[1700] Virginia Stewart Berridge. The many endings of recent epidemics: HIV/ AIDS, Swine Flu 2009 and policy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bastos:2022:NEP

[1701] Cristiana Bastos. The never-ending poxes of syphilis, AIDS, and measles. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gradmann:2022:EET

[1702] Christoph Gradmann. This is the end: Eradicating tuberculosis in modern times. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Larsen:2022:PBB

[1703] Clark Spencer Larsen and Fabian Crespo. Paleosyndemics: a bioarchaeological and biosocial approach to study infectious diseases in the past. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Webb:2022:HES

[1704] James L. A. Webb, Jr. Historical epidemiology and the single pathogen model of epidemic disease. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cohn:2022:EEB

[1705] Samuel Kline Cohn, Jr. Epidemics that end with a bang. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Kelton:2022:ESI

[1706] Paul Kelton. The end of smallpox for indigenous peoples in the United States, 1898–1903: an unnoticed finale. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pelling:2022:MEJ

[1707] Margaret Pelling. Mythological endings: John Snow (1813–1858) and the history of American epidemiology. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bailey:2022:EMT

[1708] Roderick Bailey. Ends and means: Typhus in Naples, 1943–44. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Rose:2022:CCE

[1709] Arthur Rose. Closure and the critical epidemic ending. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wigen:2022:MTE

[1710] Einar Wigen. The multiple temporalities of epidemic endings. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Apedaile:2022:BRJ

[1711] Dorothy Apedaile. Book review: Johan P. Mackenbach, A History of Population Health: Rise and Fall of Disease in Europe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Castro-Tejerina:2022:BRE

[1712] Jorge Castro-Tejerina. Book review: Erica Fretwell, Sensory Experiments: Psychophysics, Race and the Aesthetics of Feeling. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Morel:2022:BRU

[1713] Thomas Morel. Book review: Ursula Klein, Technoscience in History: Prussia, 1750–1850. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Omodeo:2022:BRA

[1714] Pietro Daniel Omodeo. Book review: Andrea Strazzoni, Dutch Cartesianism and the Birth of Philosophy of Science: From Regius to ’s Gravesande. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Prevost:2022:BRL

[1715] Jean-Guy Prévost. Book review: Laurent Mazliak and Rossana Tazzioli (eds), Mathematical Communities in the Reconstruction After the Great War 1918–1928: Trajectories and Institutions. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1): ??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Wisselgren:2022:BRK

[1716] Per Wisselgren. Book review: Kirsti Niskanen and Michael J. Barany (eds), Gender, Embodiment, and the History of the Scholarly Persona: Incarnations and Contestations. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(1):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Sander:2022:RMV

[1717] Christophe Sander. Rendering magnetism visible. diagrams and experiments between 1300 and 1700. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bardi:2022:CYH

[1718] Alberto Bardi. The cosmos in your hand. A note on Regiomontanus’ astrological interests. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Savoia:2022:KNS

[1719] Paolo Savoia. Knowing nature by its surface: Butchers, barbers, surgeons, gardeners and physicians in Early Modern Italy. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Carman:2022:TCC

[1720] Christián C. Carman. On Tycho’s calculation of the coordinates of Hamal, the fundamental star of Tycho’s Catalog. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

LiVigni:2022:HSF

[1721] Fabrizio Li Vigni. Hayek at the Santa Fe Institute. Origins, models and organization of the cradle of complexity sciences. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Brookhuis:2022:TBS

[1722] Hein Brookhuis. Transforming big science in Belgium management consultants and the reorganization of the Belgian Nuclear Research Center (SCK CEN) 1980–1990. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Krasnodebski:2022:LGC

[1723] Marcin Krasnodebski. Lost green chemistries: History of forgotten environmental trajectories. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Morales:2022:BRP

[1724] Angélica Morales. Book review: Paula S. De Vos. Compound Remedies: Galenic Pharmacy from the Ancient Mediterranean to New Spain. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Neufeld:2022:BRD

[1725] Michael J. Neufeld. Book review: David P. D. Munns and Kärin Nickelsen. Far Beyond the Moon: a History of Life Support Systems in the Space Age. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schechner:2022:BRJ

[1726] Sara J. Schechner. Book review: Josefina Rodríguez-Arribas, Charles Burnett, Silke Ackermann, Ryan Szpiech (eds). Astrolabes in Medieval Cultures. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Zaragoza:2022:BRR

[1727] Juan M. Zaragoza. Book review: Rita Felski and Stephen Muecke. Latour and the Humanities. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Barnes:2022:BRH

[1728] Joel Barnes. Book review: Hoffenberg,Peter H. A Science of our own: Exhibitions and the rise of Australian public science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(2):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hochadel:2022:SZI

[1729] Oliver Hochadel. Science at the zoo. An introduction. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(3):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cowie:2022:TTA

[1730] Helen Cowie. A tale of two anteaters: Madrid 1776 and London 1853. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(3):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Reiss:2022:KSM

[1731] Christian Reiß. Knowing is surviving. The Mexican axolotl, two Paris zoos and the role of practical knowledge in keeping unknown animals, 1850–1876. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(3):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hochadel:2022:HGP

[1732] Oliver Hochadel. How to get into the pouch. Solving the riddle of the kangaroo birth (1826–1926). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(3):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Larsson:2022:HTF

[1733] Elle Larsson. “Here they are in flesh and feather”: Walter Rothschild’s “private zoo” and the preparation and taxonomic study of cassowaries. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(3):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

McLeod:2022:HDH

[1734] Katherine McLeod. How to display a hoatzin: Animal care and conservation science in the early 20th Century. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(3):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pouillard:2022:ANA

[1735] Violette Pouillard. Animal nutrition, animal experiments, and the zoo as a laboratory, London Zoo — Paris Ménagerie, 19th century–ca. 2000. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(3):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Szczygielska:2022:UER

[1736] Marianna Szczygielska. Undoing extinction. The role of zoos in breeding back the tarpan wild horse, 1922–1965. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(3):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Carandell:2022:MDM

[1737] Miquel Carandell. Mediterranean dolphins from Miami: Knowledge and practices in the Barcelona Zoo “Aquarama” in the 1960s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(3):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Vermeir:2022:E

[1738] Koen Vermeir. Editorial. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Gavroglu:2022:SFH

[1739] Kostas Gavroglu. The Sisyphean fate of history of science: Unmoved scientists, unresponsive bureaucrats, unimpressed politicians (2020 Gustav Neuenschwander Prize Lecture). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Schaffer:2022:ODH

[1740] Simon Schaffer. Orientations and disorientations in the history of science: How measures made a difference at the Imperial Meridian (2022 Gustav Neuenschwander Prize Lecture). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VandenBroecke:2022:NTM

[1741] Steven Vanden Broecke. Negotiating theology and medicine in the Catholic Reformation. The early debate on Thomas Fienus’s embryology in the Spanish Netherlands (1620–1629). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4): ??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Prevignano:2022:PMD

[1742] Emma Prevignano. A passport for the metre: The diplomatic recognition of the metric system in a changing international order (1785–1799). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Katsiampoura:2022:ACS

[1743] Gianna Katsiampoura. An appraisal of the current status of research on Byzantine sciences: an essay review of Lazaris Stavros (ed.), A Companion to Byzantine Sciences. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VanDyck:2022:GCE

[1744] Maarten Van Dyck. Galilean challenges: an essay review of Jochen Büttner, Swinging and Rolling. Unveiling Galileo’s Unorthodox Path from a Challenging problem to a New Science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64 (4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Janiak:2022:WWN

[1745] Andrew Janiak. What was Newtonianism in Enlightenment Europe? An essay review of: Helmut Pulte and Scott Mandelbrote (eds), The Reception of Isaac Newton in Europe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cozzoli:2022:AMB

[1746] Daniele Cozzoli. The age of molecular biology: an essay review of Soraya de Chadarevian, Heredity Under the Microscope: Chromosomes and the Study of the Human Genome; Michel Morange, The Black Box of Biology: a History of the Molecular Revolution; Robin Wolfe Scheffler, A Contagious Cause: The American Hunt for Cancer Viruses and the Rise of Molecular Medicine. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Birn:2022:PHM

[1747] Anne-Emanuelle Birn. Pandemic histories: Making meaning or embeddig bromides? An essay review of: Juan Del Llano and Lino Camprubi, Sociedad Entre Pandemias; John Fabian Witt, American Contagions: Epidemics and the Law from Smallpox to COVID-19; Alex de Waal, New Pandemics, Old Politics: Two Hundred Years of War on Disease and its Alternatives; Frank M. Snowden, Epidemics and Society: From the Black Death to the Present. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Felten:2022:BRW

[1748] Sebastian Felten. Book review: Wolfgang Lefèvre, Minerva Meets Vulcan: Scientific and Technological Literature, 1450–1750. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Acosta:2022:BRR

[1749] Benedicto Acosta. Book review: Richard J. Oosterhoff, José Ramon Marcaida, and Alexander Marr (eds). Ingenuity in the Making: Matter and Technique in Early Modern Europe. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Spary:2022:BRE

[1750] Emma Spary. Book review: Elizabeth A. Williams, Appetite and Its Discontents: Science, Medicine, and the Urge to Eat, 1750–1950. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Fernandez-Garrido:2022:BRE

[1751] Sam Fernandez-Garrido. Book review: Elizabeth Reis, Bodies in Doubt. An American History of Intersex. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 64(4):??, ???? 2022. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Linte:2023:MUS

[1752] Guillaume Linte and Paul-Arthur Tortosa. ‘The most unhealthy spots in the world’: Thinking, dwelling in and shaping pathogenic environments. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(1):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Linte:2023:SSV

[1753] Guillaume Linte. “The salvation of the seamen”: Ventilation, naval hygiene and french overseas expansion during the Early Modern Period (c. 1670–1790). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(1):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tortosa:2023:ABF

[1754] Paul-Arthur Tortosa. Aetiologies of blame. fevers, environment, and accountability in a war context (France and Italy, c. 1800). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(1):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Omes:2023:ASC

[1755] Marco Emanuele Omes. “In aria sana”. Conceptualising pathogenic environments in the popular press: Northern Italy, the 1820s–1840s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(1):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bonelli:2023:STA

[1756] Costanza Bonelli. “Some typically African risks”: Safeguarding the health of Italian settlers during the Fascist Empire (1935–1941). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(1):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Delmaire:2023:LHH

[1757] Lea Delmaire. Locating the health hazard, surveilling the Gecekondu. The tuberculosis control pilot-area of Zeytinburnu, Istanbul (1961–1963). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(1):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Grunfeld:2023:CKT

[1758] Martin Grünfeld and Karin Tybjerg. Collections, knowledge and time. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(2):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Tybjerg:2023:AMC

[1759] Karin Tybjerg. The anamnesis of medical collections: The role of past collections in diagnosis and prognosis. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(2):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Nall:2023:LMU

[1760] Joshua Nall and Boris Jardine. The lab in the museum, or: Using new scientific instruments to look at old scientific instruments. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(2):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

VanAllen:2023:ETC

[1761] Adrian Van Allen. Entangled timelines: Crafting types of time through making museum specimens. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(2):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Keck:2023:FCG

[1762] Frédéric Keck. Filling China’s gaps: Viral banks and bird collections as museums for pandemics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(2):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chacko:2023:LSC

[1763] Xan Chacko and Jenny Bangham. Lively stasis: Care and routine in living collections of flies and seeds. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(2):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Beltrame:2023:MDI

[1764] Tiziana Nicoletta Beltrame. A matter of dust: From infrastructures to ‘infra-thin’ in museum maintenance. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(2):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Grunfeld:2023:MLR

[1765] Martin Grünfeld, Adam Bencard, and Louise Whiteley. From mausoleum to living room: Practicing metabolic carpentry in the museum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(2):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Arnold:2023:ERP

[1766] Ken Arnold. Epilogue: Redeeming the past, present and future. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(2):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Podgorny:2023:I

[1767] Irina Podgorny and Nathalie Richard. Introduction. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(3):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Espejel:2023:NPG

[1768] Claudia Espejel. Nation in pieces. The gathering of an archaeological collection in Late Nineteenth-century Mexico: The first collection of Francisco Plancarte y Navarrete. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(3):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Achim:2023:COM

[1769] Miruna Achim. Carving an origin for Mexico’s ancient cultures: Jade artifacts and the question of their provenance in Nineteenth-Century science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(3):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Bertran:2023:PPV

[1770] Xavier Ulled Bertran and José Pardo-Tomás. From private to public (and vice versa). Scientific collections in Barcelona, 1830–1880. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(3):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lopes:2023:OCM

[1771] Maria Margaret Lopes. Objects and contradictions on the move: From private collections to provincial Brazilian museums. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65 (3):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Reubi:2023:HDO

[1772] Serge Reubi. How do objects enter and exit collections? Exchanging material culture over the Atlantic, 1920–1940. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(3): ??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chazaro-Garcia:2023:HDA

[1773] Laura Chazaro-Garcia. How do anatomies become exhibition objects? From provincial museums to (inter)national collections. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(3):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Garcia:2023:SAL

[1774] Susana Valeria García. South American landowners, herbaria and scientific sciability networks (1880s–1910s). Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(3):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ganger:2023:FMW

[1775] Stefanie Gänger. ‘The finest in any museum in the world’. Collecting preconquest antiquities in the Southern Andes, c. 1850–1911. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(3):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Corsi:2023:NLB

[1776] Pietro Corsi. Of names, labels and books in Nineteenth-Century Tuscan paleontological collections. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 65(3):??, ???? 2023. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ackermann:2014:SID

[1777] Silke Ackermann, Richard L. (Richard Lynn) Kremer, and Mara Miniati, editors. Scientific instruments on display, volume 46 / 4 of History of science and medicine library; Scientific instruments and collections. Brill, Leiden, The Netherlands, 2014. ISBN 90-04-26439-6 (hardcover), 90-04-26440-X (e-book). ISSN 1872-0684. xxxiv + 231 pp. LCCN Q185 .S39 2014.

Appleby:2013:SKN

[1778] Joyce Oldham Appleby. Shores of knowledge: new world discoveries and the scientific imagination. W. W. Norton & Co., New York, NY, USA, 2013. ISBN 0-393-23951-9 (hardcover). 308 pp. LCCN E101 .A66 2013.

Arianrhod:2012:SLE

[1779] Robyn Arianrhod. Seduced by logic: Émilie du Châtelet, Mary Somerville, and the Newtonian revolution. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-993161-5, 0-19-993162-3 (e-book). 338 pp. LCCN Q141 .A725 2012.

Ball:2013:SRS

[1780] Philip Ball. Serving the Reich: the struggle for the soul of science under Hitler. The Bodley Head Ltd, London, UK, 2013. ISBN 1-84792-248-1. ix + 303 pp. LCCN ????

Boantza:2013:MML

[1781] Victor D. Boantza. Matter and method in the long chemical revolution: laws of another order. Science, technology and culture. Ashgate, Farnham, Surrey, UK, 2013. ISBN 1-4094-1867-7 (hardcover), 1-4094-1868-5 (e-book), 1-4724-0398-3 (e-pub). ISSN 1700-1945. xiv + 266 pp. LCCN QD14 .B63 2013.

Boner:2013:KCS

[1782] Patrick Boner. Kepler’s cosmological synthesis: astrology, mechanism and the soul, volume 39 of History of science and medicine library, Medieval and early modern science (volume 20). Brill, Leiden, The Netherlands, 2013. ISBN 90-04-24608-8 (hardcover), 90-04-24609-6 (e-book). ISSN 1872-0684. x + 187 pp. LCCN QB36.K4 .B638 2013.

Brock:2016:HCV

[1783] W. H. (William Hodson) Brock. The History of Chemistry: a Very Short Introduction, volume 456 of Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-871648-6 (paperback). xvi + 151 pp. LCCN QD11 .B758 2016.

Buckland:2013:NSF

[1784] Adelene Buckland. Novel science: fiction and the invention of nineteenthcentury geology. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-07968-6 (hardcover), 0-226-92363-0 (e-book). 377 pp. LCCN PR468.G38 B83 2013.

Camprubi:2014:EMF

[1785] Lino Camprubí. Engineers and the making of the Francoist regime. Transformations: Studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02717-8 (hardcover). xiii + 298 pp. LCCN TA87 .C35 2014.

Carson:2010:HAA

[1786] Cathryn Carson. Heisenberg in the atomic age: science and the public sphere. Publications of the German Historical Institute. German Historical Institute, Washington, DC, USA, 2010. ISBN 0-521-82170-3 (hardcover). xvi + 541 + 8 pp. LCCN QC16.H45 C37 2010.

Chapman:2014:MSW

[1787] Allan Chapman. Mary Somerville and the world of science. Springerbriefs in history of science and technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-09398-3, 3-319-09399-1 (e-book). ???? pp. LCCN Q143.S7. URL http:

//www.loc.gov/catdir/enhancements/fy1412/2014947132-d.html;
 http://www.loc.gov/catdir/enhancements/fy1412/2014947132-t.
html.

Coen:2013:EOD

[1788] Deborah R. Coen. The earthquake observers: disaster science from Lisbon to Richter. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-11181-4 (paperback), 0-226-11183-0 (e-book). 348 pp. LCCN QE539 .C64 2013.

Daipha:2015:MUW

[1789] Phaedra Daipha. Masters of uncertainty: weather forecasters and the quest for ground truth. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-29854-X (hardcover), 0-226-29868-X (paperback). 271 pp. LCCN QC983 .D26 2015.

Drucker:2014:CSA

[1790] Donna J. Drucker. The classification of sex: Alfred Kinsey and the organization of knowledge. University of Pittsburgh Press, Pittsburgh, PA, USA, 2014. ISBN 0-8229-6303-5. ix + 244 pp. LCCN HQ18.32.K56.

Erickson:2013:HRA

[1791] Paul Erickson, Judy L. Klein, Lorraine Daston, Rebecca M. (Rebecca Maura) Lemov, Thomas Sturm, and Michael D. Gordin. How reason almost lost its mind: the strange career of Cold War rationality. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-04663-X (hardcover), 0-226-04677-X (e-book). 259 pp. LCCN D843 .E69 2013.

Finkelstein:2013:EDB

[1792] Gabriel Ward Finkelstein. Emil du Bois-Reymond: neuroscience, self, and society in Nineteenth-Century Germany. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-01950-7 (hardcover). xviii + 362 pp. LCCN QP26.D8

F56 2013. URL http://muse.jhu.edu/books/9780262314848/.

Fors:2015:LMC

[1793] Hjalmar Fors. The limits of matter: chemistry, mining, and Enlightenment. Synthesis. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-19499-X, 0-226-19504-X (e-book). LCCN BD646 .F67 2015.

URL http://public.eblib.com/choice/publicfullrecord.aspx?p=
1835294;   http://site.ebrary.com/id/10984129.

Forstner:2007:QIK

[1794] Christian Forstner. Quantenmechanik im Kalten Krieg: David Bohm und Richard Feynman. (German) [Quantum mechanics in the Cold War: David Bohm and Richard Feynman]. Verl. für Geschichte der Naturwiss. und der Technik, Diepholz, Germany, 2007. ISBN 3-928186-81-7 (paperback). 244 pp. LCCN ???? EUR 30.00EUR 30.00. URL http://paperc.

de/4165-quantenmechanik-im-kalten-krieg-9783928186810.

Fox:2016:SFC

[1795] Robert Fox. Science Without Frontiers: Cosmopolitanism and National Interests in the World of Learning, 1870–1940. The OSU Press Horning Visiting Scholars publication series. Oregon State University Press, Corvallis, OR, USA, 2016. ISBN 0-87071-867-3 (paperback). xvi + 160 pp. LCCN Q223 .F698 2016.

Gal:2013:BS

[1796] Ofer Gal and Raz Chen-Morris. Baroque science. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-92398-3 (hardcover), 0-226-92399-1 (e-book). xiv + 333 pp. LCCN Q127.E85 G35 2013.

Greene:2014:GUM

[1797] Jeremy A. Greene. Generic: the unbranding of modern medicine. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1493-7 (hardcover), 1-4214-1494-5 (electronic). xii + 354 pp. LCCN RM301 .G74 2014.

Harman:2013:OSR

[1798] Oren Solomon Harman and Michael R. Dietrich, editors. Outsider scientists: routes to innovation in biology. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-07837-X (hardcover), 0-226-07840-X (paperback), 0-226-07854-X (e-book). 374 pp. LCCN QH26 .O98 2013.

Hayton:2015:CCA

[1799] Darin Hayton. The crown and the cosmos: astrology and the politics of Maximilian I. University of Pittsburgh Press, Pittsburgh, PA, USA, 2015. ISBN 0-8229-4443-X. xiii + 312 pp. LCCN BF1676 .H39 2015.

Hoffmann:2013:EBF

[1800] Dieter Hoffmann. Einstein’s Berlin: in the footsteps of a genius. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-1040-0 (paperback). xiii + 175 pp. LCCN QC16.E5 H64413 2013.

Jackson:2015:GGP

[1801] Myles W. Jackson. The genealogy of a gene: patents, HIV/AIDS, and race. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2015. ISBN 0-262-02866-2 (hardcover). xii + 336 pp. LCCN RA643.8 .J34 2015.

Kargon:2015:WFE

[1802] Robert H. (Robert Hugh) Kargon, Karen Fiss, Morris Low, and Arthur P. Molella. World’s fairs on the eve of war: science technology and modernity, 1937–1942. University of Pittsburgh Press, Pittsburgh, PA, USA, 2015. ISBN 0-8229-4444-8 (hardcover). viii + 206 pp. LCCN T395 .K37

Kisacky:2017:RMH

[1803] Jeanne Susan Kisacky. Rise of the Modern Hospital: an Architectural History of Health and Healing, 1870–1940. University of Pittsburgh Press, Pittsburgh, PA, USA, 2017. ISBN 0-8229-4461-8 (hardcover). vii + 448 pp. LCCN RA967 .K53 2017.

Krischel:2014:UNS

[1804] M. Krischel. Urologie und Nationalsozialismus: eine Studie zu w Medizin und Politik als Ressourcen füreinander, volume 1 of Schriften zur Urologiegeschichte. Franz Steiner Verlag, Stuttgart, Germany, 2014. ISBN 3-515-10849-1 (paperback). 220 pp. LCCN RC870.8 .K75 2014.

Lawlor:2012:MPH

[1805] Clark Lawlor. From melancholia to prozac: a history of depression. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-958579-2 (hardcover). xi + 265 pp. LCCN RC537 .L392 2012.

Lehoux:2012:WDR

[1806] Daryn Lehoux. What did the Romans know?: an inquiry into science and worldmaking. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-47114-4 (hardcover). xi + 275 pp. LCCN Q124.95 .L44 2012.

Malcolmson:2013:SSC

[1807] Cristina Malcolmson. Studies of skin color in the early Royal Society: Boyle, Cavendish, Swift. Literary and scientific cultures of early modernity. Ashgate, Farnham, Surrey, UK, 2013. ISBN 0-7546-3778-6. xii + 233 pp. LCCN Q41.L86.

Margocsy:2014:CVS

[1808] Dániel Margócsy. Commercial visions: science, trade, and visual culture in the Dutch Golden Age. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-11774-X (hardcover), 0-226-11788-X (e-book). x + 319 + 32 pp. LCCN Q127.N2 M37 2014.

Munns:2017:EEP

[1809] David P. D. Munns. Engineering the Environment: Phytotrons and the Quest for Climate Control in the Cold War. University of Pittsburgh Press, Pittsburgh, PA, USA, 2017. ISBN 0-8229-4474-X (hardcover), 0-8229-8276-5 (e-book). xxiii + 334 pp. LCCN QK715.5 .M86 2017. URL

http://www.jstor.org/stable/10.2307/j.ctt1r33q5x.

Nielsen:2012:SSF

[1810] Kristian H. Nielsen, Michael Harbsmeier, and Christopher J. Ries. Scientists and scholars in the field: studies in the history of fieldwork and expeditions. Aarhus University Press, Aarhus, 2012. ISBN 87-7124-014-4. 476 pp. LCCN Q181 .S385 2012.

Nye:2011:MPH

[1811] Mary Jo Nye. Michael Polanyi and his generation: origins of the social construction of science. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-61063-2 (hardcover). xxi + 405 pp. LCCN B945.P584 N94 2011.

Pickering:2010:CBS

[1812] Andrew Pickering. The cybernetic brain: sketches of another future. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-66789-8 (hardcover). x + 526 pp. LCCN Q310 .P53 2010.

Powers:2012:ICH

[1813] John C. Powers. Inventing chemistry: Herman Boerhaave and the reform of the chemical arts. Synthesis. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-67760-5, 1-280-12635-3, 0-226-67762-1 (e-book). viii + 260 pp. LCCN QD15 .P69 2012eb.

URL http://public.eblib.com/EBLPublic/PublicView.do?ptiID=
867815;   http://search.ebscohost.com/login.aspx?direct=true\
%26scope=site\%26db=nlebk\%26db=nlabk\%26AN=438283;   http:
//site.ebrary.com/id/10537830;   http://site.ebrary.com/lib/
alltitles/Doc?id=10537830;   http://www.myilibrary.com?id=353021;
 http://www.myilibrary.com?id=353021\%26ref=toc.

Riordan:2015:TVR

[1814] Michael Riordan, Lillian Hoddeson, and Adrienne W. Kolb. Tunnel visions: the rise and fall of the Superconducting Super Collider. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-29479-X (hardcover), 0-226-30583-X (e-book). xiii + 448 pp. LCCN QC787.S83 .R56 2015.

Rossiter:2012:WSA

[1815] Margaret W. Rossiter. Women scientists in America: forging a new world since 1972. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0233-5 (hardcover), 1-4214-0363-3 (paperback). xx + 426 pp. LCCN Q130 .R6835 2012.

Ruse:2013:GHS

[1816] Michael Ruse. The Gaia Hypothesis: science on a pagan planet. Science.culture. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-73170-7 (hardcover), 0-226-06039-X (e-book). xiv + 251 pp. LCCN QH331 .R8785 2013.

Schwartz:2012:WWR

[1817] Janelle A. Schwartz. Worm work: recasting Romanticism. University of Minnesota Press, Minneapolis, MN, USA, 2012. ISBN 0-8166-7320-9 (hardcover), 0-8166-7321-7 (paperback). xxv + 277 pp. LCCN PN603 .S27 2012.

Smith:2015:SLP

[1818] Albert Mark Smith. From sight to light: the passage from ancient to modern optics. University of Chicago Press, Chicago, IL, USA, onlineausg. edition, 2015. ISBN 0-226-17476-X (hardcover), 0-226-17493-X (e-book). xi + 457 pp. LCCN QC352. URL http://site.ebrary.com/

lib/alltitles/Doc?id=10985877.

Sorlin:2013:SGC

[1819] Sverker Sörlin, editor. Science, geopolitics and culture in the polar region: Norden beyond borders. The Nordic experience. Ashgate, Farnham, Surrey, UK, 2013. ISBN 1-4724-0969-8 (hardcover), 1-4724-0970-1 (e-book), 1-4724-0971-X (e-pub). xiv + 443 pp. LCCN G877 .S36 2013.

Terrall:2014:CNA

[1820] Mary Terrall. Catching nature in the act: Réaumur and the practice of natural history in the Eighteenth Century. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-08860-X (hardcover), 0-226-08874-X (e-book). 275 pp. LCCN QH15 .T47 2014.

Tresch:2012:RMU

[1821] John Tresch. The romantic machine: utopian science and technology after Napoleon. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-81220-0. 449 pp. LCCN T26.F8 T74 2012.

Truitt:2015:MRM

[1822] Elly Rachel Truitt. Medieval Robots: Mechanism, Magic, Nature, and Art. The Middle Ages series. University of Pennsylvania Press, Philadelphia, PA, USA, 2015. ISBN 0-8122-4697-7, 0-8122-2357-8. x + 255 + 24 pp. LCCN TJ211 .T74 2015.

Turchetti:2012:PAC

[1823] Simone Turchetti. The Pontecorvo affair: a cold war defection and nuclear physics. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-81664-8 (hardcover). 292 pp. LCCN QC774.P66 T8713 2012.

Vertesi:2015:SLR

[1824] Janet Vertesi. Seeing like a Rover: how robots, teams, and images craft knowledge of Mars. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-15596-X (hardcover), 0-226-15601-X (e-book). xi + 318 pp. LCCN TL799.M3 V47 2015.

Vidal:2011:SSE

[1825] Fernando Vidal and Saskia Brown. The sciences of the soul: the early modern origins of psychology. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-85586-4. xiv + 413 pp. LCCN BF101 .V5313 2011.

Westman:2011:CQP

[1826] Robert S. Westman. The Copernican question: prognostication, skepticism, and celestial order. University of California Press, Berkeley, CA, USA, 2011. ISBN 0-520-25481-3 (hardcover). xviii + 681 pp. LCCN QB29 .W47 2011.

Wilding:2014:GIG

[1827] Nick Wilding. Galileo’s idol: Gianfrancesco Sagredo and the politics of knowledge. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-16697-X (hardcover). 200 pp. LCCN Q143.S17 W55 2014.

Yalcinkaya:2015:LPD

[1828] M. Alper Yalçınkaya. Learned patriots: debating science, state, and society in the nineteenth-century Ottoman Empire. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-18420-X (hardcover), 0-226-18434-X (e-book). 314 pp. LCCN Q175.52.T9 Y35 2015.

Lykknes:2012:BWC

[1829] Annette Lykknes, Donald Opitz, and Brigitte van Tiggelen, editors. For better or for worse?: collaborative couples in the sciences, volume 44 of Science networks. Historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2012. ISBN 3-0348-0285-4 (hardcover), 3-0348-0286-2 (e-book). xiv + 319 pp. LCCN Q125 .F5935 2012. URL http://www.loc.gov/

catdir/enhancements/fy1304/2012940246-d.html;   http://www.
loc.gov/catdir/enhancements/fy1304/2012940246-t.html.

Schot:2010:TMN

[1830] Johannes Willem Schot, Arie Rip, and Harry Lintsen, editors. Technology and the making of The Netherlands: the age of contested modernization, 1890–1970. Walburg Pers, Zutphen, The Netherlands, 2010. ISBN 0-262-01362-2 (MIT Press), 90-5730-630-1 (Walburg Pers). 635 pp. LCCN

???? URL http://catalogue.bnf.fr/ark:/12148/cb422097099.